A Complete Bibliography of Publications in Isis, 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.15

Title word cross-reference

⊃ [521].

1 [511]. 1050 [362]. 10th [521]. 11th [1186, 521]. 125th [737]. 1350 [1250]. 1485 [566]. 14th [1409]. 1524 [1554]. 1528 [1484]. 1537 [660]. 1561 [794]. 15th [245]. 1600 [983, 1526, 261]. 1617 [528]. 1632 [805]. 1643 [1058]. 1645 [1776]. 1650 [864]. 1660 [1361]. 1671 [372]. 1672 [1654]. 1674 [1654]. 1675 [88]. 1680 [889]. 1687 [1147]. 1691 [1148]. 1692 [888, 371]. 1695 [296]. 16th [1823]. 1700 [864]. 1700-talets [890]. 1704 [476]. 1708 [265]. 1713 [1415]. 1733 [756]. 1741 [1494]. 1751 [1197]. 1760 [1258]. 1774 [1558]. 1777 [1909, 572]. 1780 [314, 663]. 1792 [269]. 1794 [266]. 1796 [1195, 840]. 1799 [128]. 1799/1804 [128]. 17th [1256, 623, 1813]. 1800 [1641, 100, 1343, 1044, 1655, 248, 1331]. 1802 [127, 437]. 1803 [405, 1778]. 1804 [128]. 1807 [625]. 1814 [668]. 1815 [1777]. 1820 [1660]. 1826 [1857]. 1832 [668]. 1841 [1362]. 1844 [1913, 946]. 1848 [1708]. 185 [1327]. 1850 [1230, 1391]. 1855 [442]. 1860 [301, 1232, 1917, 1367]. 1865 [445, 1263].

1866 [253, 71]. 1868 [1019]. 1870’s [674]. 1875 [1364]. 1878 [25]. 1880 [1427, 807, 1894]. 1882 [381]. 1889 [1428]. 1893 [1588]. 1894 [1921]. 1895 [896]. 1897 [762]. 18th [1013, 1256, 623]. 1900 [170, 692, 649, 1047, 28, 1423, 1500, 1421, 482, 1004, 348, 787]. 1904 [1785]. 1905 [1199, 1565]. 1906 [470]. 1907 [1786]. 1909 [105]. 1910 [132, 1855]. 1911 [1623]. 1912 [1569, 1812]. 1913 [1125, 689, 1395]. 1914 [1924, 1202, 1282, 1154, 1068, 1795]. 1915 [989]. 1917 [1313]. 1919 [669]. 1920 [1748, 1619, 1849, 1862]. 1922 [1069]. 1925 [1310]. 1929 [1431]. 1930 [632, 490, 898]. 1932 [676]. 1936 [1691]. 1940 [447]. 1941 [1860]. 1945 [324]. 1947 [1024]. 1947/1952 [107]. 1950 [611, 949]. 1955 [134, 153, 636]. 1960 [305, 41]. 1962 [1071]. 1965 [581, 328]. 1966 [526]. 1968 [322]. 1969 [1226, 312, 17, 883, 978, 1249, 620, 741, 759]. 1970 [1431, 847, 637, 1119, 250, 583]. 1971 [403, 1057, 703, 847, 1098, 952]. 1972 [1227]. 1973 [961, 999, 1332, 1010]. 1974 [1781, 1087, 1430]. 1975 [1324, 1281, 1466, 1774, 1375]. 1976 [1388, 1168, 1886, 1508, 1605, 1645, 1405, 1574]. 1977 [1680, 1521, 1520, 1809, 1450, 1732]. 19th [1823, 673, 760, 1498, 245, 8]. 19th-Century [760]. 1er [1142].

2 [1799]. 200 [1589]. 20th [1590]. 227 [1451].

3 [972]. 37 [375].

4002 [1059].

500th [1010]. 50th [1574].

65 [922].

80 [1937, 1936].

93e [695].

êtres [1063].

A-Z [702]. A.5.2 [258]. A.D [864, 946, 520, 794]. A.D. [1327, 520]. A13 [1900]. Aaron [846]. Åsberg [1494]. Abū [1522]. Abū’l [778]. Abano [79]. ’Abd [259, 546]. Abdus [771]. Abraham [75]. Abrahams [717]. abril [741]. Absorption [290]. abstrakten [106]. Abu [364, 1903, 521, 1649]. Abyss [1738]. ac [750]. Academic [1560, 167]. Academies [1256]. Academy [1815, 312, 405, 1783, 1457]. Accelerated [286]. Accelerators [228]. accidentium [1761]. Accomplishment [1050]. Account [544, 1238, 269, 115]. Accounts [251]. Accuracy [1323, 1325]. Acheson [584]. Achievement [982, 1311]. Achievements [1622]. Achilles [260, 1299].

Achinstein [59, 739]. Acidity [588, 1079]. Ackerknecht [357, 218]. Ackermann [1333]. Acontius [750]. Acorn [1778]. Acosta [1702]. Acoustics [1740]. Actas [741]. Actes [1142, 550]. Actinium [201]. Action [115, 721]. Active [855]. Activist [502]. Ad [532, 1490]. Adair [1750]. Adam [891]. Adams [1229]. Adaptation [487]. Add [1059]. Addison [630]. Addresses [1119, 618]. Adelaide [844]. Adelin [535]. Adelmann [1557]. Adept [1050]. Adherents [1635]. Adjacent [756]. Adkins [793]. Administration [909]. Administrator [1160]. Adolf [1021]. Adriaan [782]. Adrian [69]. Advancement [224, 1367, 1119, 1171]. Advances [662, 1643]. Advent [1817]. Adventure [951]. Adventures [1456, 303, 581]. Adversary [1132]. Adviser [387]. Advisor [395]. Advisors [1266]. Advocacy [1199]. Aepinus [334, 1033]. Aerial [1164]. Aeronautics [909]. Aesthetic [940]. Aesthetics [1821]. Aether [632, 579]. Aether-Drift [632]. af [1741]. Affair [277]. Affinity [445]. Affirmation [596]. Africa [871]. African [871, 627, 751]. Africanus [627]. after [540, 864, 858, 1445, 1779]. against [1412, 417, 1402, 367]. Agassi [1591, 436]. Agassiz [36, 1118, 1582]. Age [402, 1150, 192, 1740, 1614, 1012, 1852, 921, 439, 601]. Agents [1721, 1347]. Ages [1898, 1631, 1181, 1297, 982, 1609, 31, 742, 354, 424, 1475]. Agostino [1496]. Agricultural [1427]. Agriculture [1428, 1600]. Agrimensores [706]. agronomes [156]. Agronomie [156]. Ahmad [933, 1479, 1105, 1903, 1649]. Aichi [1106]. Ainsworth [1399]. Air [1172]. Airship [1185]. Aitken [1404]. Akadémia [1457]. Akademie [1714, 1929, 1617, 1558, 1430]. Akademiemitglied [1857]. akademii [1256]. Akademiya [1783]. Akhundov [923]. Akten [1414]. Al [1460, 1252, 1903, 259, 1186, 1479, 1761, 521, 1497, 1551, 1409, 1251, 409, 1649, 1252, 1810, 234, 732, 546, 1903, 259, 309, 564, 933, 1252, 748, 1105, 1054]. al-Amakin [1186]. al-a’rad [1761]. al-Baġdādī’s [546]. al-Bagdadi’s [259]. Al-Bahir [933, 1105]. al-Biruni [1649, 1054]. Al-Bitruji [748]. al-Buzajani [521]. al-Dīn [234, 732]. al-Fārābī [1810]. Al-Fusul [1252]. al-Haitham [309]. al-handasiyya [1251]. al-Hasan [1903]. al-Hindi [1252]. al-hiyal [1251]. al-Jazari [1251]. al-Karaji [521]. Al-Kindi [1460]. al-Latif [259, 546]. al-Quff [564]. al-Rayhan [1649]. al-Razzaz [1251]. al-Shātir [732]. al-Shatir [1409]. al-Tūsī [234, 732]. al-Uqlidisi [1252, 1903, 1252, 1903]. al-Wafa [521]. Alain [1416]. Alan [1307, 1532, 1685, 72, 605]. Alani [747]. Albani [1703]. Albani-Boncompagni [1703]. Albany [919]. Albert [226, 1884, 193, 1800, 74, 134, 720, 1206, 636, 634, 1476, 438, 34, 36, 1063, 271, 1154, 990, 1231, 1676, 729, 723]. Alberti [1878]. Alberto [1647]. Albion [1260]. Albrecht [1853, 1477, 776, 1704, 1909, 806]. Albucasis [1298]. Album [1859]. Alchemical [1050]. alchemistische [1899]. alchemistischen [1187, 1482]. Alchemists [1606, 176]. Alchemy [1050, 32, 1163, 1225, 1606]. Alchimie [200]. Alcmaeon [836]. Alderotti [1213]. Aldgate [310]. Aldrovandi [1555, 1555]. Aleksander [694]. Aleksandrovich [55]. Alexander [1657, 1244, 1473, 1295, 99, 1313, 358, 1875, 1363, 1420].

Alexandra [1286]. Alexandre [873, 128]. Alexandria [307, 168, 1902]. Alfred [1490, 492, 70, 578, 728, 504, 327, 70, 905]. Alfredo [132]. Algebra [644, 1867, 1841, 1112, 1094, 29]. algèbre [1396, 1105, 933]. Alger [388]. Alice [384]. Alienation [699]. Aliis [1442]. Alimberto [1412]. Alkalis [687]. Alkindi [233]. alla [1846, 1485, 1301, 1910]. Allamand [858, 858]. Allan [298, 1475, 1706]. állandó [1457]. Allaud [1661]. alle [1878, 1703, 936]. Allemagne [1488]. Allen [652, 624, 1833, 1048, 1395, 1346, 1011, 1357, 633, 1875, 1733, 471, 1418]. Allgemeine [1607]. allgemeinverständlichen [1413]. Allibone [988]. Allied [635, 363, 1540]. Alma [884]. Almagest [1480, 1245]. Almanus [566]. Almost [1862]. Alois [750]. Alpha [1036]. Alphonso [685]. Alsopp [219]. alten [1293]. älteren [835]. ältesten [1548]. altorientalischen [1182]. Altschule [1538]. Alwayn [492]. am [8]. Amakin [1186]. Amann [1687]. Amazons [382]. Ambiguous [1274]. Ambrose [1634]. America [667, 355, 71, 443, 18, 100, 756, 807, 1917, 1569, 343, 760, 41, 1462, 1660, 1895, 1443, 1382, 1444, 816, 159, 880, 1008, 1470, 222, 1055, 1663, 930, 1916, 1699, 1362]. American [1069, 1199, 1310, 339, 1286, 490, 1204, 370, 1771, 525, 797, 1302, 1778, 1568, 1367, 266, 1643, 1119, 1200, 1156, 1646, 1009, 668, 1179, 1614, 697, 613, 1751, 118, 1341, 1711, 1118, 1501, 1418, 1862, 1860, 328, 1540, 502, 402, 141, 1037, 777, 589, 170, 542, 1322, 1341, 743, 673, 1855, 1419, 1923]. Americanarum [568]. Americans [1427]. Amerika [1066]. amerikanischer [170]. Ames [581]. Ami [1193]. Ammines [1320]. Ammonium [504]. Amos [1774]. Ampère [1635]. Amprère [927]. Anaesthetic [1697]. Anaesthetists [1697]. Análisis [1355]. Analogy [916]. analyse [1108]. Analysen [1548]. Analyses [1518]. Analysis [1069, 247, 1288, 1052, 165, 1823, 21, 950, 336, 47, 974, 1272, 942]. Analytical [819]. analytique [1615, 432]. anatomía [1707]. Anatomical [640]. Anatomie [564]. Anatomy [405, 1678]. anciens [1731]. Ancient [1596, 1604, 614, 701, 1756, 1542, 608, 1132, 975, 1757, 422, 1050, 1829, 1103, 836, 96, 1677, 1869, 491, 1051, 834, 1327, 7, 1395, 1347]. Ancients [833]. Anderson [1692]. Andre [375, 561, 1615, 156, 622]. André-Félix [622]. Andreae [1191]. Andreas [1413]. Andrew [583, 878]. Andrews [1206, 1023]. Andronikos [1473]. Anfänge [621]. Anfängen [1097]. Angelo [936]. Angle [187, 1487]. Anglo [1722]. Anglo-Saxon [1722]. Anglo-Saxons [1722]. angulaire [1306]. Angus [1920]. Anha [1761]. Animal [1739, 527, 1169, 1230, 1729, 48, 39, 1374]. animala [1146]. animalium [1897, 1355]. Animals [661, 751, 795, 792]. Ann [1425, 1158]. Anna [839]. années [1058]. Annette [1892, 622]. Anniversary [1574, 995, 737]. Anniversay [1010]. Annotated [1290, 132, 1610, 1561, 187, 1711, 648, 305]. Annual [1644, 1605, 1223, 1449]. Annus [526]. Anonymous [1276]. ans [561]. Antall [1228]. Ante [1200, 816]. Ante-Bellum [1200, 816]. Antebellum [1614, 1444]. Anthology [658]. Anthony [1814, 661, 1470, 869, 1646, 878, 879, 1287]. Anthropological [1406]. Anthropologie [759]. Anthropology [1662, 861, 1356, 1214, 1382, 73, 1099, 1591, 304, 26, 807]. Anti [998]. Anti-Lyellian [998]. Anticipated [1790]. Anticipations [857]. Anticlaudianum [747]. Antike [214]. antiken [360, 707]. antiker [358]. Antikythera [1247, 1937, 1936]. antique [1472]. antiquité [1547]. Antiquity [152, 1740]. Antivivisection [1562]. Antoine [957, 1102, 1484]. Antoinette [188]. Antología [57]. Anton [1877]. Antonina [1878]. Antonio [895, 839]. Antony [1729]. Anwendungen [1847]. Aphorismi [1348]. Aphrodisias [1473]. apologista [570]. Apostle [154]. Apothecaries [328]. Apotheker [1849]. Apparatus [1697]. Appeal [11]. Appearance [241]. Appendektomie [1851]. Application [1019, 289, 1108]. Applications [223, 1566]. Applied [141]. Apply [1466]. Appraisal [1565, 697]. Appreciation [194]. Approaches [1166, 1840]. April [1168, 1654]. Aquatic [1880]. Aquin [1899]. Arab [1409]. Arabic [1479, 1298, 363, 1550, 1648, 12, 1081, 521]. arabischen [1187, 1482]. arabischer [1403]. Arabs [1050]. a’rad [1761]. Arbeitsweise [43]. Arboretum [513]. Archaeoastronomy [1470]. Archaeological [1504, 1435]. Archaeologist [491]. Archaeology [1468, 553, 1009, 792]. Archeology [883]. Archibald [668]. Archimède [744]. Archimedean [964]. Archimedes [1898]. Architectural [219, 28, 1239]. architettura [839]. Archiv [1330, 8]. Archival [11]. Archive [8]. Archives [598, 737, 1783, 1171, 437]. Arctic [1712]. Area [883]. Arechaga [1707, 1782]. Argand [1800, 1193, 758]. Argüelles [940]. Argyll [1214]. Arié [1761]. Ariga [1567]. Arising [415]. Aristotele [745, 519]. Aristóteles [1407]. Aristotelian [423]. Aristotelianism [620]. Aristotelicum [1837]. Aristotelis [79]. Aristotelismus [1473]. Aristotle [1180, 1104, 983, 1672, 1445, 1081, 1581, 154, 983, 123, 1837, 1897, 799, 154]. Arithmetic [1252, 1216, 1903, 521, 778, 962, 396]. Arithmetical [887]. Arithmetics [644]. arithmetischer [98]. Arkhivakh [1783]. Armament [223]. Armand [1776]. Armas [1293]. Armenian [169]. Armillary [1007]. Armin [74, 530, 1461, 1508, 702, 1572]. Armory [1429]. Armstrong [898]. Arnaldi [1348]. Arnaldo [1410]. Arnold [74, 1395, 272, 276, 513, 893, 1241, 1817, 1571]. Arrangement [1775, 349]. Art [1190, 1034, 1214, 1244, 1419, 66, 793, 1869, 180, 1050, 1606, 751]. arte [86]. Artelt [928]. Arthur [907, 1592, 919, 69, 791]. Articles [1845]. Artillery [646]. artistes [1731]. Artists [1599]. artium [750]. Arts [1047, 176, 215, 96, 1228, 1250]. Arturo [51]. Arzneibereitung [1182]. Arzneimittel [1066]. Arzneimittel-Standardisierung [1066]. Arzt [260, 1299, 67, 1652, 802]. ärzte [947]. as-Salah [1480]. As-Samaw’al [933]. Ascent [971]. Ash [1587]. Ashforth [438]. Ashmole [1145]. Ashworth [1779]. Asia [1800, 1764]. Asit [556]. Aslam [1522]. Aspect [630, 963]. Aspects [1623, 881, 227, 771, 1246]. Aspetti [433]. Assessment [326, 1879, 1645, 453]. Assessments [893]. Assistant [1686]. Association [1367, 1119, 1697, 1234, 1171, 1632]. Associative [1867]. Assunta [1690]. Astrolabe [1007]. Astrolabes [1759]. Astrologers [356]. Astrological [46, 509, 125]. astrologicum [1471]. Astrologie [256, 886]. astrologische [214]. Astrologumena [214]. Astrology [544]. Astronom [1652]. Astronomer [1409]. Astronomers [348]. astronomiae [1647]. astronomica [1910]. astronomicae [1490]. Astronomical [145, 959, 1350, 1517, 1741, 873, 749]. Astronomie [549]. Astronomies [1756]. astronomique [885]. astronomiques [1484]. astronomiske [1741]. Astronomy [822, 4, 287, 20, 1926, 1340, 1871, 123, 537, 1322, 1051, 669, 1596, 225, 1738, 748, 1646, 300, 1137, 1926, 245, 831]. Astrophysical [820]. Astrophysics [326]. Athanasius [4]. Athanese [1530]. Atom [689, 1155]. atome [811]. Atomic [107, 1571, 957, 1264, 558, 1719, 143, 107, 119]. Atomism [301, 1866]. Atoms [272]. Atomzeitalters [302]. Attanasio [215]. Atti [179]. Attitudes [685, 482]. Aub [383]. aube [1907, 631]. Aubrey [1359, 1835]. Auctions [1469]. Audrey [889, 271, 864]. Aufbau [1684]. Aufklärung [1413, 341]. Aufsätze [1332]. Augsburg [566]. August [1375, 1324, 573, 1521, 403, 999]. Auguste [1909, 1368]. Aurel [1504]. Aurigemma [1547]. Aurorae [1677]. Aurousseau [444]. Auseinandersetzung [752]. Australia [1425]. Australian [312]. Author [1559]. Autobiographia [532]. Autobiography [1375, 1491, 449, 197]. Autograph [1059, 86]. Autoren [841]. autrichiens [622]. Auvergne [1477]. Aux [1233]. Aveni [1470, 1646]. Averill [1538]. Averlino [839]. Avery [1622]. Avogadro [957, 1635]. Awakening [187, 1137]. Award [1388, 1450]. Awarded [403, 507, 823]. Axel [1877]. Axiome [61]. Ayala [1138]. Ayer [37].

B [1713, 1292, 1318, 717, 184, 1746, 468, 370, 889, 1157, 1570, 1750, 1557, 756, 1168, 132, 1561, 70, 874, 754, 764, 1534, 1424, 1759, 513, 1818, 31, 634, 1749, 354, 1023, 1076, 270, 1024, 950, 709, 534, 760, 887, 182, 477, 347, 466, 1575, 1137, 952, 1155, 876, 620, 1649]. B.C [794]. B.C. [1937, 1936, 1406, 1247]. Baġdādī’s [546]. Babai [1818]. Babbage [1787]. Babylon [1454]. Back [44, 76, 110, 135, 139, 163, 198, 229, 279, 283, 330, 389, 450, 493, 497, 538, 585, 638, 679, 683, 726, 772, 813, 848, 911, 953, 993, 1025, 1029, 1074, 1121, 1161, 1207, 1267, 1315, 1376, 1436, 1512, 1577, 1624, 1666, 1716, 1803, 1863, 1931, 1935]. Background [1565, 1076, 559]. Backgrounds [184]. Bacone [1301]. Baconian [1614]. Bacteria [1402]. Bad [1918, 1761]. Badash [103, 1259]. Baden [1709]. Badger [312]. badische [1795]. Baer [1783]. Baers [1484]. Bagdadi’s [259]. Bagdedinus [291, 595]. Baglivi [1194]. Bahir [933, 1105]. Bahnbrecher [302]. Bailyn [41]. Bainbridge [1620]. Baker [1312, 120, 1143]. Balaguer [1355]. Baldi [1703, 1485, 838, 781]. Baldiego [838]. Baldry [417, 1402]. Balfour [1166]. Ball [1089]. Ballester [1782, 1766, 798, 1893, 1551, 884]. Balmer [1590, 1704]. Band [1015, 522, 1765, 1414, 1799, 33, 1612, 511, 517]. Banesh [723]. Banister [371]. Bankers [18]. Banneker [523]. Baptiste [1063]. Barbara [1196, 1861, 342, 102, 878, 1010]. Barber [325]. Barclay [751]. Barker [670, 59]. Barlaam [1552]. Barnes [1394]. Barometers [19]. Baron [756, 160, 397]. Barr [1003]. barre [1839]. Barrett [1065, 1312, 254, 1065, 1914]. Barriers [1466]. Barrow [529]. Barry [1394]. Bartlett [1375]. Bartolomeo [1213]. Barton [1382]. Bas [622]. Basalla [478, 89]. Based [625, 449, 1697, 380]. Basis [7]. Bassett [1019]. Bastin [415]. Bates [382, 484]. Bathybius [997]. Battelli [1930]. Battle [417, 1402]. Baum [1086]. Bayan [1761]. Bayle [986]. Be [1507, 1089]. Beagle [72]. Beaglehole [984]. Bear [944]. Beast [527]. Beast-Machine [527]. Beasts [795, 661]. Beaton [182]. Beatty [1187, 1482, 880, 742, 1183, 1829]. Beaulieu [1776]. Beaune [1487]. Beche [1794, 998]. Beck [1349]. bedeutender [1799]. Bedeutung [1607]. Bedient [1533]. Bedini [1302, 523]. Beetles [18]. before [1727, 1343, 1232, 455, 610, 1250, 1445]. Begegnung [752]. Beginnings [64, 469, 471, 100, 1244, 1715, 1619]. Begriff [1831]. Begründung [273]. Begrüssungsansprachen [1414]. Behavioral [1748, 1695]. Behind [241]. Behnke [633]. Being [225, 712, 1050, 303, 716, 337]. Beit [1761]. Beitrag [1714, 1112, 1061]. Beitragen [1875]. Bel [1261]. Belief [1005, 1331]. belka [1883]. Belkin [1905]. Bell [1393, 1420]. Bellum [1200, 816]. Belmont [961]. belysande [262]. Ben [1350]. benannte [858]. Bendix [574]. Benjamin [223, 523, 215, 510, 1867, 1382, 1448, 64, 715, 1559, 1060]. Bennett [69]. Benoît [34]. Benson [1008]. Benz [752]. Ber [1783]. Bera [1783]. Bereano [1734]. Bereich [255]. Berendzen [1737]. Berg [1115]. Bergman [966, 235, 967, 545, 783]. Bergson [130, 722]. Berichte [1414]. Berichten [295]. Berkeley [753, 99, 1055, 99, 1055, 1303]. Berkson [1062]. Berlant [1334]. Berliner [1017, 759, 1617, 1558, 759]. Berman [1913]. Bern [1031, 1388]. Bernard [225, 1115, 251, 1189, 843, 1169, 812, 302, 748, 1511, 1350, 1708, 1842, 1058, 41, 325, 1539, 837, 427, 920]. Bernardina [838]. Bernardino [1703, 781, 1485, 838]. Bernardus [1875]. Bernhard [1057, 1929, 157, 1332, 1521]. Bernice [1232]. Bernier [1233]. Bernoulli [1612, 511, 511]. Bert [1467]. Berthold [670]. Berthollet [1079]. Bertrand [1832, 1261]. Berufsbildes [1096]. Beryl [1888]. Beschreibung [1187]. Bessie [669]. Besso [636]. Bestehen [759]. Beta [1036]. Betrachtungsweise [360]. Betsy [314]. between [301, 1083, 775, 857, 241, 1336, 1173, 34, 520, 170, 1580, 948]. Bewegung [255]. Beyerchen [1685]. Beyond [276, 415]. Beziehungen [548]. Bharatha [1858]. Bialas [522, 1112]. Bibby [896, 576]. Bibliografia [409, 1045]. Bibliograficheskii [1589]. Bibliografie [1774]. Bibliographia [572]. Bibliographic [1695, 542, 305]. Bibliographical [1589, 1561, 1527]. Bibliographie [33, 58, 517]. Bibliographies [922, 1689]. bibliographique [1642]. bibliographischen [1875]. Bibliography [398, 1343, 1588, 922, 132, 1610, 1689, 1044, 1749, 926, 138, 282, 496, 682, 851, 1028, 1210, 1439, 1669, 1934, 147, 462, 983, 648, 305, 620, 1290, 1912, 978, 825, 931, 363, 615]. Bibliothèque [720, 1487]. bibliothèques [1650]. Bicameral [1747]. Bicentenary [312, 475]. Bicentennial [1643]. Bickel [1205]. Bidloo [661]. bidrag [262]. Bienayme [1790]. Bieńkowska [1010, 342]. Biermann [1657, 1017]. Biesterfeldt [1403]. Bifocal [1192]. Big [302]. Biggs [1691]. Bijvoet [941]. Bildern [295]. Bildung [483]. Biliński [1651, 1703, 1485, 838]. Billings [979]. binären [1061]. Binet [905]. Bingham [187]. Bio [33]. Bio-Bibliographie [33]. Biochemistry [656, 541, 1044, 1744, 1579, 1739, 183, 796]. Biografii [1589]. Biographical [820, 1237, 1003, 146, 1044, 613, 1923, 864]. Biographien [1799, 702]. Biographies [1589]. Biography [1205, 131, 1451, 1506, 332, 990, 838, 937, 104, 1726, 193, 102, 1576, 211, 738]. Biologen [1799]. biológica [109]. Biological [487, 976, 633]. biologiche [745]. biologie [1233, 1889, 1642]. Biologist [384]. biologiste [133]. Biology [1574, 1138, 1524, 600, 633, 1373, 630, 480]. Biomedical [881]. Biosynthesis [1744]. Biran [446]. Bird [1888, 685]. Birdes [894]. Birds [1888]. Birk [810]. Birkenmajer [694]. Birkhoff [942]. Birth [1013, 1500, 1920, 485, 1137, 1010]. Biruni [1186, 1649, 1054]. bis [218, 273, 1473, 302, 874, 98, 1097, 1617, 1918, 1391]. Biskup [660]. Biswas [556]. Bitruji [748]. bizottságai [1457]. Björnbo [1877]. Black [1078, 1812, 1226]. Black-Body [1812]. Blackberry [693]. Blackmore [897]. Blaeu [937]. Blanche [1343]. Blood [917, 666]. Bloor [1394]. Blumberg [1372]. Blumenberg [1459]. Blumenthal [1419]. Blunt [426]. Board [1880]. Boas [1654, 372, 502]. Bodies [1672, 1631, 1383]. Body [257, 1812]. Boeren [1546]. Boerhaave [263, 1779]. Böhme [1816]. Boies [528]. Bollack [746]. Bolton [182]. Boltwood [103]. Boltzmann [1198]. Bolzán [1407]. Bolzano [1115, 1115, 1708]. Bombelli [644]. Bonaventura [955]. Boncompagni [1703]. Bonelli [1257, 465, 1113]. Bones [18]. Bonner [528]. Bonnet [567]. Bonnot [712]. Book [1596, 1601, 225, 1604, 1853, 277, 1713, 614, 254, 264, 301, 888, 1330, 1411, 1490, 1824, 712, 1656, 1773, 105, 274, 414, 1428, 1434, 1492, 1574, 1664, 1924, 1661, 933, 1252, 1837, 1897, 1524, 368, 1057, 1458, 740, 805, 652, 1359, 1650, 1281, 871, 1832, 107, 226, 278, 515, 536, 767, 1001, 1071, 1088, 1266, 1372, 1571, 1850, 1928, 1469, 1733, 1687, 1292, 1108, 616, 745, 676, 1069, 1199, 799, 1814, 1618, 743, 1927, 159, 463, 1203, 1398, 1700, 40, 1310, 1739, 1753, 1706, 1190, 1731, 1752, 1815, 675, 1408, 339, 1884, 1595, 483, 532, 1848, 218, 1427]. Book [247, 1770, 1921, 246, 385, 1115, 667, 844, 734, 1890, 902, 345, 741, 407, 1611, 1856, 650, 701, 1425, 670, 891, 1086, 1144, 352, 635, 715, 1015, 1205, 1465, 1851, 315, 64, 388, 717, 1240, 1393, 445, 632, 1062, 1789, 117, 624, 986, 1354, 93, 1100, 1237, 1538, 1590, 1748, 1756, 15, 273, 1265, 1432, 1657, 63, 464, 1150, 973, 1202, 1454, 1734, 124, 185, 1466, 1543, 379, 1777, 364, 1638, 1431, 1766, 1151, 184, 242, 1196, 1235, 1746, 1861, 1602, 131, 251, 669, 692, 719, 932, 312, 316, 832, 1562, 1779, 1896, 1477, 1758, 1888, 413]. Book [423, 374, 890, 1170, 1413, 1692, 1092, 355, 17, 25, 649, 1065, 1559, 1793, 435, 519, 627, 253, 880, 1782, 375, 71, 1189, 1013, 1185, 262, 1286, 468, 870, 1371, 1184, 601, 744, 1172, 1887, 469, 1191, 490, 1147, 1197, 1542, 446, 527, 695, 809, 843, 59, 344, 384, 1374, 219, 883, 1554, 1876, 120, 895, 1375, 1110, 560, 1152, 1422, 443, 1892, 608, 1854, 405, 1370, 481, 487, 1641, 327, 535, 1662, 1204, 1566, 1858, 1017, 1067, 1502, 1885, 193, 370, 1228, 1830, 1737, 889, 769, 810, 265, 18, 1016, 1415, 1333, 1903, 934, 1587, 1771]. Book [1833, 755, 65, 188, 192, 224, 338, 842, 892, 1769, 1653, 1451, 798, 30, 1047, 471, 434, 708, 935, 971, 1296, 28, 1373, 1727, 100, 382, 426, 525, 877, 1008, 1048, 1157, 1290, 1343, 1399, 1494, 1537, 1570, 1599, 1694, 1745, 1588, 556, 1531, 479, 1800, 1116, 1335, 430, 714, 826, 1232, 1358, 1430, 1827, 804, 1180, 1262, 1889, 661, 1750, 444, 478, 1159, 1874, 103, 1099, 1545, 1847, 1533, 711, 275, 1506, 1620, 428, 1308, 1412, 1555, 1591, 1637, 1690, 1843, 1878, 1169, 362, 654, 1898, 43, 74, 119, 134, 241, 408, 907, 908]. Book [947, 992, 1640, 1684, 1714, 1801, 486, 419, 922, 1368, 671, 1283, 1356, 297, 518, 1360, 1557, 425, 978, 1834, 797, 720, 756, 603, 1140, 1786, 514, 1796, 1768, 827, 630, 1693, 243, 1426, 1019, 1523, 1007, 1470, 1835, 1244, 95, 522, 1553, 1765, 1825, 1904, 57, 109, 132, 570, 1045, 1433, 1658, 665, 319, 351, 1302, 812, 975, 1480, 1339, 440, 1778, 1738, 1111, 1395, 753, 1473, 55, 923, 1589, 1143, 256, 1198, 1239, 152, 903, 1423, 1568, 611, 1307, 1256, 1194, 1309, 329, 350, 1491, 149, 304, 807, 1102, 647, 1288, 904, 1917, 1503]. Book [1565, 1056, 875, 1334, 154, 255, 746, 1295, 272, 569, 637, 725, 1139, 150, 646, 656, 666, 833, 1053, 1134, 1254, 1355, 1610, 1757, 1886, 27, 92, 259, 309, 1742, 410, 625, 938, 1142, 1416, 1177, 472, 1300, 1547, 1901, 925, 1052, 982, 924, 62, 326, 1452, 910, 1775, 106, 674, 1822, 88, 126, 306, 467, 492, 302, 530, 557, 678, 771, 723, 752, 898, 985, 1003, 1230, 1361, 436, 989, 1178, 1461, 295, 1529, 1592, 387, 267, 381, 1153, 1561, 509, 1186, 748, 699, 739, 768, 788, 790, 828, 829, 830, 945, 1176, 811, 162, 735]. Book [1802, 342, 660, 694, 1479, 1005, 664, 161, 1303, 217, 1489, 1367, 920, 703, 458, 1060, 1429, 1823, 158, 365, 276, 1096, 584, 704, 1774, 1729, 789, 1826, 70, 183, 796, 847, 1600, 1798, 1504, 1187, 1482, 422, 1754, 653, 867, 1705, 626, 153, 574, 825, 1282, 1556, 1642, 1689, 1728, 1877, 99, 1500, 20, 102, 125, 222, 314, 1510, 1919, 121, 550, 750, 874, 1560, 1050, 1761, 1838, 35, 191, 244, 266, 268, 317, 318, 429, 447, 461, 488, 754, 787, 899, 1145, 1206, 1227, 1654, 1929, 24, 196, 299, 349, 449, 459, 1041, 1242, 1261]. Book [1336, 1397, 1532, 1112, 1304, 1085, 1195, 1257, 1511, 1643, 981, 1066, 721, 869, 951, 1686, 1546, 260, 310, 521, 524, 673, 1105, 1299, 1350, 578, 1264, 1785, 1915, 636, 1812, 1685, 764, 1014, 1859, 800, 37, 558, 1528, 1593, 600, 1044, 1509, 1119, 1200, 1284, 1534, 1882, 1911, 21, 441, 942, 1094, 1708, 533, 1424, 1621, 91, 489, 651, 1156, 1730, 759, 882, 1414, 1842, 360, 885, 1248, 1552, 1759, 1467, 1609, 513, 1098, 353, 417, 359, 1704, 346, 1249, 1818, 1117, 1772, 1712, 1468, 1018, 512, 553, 1258, 1732, 940, 868, 1922, 1464, 1181, 1784, 929]. Book [38, 1845, 1883, 128, 1289, 1608, 1173, 1487, 1655, 1683, 1820, 1930, 893, 31, 629, 580, 98, 617, 1616, 187, 240, 1253, 1478, 1740, 386, 1146, 32, 949, 1417, 609, 1474, 1549, 1174, 1055, 1881, 1305, 634, 742, 1597, 1508, 1114, 1460, 571, 1613, 1651, 1703, 1844, 1406, 1483, 1749, 554, 1569, 1879, 1497, 26, 298, 1619, 1603, 1710, 1893, 1755, 1183, 1536, 227, 211, 307, 322, 324, 354, 439, 475, 931, 980, 1101, 1698, 1799, 411, 1691, 1849, 1476, 1646, 424, 1836, 948, 321, 1023, 1072, 438, 480, 808, 896, 1914, 1463, 1061, 1623]. Book [258, 311, 747, 1351, 1551, 1009, 1435, 1925, 738, 1594, 1644, 221, 1909, 19, 250, 939, 724, 1743, 1021, 801, 1484, 270, 376, 897, 1364, 561, 1104, 529, 223, 1193, 1090, 1563, 1663, 707, 1136, 176, 213, 406, 559, 622, 668, 926, 1058, 1097, 1342, 1472, 1496, 1567, 1659, 1688, 1776, 1797, 1763, 1346, 1179, 58, 552, 737, 1913, 1255, 523, 872, 1241, 1314, 1605, 1926, 194, 465, 526, 566, 1247, 1614, 1855, 1419, 1420, 1348, 1530, 1622, 1024, 1456, 757, 437, 516, 1201, 1781, 1783, 1846, 573, 34, 303, 1564, 36, 1285, 383, 1238, 1345]. Book [1402, 1920, 1617, 1788, 1046, 1741, 1894, 108, 343, 950, 761, 1311, 418, 220, 575, 1063, 1338, 97, 186, 709, 1906, 1829, 214, 195, 1645, 905, 181, 369, 604, 1291, 1744, 979, 1421, 1507, 1182, 1548, 1607, 655, 623, 697, 420, 485, 927, 1020, 1918, 534, 760, 583, 887, 1535, 613, 1344, 1900, 101, 271, 127, 371, 367, 1767, 1821, 1880, 1404, 1751, 548, 41, 118, 263, 325, 462, 555, 930, 1154, 1394, 1526, 1699, 1817, 1697, 1012, 66, 1405, 482, 182, 1905, 377, 894, 562, 884, 1103, 1707, 1831, 793, 906, 1365, 1791, 1787, 1341]. Book [1369, 1709, 1792, 67, 710, 1485, 33, 873, 1306, 1459, 442, 1539, 1916, 749, 1352, 795, 1455, 1260, 610, 1226, 577, 766, 1234, 68, 151, 216, 296, 477, 864, 1011, 1312, 1337, 1544, 602, 1794, 976, 1138, 1525, 22, 1340, 1409, 1471, 1839, 252, 1233, 835, 1049, 1246, 1298, 1347, 1403, 1481, 839, 1902, 567, 1251, 1606, 130, 722, 1639, 612, 179, 215, 366, 409, 619, 865, 936, 983, 1188, 1652, 941, 1022, 29, 248, 572, 700, 716, 758, 1453, 42, 1004, 1735, 60, 1396, 943, 1790, 946, 372, 1059, 1109, 1148]. Book [1615, 1908, 1331, 836, 313, 433, 663, 713, 1113, 1301, 1495, 1910, 1711, 347, 421, 1118, 1229, 1780, 466, 189, 1313, 628, 919, 69, 618, 648, 1106, 1135, 1141, 1353, 1701, 61, 1363, 528, 1488, 1089, 257, 308, 363, 473, 520, 564, 1357, 1550, 1648, 96, 155, 1091, 878, 1158, 1243, 1392, 1501, 1575, 1695, 1263, 1702, 1462, 806, 1576, 412, 160, 762, 1499, 736, 156, 474, 565, 802, 1293, 1391, 1418, 1660, 1923, 658, 718, 901, 1064, 706, 1527, 633, 1862, 1895, 357, 606, 657, 1054, 1541, 269, 320, 470, 484, 531, 693, 803, 879, 1860]. Book [340, 341, 1612, 568, 157, 197, 228, 249, 415, 765, 1498, 615, 94, 579, 581, 909, 990, 549, 1736, 1297, 1475, 1899, 1236, 1457, 1841, 974, 1231, 1907, 56, 89, 460, 984, 1171, 305, 328, 380, 448, 751, 763, 792, 1192, 1400, 1598, 1696, 1828, 1042, 659, 977, 1760, 23, 123, 245, 300, 323, 348, 356, 358, 838, 886, 1137, 1245, 677, 991, 1073, 1649, 582, 702, 952, 988, 1087, 1155, 1259, 1332, 1572, 1540, 1006, 1093, 432, 511, 1410, 1558, 576, 987, 1068, 1795, 900, 1294, 39, 517, 607, 261, 1095, 1493]. Book [1857, 133, 178, 876, 881, 1160, 1505, 1349, 551, 1175, 1107, 1362, 672, 840, 1813, 662, 361, 620, 972, 1250, 129, 212, 373, 416, 537, 791, 831, 937, 1043, 1287, 1366, 563, 621, 928, 1875, 190, 1891, 1816, 696, 837, 866, 1407, 1647, 104, 1573, 427, 431, 457, 476, 1726, 1912, 1010, 1486, 1665, 73, 122, 177, 944, 1070, 1852, 786, 1225, 378, 846, 510, 841, 770, 921, 1002, 1149, 16, 75, 1715, 845, 1747, 180, 337, 1819, 491, 1764, 90, 1401, 1762, 1840, 1120, 72, 698, 631, 605, 705, 794, 1051, 834]. Books [649, 1343, 826, 154, 462, 864, 1648]. Booth [314]. Borek [448]. Borel [775]. Borelli [1355, 873]. Borghetty [66]. Boris [461]. Born [134, 302, 134]. Borscheid [1709]. Borst [1292, 1292]. Boruch [16]. Boss [840]. Botanical [1343, 1599, 864]. botánico [570]. Botaniker [1849]. Botanist [1796]. Botanists [1599, 864]. Botany [1130]. Bourbaki [700]. Bourde [156].

Bowden [1289]. Bowers [1643, 1364, 90]. Bowl [1681, 23]. Bowler [1597]. Boxed [1588]. Boy [886]. Boyd [669]. Boydston [1158]. Boyer [1168, 1608]. Boyle [191, 1772, 1276]. Bozeman [1614]. Bradwardine [503]. Bragg [1850]. Brahe [262, 1679]. Brain [640, 832, 487, 152]. Branca [1784]. Brass [1681]. Braun [94]. Bray [375, 1426, 1426]. Brazilian [1619]. Breakdown [1747]. Breathing [115]. Breck [633]. Brednow [376]. Breguet [1190]. Breidert [753]. Breuers [1853]. Brevans [979]. Brew [26]. Brian [1198, 1503, 1500, 1364, 1337, 1006, 837]. Bridge [1922]. Bridges [1608]. Brief [228]. Briefe [74, 1909, 1330, 1067]. Briefwechsel [1170, 1844, 572, 1558, 1330, 1299, 33, 1657, 134]. Briefwechsel. [74]. Brieger [667, 1540]. Briguglio [902]. Brinner [1789]. Britain [1924, 1281, 1571, 1733, 1777, 1659, 437, 1697, 598, 1865, 1835, 306, 1719, 1334, 1023, 1072, 1926]. Britannia [1922]. britannici [1497]. British [1469, 117, 25, 1833, 192, 1599, 908, 756, 1835, 1886, 1178, 222, 583, 1154, 657, 470, 1171, 1516, 334, 1379, 1343]. Brittain [1751]. Brittan [1906]. Broberg [1356]. Brock [898]. Brodrick [792]. Brody [16]. Broglie [1387]. Broneer [60]. Bronislaw [1651, 1703, 1485, 838]. Brook [1303]. Brooke [1771]. Broun [1160]. Brown [1088, 1798, 223, 1286, 1335]. Browning [1312]. Brozek [1746, 320]. Bruce [219, 1566, 1339, 321, 668, 1089, 1263, 901, 1420]. Bruijn [1493]. Brunelleschi. [311]. Brunés [608]. Bruno [1531, 934, 188]. Brush [652, 515, 754, 1498]. Bruylants [1063]. Bryden [1004]. Bryn [1697]. Bryozoa [1019]. Bubonic [117]. Buchanan [925]. Buchdahl [604]. Bücher [98]. Buchillustration [517]. Buck [531]. Buffon [1124, 856, 804]. Buffon-Linnaeus [1124]. Bugge [101]. Building [854, 452]. Bukharin [460]. Bulferetti [902]. Bullough [178]. Bunge [924]. Bunning [1849]. Bunt [1533]. Burchardt [1684]. Burchfield [1150]. Burgers [941]. Burgess [864]. Bürgi [1112]. Burhop [767]. Burke [1768, 1852]. Burkert [834]. Burkhardt [1524]. Burks [1592]. Burl [1835]. Burmeister [260, 1299, 33]. Burner [1193]. Burnett [803]. Burnham [1115, 343]. Burning [1550]. Burrow [1710]. Burt [502]. Busard [1648]. Bush [721]. Busvine [1598]. Bütow [564]. Buttmann [1785, 104]. Butts [1769, 369, 575]. Buzajani [521]. Bykov [1883, 1883, 1535]. Byzantine [1826]. Byzantium [1836].

C [1434, 1088, 1927, 385, 670, 315, 1756, 185, 251, 1758, 1286, 490, 446, 843, 1858, 1833, 1769, 708, 1296, 1157, 1399, 1570, 756, 1426, 152, 1307, 1610, 982, 1005, 789, 867, 314, 1546, 1467, 1460, 1844, 1406, 223, 1530, 1846, 1285, 383, 343, 1443, 1584, 66, 182, 1790, 836, 204, 1576, 412, 320, 531, 803, 984, 1171, 900, 361, 1573, 73, 1852, 1554, 798]. C. [999]. côté [809]. ca [1247, 1937, 1936]. Cairo [521]. Cajus [835]. calcul [1839]. Calculus [501]. Calendar [1247, 1711, 1936, 1937]. Calendars [1398]. Calev [1905]. California [1206]. Calinger [1842]. Call [1498]. Called [1879]. Caloric [1338]. Cambridge [17, 1227, 1059, 582]. Campanella [986]. Campanus [510]. Campo [747]. Camps [144]. Can [1507]. Canada [1928, 48, 1880]. Canadense [209, 48]. Canadian [1022, 1089]. canalizzazione [713].

Canberra [312]. Canguilhem [827, 512]. Cannon [1876]. Canterbury [952]. Cantor [1630, 238]. Cantore [119]. Canyon [73]. Capaldi [16]. Capek [722, 1639]. Capital [1421]. Caplan [760, 760]. Capra [1336]. Captain [1496, 528, 984, 312]. Carbon [506]. Carbon-14 [506]. Cardano [1652]. Carder [1095]. Cardiac [153]. Cardini [519]. Cardwell [704, 439, 220]. Care [929, 1861, 1184]. Career [1673, 1157, 1570, 1517, 786]. Cargill [1802]. Caricature [998]. Carinthia [1648]. Carison [1539]. Carl [1657, 1431, 1494, 430, 1356, 319, 1174, 931, 1644, 96, 590, 1168, 193, 67]. Carlo [1354, 1202, 1012, 567, 179]. Carlos [1768]. Carmen [1471]. Carnis [656]. Carnot [315, 1858]. Caroe [1850]. Carol [1749]. Carolyn [1919]. Caroti [1647]. Carozzi [1800, 488, 34, 36, 271, 1231]. Carroll [355, 877, 1711]. Carson [795, 1462]. Cartelli [1410]. Carter [1567]. Cartesian [592]. Cartesianism [592]. Cartographer [937]. Cartography [655]. Cartwright [1827]. Carus [170]. Caryn [1842]. Case [1144, 455, 1579, 1379, 997]. Casimir [133]. Casimir-Joseph [133]. Casini [313]. Caspar [1486]. Cassedy [100]. Cassirer [413]. Casson [422]. Castellani [1846, 567]. Catalog [91, 552, 118, 373, 1259]. Catalogue [1296, 1116, 1825, 1728, 1227, 864, 1544, 245, 886, 252]. Catalogues [1145]. Catesby [1462]. Cathedral [258]. Catherine [807, 73]. Cathleen [1180]. Catholique [1525]. Caton [935]. Cauchy [1499]. Causality [80, 696]. Cause [502, 1592]. Causes [306, 1761, 488, 1583]. causis [1761]. Cavalieri [955]. Cavendish [1087, 172]. Cayeux [488]. Cecco [1412]. Cecil [767, 62, 1536]. Celebration [995]. Celebrations [1886]. Celestial [1325, 1807]. celi [472]. Celina [1407]. Cell [1374]. Celli [1703]. Cellular [380]. celular [1782]. Cena [1587]. Ceneri [1587]. Cenkalska [1010]. Census [749, 1352]. Cent [561]. Centenaire [1175]. Centenary [1886, 322]. Centennial [952, 70, 1261]. Center [951, 118, 581]. Centers [1002]. centesimum [532]. Central [1229]. Centre [1525]. Centry [1357]. Centuries [1753, 1457, 245, 1813, 921, 306, 1911, 623, 1875]. Century [1601, 159, 667, 1425, 624, 1590, 1265, 890, 253, 1013, 1371, 487, 265, 18, 1373, 1008, 630, 1631, 1256, 904, 678, 1186, 458, 521, 673, 1509, 187, 480, 1925, 250, 1255, 1456, 757, 1920, 1788, 760, 1154, 1341, 1369, 818, 1916, 577, 69, 473, 1923, 328, 659, 987, 686, 1107, 662, 837, 431, 1322, 1762, 288, 402, 1216, 1078, 338, 1271, 777, 1386, 1378, 8, 81, 167, 731, 596, 962, 3, 1317, 453, 1272, 1583, 333, 396, 1604, 1773, 1370, 1566, 1768, 152, 569, 1479, 1823, 1859, 1621, 513, 1597, 1114, 1238, 68, 1409, 1498, 1486]. Century [241, 1085, 762, 579]. Ceramics [27]. Cerebral [487]. Cesare [1188]. Chûgoku [1353]. Chaff [1749]. Chain [314]. Chaining [1022]. Chains [1088]. Chair [1069, 1125]. Challenge [650, 1429]. Challenger [808]. Challenging [633]. Chalmers [1469, 1879]. Chalmers-Hunt [1469]. Chamber [1815]. Chamberlin [1584]. Chambers [93]. Chance [1736, 380, 1592]. chang [98, 1107, 1406]. Change [1913, 930, 56, 463, 1429, 1730, 1772, 1174, 698, 1255]. Changing [1273, 1135]. Chappert [1615]. Chapters [616, 319, 1905]. Character [803]. Chardin [109, 1433]. Chargaff [1727]. Charles [315, 388, 1789, 464, 379, 669, 744, 1554, 608, 535, 224, 1451, 1588, 1361, 1561, 161, 99, 1242, 37, 128, 1616, 1114, 1914, 1364, 1697, 1787, 1312, 567, 1711, 432, 620, 416, 537, 177, 944, 82, 141, 1872, 543, 1851, 1658, 387, 381, 940, 602]. Charlton [154]. Chart [1911]. Chaucer [125, 659]. Chauliac [425]. Chauncey [125, 1238]. Chebyshev [1921]. Check [709]. Check-list [709]. Checklist [1230, 187, 305]. Chemical [445, 1518, 203, 499, 83, 196, 349, 1483, 1319, 1341, 1866, 1875]. Chemie [1655, 1097]. Chemiker [1096, 1849]. Chemistry [1739, 1375, 889, 1078, 83, 70, 183, 847, 24, 349, 449, 600, 1424, 1883, 1483, 559, 1082, 1341, 391, 466, 289, 120, 1169, 272, 1254, 70, 299, 406]. Chemists [1254, 1341]. Chester [1187, 1482]. Cheyne [1539]. chez [920]. Chicago [1431, 676, 1431]. Child [1861]. Children [216]. Chimie [730]. chimique [1079, 622, 1063]. Ch’in [1107]. China [617, 1406, 1483, 1763, 618, 506, 69]. Chinese [1874, 1406, 648, 1106, 1141, 1353, 1762, 1681, 1163, 1272, 7, 1874, 32, 1107]. chinesisch [98]. chinoise [1701]. Chirurgie [564]. chirurgien [631]. Chissick [1573]. Chittick [1610]. ch’iu [1106, 1107, 1107, 98]. chiu-chang [1107]. Chiu-shao [1107]. Chlorinity [1342]. Chlorinity/Salinity [1342]. Choix [1801]. Cholera [1618]. Chouraqui [887]. Chr [98]. Chr. [1473]. Chreias [257]. Christchurch [952]. Christiaan [1383]. Christian [1309, 92, 800]. Christianson [1738]. Christina [1010]. Christine [1923]. Christman [634]. Christoph [571, 22]. Christopher [627, 314, 1090, 1255, 356]. Chromatic [914]. Chromosomes [1514]. Chronicle [310]. Chronology [1519]. Chronos [1741]. Chrysippus [335]. Chu [1106, 1106]. Chu-Ni [1106, 1106]. Chugoku [1141]. Church [429]. Churchill [635]. Chyle [917]. Chymici [1546]. ciencia [1766, 798, 57, 570, 1045, 1613]. Ciges [1893]. Cinnabar [1483]. Cinq [5]. Cinquecento [1188]. cinquencento [1555]. Cipolla [1354, 1012]. circa [112, 1560, 566]. Circle [1793, 913]. Circles [1835]. Circular [472, 333]. Circulation [666, 889]. Cirkelkwadratuur [362]. Citations [1519]. City [899, 1421]. Civic [409]. Civil [1216, 1286, 755, 617, 613]. Civilisation [617, 1483, 1763]. Civilization [1244, 1406, 882, 655]. Clagett [1898]. Claim [859]. claires [1770]. Claparède [763]. Clark [1595, 1600, 1781, 377, 1418, 943]. Clarke [832, 152, 318, 318]. Clash [1271]. clásicos [57]. Class [173]. Classic [251, 1750]. Classical [1858, 942, 604, 696, 349]. Classics [70, 1424, 1893]. Classification [729, 1036, 1233]. Claude [530, 920, 1169]. Claudia [871]. Claus [1098, 517]. Clausius [439, 9]. Clavelin [65]. clavicula [1053]. Clavis [337]. Clemens [40]. Clergyman [1796]. Clerks [618]. Climatology [303]. Climbing [968, 703]. Clinical [297]. Clinton [1157]. Clive [1455]. Clocks [1752]. Cloths [66]. Cloud [303]. Cloyd [803]. Club [529]. Coast [64]. Cobalt [1320]. Cobalt-Ammines [1320]. Cochrane [921]. Cockatrice [1673]. Coconut [1386]. Codebreakers [95]. Codex [1408, 566]. Codices [1134, 1546]. Coe [1500]. Coelestis [337]. Cognitive [1244]. cognoscendi [619]. Cohen [527, 479, 1528, 1601, 1189, 427]. col [842]. Colding [718]. Coleman [375, 478, 1426, 480, 1337, 39]. Coley [1739]. Colin [1565, 300, 129, 831]. Collard [1400]. Collected [1116, 812, 1901, 1249, 1914, 223]. Collection [1295, 703, 1206, 1697, 864, 657, 791]. Collections [1815, 1728, 1797, 420, 118, 245]. Collective [734]. Collectors [1599, 462]. College [669, 518, 1768, 1544, 1141, 328, 1125, 777, 826, 1200]. Colleges [607]. Collegeville [13]. Colliding [1383]. Collinder [348]. Collingwood [768]. Colloques [550, 1525]. Colloquium [415]. Colodny [345, 650]. Colonial [1286]. Colonies [222]. Color [249]. Colorado [73]. Colors [290]. Columbia [1900]. Columbian [1470]. Combe [878]. Combustion [394]. Comedy [1335]. Comet [1679]. Commandino [291]. commensurabilitate [472]. Comment [966, 861, 783]. Commentaries [983]. Commentariolus [367]. Commentary [1186, 1478, 186, 473, 1054, 1298]. Commentationes [1490]. Comments [690, 1329]. commentum [747]. Commercium [1170]. Commission [107, 442]. Commitment [1637]. Committees [1457]. Common [265, 816, 66]. Communications [64]. Communis [361]. Communities [607]. Community [246, 1367, 1685, 1663, 1319, 1421, 760, 908]. Comp [298]. Companion [1100]. Company [528]. Comparative [1153, 340, 1090]. Comparison [1441]. Compass [1726]. Competition [908]. Compilation [1644]. Compleat [426, 1911]. Complete [1656, 886, 1588]. Completeness [288, 958]. complètes [1499]. Composition [292, 687, 1369]. Compound [1672, 503]. Compounds [855, 1883, 1369]. Comprehensive [1588]. Comprising [1343]. Comptes [695]. Compton [907, 1665]. Compulsory [1862]. Computer [1247, 1093, 1936, 1937]. Computers [1006]. Comroe [1602]. Comte [1368]. comune [1497]. Concept [413, 6, 1356, 392, 267, 830, 1342, 1234, 22, 1339, 1828]. Conception [992]. Conceptions [977]. Concepts [1858, 1631, 609, 1639, 1231, 734, 557]. Conceptual [1317]. Concern [760]. concerning [1447, 238, 878]. Condenser [475]. Condillac [712]. Condit [1431, 1431]. Condition [319]. Condorcet [1108, 1143]. Cone [335]. Conference [1227]. Conferences [1623, 770]. Configuration [327]. Confirmation [587]. Conflict [301, 1372, 1580]. Congrès [695, 550]. Congreso [741]. Congress [460]. Conjectures [867]. Conjunction [46]. Conor [88]. Conquest [1420, 1073]. Conquests [1621]. Conrad [1617, 765]. Conry [1151, 1616]. Conscience [162, 869]. Consciousness [1747]. Consequences [306]. Conservation [301, 718, 1064]. Conservative [1378]. conservé [720]. Considerations [1412]. Constance [194]. Constellations [867]. Constitutive [1721]. Constructed [1858]. Construction [626, 979, 979]. constructions [758]. Consultant [635]. Contemporary [345, 1538, 830]. Content [235, 187, 1344]. Context [352, 487]. Contexts [1675]. Conti [1875]. Continuity [923, 1255, 879]. Contraction [656]. Contribution [53]. Contributions [1748, 326, 1130, 504, 453, 1588, 1662, 1101]. Contributo [1910]. Contributors [1279, 1392]. Control [1920, 305, 1798]. Controversy [1214, 172, 381, 142, 1515, 1124, 1860, 151, 846]. convegno [179]. Conversations [34, 276]. Cook [1428, 312, 1588, 984]. Cooke [518]. Cooper [1687]. Coordination [70, 855, 1424]. Cope [1273]. Copernican [367, 913]. Copernicanism [500]. Copernico [781, 1485, 1651]. Copernicu [886]. Copernicus [1553, 1765, 367, 873, 1331, 838, 1010, 1653, 781, 886]. Cord [152]. Cordemoy [878]. Core [1806]. Cornelio [1843]. Cornélis [1058, 1058]. Corner [314]. Coronary [353]. Coronelli [1325]. Corporate [1699]. Corpus [1548, 1607]. Corpuscles [630]. corpusculaire [1615]. Correction [598, 1278, 545]. Correspondance [636, 1616, 1058, 1776]. Correspondance. [1801]. Correspondence [717, 1415, 1116, 4, 1557, 1194, 1654, 1783, 485, 372, 1259, 662, 476, 8]. Corresponding [1171]. Cortisone. [385]. Cosmographical [1904]. Cosmological [1775]. Cosmology [934, 303, 1549]. Cosmos [675]. Coss [1112]. Costabel [888, 1487]. Coterie [1307]. Cotta [8, 8]. Cotta-Archiv [8]. Cotter [669]. Cottons [66]. Cottrell [1180]. Coughlan [1501]. Coulomb [755, 458]. Coulter [1465]. Count [223]. Country [1559, 125, 73]. Counts [871]. Court [944]. Courtès [1488]. Cousins [180]. Coutant [1474, 423]. Cowan [1604]. Cowles [903]. Cowpox [1696]. Cracow [732]. Cradle [1406]. Crafts [96]. Craftsmen [618]. Craig [754]. Crane [607]. Cranz [983]. Crapulli [366]. Craven [1897]. Cravens [1860]. Crawford [1825]. Creation [1297, 1674, 1042, 1568]. Creative [1312]. Creativity [1319]. Creator [723]. Creed [356]. Crellin [657, 27]. Crescas [49]. Crescini [936]. Créteil [1142]. Criminological [1123]. Crisis [1199]. Crispini [1843]. Cristiana [713]. Cristofano [1012]. Critical [1805, 398, 922, 1565, 1295, 625, 306, 945, 1689, 893, 138, 282, 496, 682, 851, 1028, 1210, 1439, 1669, 1934, 147, 47, 620, 1762, 768]. Criticism [151, 605]. critico [1308]. Critics [381]. Critique [1083, 590, 965, 233]. critiques [1705]. Cronstedt [819]. Crookes [1035]. Crop [610]. Crosland [1304, 516, 101]. Cross [1594, 460, 1191]. Cross-Disciplinary [1594]. Crowe [21]. Crowther [344, 275, 1087]. Crucial [329]. Crucible [1191]. Cruelty [795]. Crusade [1865]. Cruz [1619]. Crystal [1033, 204, 1536]. Crystals [941, 1373, 1369]. Cullen [1078]. cultura [409, 865, 1188]. Cultural [1793, 653, 138, 282, 496, 682, 851, 1028, 1210, 1439, 1669, 1934, 60]. Culture [402, 1374, 1876, 589, 1419, 655, 1526, 1118, 658, 871, 514, 1091, 736]. cumulatif [1058]. Cumulative [398, 922, 1689]. Curie [1506, 1801, 1001]. Curious [822, 1089]. Currents [285]. Curriculum [1018]. Cutright [377]. Cutting [1349]. Cybernetics [231]. Cycles [1042]. Cynthia [1569]. Cyril [1053, 896, 520, 576]. Cyrurgie [425]. Czartoryski [886]. Czechoslovak [147].

D [635, 1393, 1538, 1150, 185, 184, 375, 1375, 1152, 471, 1533, 603, 1778, 55, 923, 1239, 152, 1491, 807, 904, 1295, 703, 365, 704, 1728, 1761, 1685, 1534, 1478, 1597, 322, 439, 311, 221, 220, 1341, 710, 766, 1246, 248, 716, 1004, 1148, 61, 633, 123, 1857, 73, 1070, 798, 750, 1297]. D. [1450]. Dīn [234, 732]. Daele [1816]. Dagognet [1537, 252]. Dahl [718]. dai [1106]. Daigaku [1106]. Dain [253]. Daisy [484]. Dakota [1426]. dal [409, 1497]. d’Alembert [189, 1083]. d’Alep [1650]. Dales [1758, 982]. Dalla [1301]. Dalton [893, 220]. Dams [979]. Daniel [1748, 375, 481, 1737, 1552, 1663]. Daniels [1190, 402, 1144]. Daniloff [675]. Dann [1345]. dans [1650, 1889, 1547, 622, 1472, 758]. D’Arcy [1311]. dargestellt [1291]. Darnell [1099]. Dartmouth [1088]. Darwin [1793, 1658, 455, 381, 1561, 1616, 1914, 1312, 1711, 774, 1131, 392, 854, 642, 1217, 1331, 1014, 1914, 1706, 1065, 486, 1569, 72]. Darwinian [1019, 38]. Darwinism [481, 818, 1153]. Darwinisme [1151]. d’As-Samaw’al [1105]. Dass [1403]. Data [1044, 235, 10]. Date [292]. Dating [506, 838]. datis [336]. Daubenton [856]. d’Augustin [1499]. Daumas [354, 623]. Davaine [133]. Davenport [658]. David [277, 1733, 1595, 1848, 715, 1432, 364, 649, 1067, 708, 1296, 1693, 95, 306, 1282, 266, 1195, 558, 1284, 1018, 1289, 1549, 1460, 1836, 1346, 1926, 383, 1394, 1526, 1699, 894, 749, 1352, 1234, 976, 1471, 470, 249, 56, 1073, 361]. Davies [1687, 25]. Davis [889, 369]. Davy [440]. Dawn [48]. Dawnseekers [1179]. Dawson [379]. Day [1797, 1218, 370]. Days [582]. Dazzi [1890]. DDR [1714]. Deacon [692]. Death [310]. Debate [1278, 674, 1212, 624]. Debates [592]. Debus [1875, 624, 1011, 1357]. débuts [622]. Decay [25]. December [1203, 1680, 919, 1809]. décembre [1142]. Décembre. [886]. Dechend [122]. décisions [1108]. Decker [1786]. Declaration [1492]. Decline [407, 592, 882, 1049]. Dedekind [1502]. Dedicated [1874]. Dedication [1031]. Dee [1300, 291, 595, 1807]. Deely [866]. Deep [1753]. Definition [914]. Definitive [1298]. degli [409]. degrees [566]. Del [1690, 254, 741, 1843, 756, 1355, 1707, 179, 1188, 1495]. Delegates [460]. Delivered [716]. Dell [1243]. della [1890, 1690, 1843, 1878, 1658, 465, 1846, 179, 409, 936, 313, 433, 1495]. dell’anno [1485]. Dell’Astronomia [822]. dell’illuminismo [665]. dell’Italia [865]. dell’opera [1910]. Dellow [1659]. Delno [1467]. Demandt [358]. Democritean [335]. Demography [100]. Demonstration [1126]. Demotic [705]. Denken [806]. Denkweisen [1332]. Denys [1875]. Deo [736]. d’Épicure [746]. déraison [1611]. Derangement [878]. Derbyshire [1070]. Derek [1281, 1454, 1594, 1247]. Derivations [149]. Descartes [527, 434, 524, 604, 1083, 1215, 1770, 935]. Descent [1019]. Described [101]. Description [817]. Desideratum [1421]. Design [1604, 1699]. Designing [1720]. Desmond [1599]. Destruction [1318]. Details [979]. Determination [1514, 687]. Deterrence [1571]. detto [839]. Deutsche [43, 1698]. deutschen [1785, 8, 1698, 418]. Deutscher [947, 170]. Deutschland [562, 1068, 932, 1684]. Deutschlands [1066]. Development [247, 1062, 1538, 1432, 1381, 1339, 272, 1417, 1623, 1193, 1342, 1241, 1741, 1697, 762, 56, 1400, 1272, 1813, 656, 441, 609, 1136, 1639, 531, 646, 489]. Developmental [1373]. Device [234]. Devices [223, 1251]. Dewey [1501, 1158]. Dewhurst [832, 297]. d’explication [1615]. d’Histoire [550, 601, 694, 512, 1525, 700]. D’Holbach [1307]. Diabetestraktat [259, 546]. Diagram [337]. Diagrams [1904]. Dialectical [1271, 949]. dialektika [55]. Dialog [563]. Dialogue [1412, 1277]. Diamond [1100]. Diana [607]. Diane [1160]. Diaries [1789]. Diary [635, 1659, 485]. Diaz [1128]. Dibble [1260]. Dibner [1031, 1388]. Dickinson [351]. Dickreiter [1056]. Dicks [123]. Dickson [23]. Dictionary [613, 864, 1106, 1106, 1599, 211, 738, 1453]. Didactic [1254, 1126]. Diego [745]. dien [617]. Dieter [1015, 1590, 562, 1902]. Dietlinde [1097, 1182, 1899]. Dietrich [1655, 376]. Dietz [374]. Differential [290]. Differenzierung [1563]. Diffraction [941, 47]. Diffusion [1922, 914, 607]. Digital [1006]. Dijksterhuis [261]. Dilemma [774, 335, 162]. Dilke [706]. Dimension [1164, 1580]. Diminution [34]. Din [1479]. Dinosaur [1720]. Dinosaurs [853]. Diocles [1550]. Dionysius [979]. Diophante [1396]. Diplomacy [141]. direction [1770]. Directions [633]. Directorships [669]. Disciplinary [1594]. Discipline [452]. Disciplines [1819]. Discontinuity [923, 1812, 879]. Discorso [1555]. Discours [775]. Discourse [36, 878]. Discourses [225]. Discoverer [203, 765]. Discoveries [853]. Discovering [300]. Discovers [1737]. Discovery [1601, 1514, 242, 1196, 1165, 855, 201, 578, 1883, 1629, 1483, 1719, 1763, 1456, 412, 3, 1064, 416, 1602]. Discussione [842]. Discussions [415]. Disease [1073, 353, 417, 1192]. Diseases [1052, 1539]. d’Isis [1671]. Disordered [253]. Disorders [760]. Dispersion [914]. Disputes [1582]. Dissemination [1386, 334]. Dissertations [435]. Disturbances [1309]. Divergent [856]. diversis [258]. Diversity [1401]. Divided [1465]. Divine [1050]. Division [112, 800]. Divisionibus [291, 595]. djen [1483, 1763]. DNA [1601, 1622]. do [342, 660]. Dobzhansky [1138]. Doctor [761, 1055]. Doctorates [1226]. Doctors [1466, 742, 864]. Doctrine [1487]. Doctrines [1052, 308, 827]. documentaires [1705]. Documentary [1425, 1172]. Documents [537]. Dokumente [1929]. Dollars [159]. Domain [650]. Domandl [621, 621]. Domestic [19]. Dominick [629]. Don [807, 1818, 73]. Donald [82, 1694, 1014, 1156, 108, 41, 1251, 1736, 177, 944]. Donaldson [475]. Donato [151]. Donauraum [932]. Donder [720]. Donders [320]. Donkersloot [937]. Donkersloot-de [937]. Donna [1373]. Donnet [1552]. Donovan [1511]. Dörflinger [1489]. d’organisation [1889]. Doris [999]. Dorotheus [1471]. Dorothy [719, 1194, 656, 99, 1055, 990]. Double [1543, 1120]. Douglas [1185]. Dowling [1621]. Dozenten [1017]. Dr [314]. Dr. [99]. Drafts [1793]. Drag [780]. Drake [375, 1726, 264, 1412, 964]. Drawing [817]. Drawn [1119]. Drazen [124]. drei [43]. Dressendörfer [1892]. Drew [1432]. Drift [632]. Drinker [1273]. Dritte [1330]. Drosophila [956]. Drug [1798]. Dualistic [50]. Duane [184, 603, 1728]. Dubbey [1787]. Dublin [265]. Dubos [1622]. Duc [892]. Dudley [1149]. Dugac [1502]. Dugès [132]. Duhem [1507]. Duhem-Quine [1507]. Dukas [723]. Duke [1214]. Dukes [921]. Dumas [1063, 1635]. d’un [227]. Dunbar [1915]. Duncan [107, 1071, 1553]. d’une [410, 1839, 631]. Dunn [1579]. Durbin [1644]. durch [572]. during [674, 1041, 187, 1597].

Durkheim [945, 629, 845]. Dutrochet [1262]. Dwight [1614]. Dyadik [1061]. Dyeing [66]. Dyers [66]. Dying [1235]. Dykhuizen [1158]. Dynamic [242, 854]. Dynamics [888, 338]. dynamique [1615]. Dynamische [360]. Dynamism [736]. Dyos [899]. dzialalno’sci [660].

E. [1318]. Eagle [1569]. Earle [1905]. Earlier [693]. Earliest [853, 1343, 84, 838]. Early [1773, 290, 1281, 1393, 1518, 1865, 1286, 957, 1876, 100, 1131, 1164, 1302, 1806, 1176, 1600, 1320, 1156, 817, 1406, 439, 1517, 456, 1744, 215, 289, 1527, 1519, 1294, 1349, 840, 696, 64, 1406, 355, 1771, 826, 132, 1340, 941, 123, 791]. Earth [1777, 25, 1597, 303, 590, 1441, 815, 1150]. Earthquakes [6]. Earths [819]. East [1406, 528, 1764, 1752]. Eastern [253, 1336, 1836, 1167]. Eaton [910]. Ebenezer [544]. Eber [84]. Eberhard [1111, 1096]. Ebied [363]. Ebison [1532]. Eclipse [129]. éclipses [1552]. École [1175]. Ecological [1092]. Ecology [1694, 1745]. economia [1658]. Economic [797, 461]. Economy [1694]. ed [1658]. Eddy [1312]. Edelstein [66]. Edge [1926, 1349]. Edgerton [1253]. Edifice [1775]. Edinburgh [80, 167, 1825]. Edited [303, 521]. Edith [928, 761]. Edition [1588, 625, 1478, 1298, 307]. Editions [187, 983]. Editor [821, 1389, 597, 1585, 1039, 1724, 784, 862]. Editorial [1626]. Edmond [1720, 1145, 129]. Edmund [1426, 99, 1055, 1345, 1217]. Edouard [763]. Edsall [383]. Eduard [24, 171]. Education [624, 904, 868, 1855, 576, 1444, 185]. Educationalist [1796]. Edward [1372, 385, 1273, 635, 131, 1793, 1588, 1019, 455, 472, 828, 1248, 424, 367, 894, 442, 983, 1118, 987, 1840, 1443, 370, 584, 1916, 1696]. Edwin [832, 1197, 152, 552, 930, 581, 73, 834, 782, 1160]. Edythe [761]. Effect [1665]. Egg [1807]. Egypt [1604, 1103, 491]. Egyptology [470]. Ehrenfest [158]. Ehrenfried [1110, 1844]. Ehrenwald [1463]. Eichenlaub [1474]. Eicke [1590]. eigenen [295]. Eight [183]. Eighteenth [1773, 890, 1768, 569, 458, 1911, 167, 68, 987, 662, 1078]. Eighteenth-Century [458, 987]. Eighth [682]. Eiko [1567]. Eileen [354]. Einf [1345]. Einführung [1640, 1841]. Einheit [806]. einigen [841]. Einschluss [1291]. Einstein [74, 134, 723, 874, 636, 927, 1676, 1385, 1062, 408, 1020]. Eisele [1919]. Eisenhower [1686]. Eleatics [1346]. Electric [285]. Electrical [388, 1034, 1751, 1337, 204]. Electricians [334]. Electricities [1032]. Electricity [467, 166, 1033, 1337, 1448, 1230]. Electrochemical [1675]. Electromagnetic [641, 113, 958]. Electron [64]. Electronics [615]. Electrophoresis [53]. Elektrizität. [752]. Elementary [1533, 609, 915, 396]. Elements [445, 203, 1300, 1919, 196, 1648, 1294, 302, 1758, 186, 349, 58, 700]. Eleventh [596]. Eleventh-Century [596]. Elgood [520]. Elias [1145, 1613]. Elihu [717]. Elio [842]. Elisa [551]. Elisabeth [1098, 1780]. Elisée [1915]. Elite [1735]. Elixirs [1483]. Elizabeth [1861, 1643, 1312, 484, 961]. Elizabethan [46]. Elkana [1817, 1173, 1064]. Elkin [1159]. Ellen [1230, 1505].

Ellenberger [242, 1907]. Ellery [537]. Elliot [1159]. Elliston [1733]. Ellsworth [991]. Elmer [796, 628]. Eloge [1809, 1446, 82, 779, 594, 1868, 961, 1168, 1275, 1384, 1680, 171, 1324, 505, 999, 1521, 397, 1520, 1128, 1634, 236]. Elusive [872, 1829]. Embriologia [1846]. Emergence [1427, 890, 1169, 1176, 1304, 1926, 943, 1819]. Emil [1490]. Emile [1800, 629, 945]. Emilio [278, 1355]. Eminence [167]. Eminent [1927, 1145, 1392]. Emmanuel [1484]. Emmerson [868]. Empedocles [815, 115]. Emphasis [1597]. Empire [756, 190, 365]. Empirical [879]. Empirismo [1878]. Empty [596]. Encounter [1463]. Encounters [276]. Encyclopédie [1489]. Endless [1278, 1212]. Endocrinology [670]. Endowment [1279, 1443]. Enemies [1829]. Energieverteilung [273]. Energism [85]. Energy [107, 1571, 1566, 908, 718, 901, 1064, 988, 1372, 1339]. enfermedad [884]. Engelhardt [1655]. Engine [1833, 475]. Engineer [628]. Engineering [1393, 1479, 458, 617, 322, 388, 755, 1757, 868]. Engineers [1923, 613, 1341, 613]. Engines [355, 1858, 833]. England [889, 1343, 1911, 1193, 1920, 761, 685, 1292, 1753, 904, 1018, 1417, 948, 1255, 1357, 659, 431]. English [649, 1422, 88, 989, 20, 1772, 1478, 1351, 1539, 1298, 1106, 473, 19]. Engrav’d [756]. Enlightenment [1307, 1258, 296, 803, 1016, 189]. Enquiry [1503]. Enrico [119, 278]. enseignement [1102]. Ensign [384]. Enterprise [1310, 92, 1791]. Entitled [337]. Entomology [877]. entre [1616]. Entstehung [341]. Entstehungsdatum [548]. Entwicklung [1096, 874, 1929, 418, 1068]. Environment [760, 1860]. Environmental [1604]. Enzyme [541]. Eozoön [209, 48]. Epicurean [1808]. Epidemic [1618]. Epidemics [1829]. Epilepsy [469]. Episode [1741]. Episodes [875]. epistolae [1490]. Epistolary [1783]. Epistolicum [1170]. Epistolyarnogo [1783]. Epitomised [337]. Epoch [1193]. Epochen [43]. Epstein [1180, 1180]. Equivalent [286, 235, 1129]. Equivalents [966, 967, 545, 783]. Era [101, 50, 1710, 215]. Erasmus [1014]. Eratosthenes [12]. ère [631]. Erfahrung [1332]. Erhard [950]. Eri [1284]. Eric [1411, 1281, 1013, 671, 938, 808, 1659, 1539, 56, 662, 921]. Erich [6, 1909, 271]. Erika [1066]. Erkenntnis [1332]. Erkenntnisgrund [619]. Erna [1656, 483, 891, 886]. Ernest [829, 1882, 1249, 878, 448, 845, 1155]. Ernst [481, 1503, 1139, 752, 1818, 1783, 328, 841, 1401, 170, 413, 479, 897]. Erratum [966]. Error [1503]. Errors [83]. Erwin [218, 1727, 1180, 1849, 357, 1387]. Erziehung [621]. espagnole [128]. España [741, 57, 1045, 1551, 1782]. española [1707, 570]. esprit [1770]. Esquisse [410]. Essai [1079, 561, 1063, 1108, 758]. Essay [712, 755, 935, 1693, 1395, 272, 739, 1501, 122, 542]. Essays [345, 352, 1543, 1184, 297, 243, 1886, 492, 768, 829, 1729, 574, 1643, 1528, 600, 1911, 1289, 554, 1507, 1011, 1135, 1382, 736, 415, 1813, 1295, 1901, 678, 1461, 357, 1401, 15, 1554, 1603, 1743, 1688]. Essences [643]. Essential [1368, 1730]. est [750]. Establishment [879]. Establishments [1560]. estestvenno [1813]. estestvenno-nauchnykh [1813]. Estimate [969]. estudio [740]. Eternal [1042]. Ethel [864]. Ethereal [632]. Ethnologie [759]. Ethnology [357, 1464]. Etienne [712, 1615]. Etnografii [1236]. Etude [252, 1839]. Etudes [1705, 694, 601, 512]. Euan [1835]. Eucharius [1905]. Euclid [1300, 186, 1648, 233]. Euclidean [690, 1294, 393]. Euclides [61]. Eugene [771, 882]. Eugenics [454, 1395, 1924, 108]. Eugenio [151, 409]. Euler [247, 1170, 432, 662, 73]. Euler-Mayer [662]. Euleri [1490]. Eulers [1558]. Eulner [928, 1875, 418]. Eupalinus [168]. Europa [1851]. Europe [1783, 41, 56, 1304, 1527, 1683]. European [1887, 1123, 1752]. Evaluation [1432, 1054, 722]. Evans [384, 1746, 384, 236]. Evelyn [1745]. Evening [225]. Events [306]. Everett [881]. Evidence [486, 689, 1592]. evolución [109]. Evolution [956, 1434, 734, 242, 1080, 861, 297, 458, 24, 558, 21, 996, 1597, 303, 482, 766, 130, 1022, 1860, 1294, 537, 866, 1401]. Evolutionary [1524, 1273, 1214, 866]. Evolutionist [1434]. Evropy [1783]. évszázada [1457]. Ewan [371, 371]. ex [736]. Exact [749, 1352]. Examination [655]. Examined [1305]. Excellence [1024, 1204]. Excitement [1927]. Exhaustive [1649]. Exhibition [929]. Exiled [88]. Existence [392]. Expedition [71]. Expeditions [1426, 177, 944]. Experience [1221, 430, 583]. Experiment [690, 1257, 593, 393, 1126, 1218]. Experimental [587, 1320, 1836, 964, 920, 93]. Experiments [290, 632, 1265, 428, 329, 394, 308, 1052]. Expertise [1791]. Explanation [1672, 739, 789, 696, 414]. Explication [337]. Exploration [1336]. Explorations [1896, 251]. Explorer [1712, 1504]. Explorers [1880]. Exploring [251]. Expositio [79]. Extant [84]. exterendum [1442]. External [89]. Extra [1076]. Extra-Sensory [1076]. Extractions [165]. Extracts [485]. Extramission [1810]. Extraordinary [896]. Extrascientific [392]. Exzerpte [835]. Eye [1532]. Eyles [1634].

F [1595, 64, 379, 242, 446, 59, 1858, 28, 877, 1827, 804, 444, 1470, 1825, 1886, 306, 678, 1153, 1761, 1206, 1643, 1621, 1930, 1879, 1646, 819, 1285, 1699, 894, 1544, 983, 1022, 1575, 718, 320, 579, 974, 23, 677, 476, 75, 1051]. F. [1579, 1019]. F.R.S [1364]. F.R.S. [191, 160]. Fārābī [1810]. Fa [1762, 396]. Fabricius [835, 1606]. Fabrics [1373]. Face [1543]. facie [563]. Facsimile [598, 307, 437, 1059, 886, 337]. Factors [89]. Faculty [1742]. Fagen [1393]. Failure [205]. Faith [298, 92]. Faksimile [1765]. Falconry [124]. Fall [443, 55]. Falling [469]. Fallstudie [1563]. Familie [858]. Family [703, 968, 858]. Fantasies [298]. Faraday [1062, 1675, 1129, 436]. Farber [24, 171]. Faris [676]. Farming [1753]. Fascination [1927]. Fate [1632]. Father [1217, 988, 1363]. Faye [1876]. Featherstonhaugh [668]. Federico [902]. Federn [1139]. Feedback [420]. Feisenberger [1145]. Feld [812]. Feldberg [1886]. Félix [622]. Fellmann [1490, 1147]. Fellows [804]. Fenichel [1749]. Ferdinand [621]. Ferguson [105]. Fermat [786]. Fermi [278]. Fernando [109]. Ferrari [1410]. Ferrier [487]. Ferry [1429]. Fertey [375]. Festgabe [928]. Festschrift [1824, 1892, 759, 1098]. Feudal [90]. Feuer [1020]. fi [1761, 1251]. Fichant [346]. Field [830, 1633, 1089]. Fields [1373, 1062]. Fiery [46]. Fierz [1652]. Fifteenth [731]. Fifteenth-Century [731]. Fifth [138]. Fiftieth [995]. Fifty [1573, 275]. Fighting [1621]. fil [1252]. fil-Hisab [1252]. Filarete [839]. filosofi [663]. Filosofia [665, 313]. filosofica [895, 1690]. filosofo [1495]. Final [1887]. Finding [1013]. Fine [886, 960]. Finoli [839]. Fire [1727, 165, 423]. Firenze [465]. Firmament [20]. First [820, 669, 1793, 253, 18, 1244, 1742, 196, 1686, 394, 513, 187, 1836, 1179, 1210, 1605, 186, 960, 506, 1871, 1059, 473, 399, 328, 1362, 337]. Fischer [713]. Fisheries [1008, 1880]. Fishes [1797]. Fisica [1407]. físico [1355]. físico-matemático [1355]. Fisiologia [1846]. Fitch [1491]. Fixed [1447, 687]. Flacius [619]. Flack [1421]. Flag [944]. Flamsteed [1411]. Fleming [41]. Fletcher [1579]. Flew [1729]. Florence [1312, 921, 892, 921]. Florentines [921]. Florenz [1877]. Florey [1205]. Florimond [1487]. Florkin [796, 1744]. Flowers [1706, 1629]. fluviale [713]. Fluxions [501]. folgen [1403]. folie [1891, 1611]. Followers [1140]. fondements [1502]. fonds [720]. Fontenelle [1416]. fonti [179]. Food [1798]. Forberger [1875]. Forbes [1281, 1411, 1582, 1013, 938, 126, 985, 310, 127, 662]. Force [635, 1062, 1172, 456, 583, 1002, 338]. Forces [1721, 1082, 1583]. Ford [1070]. Förderung [1684]. Foremost [761]. Forgotten [921]. Form [870, 1536, 480, 39]. Forman [1560]. Formation [1367, 940]. Formative [1069, 1500, 1855]. formazione [895, 1555]. Formed [1689, 922]. Forming [514]. Forrest [1152]. Forscher [1857]. Forschung [750]. Forschungsförderung [43]. Forschungsgemeinschaft [43]. Forster [1358]. Fort [1375]. Fortschritts [1816]. Fossil [5]. Fossils [875, 1597]. Fothergill [314]. Foucault [553]. Foulkes [1503]. Found [661]. Foundation [1278, 1858, 847, 637, 847, 1212, 1510]. Foundations [247, 1748, 1221, 132, 1806, 1178, 1123, 143, 1573, 1225]. Founder [70]. Founding [1275]. Four [669, 28, 319]. Fourier [625, 1201]. Fourteenth [659]. Fourth [1035, 1218, 1934]. Fowler [807, 73, 807, 73]. Fowlers [534]. Fowles [894]. Fox [1338]. fracturi [1902]. Fragments [1003, 169]. Fraisse [93]. Frame [122]. France [1151, 625, 1195, 1616, 561, 1193, 757, 1125, 101, 127, 156, 1719, 1102, 458]. Frances [1016]. Francesco [1875, 1301]. francese [665]. Francis [107, 1071, 1554, 1422, 428, 1488, 510, 454, 1281, 1152, 88, 1797]. Franciso [1138]. Franco [1843, 362]. Francois [58, 560, 1537, 252, 1841, 1907]. Frängsmyr [890]. Frank [1656, 767, 1692, 1491, 35, 1836, 948, 311, 710, 380, 1068, 791, 180, 1656]. Franke [1761]. Frankenstein [216]. Frankfurt [949]. Frankl [28]. Franklin [1601, 1380, 1060, 1263, 1475, 1448]. Franklinist [1380]. Frankreich [562]. Frankreichs [1066]. Franz [522, 474, 502]. französischen [1413]. Fraser [1285]. Fráter [1457]. Fraunhofer [1857]. Fred [1100, 1464, 1160]. Frederic [1169, 1930, 858, 1506]. Frédéric-Louis [858]. Frederick [1827, 1264, 1740, 1788, 1046, 94]. Freeman [715, 1561]. Frémont [177, 944, 177]. Frémont. [944]. French [527, 1335, 756, 1608, 1419, 34, 1369, 943, 1562, 1319, 960, 1671]. Fresh [1402]. Fresnel [780]. Freud [810, 1312, 906, 1235, 1140, 1786, 1365]. Freudian [769]. Friday [225, 1116, 1171]. Friedrich [1657, 316, 193, 1177, 302, 1101, 590, 1785]. Friendship [314]. Frier [878]. Frihetstida [890]. Frisi [713]. Frisinger [1641]. Frison [244]. Frithjof [531]. Fritjof [1336]. Fritz [1057, 1384, 360, 707]. Froderberg [193]. Frolov [55]. Front [1, 45, 77, 111, 137, 140, 164, 199, 230, 281, 284, 331, 390, 451, 495, 498, 539, 586, 639, 681, 684, 727, 773, 814, 850, 852, 912, 954, 994, 1027, 1030, 1075, 1122, 1162, 1209, 1211, 1268, 1316, 1377, 1438, 1440, 1513, 1578, 1625, 1668, 1670, 1717, 1804, 1864, 1933]. Frontier [105, 1278, 1212]. frühen [255, 1112, 98]. Frühgeschichte [63]. Fruton [600, 1044]. Fryer [69]. Führer [572]. Full [794]. Fullmer [440]. Function [1601, 345, 480, 832]. Functional [974]. Functions [640, 1052]. Fund [1279]. Fundamenta [316]. Funding [1560]. Furley [471, 1346]. Furner [1199]. Furrow [1753]. Further [966, 544, 967, 783]. Fury [846]. Fusion [1365]. Fussell [610]. Fusul [1252]. Future [1458]. Fyfe [627].

G [1411, 1824, 1837, 652, 1281, 107, 515, 767, 1071, 1850, 502, 1739, 532, 345, 1856, 650, 624, 1202, 1861, 312, 1013, 870, 1147, 344, 1370, 1833, 1008, 1745, 1180, 661, 1159, 275, 1111, 1395, 154, 1053, 1355, 1757, 938, 326, 1822, 1230, 989, 1529, 768, 161, 789, 153, 874, 673, 1579, 1882, 1922, 1883, 1710, 1476, 1061, 939, 668, 1846, 1285, 1906, 1507, 1584, 1404, 610, 1011, 602, 1298, 941, 1357, 1550, 1498, 763, 952, 1087, 987, 1493, 1505, 662, 1250, 1875, 1852, 510]. G. [719]. Gad [31, 887]. Gadfly [715]. Gaisinovich [1713]. Galaxies [1737]. Galaxy [416]. Gale [1738]. Galen [836, 30, 1052, 257, 308]. Galenic [1678]. Galenism. [1049]. Galeno [884, 798]. Galens [835, 1403]. Galilean [859, 1723]. Galilée [892, 927, 65]. Galilei [874, 1726, 1308]. Galileo [286, 587, 1080, 1126, 643, 114, 1726, 587, 1132, 1221, 1447, 593, 963, 1218, 293, 1277, 1769, 1556, 1012, 1113, 1910, 264, 805, 1412, 429, 1726, 1149]. Gall [487]. Gallehus [548]. Gallwey [833]. Galton [1152, 1422, 454]. Game [1242]. Gamgees [1311]. Gamma [1036]. Gamow [197]. Garcia [884, 1766, 1782, 798, 1893, 1551]. Garden [99, 1418]. Garin [409]. Garrett [942]. Garrison [305]. Gartrell [1230]. Gary [1915]. Gaseous [1718, 289]. Gases [9, 237, 1338, 1498]. Gasking [961]. Gasman [481]. Gassendi [754]. Gasser [260, 1299, 1851]. Gaston [1616, 573, 908]. Gastric [370]. Gatzemeier [255]. Gaudin [957]. Gaus [690]. Gauss [193, 393, 1657]. Gautier [1907]. Gave [1205]. Gavioli [551, 551]. Gay [1635]. Gay-Lussac [1635]. Gears [1937, 1936, 1247]. Gebbett [1796]. Gebrauch [98, 1607]. Geburtstag [1098, 1892, 928]. Geck [1098]. Gedo [1365]. gedruckte [572]. Gegenwart [1831]. Geiger [689]. Geillustreerde [244]. Geist [841]. Geldsetzer [619]. Genera [525]. General [1288, 1125, 1313]. générale [1262]. Generation [674, 1922, 302]. Generations [1020]. Generazione [1846]. genèse [927]. genesi [366, 1495]. Genesis [106, 530, 1459, 470, 1297]. Genetic [1601]. Genetics [1664, 468, 697]. Genetika [55]. genevoise [1930]. Genius [129, 1152]. genombrott [890]. Genovesi [895]. Gentleman [1160]. genus [858]. Geocentricity [1807]. Geochronological [1326]. Geoffrey [1924, 880, 18, 1399, 742, 1183, 1829, 991]. Geografii [1783]. géographe [128]. Geographen [1785]. Geographical [267]. Geographie [1489]. Geographies [1289]. Geography [267, 169, 1783, 1885, 351, 1186]. Geological [774, 969, 1326]. géologie [1907]. Geologist [455, 435]. Geology [1428, 1777, 455, 1742, 1379, 668, 303, 1584, 2, 1363, 542, 602, 62, 488, 1743, 105, 1742]. Geometric [1726]. Geometrical [1321, 1288, 7]. Geometrie [1300]. géométrique [1615]. géométriques [758]. Geometry [1144, 608, 857, 1294]. Geomorphology [25]. Georg [256, 376, 1345, 1875, 1630, 238, 33]. George [1492, 1190, 1115, 1265, 1184, 1887, 1885, 382, 478, 753, 903, 329, 910, 70, 1424, 868, 411, 1743, 485, 1365, 1539, 1312, 466, 1158, 197, 615, 89, 537, 403, 507, 823, 402, 996, 452, 1428]. Georges [827, 811, 512, 1930]. Georgii [1535]. Geosciences [1597]. Geosophy [1289]. Gerald [760, 1916, 795]. Geraldine [131]. Gérard [1819, 891]. Gerardi [1270]. Géraud [878]. Gerd [604]. Gerhard [1590, 654]. Gerhardt [1844]. Gericke [217, 1841]. Germ [1318]. Germ-Layer [1318]. German [481, 1258, 182, 1679, 1271, 8, 1721, 858, 50, 170]. Germany [1786, 127, 141, 540, 1866, 1788]. Gernot [1816]. Gerson [1350]. Gerstner [929]. Gert [667, 1607, 1540, 928]. Gertrud [1368, 812]. Gesammelte [522]. Gesamtausgabe. [1765]. Gesamtinterpretationen [1414]. Geschichte [1892, 43, 255, 1414, 1755, 1061, 707, 1097, 702, 1332, 1714, 1015, 1847, 654, 217, 214]. geschichtlichen [1096]. Gesellschaft [1684, 759]. gesellschaftliche [1816]. Gesellschaftstheorie [1414]. Gestalt [1632]. Gestalten [654]. Geza [769]. gg [1713]. Ghalioungui [1103]. Ghanem [1409]. Ghiselin [38]. già [1703]. Giants [1185]. Gibbs [1204, 1700, 485]. Gibson [716, 1927]. Gifford [669, 1492]. Gilbert [292, 1584]. Gilette [1801]. Gille [1832]. Gillispie [315]. Gimpel [1609]. Gioanventura [66]. Giordano [934, 188]. Giorgio [842, 150, 122, 1128]. Giornali [1846]. giovanili [1495]. Giovanni [366]. Giove [842]. Giovio [1875]. Girolamo [1652]. Giuliano [1658]. Giuseppe [1555, 900]. Gizycki [1563]. Glacier [1582]. Glaciers [36]. Gladys [1891]. Glasgow [716]. Glass [1904, 979, 657, 657]. Gleanings [286, 587, 1080]. Glenn [674, 186, 1541, 75]. Glick [1153]. Glimpses [365]. Globe [1325, 1911]. Globi [212]. Glory [71]. Gnomonics [245, 180]. Gnudi [1496]. Goal [958, 1671]. Goals [16]. Gods [1882, 795]. Goerke [714, 67]. Goethe [764]. Gold [1763]. Goldammer [1875]. goldenen [74]. Goldhörner [548]. Goldin [1239]. Goldschmidt [1761]. Goldsmith [1506]. Goldstein [748, 1350]. Goldstine [1823, 1093, 794]. Goltz [1899, 1097, 1182]. Good [1539]. Goodison [19]. Goodman [1728]. Goodrich [584]. Goodstein [1206, 489]. Gordon [25, 904, 1177, 1009, 1435, 1894, 1906, 1544, 251]. Görgemanns [563, 563]. Gosse [1217]. Gotland [1494]. Gottfried [1330, 1844, 1842]. Government [1023]. Gowing [1571]. Gr [1530]. Grace [1239]. gradibus [1348]. Grafton [1451, 1159]. Graham [1420, 580]. Grahame [1600]. Grand [921, 133].

Grandfather [964]. Grant [1720, 472, 1248, 424]. Graph [10, 1691]. Graphic [1041]. Grassi [839]. Grattan [625, 247]. Grattan-Guinness [625, 247]. Grau [1617]. Graunt [10]. Graves [1366]. Gravity [1002]. Gray [1732]. Great [598, 714, 1334, 1361, 1659, 437, 485, 1697, 66, 1515, 729, 1277]. gréco [1547]. gréco-latine [1547]. Greece [96]. Greek [616, 1898, 1826, 1246, 836, 123, 1294, 1700, 646, 29]. Greeks [469, 1247, 1937, 1936]. Greene [1852, 1225]. Greeno [789]. Greenwich [1281, 1281]. Gregor [1171]. Gregorii [532]. Gregory [1886, 1367, 1788]. Grene [554]. Grensemann [359, 1548]. Gresham [1411]. Griechen [1473]. griechischen [256, 1182, 549]. Griesinger [81]. Grmek [124]. Grob [1916]. Grossbetrieb [1795]. Grossbritanniens [1066]. Grosse [302]. Grossen [1617]. Grosseteste [1407]. Group [1375]. Growth [463, 1110, 605, 1638, 797]. Groza [609]. Gruber [1065]. Grünbaum [1021]. Grunder [1590]. Grundlagen [753]. Grundprobleme [707]. Gründung [1684]. Gruppenbegriffes [106]. Guenter [876]. Guenther [574]. Guerlac [823, 1688]. Guglielmo [1910, 1868]. Guicciardini [1875]. Guide [1888, 1172, 1422, 737, 657, 1695, 1206]. Guido [1542, 1098]. Guillermo [1782, 1893]. Guinness [625, 247]. Gukov [1881]. Gulick [996]. Gundel [256]. Gunnar [1815, 1356]. Gunns [1206]. Gunter [1918, 928, 130, 1892, 1929, 1857]. Gunther [447, 447, 1154, 1785, 104]. Gupta [565]. Guralnick [1200]. Gustav [435, 1166]. Gustina [311, 710]. Guterman [699]. Gutt [1564]. Guy [425, 1455]. Gwei [617, 1483, 1763]. Gwei-dien [617]. Gwei-djen [1483, 1763]. Gweneth [666]. Gwinn [1372]. Gynaecology [520].

H [1601, 1434, 1661, 1458, 767, 1595, 218, 667, 1856, 402, 445, 1602, 669, 1065, 1185, 1641, 1067, 100, 1750, 478, 1533, 486, 603, 350, 1886, 326, 1775, 898, 1823, 1165, 796, 1728, 899, 1145, 764, 1597, 113, 1914, 939, 270, 566, 1285, 905, 182, 1365, 1791, 1369, 1539, 983, 1089, 1648, 1091, 658, 357, 1297, 677, 1540, 1093, 576, 1362, 672, 791, 1043, 1840, 794, 750]. H. [1222, 898, 989]. Ha-håkra’ah [1761]. håkra’ah [1761]. Haas [1818]. Haber [1285, 1226]. Haberer [246]. Habicht [1490, 1170]. Hacienda [254]. Hacking [1176]. Haeckel [481, 997]. haeckelii [997]. Hägg [941]. Hahm [1549]. Hahn [405]. Haimbach [1749]. Haitham [309]. Hakim [309]. Hale [537]. Half [1457, 1509]. Hall [719, 193, 1415, 1654, 1467, 1605, 372, 17, 1802, 524, 1597]. Haller [1380, 806, 482, 776, 1704]. Hallers [1909]. Halleux [1472]. Halley [1145, 129, 232]. Hallowell [1662]. Hals [564]. Hals- [564]. Hamarneh [606, 1054, 546]. Hamburg [399]. Hamilton [1860, 1269]. Hamlet [122]. Hammell [1428]. Hammer [522]. Hand [1542]. handasiyya [1251]. Handler [1455]. Handlist [1561]. Handschriften [1187, 1482, 1187]. Handwerker [1857]. Hankins [1083, 189]. Hanley [1462]. Hannah [1786]. Hannaway [1254]. Hannover [1414]. Hans [1590, 273, 1180, 947, 256, 106, 1061, 418, 1459, 1403, 1899, 928, 1875, 262, 259, 1520]. Hans-Heinz [418, 928, 1875]. Hanson [867, 412]. Hanzeit [98]. Hapgood [383]. Harald [1347]. Haraway [1373]. Harcourt [1335]. Harding [1507]. Hariot [368]. Harlow [42]. Harms [878]. Harold [717, 1007, 299, 951, 1305, 473]. Harpers [1429]. Harré [651]. Harriet [1817, 1735, 1287]. Harris [1465, 68, 836]. Harrison [894, 1894, 505, 1908]. Harry [1621]. Hart [1737, 1157, 1570]. Hartkopf [1714]. Hartley [299]. Hartlib [224]. Hartman [581]. Hartner [548, 403, 1824]. Harvard [669, 1375, 383]. Harvest [1532]. Harvey [666, 869, 1181, 1456]. Hasan [1903]. Hasdai [49]. Hashish [474]. Hassan [1479]. Hassler [1014]. Hauksbee [428]. Hauptschriften [1061]. Hausen [1545]. Hawaii [1232]. Hawke [715]. Hawkins [441]. Hawthorne [1053, 1925]. Hayes [1022, 323]. Headlines [534]. Headquarters [635]. Heads [534]. Headwaters [375]. Healing [1542, 1228, 1463]. Health [1354, 1184, 365, 1784, 1012, 1862, 1895, 1505, 1895]. Heaps [1672]. Hearing [308]. Heart [353, 417, 836, 153]. Heat [290, 1858, 625, 223, 1498, 1129]. Heauen [894]. Heavenly [1553]. Hebdomadon [616]. Hebrew [31]. Hector [66]. Hedwig [134]. Hegel [1655, 378]. Heidegger [464]. Heidelberg [1795]. Heilbron [1560, 989]. heilige [359]. heiligen [619]. Heilkunde [1182]. Heilsystems [1291]. Heinar [75]. Heinrich [1789, 1590, 947, 1414, 1403, 572, 1789]. Heinroth [1309, 324, 324]. Heinz [374, 714, 746, 260, 1299, 1704, 418, 67, 33, 928, 1875]. Heischkel [928]. Heischkel-Artelt [928]. Heisenberg [276, 1508, 1572, 302]. heiten [564]. Held [225, 460, 952]. Helden [1884]. Helen [1928, 844, 1232, 723, 537]. heliocentryzm [342]. heliozentrische [549]. Helix [1120]. Hellman [18, 999]. Helm [802]. Helmuth [217, 1841]. Helped [1793]. Henderson [325]. Heninger [1904]. Henkin [1530]. Henle [977, 386]. Hennessy [354]. Henrey [1343]. Henri [1731, 242, 934, 1262, 1705, 885, 1484, 561, 1907, 631, 1262, 244]. Henry [1604, 614, 1850, 463, 1203, 375, 756, 1654, 940, 1417, 438, 896, 1688, 1419, 485, 372, 1091, 1297, 823, 854, 919, 803, 382, 266, 1794]. Hensansho [1106]. Hensen [274]. Henslow [1796]. Heraclitean [1727]. herausgegeben [1844]. Herb [474]. Herbals [1692]. Herbert [1266, 769, 1917, 467, 216, 236, 766]. Herbig [326]. Herder [270]. Heredity [956, 1123, 1860]. Heredity-Environment [1860]. Heresy [1317, 443]. Heribert [1765]. Heritage [369, 1024]. Herivel [1201]. Herman [1069, 1139, 1823, 1341, 1093, 794, 263]. Hermann [1067, 530, 449, 359, 1508, 194, 1548, 1648, 74, 702, 1572, 1632]. Hermetic [1807]. Hermeticism [1781]. Hermit [1050]. Herneck [302]. Hero [307, 168]. Herpetological [132]. Herpetology [132]. Herschel [104]. Herskovits [910]. Herstmonceux [1281]. Hertha [1896, 122]. Hertwig [112]. Hertz [1789, 1789, 166]. Herwig [563]. Hesse [1593]. Hessen [461]. Heurck [244]. heutigen [1291]. Hevelius [373]. Hewlett [107, 1071]. Heyde [1790]. Heyman [755]. Hickel [1066]. Hickok [878]. Higbie [269]. High [908, 20]. Hilbert [1067, 61, 194]. Hilfstein [886]. Hill [1255, 1251, 1363]. Hilts [1422]. him [858]. Hindi [1252]. Hindle [1771].

Hintikka [1288, 1104]. Hintzsche [1909]. Hippiatry [124]. Hippocrates [154, 473]. Hippocratic [616]. Hippocraticum [1607, 1548]. Hippokrates [1902]. hippokratische [359]. Hiram [935]. Hirosige [1324]. Hirsch [909]. Hirschmuller [1853]. Hisab [1252]. Histoire [695, 1537, 856, 1616, 560, 1889, 410, 1642, 346, 561, 58, 1832, 1102, 432]. Histology [380]. Historia [741, 57, 568, 1551, 1564, 1702, 1236]. Historian [1915, 1712]. Historic [1406, 149]. histórica [1045]. Historical [1595, 1538, 1531, 1533, 514, 1339, 152, 646, 656, 788, 945, 1005, 1528, 600, 1284, 1289, 240, 609, 1483, 516, 1238, 343, 697, 420, 864, 1340, 308, 974, 1171, 356, 1540, 1896, 121, 977]. histórico [740]. historien [128]. Historiker [260, 1299]. Historiography [688]. Historische [1414, 244]. Historisches [1590]. History [1596, 1601, 1604, 1661, 1279, 1281, 107, 515, 1469, 1815, 1465, 64, 1393, 632, 117, 93, 1100, 1237, 1748, 15, 185, 1543, 184, 242, 1746, 251, 669, 832, 1692, 17, 25, 1185, 1184, 1172, 1887, 469, 398, 590, 1641, 755, 28, 1343, 1399, 1827, 83, 1381, 1099, 1034, 822, 419, 922, 518, 978, 603, 1693, 1275, 1778, 1631, 206, 903, 611, 875, 1886, 27, 1742, 1177, 62, 1822, 467, 557, 1380, 1461, 1529, 400, 509, 699, 1802, 1823, 183, 796, 1600, 1754, 153, 825, 1282, 1689, 1728, 1510, 447, 24, 196, 299, 459, 1044, 1534, 21]. History [91, 1156, 353, 1732, 1379, 1464, 929, 1883, 949, 1603, 354, 1463, 1009, 19, 1743, 270, 1193, 1663, 406, 559, 926, 1688, 1179, 138, 282, 496, 552, 682, 851, 1028, 1210, 1439, 1669, 1934, 303, 1125, 1741, 1894, 1744, 420, 1082, 1344, 271, 371, 147, 1880, 118, 1526, 1697, 1012, 894, 1365, 795, 1234, 864, 998, 612, 716, 1453, 60, 466, 618, 1135, 1141, 1243, 1695, 1263, 1462, 1527, 633, 1895, 399, 606, 1541, 803, 1498, 615, 581, 460, 305, 1696, 1828, 300, 1073, 1540, 1869, 1349, 937, 542, 770, 921, 53, 398, 1597, 1089, 1733, 1425]. History [219, 18, 1239, 704, 868, 1608, 1751, 1089, 1501, 228, 1231, 1598, 652, 339, 877, 1290, 556, 1605, 97, 181, 577, 94, 328, 876]. Hitcham [1796]. Hitchcock [455]. Hitler [1685]. hiyal [1251]. Hoare [1358]. Hobart [132]. Hobbes [1305, 1109]. Hobbins [919]. hoc [750]. Hochschule [1795]. Hochschulwesen [1795]. Hodge [401]. Hodges [614]. Hodgkin [1886]. Hodson [1051]. Hoeniger [894]. Hoff [1261]. Hoff-Le [1261]. Hoffmann [316, 723]. Hofmann [1110, 1844]. Hohlfeld [1816]. Holbrook [623]. Hollerith [146]. Holly [907]. Holmes [1169]. Home [803]. Homeopathy [443]. homine [1109]. Homo [1356, 1356]. homogène [1839]. Honest [1089]. Honor [492, 829, 1467, 1289, 1011]. Honour [1135]. Honourable [191]. Hooke [6, 317, 1145, 1274, 317]. Hopkins [1024, 1579, 1456]. Hoppen [265]. Horizontal [587]. horloge [886]. Hormone [385]. Horský [245]. Horst [564]. Horticultural [1343]. Horticulturists [1599]. Horton [627]. Hoskin [1148, 337, 221]. Hospital [253, 1239, 1455, 1362]. Hospital-250 [1455]. Hospitals [1015]. Hossein [1610, 1481]. Hostile [144]. Hostilities [1894]. Houghton [792, 1423]. House [1103]. Hovenkamp [1917]. Howard [159, 1753, 1205, 1065, 384, 1641, 1557, 757, 1179]. Howse [1281]. Howson [1565]. Hoyt [1366]. Hua [332]. Huard [190]. Hubble [782]. Hubert [900]. Hudson [883, 1468]. Hugh [869, 1193, 1404]. Hughes [250, 628, 1828].

Hugli [1175]. Hugonnard [885]. Hugonnard-Roche [885]. Hui [1762]. Huling [659]. Hull [976]. Human [712, 1888, 1335, 319, 152, 703, 1397, 553, 216, 734, 881]. Humanism [1365]. Humanist [1299, 260]. Humanists [1556]. Humboldt [1657, 128, 531]. Humphreys [867, 412]. Humphry [440]. Hund [1177]. Hundertjährigen [759]. Hundred [820, 304, 1742, 196, 1041, 26, 1210, 1439, 1669, 1934]. Hundredth [1028]. Hungarian [1457]. Hunt [1664, 864, 1469, 1740]. Hunter [670, 1405, 1359, 385]. Hunting [1225]. Huntington [991]. Hutchinson [1600, 1745]. Hutton [1441]. Huxley [1856, 1886, 896, 997, 576, 438]. Huygens [232, 1383, 10]. Hwasser [671]. Hydraulics [1405]. Hydro [1754]. Hydro-Technology [1754]. Hydrology [556]. Hygiene [1598]. Hymes [1243]. Hypochondriack [1539]. Hypothesis [957, 1568, 1635, 337]. Hysterick [1539].

I-II [598, 437]. I. [1790]. Ian [1664, 1176, 1072]. Iatromagie [1831]. Ibn [1480, 1050, 732, 1409, 1839, 1251, 564, 1522, 1903, 1396, 309]. Ibrahim [1903]. Idea [78, 21, 1597, 366, 870]. Ideal [1614, 1488]. Idealism [1241]. Ideas [454, 1164, 392, 1176, 24, 1379, 1406, 1241, 1453, 1527, 1813, 803, 1100, 150, 1234, 1453]. idée [410]. Ideen [874]. Identification [83, 12]. Identifications [1767]. Identity [205]. idéologie [1488]. Ideologies [1271]. Ideology [490, 540, 1125]. Idiocy [442]. idraulica [713]. Idyll [251]. Ier [720]. igakushi [1141]. Ignace [573]. igne [423]. ihm [858]. ihre [1017, 214]. ihrer [1847, 1714]. II [598, 1927, 1203, 695, 1898, 1424, 1414, 223, 1346, 1348, 437, 127, 960, 864, 1527, 1137, 432, 944, 154, 749, 107, 635, 1544, 941, 1054]. III [741, 1015, 1689, 1299, 617, 223, 33, 1558, 518, 985, 1352, 1612]. ikonographischen [548]. illiministi [1497]. Illinois [306]. Illness [1312]. Illuminations [1826]. illuminismo [409]. Illustrated [832, 1692, 152, 801, 1193, 1481]. Illustrating [305]. Illustration [544, 1693]. Illvricus [619]. im [1853, 374, 947, 1684, 1480, 1473, 255, 752, 1096, 1066, 8, 1698, 1548, 1607, 806, 1293, 1558]. Imad [1409]. Image [776, 1418]. Imagery [1394, 125]. Images [899, 1310, 1731]. imaginaires [758]. Imaginary [593]. Imagination [949, 1337]. Immortality [1763]. Impact [1215, 1802]. Impenetrability [1631]. Imperialismus [1617]. Importance [1447]. Important [1883]. Imre [605]. Ina [1640, 1594]. Ina-Susanne [1640]. Including [1002, 1599, 1588]. incommensurabilitate [472]. Incommensurable [1294]. Independence [1492, 1571]. Independent [583, 1166]. Index [1588, 1423, 1003, 1058, 1923]. Indexed [132]. India [528, 97, 709, 565]. Indian [375, 1797, 34, 1130]. Indiana [552]. Indians [1426]. Indias [1702]. Indicators [1817]. Indices [1391]. Indigenous [1406]. indischen [549]. Induced [486]. Induction [1176, 1522]. Industrial [883, 306, 1609, 1597, 439, 1659, 56]. industriale [1202]. Industrialist [584]. Industrialization [1833]. Industrie [1709, 622]. Industry [1310, 1687, 1468]. Inertia [587, 1475]. Infection [1621]. Infeld [1928]. Inference [975, 1176, 1593]. Inferiority [482]. Infinite [188, 501]. Infinities [49]. Influence [454, 237, 208, 644, 50, 840, 897]. Influences [1793, 138, 282, 496, 682, 851, 1028, 1210, 1439, 1669, 1934]. Informal [1874, 1886, 1510, 197]. Information [136, 280, 494, 680, 849, 1026, 1208, 1437, 1667, 1932]. informes [827]. Ingeneis [710]. ingénieure [892]. Ingenious [1496, 1251]. Ingenuity [1241]. Inka [981]. Innovation [306, 881]. Innovations [757]. Inquiry [1144, 1592, 1406, 412]. Insanity [442]. insbesondere [564]. Inschriften [548]. insecabilibua [519]. Insects [1629, 1598]. Insights [1602]. Instauration [1361]. Institut [206, 625]. Institute [27, 1742, 306, 1206, 1579, 949, 1622, 118, 864, 1263, 306, 205, 399]. Institution [225, 1281, 1220, 1913, 737, 405, 737]. Institutional [1379]. Institutionalization [879]. Institutions [847, 1024, 537, 561, 223]. Instructions [66]. Instrument [1559, 1004]. Instrument-Maker [1559]. Instrument-Makers [1004]. Instruments [1815, 1452, 91, 939, 623, 1298, 245, 791, 1731, 91, 1287]. Insulin [1483]. Insurance [1862]. Integration [441]. Intellectual [805, 352, 711, 102, 1772, 1114, 1569, 1421, 41, 1365, 1501, 531]. Intellectuel [730]. Interaction [1173]. Interchanges [1419]. Interest [187, 743]. Interests [303]. Internal [782, 729, 89]. International [550, 144, 1818, 460]. Internationalen [1414]. Internationales [1057]. Internationalism [540]. internazionale [179]. Interplay [600]. Interpretation [1278, 1212, 114, 39, 913, 1814]. Interpretations [1767, 554]. interpretazione [1910]. Interrelations [1822, 1529]. interrelationships [170]. Intimate [1073]. introducción [1782, 57, 1355]. Introduction [956, 1745, 119, 788, 271, 147, 69, 412, 706, 1527, 1054, 1340, 1399, 1642, 427]. Introductions [1478]. Intromission [1810]. Invariance [1882]. Inventaire [1487, 720]. Inventaris [244]. Invention [1884, 306, 1483, 354, 1763, 1226, 91]. Inventions [306, 311]. Inventor [584, 628]. investigandarum [750]. Investigations [1750]. Invisible [607]. Inward [1181]. Inyokern [634]. Ireland [1343, 1697]. Irene [881, 1801]. Iris [269]. Irish [1599]. Irmer [1902]. Irregular [1874]. Irrenhauses [562]. Irving [1662]. Isaac [1415, 35, 221, 526, 1059, 1148, 1908, 476, 1189, 1495]. Isis [922, 1689, 922, 1689, 1671]. Islamic [1610, 980, 1481]. Isles [1469, 1835, 117]. Isolation [996]. Isotopes [578]. Israel [671]. Issac [268]. Issue [1574]. Issues [1371]. Istanbul [521]. istorii [1883]. Istoriia [1535]. Istoriya [1883, 1883]. Istruzione [1202]. István [1847]. Italia [1202]. Italian [1556, 409]. italiana [179]. italienisches [1900]. Itself [1882]. Itzkoff [413]. Iudif [1256]. iuvenulia [532]. IV [617, 939, 223, 1744, 476, 1530]. Ivan [55]. Iversen [271]. Ivor [247]. IX [1490]. Iz [1783].

J [1604, 254, 1281, 1469, 1687, 616, 1618, 532, 1086, 64, 717, 986, 1692, 1172, 344, 1374, 1222, 1204, 1833, 1769, 471, 479, 1180, 661, 1750, 275, 362, 419, 630, 243, 1019, 1825, 57, 1045, 1309, 1503, 875, 1757, 625, 62, 910, 1775, 126, 985, 989, 267, 1303, 1060, 158, 276, 789, 20, 102, 1761, 429, 899, 1264, 37, 1911, 21, 353, 1772, 1289, 996, 1930, 1478, 1597, 571, 26, 1536, 322, 1691, 1023, 1061, 1076, 1663, 559, 1342, 1797, 1346, 1926, 566, 1622, 1781, 1285, 950, 979, 1421, 655, 583, 127, 1404, 325, 462, 766, 216, 864, 1794]. J [1525, 22, 941, 248, 1004, 1790, 347, 1229, 61, 1550, 155, 156, 657, 157, 1899, 984, 1598, 977, 991, 1087, 261, 1493, 937, 1043, 866, 1407, 476, 401, 510, 770, 75]. J. [1787, 204]. Jaager [320]. Jaakko [1104]. Jaako [1288]. Jack [1601, 1533, 485]. Jackson [177, 402, 322, 944]. Jacob [1047, 29, 320, 560, 1772, 395]. Jacobellus [124]. Jacobus [750]. Jacques [973, 755, 1360, 318, 927, 448]. Jacquin [568]. Jaffe [302]. Jagdish [992, 1623]. Jaggi [97]. Jahren [1617]. Jahrhundert [1473, 1066, 1900, 8, 218, 374]. Jahrhunderts [1590, 1617, 1408, 1780, 841]. Jahrhundertwissenschaft [1572]. Jahrundert [752]. Jaki [1775, 872, 1042]. Jakob [1612, 511, 50]. Jakov [124]. James [1856, 1538, 1197, 28, 100, 1116, 430, 903, 664, 1686, 1710, 475, 979, 1751, 1226, 919, 1089, 984, 1171, 1441, 1588, 37, 803]. Jameson [1854]. Jammer [557]. Jan [1115, 1774, 1463, 747]. Jánosné [1457]. Jansen [661]. Jansz [937]. January [1324, 1279, 138, 282, 496, 682, 851, 1028, 1210, 1439, 1669, 1934]. Janus [1543]. Jaouiche [1839]. Japan [1284, 391, 90]. Japanese [1106, 207]. Jaquel [1705]. Jaques [631]. Jaques-Louis [631]. Jarcho [1603]. Jaroslav [1094]. Jarrell [1089]. Jarvis [1916, 442]. Jasenas [926]. Jasmin [843]. jawab [1761]. Jay [949, 1226]. Jaynes [977, 1747]. Jazari [1251]. Jean [1428, 93, 973, 809, 1360, 1796, 746, 1705, 533, 1609, 1732, 1063, 133, 1175, 1083, 189]. Jean-Baptiste [1063]. Jean-Henri [1705]. Jean-Jacques [973, 1360]. Jeanne [1373, 1920]. Jeannette [1504]. Jeeves [92]. Jefferson [1852]. Jeffrey [1334, 789]. Jellyfish [1371]. Jenkins [1310]. Jenner [1696]. Jeon [1452]. Jeremy [1009, 1527, 266]. Jerome [1861, 429, 555]. Jerry [908]. Jerusalem [1050]. Jetter [1015, 562]. Jewish [975, 363, 860]. Jewkes [306]. Jih [396]. jiten [1106]. Jo [533, 1158]. Joachim [1590, 33, 1899]. Joan [118, 751, 537]. Jochem [1755]. Joe [1150]. Joel [1830, 1645]. Johan [320]. Johann [1656, 1191, 842, 1358, 1309, 1306, 572, 336, 295]. Johanna [1789, 1905]. Johannes [522, 1056, 1489, 196, 1112, 917, 1606, 802, 86, 522, 802, 373]. Johannisson [1146]. John [1411, 1196, 1559, 1191, 1333, 708, 486, 1239, 1491, 807, 1053, 27, 1300, 788, 314, 1560, 1838, 318, 1145, 1643, 1712, 1417, 1055, 571, 897, 529, 1201, 343, 369, 1344, 462, 482, 793, 1365, 1544, 612, 1908, 919, 1091, 1158, 1501, 160, 1418, 633, 470, 909, 1192, 977, 866, 104, 73, 177, 944, 1852, 378, 75, 90, 996, 291, 1576, 595, 1807, 1359, 670, 1116, 102, 893, 220, 371, 69, 361]. Johns [1024, 1456]. Johnson [153, 477, 1712, 1160]. Johnstone [1880]. Joja [1530]. Joliot [1506]. Joliot-Curie [1506]. Jolly [1117]. Joncich [131]. Jones [669, 1533]. Jope [1600]. Joravsky [277]. Jordanus [336, 1759]. Jørgensen [670]. José [740, 570, 1275, 940, 1138, 1702, 269]. Josef [1853, 532, 1746, 320].

Joseph [1814, 1203, 246, 375, 1110, 1262, 1889, 1591, 243, 436, 183, 1600, 1511, 600, 1044, 617, 1844, 1483, 1763, 383, 371, 919, 618, 1135, 1701, 50, 909, 133, 1446, 1078, 1426, 625, 1201, 1857]. Josephi [568]. Joshua [1817]. Josiah [1204]. Jost [1112]. Joule [1129]. Journal [820, 1874, 266, 668, 1540]. Journals [375, 1116, 1426, 161, 127, 528, 1003]. journée [1142]. Journey [1494]. Journeys [127]. jours [561]. Jozsef [1228]. Jr. [552, 736, 1467]. Juan [1782, 1707]. Judaeus [145, 591]. Judd [1232]. Judith [93, 1206, 1821, 1895, 881]. Judy [1476]. Juif [885]. Juli [1414]. Julia [1069]. Julian [977, 1747]. Julius [1602, 901]. July [460, 689, 952, 1809, 689]. June [883, 440, 1654, 1680, 460, 1520]. June-September [883]. Jungk [302]. Jürgen [259, 341]. Juskevic [1170, 1558]. jusqu’à [1488]. Justice [502]. Justification [1174]. Justus [1427]. Juvamentis [1678].

Kāmil [1522]. Kádár [1826]. kagaku [1353]. Kagan [1861]. Kahn [95]. Kaiser [1684]. Kaiser-Wilhelm-Gesellschaft [1684]. Kaiserreich [1918]. Kallinich [1892]. Kames [803]. Kämmerer [841]. Kangro [273, 1520]. Kant [464, 1805, 604, 1488, 267, 1906]. Kapitsa [226]. Karaji [521]. Karel [244]. Kargon [478, 1119, 1791]. Karin [1545, 1146, 1841]. Karl [1057, 1333, 1433, 260, 1299, 1783, 1831, 33, 60, 876]. Karla [1783]. Karlfried [1590]. Kasimir [810]. Katalog [1187, 1482]. Kataye [1650]. Katharina [1408, 324]. Kauffman [70, 70, 1424, 466]. Kaufman [443, 1855]. Kay [1023]. Kean [1750]. Keeling [528]. Keene [105]. Keepers [1154]. Keir [1157, 1570]. Keith [1143]. Keller [1100]. Kelvin [1150]. Kendall [385]. Keng [332]. Kennedy [1409, 1186, 703, 1649, 900]. Kenneth [832, 1588, 1750, 297, 387, 703, 1468, 697, 1880, 1341, 579, 1809]. Kentmann [802]. Kepler [1057, 842, 522, 295, 1112, 873, 522, 1056, 1460, 1306]. Kepler-Symposium [1057]. Kerslake [1544]. Kerwood [919]. Ketner [1588]. Keuning [937]. Kevles [1663]. Key [1706, 1053]. Khaimovna [1256]. Khalid [1050]. Khalil [1839]. Khatri [1923]. khimii [1883, 1535]. Kielmeyer [590]. Kieser [376]. Killian [1686]. Kilwardby [1476]. Kind [1498]. Kindi [1460]. Kindness [795]. Kinematics [472]. Kinetic [9, 237, 1498]. King [1050, 1697, 1773, 652, 352, 316]. Kipphardt [75]. Kircher [4, 208]. Kirchhoff [1166, 1166]. Kirk [80]. Kirkpatrick [1421]. Kirtland [1289]. Kisch [1531]. Kish [1885]. Kisiel [243]. Kitab [1186, 1251]. Kiyoshi [1353]. Kiyosi [507]. Klaus [1698]. Klein [184, 1761, 158, 29, 846]. Klein-Franke [1761]. Kleinert [1413]. Klepsydra [115]. Kline [701]. Knidische [1548]. knidischen [1548]. knidischer [1548]. Knight [649, 1693, 1282, 1526]. Knizních [1774]. Knobloch [1111]. Knorr [1294]. Know [725]. Knowledge [712, 413, 1286, 1818, 1441, 555, 1394, 1251, 607, 1394, 1922, 605, 1503]. Knowles [91]. Kockelmans [243]. Koelbing [357]. kogi [1141]. Köhler [386, 386]. Kohlstedt [1367]. Kolakowski [699]. Kolb [1486]. Kolonialzeit [1015]. Kombinatorik [1111]. Komenian [1774]. Komensky [1774]. Kommentar [43, 1902]. końca [342]. Kongresses [1414, 1918]. König [1820]. Kónya [1457]. Konzepte [1831]. Koordinatenüberlieferung [1480]. Kopal [1095]. Kopelevich [1256]. Kopernik [342]. Kopernika [660, 838]. kopernikanischen [1459]. Kordig [1174]. Korea [1452]. Körpers [1403]. Korrespondenz [8]. Korrespondenzen [1558]. Kors [1307]. Kosmos [1741]. Koyré [873]. Krafft [1057, 1851, 360]. Kräfte [1403]. Krankenpflege [1851]. Krankheit [359]. Krankheitsbezeichnungen [1607]. Kranzberg [658]. Krassilnikoviae [173]. Krausz [768]. Kremers [1541]. Kremlin [675]. Kreutterbuch [1905]. Kriss [173]. Kristian [1741]. Kritik [1480]. Kröber [1929]. Kroeber [327, 327]. Krohn [1816]. Krönig’s [237]. Kropp [654]. Krupp [1756]. Kuchler [1374]. Kuhn [1730]. kulturze [342]. Kunitzsch [1480]. Kunst [1698]. Künsten [750]. Kurt [1657, 932, 1017, 43, 1414, 98, 707, 1900, 1875]. Kurt-R [1657, 1017]. Kurz [1752]. Kuzakov [1813]. Kuznecov [874].

L [301, 1837, 652, 1292, 676, 1700, 1465, 64, 388, 1265, 1734, 131, 251, 1083, 25, 375, 1887, 1422, 1067, 1833, 430, 1169, 1356, 630, 1019, 975, 1568, 329, 1334, 1886, 1775, 989, 704, 153, 1560, 1532, 1859, 1284, 489, 1156, 1930, 1474, 1749, 439, 1691, 1659, 872, 220, 462, 1298, 189, 1648, 1527, 1862, 1895, 803, 249, 1042, 1137, 178, 1505, 834, 1473]. L. [325]. l’Abbé [567]. Laboratory [1087]. Lacaita [1202]. LaCapra [629]. Lachesis [319]. Ladislao [779]. Ladle [815]. Ladrière [1525]. l’agronomie [1537]. Laid [1858]. Laine [1588]. Lakatos [605]. Lam [1762]. Lamarck [1524, 1131]. Lambert [1775, 1705, 572]. Lambertiana [572]. l’Amérique [128]. Lampsakos [255]. Lancelot [594]. Land [706]. Lande [492]. Landels [1757]. Landes [56]. Landesbibliothek [1330]. Landmark [1265]. Landsborough [1509]. Landucci [663]. Lange [1929]. Langenstein [1297]. Langford [429]. Language [986, 1793, 1748]. Languages [151]. Lanza [745]. Lapidibus [1838]. Laplace [1568, 1519]. Lardner [979]. lärdomshistoria [1283]. Large [731]. Larry [1638]. Larson [430]. L’art [1396]. Last [1157, 1570]. lat [660]. Late [1768, 948, 167]. Later [857]. Latest [292]. Latif [259, 546]. Latin [983, 1648, 1448]. latine [1547]. Latitudes [1270]. Laudan [1638]. Laura [1415]. Laureates [1735]. Laurens [878]. laus [1799]. Lausanne [1175]. Lavoisier [1338, 1792, 588, 394, 391, 289]. Law [594, 1568, 916, 731, 1274, 739, 1166]. Lawrence [1172, 103, 1556, 1259]. Laws [1675]. Lay [1762]. Layer [1318]. Layton [552, 1018, 930]. Lazare [315]. Lazzaro [1846]. League [481]. Leake [1238]. Learned [1286, 878]. Learning [1213, 1359, 224]. Leavitt [1895]. Leben [1853, 274, 295, 1785, 376]. Lebesgue [441]. L’école [1930, 1175]. Lecons [1063, 927]. Lecture [716]. Lectures [1411, 183, 847, 1059, 618, 1141, 388]. Lederberg [1817]. Lee [1050, 177, 944]. leeren [213]. Leeuwenhoek [661]. Lefebvre [1493]. Leffel [979]. Leffler [206]. Left [1928, 769]. Legacy [1465, 59, 1783, 1260, 537]. Lehmann [803]. Lehre [750, 1548]. Leib [841]. Leib-Seele-Geist-Problem [841]. Leibniz [1330, 1147, 1111, 1414, 1844, 1061, 181, 174, 142, 1842, 888, 1110].

Leibniz-Archiv [1330]. Leibniz-Kongresses [1414]. Leibnizian [172, 592]. Leibowitz [1761, 353]. Leichhardt [444]. Leiden [521, 987]. Leith [387]. Lemaître [811]. Lemaine [1819]. Lémery [730]. Lempriere [756]. Lenins [1929]. Lennard [1205]. Lenoble [410]. Lenon [421]. Lenzen [64, 1384]. Lenzer [1368]. Leo [812]. Leon [1530, 1102]. Leonard [1088, 1237, 161, 602]. Leone [1846]. Leonhard [1170, 1558]. Leonhardi [1490, 1490]. Leonora [527]. Leopardi [822]. Leopold [1928, 566]. L’équilibre [843]. Lerner [55]. Lértora [1407]. Lesky [891, 1656, 483]. Leslie [305]. Lester [1773, 352, 316]. Leszek [699]. Let [1089]. L’età [409]. Lethal [1829]. L’Étrenne [1306]. Letter [821, 784, 84, 367, 661]. letterario [1308]. Letterio [902]. Letters [1713, 1789, 527, 444, 103, 1426, 1775, 989, 314, 1783, 597, 1039, 862, 1711, 1448, 1925, 1314]. lettre [746]. Lettres [1801, 567, 763]. letzten [806]. Leuckart [1799]. leur [1731]. Level [1687]. Levere [1089, 445, 939]. Levi [1350]. levnadsteckning [262]. Lëvshin [1713]. Levy [227]. Lewis [1928, 1418, 672, 1298, 377]. Lexikon [702]. Leyden [1779]. lezioni [1495]. l’hérédité [560]. l’homme [1109]. L’hypothèse [811]. Liang [1106]. Libbrecht [1107]. Liber [258, 1760]. Liberalism [1699]. Libraries [1145, 1227, 462, 1728]. Library [225, 1197, 826, 978, 1825, 1778, 1194, 1187, 1781, 864, 1059, 1908, 1711, 13, 1486, 1482, 864]. libri [1653, 960]. Libro [212]. libros [1407]. Lieber [1761]. Liebig [1427, 1792, 610]. Liedman [671]. Life [1428, 226, 1372, 40, 1205, 379, 375, 870, 1371, 1554, 1152, 1727, 426, 1290, 4, 408, 238, 1309, 625, 126, 985, 989, 945, 183, 191, 310, 1119, 893, 554, 896, 1925, 939, 897, 1193, 523, 1622, 127, 1103, 1409, 409, 1229, 1158, 765, 984, 448, 991, 1493, 1401, 600, 1575, 1715, 1085, 160]. Lifting [980]. Light [232, 859, 434, 240, 223, 914, 990, 1088]. Liliana [839]. Liliane [720]. Limit [1229]. Limitations [641]. Limits [590, 501]. Lin [1106]. Linacre [1554]. Lincoln [258]. Lind [1232]. Linda [1923]. Lindberg [708, 1296, 1460, 361]. Lindroth [1732, 1283]. Lindsay [973, 1566, 1339, 321, 901]. Line [88, 197]. Linear [1253, 1867]. lineis [519]. Linens [66]. Ling [617]. Linguistic [531]. Linguistics [1243]. Linklater [808]. Linn [383]. Linnaeu [1356]. Linnaeus [1494, 1124, 426, 319, 1494, 714]. Linné [843, 67, 430]. Linnés [1356]. Linton [535, 535]. L’introduction [1151]. Lionel [422]. Liquors [979]. Lisa [1789]. List [1112, 709]. Litchfield [1567]. literarischen [548]. literas [619]. Literatur [1430, 214]. Literature [986, 1343, 1335, 1258, 31, 1749, 1419, 1337, 1633, 860, 1327, 305, 751, 659, 667, 975, 20]. Little [1053]. Littlenesse [501]. liv [262]. Liver [661]. Lives [1154, 1312]. Living [1235]. Livingston [228, 990]. livre [1839]. Lloyd [1837, 1691, 751]. Lo-Keng [332]. Loadstone [860]. Loan [791]. Lobachevsky [857]. Localization [487, 729]. Located [552]. Lock [767]. Locke [712]. Lodge [1117]. Lodovico [1410]. L’oeuvre [1433, 885]. Logan [1197, 1418, 664]. logarithmes [432]. Logic [1793, 1858, 688, 1249, 181, 489, 1530]. Logical [59, 950]. logique [560, 724]. Logos [800]. Lohne [157]. Lohr [983]. Lokken [664]. Lomask [1510]. Lomax [1861]. Lomonosov. [1845]. London [518, 756, 1920, 1544, 460, 1442, 310, 1349]. Long [1753]. Longitude [1013].

Longo [747]. Look [1402]. López [740, 1782, 57, 1045]. Lord [803, 1150, 321]. Lore [834]. Loren [580]. Lorentz [113]. Lorenzo [983]. Lorna [1571]. Losee [788]. Lost [1550]. Lothar [1684]. Löttechnik [1755]. Lotze [1632]. Louis [1661, 569, 36, 1226, 1615, 1118, 1387, 858, 631]. Louise [347]. Louvain [1484, 1525]. Lovell [225]. Löw [1892, 1792]. Lowell [1366]. Lowenthal [1289]. Lowie [677]. Loyd [632]. Lu [617, 1483, 1763]. Lubos [1094]. Lucas [1533]. luce [1485]. Lucien [488]. Lucius [1798]. Ludmerer [697]. Ludvig [718]. Ludwig [444, 1198]. lui [1416]. lui-même [1416]. Luigi [902, 1547, 1497]. Luik [362]. Luis [1766, 1782, 798, 1893, 1551, 884]. Luisa [1257, 465, 1113]. Lukes [945]. Lukina [1783]. Lumb [864]. lumière [232, 1615]. Lunacy [442]. lunae [563, 563]. Lunar [298, 662, 1802]. l’Université [1525, 1175]. L’universo [313]. L’universo-macchina. [313]. Luqa [1396]. Lurie [1118]. Lussac [1635]. Lutz [619]. Lutzker [761]. Lyceum [915]. Lyell [161, 969, 2, 1326, 602]. Lyellian [998]. Lyle [669]. Lymph [917]. Lynn [1901, 1467, 736, 972]. Lyon [1225]. Lysenko [277, 55].

M [740, 767, 1088, 1850, 1469, 1687, 1292, 1108, 1398, 1393, 1354, 1514, 464, 1861, 312, 649, 883, 608, 487, 1116, 444, 1878, 362, 419, 518, 1019, 1553, 1765, 57, 132, 1045, 923, 1194, 656, 1600, 1728, 1761, 1838, 1195, 1511, 521, 1014, 1200, 322, 221, 757, 1783, 1920, 697, 1526, 1012, 66, 1787, 1341, 1226, 864, 1298, 567, 941, 1148, 919, 565, 357, 228, 1736, 1171, 305, 1647, 401, 1715, 236]. M.D [761]. même [1416]. ma [1251]. maître [885]. Ma’amar [1761]. Maanen [782]. Maarten [320]. MacAdam [249]. Maccagni [179]. macchina. [313]. Mach [479, 1503, 897, 170, 735]. Machamer [1822, 1529]. Machina [656, 736]. Machine [527, 1609, 302]. Machines [1479, 1496, 1240]. Machometus [291, 595]. Macintosh [1159]. Mackay [1532, 1736]. Mackenzie [1314]. Mackenzie-McNaughton [1314]. Mackey [931]. MacKie [1835]. MacKinnon [828]. Macksey [151]. MacLeod [1116, 1171, 1819]. Macromolecules [449]. Mad [1925]. Maddison [1554, 888, 934, 191, 873]. Made [1738, 1055]. Madeleine [534, 886]. Madrid [1134]. Maeyama [1824]. Magda [922, 1689, 449]. Magdeburger [213]. Magia [1301]. Magic [1463, 407]. Magical [1103]. Magisteries [1763]. Magisterium [1050]. Magnetic [1865, 916, 456]. Magnetism [1230, 1033, 467]. magnetismen [1146]. Magnetisörernas [1146]. Magneto [3]. Magneto-Optical [3]. Magnitudes [1294]. Magno [1647]. Magyar [1457]. Maillet [34]. Maimonides [1761]. Maine [446]. Mainstreams [352]. Mainz [1430]. Majno [1542]. Major [703]. Maker [1559]. Makers [351, 1004, 19]. Making [1777, 158, 1911, 1571]. Malady [486, 1539, 1312]. Malaria [1894]. Malcolm [92]. Malerei [1780]. Malgorzata [1236]. Malpighi [1557]. Malus [1615]. Mamiami [1495]. Man [1850, 1856, 1748, 775, 1542, 527, 1159, 1244, 703, 728, 1055, 5, 298, 321, 1201, 1894, 655, 577, 151, 816, 251, 968, 1875, 543, 413, 1065, 903, 1729, 524, 1092, 1737, 971, 1754]. Man-Machine [527]. Manager [437, 598]. Manchester [322, 1791, 864]. Mandelbaum [577]. Mandeville [1539]. Manfred [1187, 1482]. Manier [1793]. manière [758]. Manipulation [540, 1163]. Mankind [1829]. Mann [188, 1918, 928]. människolära [1356]. Manoscritti [1703, 1485]. Mansfeld [616]. Manual [825]. Manuel [35]. Manufacture [979, 155]. Manufacturing [1567]. Manuscript [1479, 12, 258, 566, 118, 336, 473, 13, 886, 595]. Manuscripts [1220, 587, 807, 552, 709, 864, 1296, 1826]. manuscrits [1650]. Manuskriptes [1765]. Map [756, 1911, 1055, 177]. Mappae [1053]. Mapping [1095]. Maps [655]. Maqalah [1761]. Maraghah [1871]. Marbach [8]. Marbode [1838]. Marc [957]. Marcel [796, 1930, 1744, 1312]. Marcello [1890, 1557]. March [1781, 999, 350]. Marcia [1728]. Marcohandschriften [1877]. Marcus [1761]. Marcuse [769]. Margaret [1571, 1427, 692, 1554, 425, 1798, 187, 258, 257]. marginadas [1766]. Marginalia [1147]. Maria [519, 1690, 1257, 465, 839, 1113, 1780, 551, 1275]. Marian [660]. Mariano [1782, 269, 710]. Marie [1494, 1801, 1654, 372, 1109, 1001]. Marijke [937]. Marin [892, 1058, 1776, 955]. Marine [251, 692, 1019]. Marinus [800]. Mario [745, 924]. Marion [717]. Marius [1758]. Marjorie [1608, 554]. Mark [1538, 1462, 582]. Marks [880, 742, 1183, 1829]. Markus [1652]. Marqueses [254]. Marquis [1108]. Mars [1323, 1366]. Marsden [689]. Marshack [1244]. Marshall [1898, 844]. Martha [1897, 375, 1426, 1112, 1496]. Martial [944, 944]. Martin [1559, 443, 875, 790, 158, 949, 1855, 1661, 991]. Martínez [1707]. Martinus [939, 126, 985, 127, 1493]. Marton [64]. Martyn [1289]. Marum [126, 985, 939, 127, 1493]. Marx [1528]. Mary [1199, 1466, 1371, 384, 1825, 1593, 533, 1818, 386, 623, 1312, 977, 73, 177, 944, 844]. másfél [1457]. Masha’allah [509]. Ma’shar [364]. Masotti [1410]. Massachusetts [1742, 442]. Master [528, 990]. Masterbuilders [1604]. matematica [1410]. matematici [1703, 1485]. matemático [1355]. Mateos [1782]. Materia [1054]. Material [820, 1640]. Materialien [1820]. Materialism [192, 1271, 1788]. matérialisme [1262]. materialov [1845]. Materials [206, 238, 1845, 11, 1588, 24]. materialy [660]. Mathematica [1147, 174, 716, 457]. Mathematical [1596, 247, 1110, 1204, 1186, 489, 221, 194, 143, 181, 1787, 29, 716, 1148, 786, 1522, 1762, 705, 701, 31]. Mathematicall [1300]. Mathematicians [1556]. mathématicien [892, 227]. Mathematics [1086, 1628, 1630, 206, 1269, 1556, 1882, 1759, 609, 248, 716, 453, 1869, 1107, 1533, 1143, 1919, 647, 825]. Mathematik [1017, 654, 753, 571]. Mathematiker [1652]. Mathematikern [1844]. Mathématique [1108]. mathématiques [700, 1502]. mathématisation [827]. Mathematische [1844]. Mathematischen [1877, 1017, 1111]. Mathematischer [1330, 932]. Mathesis [366]. Mathilde [1789]. Matrix [1630]. Mats [1767]. Matter [44, 76, 110, 135, 139, 1, 45, 77, 111, 137, 163, 198, 229, 279, 283, 140, 164, 199, 230, 281, 330, 389, 450, 493, 497, 284, 331, 390, 451, 495, 538, 585, 638, 679, 683, 498, 539, 586, 639, 681, 726, 772, 813, 848, 684, 727, 773, 814, 850, 911, 953, 993, 1025, 1029, 852, 912, 954, 994, 1027, 1074, 1121, 1161, 1207, 1030, 1075, 1122, 1162, 1209, 1267, 1315, 1376, 1436, 1211, 1268, 1316, 1377, 1438, 1512, 1577, 1624, 1666, 1440, 1513, 1578, 1625, 1668, 1716, 1803]. Matter [1863, 1931, 1935, 1670, 1717, 1804, 1864, 1933, 445, 1035, 272, 1529, 516, 204, 952, 75, 924]. Matter-Theory [272]. Matthias [255, 619]. Maturing [1119]. Maturity [1110]. Maupertuis [1142, 1335]. Mauri [1412]. Maurice [1661, 93, 65, 1506, 653, 1532, 1304, 354, 516, 623, 101, 577, 1698]. Maurin [1109]. Maurizio [1495]. Maurus [473]. Mauryan [365]. Mauskopf [1369]. Max [1235, 134, 302, 557, 574, 572]. Maxwell [302, 641, 166, 231, 958]. May [312, 1809, 267, 257]. Mayer [901, 662, 938]. Mayotte [746]. Mayr [1240, 63, 1132, 420, 1401]. Maze [1241]. Mazzeo [1814]. McCance [1886]. McCann [1511]. McCartney [1794]. McClain [1882, 1882]. McCormack [1503]. McCormmach [121]. McCown [703]. McDowell [322]. McGehee [1456]. McGrath [387]. McGucken [762]. McGuinness [1503, 1198]. McGuire [1781]. McIntosh [356]. McKee [73]. McKinney [672]. McLaren [1920]. McLeish [1344]. McNaughton [1314]. McVaugh [1348]. Mea [1706]. Meadows [1281, 20]. Mean [12]. Meaning [112, 875]. Measurement [202]. Measurements [456]. mécaniste [1262]. Mechanica [1490]. mechanicae [1490]. Mechanical [1479, 1251, 1129, 428]. Mechanics [592, 1263, 1573, 333, 626, 950, 15]. Mechanik [360]. mechanischen [1847, 1698]. Mechanism [1247, 1936, 1937, 192]. Mechanisms [626, 420]. Mechnikov [1713]. med [262]. Medal [403, 507, 823, 1388, 1450]. médecin [133]. médecine [190]. Mediaeval [363]. Medica [1054, 1348]. Medical [1656, 352, 1465, 1354, 185, 1466, 1602, 1562, 1037, 1184, 443, 889, 1768, 1509, 929, 81, 1619, 1855, 1024, 1456, 1920, 709, 1103, 1916, 296, 864, 1049, 1213, 520, 305, 1696, 90, 667, 27, 1195]. Medically [486]. médicaux [1650]. Medicina [741, 57, 1355, 1551, 1766]. medicinae [316]. Medicine [1773, 743, 352, 1896, 880, 1184, 1750, 297, 978, 1334, 1361, 365, 1643, 1784, 1749, 1710, 801, 383, 864, 1011, 836, 1141, 363, 520, 1895, 659, 1540, 1875, 978, 1543, 30, 1827, 27, 1603, 1183, 1351, 1246, 305, 1830, 1874, 1357, 357]. Medicines [1518, 503]. medicinis [258]. médicos [57]. Medieval [124, 1296, 975, 1053, 1248, 1609, 1249, 1606, 474, 1527, 972, 696, 510, 503, 287, 145, 1806, 685, 983, 860, 1901, 1608, 1351]. Mediterranean [1167]. Medizin [1097, 1548, 1831, 1918]. Medizinalbuch [1408]. Medizingeschichte [928]. medizinischen [418]. medizinischer [1391]. Medvedev [55]. Meeting [1223, 1449]. Meetings [598, 437]. Megalithic [1292]. Megara [1300]. Mehra [992, 1623]. Meites [1511]. Mel [1314]. Melanchthon [913]. Mellor [558]. Melville [910]. Melvin [658]. Members [1783]. Membrorum [1678]. Memoir [755, 1686]. Memoirs [385, 1789]. Memorie [902]. Memory [799]. Men [1779, 1422, 1302, 1738, 150, 16, 795]. Menabrea [902]. Menard [463]. Mendel [532, 830]. Mendelsohn [881]. Mendoza [1407]. Menninger [60]. Menon [767]. Menschenbild [621]. mensurandi [86]. Mental [1309, 760, 878, 320]. Mentioned [1219]. Merkmale [1831]. Merleau [927]. Merleau-Ponty [927]. Merriam [1157, 1570]. Merritt [1429]. Mersenne [955, 1058, 1776, 892, 500, 1038].

Merton [1817, 431]. Merwe [492]. Merzbach [942]. Meseguer [740]. Mesopotamien [1293]. Messungen [273]. metafisica [1878, 1843]. Metals [1183, 235]. Metamorphosis [165]. Metaphor [514]. Metaphors [1373]. Metaphysics [1269, 604]. métaux [1472]. Meteorological [91]. Meteorology [1641]. Method [93, 640, 1288, 829, 7, 93, 38, 575, 1565, 924]. Methode [750]. Methodo [750]. Methodological [369, 1528]. méthodologie [724]. Methodologies [1584]. méthodologique [252]. Methodology [1870, 1530]. Methods [1479, 24, 819, 16, 1879]. metodo [1495]. metodologico [936]. Metric [1817]. Mettrie [527]. Mexican [132]. Meyer [1057, 467]. Meylan [561]. Meynert [81]. Michael [1359, 1358, 1768, 1825, 55, 904, 1056, 768, 899, 21, 38, 1193, 926, 1926, 1348, 528, 378, 337, 1819]. Michaelis [869]. Michel [934, 346, 553, 1731]. Michele [636]. Michelmore [536]. Michelson [632, 990, 990]. Michigan [552]. Microbes [448]. Microbiology [1400, 674]. Microfiche [1588]. Microfilm [13]. Microorganisms [173]. Microphysics [119]. Microscope [1037, 64]. Microscopen [244]. Microscopes [1156]. Microscopy [141]. Mid [1920, 1486]. Mid-Sixteenth [1486]. Mid-Victorian [1920]. Middle [805, 890, 1898, 1631, 1256, 982, 589, 1609, 1181, 31, 424, 1297, 1475]. Middleton [91]. Migration [355, 41]. Miguel [1893]. Mikolaja [660, 838]. Mikulas [1135]. Miles [1341]. Milestone [855]. Milic [722, 1639, 569]. Milieu [730, 1123, 845]. Militärarzt [374]. Military [1726]. Milky [872]. Mill [979, 122]. Millás [1275]. Millburn [1559]. Miller [632, 375, 948, 159, 71, 690, 1328]. Millikan [1206, 1378]. Milliken [804]. Milton [1510]. Minar [834]. Mind [1837, 986, 100, 1289, 554, 1118, 1158, 411, 1747, 487]. Minds [253, 1312]. Mineral [1905]. Mineralnamen [1097]. Mineralogical [668]. Mineralogy [542]. Minerals [1905, 1852]. Minguet [128]. Minime [1058, 1776]. Mining [421]. Minio [983]. Minio-Paluello [983]. Minkowski [1067]. Minnesota [13]. Minor [1510]. minoría [1551]. minorías [1766]. Minutes [598, 437, 1139]. Miolati [51]. Mirabilis [526]. Miracle [1510]. Mirko [124]. Mirrors [1550]. Mirsky [1504]. Misattributed [293]. miscellanea [1490]. Mischungen [1403]. Misconception [777]. Mispar [887]. Missing [238]. Missionary [34]. Mississippi [375]. mitad [1707]. Mitcham [931, 1644]. Mitchell [725, 1055]. mitt [890]. Mittag [206]. Mittag-Leffler [206]. Mittelstrass [341]. Mittler [877]. Modality [1104]. Mode [1378]. Model [924, 1326]. modelo [1355]. Modern [844, 701, 184, 469, 714, 351, 1738, 329, 149, 1565, 1806, 306, 626, 1336, 1284, 1094, 882, 1467, 1784, 1893, 1663, 1688, 1123, 383, 216, 722, 215, 1106, 1527, 89, 790, 793]. moderna [57, 1355, 1878, 865, 936]. moderne [1690, 1907, 1545]. modernen [74]. Modernity [569]. Modest [1089]. Moerbeke [1898]. Moesgaard [1741]. Mogens [1340]. Mohammed [309]. Mohawk [883]. Mohawk-Hudson [883]. Moisil [1530]. Molecular [1077, 1369, 1329, 1576, 533]. Molecules [600]. Molella [919]. moment [1839]. Monastic [13]. Monboddo [803]. Mond [981]. mondes [775]. Mondfinsternisse [358]. Mondgesicht [563].

Mone [887]. Monist [481]. Monna [974]. Monnier [1930]. Monod [448]. Monograph [625]. Monopoly [1334]. Montgomery [1191]. Monthly [668]. Moody [1249]. Moon [1189, 1095, 1155]. Moone [298, 1196]. Moons [794]. Moorat [864]. Moore [446, 698]. Moorehead [72]. Moormann [1433, 1433]. moral [1702]. Morales [740]. Morality [845]. Moraux [1473]. Moravia [665]. Morgan [1664]. Morgenbesser [829]. Morienus [1050]. Morion [257]. morisca [1551]. Moriscos [1766]. Morley [632]. Morley-Miller [632]. Morris [701, 1913, 473, 1618]. Morrow [186]. Morse [1715]. Morsier [1930, 763]. Morton [384, 829, 305]. Moseley [989, 302, 52]. Moses [1761]. Mosquitoes [1894]. Mosse [1887]. Most [1883, 1829]. Motion [1383, 1189, 499, 286, 472, 1581, 643, 1498, 333, 1529]. Motions [782]. Motivations [776]. Motive [1831]. Mott [1750]. motu [1897, 1080, 1355]. motuum [472]. Mountain [854]. Mourelatos [1295]. Moyal [1425]. Moyen [601]. Mozans [1222, 1043]. Moziño [269]. Mozley [1425]. Mr [712]. MS [521, 258, 1059]. MSS [521]. Mu’awiyya [1050]. Muctazilite [596]. Mugler [744]. Muhammad [1649]. mulhousien [1705]. Mulkay [1819, 1926]. Muller [216, 1091, 374, 1414, 917, 564, 981, 1285]. Multiplikation [1899]. Mundo [1305, 292]. Muriel [368]. Muris [86]. Murphy [1778, 677]. Muscular [656]. Museo [465]. Muses [96, 606]. Museum [18, 244, 91, 8, 929, 939, 420, 1154, 245]. Museums [606]. Musgrave [605]. Music [1882, 1337]. Musiktheoretiker [1056]. Muslim [474]. Musson [56]. musulmana [1551]. My [302, 449, 1575, 693, 197]. Mycology [1399]. Mysticism [1257, 1336]. Myth [690, 1882, 393, 122, 1696, 837]. Myths [565].

N [1921, 1748, 375, 1833, 877, 1159, 1533, 1426, 664, 653, 1883, 1691, 195, 181, 1916, 1091, 537, 866, 98]. nach [858, 564]. Nachum [975]. Nacional [741]. Nagatsune [1567]. Nagel [829]. Naissance [1891]. Najstarszy [838]. Nakayama [1284]. named [858]. Namen [1902]. Names [1767, 1237]. Nancy [1619, 1363, 1250]. Napier [608]. Narratio [367]. Nasīr [732]. Nasen [564]. Nasen- [564]. Nash [530]. Nasir [234]. Naslediya [1783]. Nasr [1610, 1481]. Natale [1875]. natalem [532]. Nathan [1203, 1834, 797, 1922, 32, 919, 1660, 1764]. Nation [1588, 483]. National [1278, 71, 490, 481, 978, 1510, 420, 118, 803, 909, 245, 695, 1212, 205, 8]. Nationalmuseum [8]. Native [1646]. Natsu [1141]. Natur [1655]. natura [1555, 1495, 1658]. Natural [1706, 640, 1793, 18, 1078, 192, 1116, 1335, 430, 241, 243, 1143, 1568, 436, 318, 1528, 1712, 1581, 668, 271, 371, 1217, 1226, 633, 672, 401, 1702, 1088, 1469, 649, 1462]. naturale [1555]. naturalis [457]. Naturalism [948]. Naturalist [1733, 382, 426]. Naturalists [1157, 1570, 8, 377]. Naturanschauungen [1477]. Nature [78, 345, 423, 1092, 1356, 690, 1886, 166, 1592, 1729, 685, 349, 1915, 240, 1879, 113, 393, 285, 1526, 68, 1213, 1192, 1875, 843, 1848, 1856, 1727, 992, 410, 378, 1694, 1118]. naturelle [65, 856]. Naturforscher [947, 302, 260, 1299, 67, 802]. Naturforschern [8]. Naturforschung [947].

Naturkenntnisse [1477]. Naturphilosoph [1652]. Naturphilosophie [255]. naturuppfattning [1356]. Naturwissenschaft [752, 1918, 1709]. naturwissenschaftlich [1391]. naturwissenschaftlich-medizinischer [1391]. naturwissenschaftlichen [1332]. Naturwissenschaftlicher [1330]. Naturwissenschaftsgeschichtliche [1824]. Nauchno [1881]. Nauchno-tekhnicheskaia [1881]. Nauchnykh [1589, 1256, 1813]. Nauheim [1918]. Nauk [1783]. Nautical [669]. Nautics [617]. Naux [432]. Naval [1712, 634]. navalem [1490]. Navigation [1013]. Navigator [312]. Navy [1071, 634]. Near [1752]. Nebulae [202, 782, 1718]. Nebular [1568]. Necessity [1104]. Neckar [8, 8]. Nectar [1629]. Nectaries [1629]. Need [1466, 1218]. Needham [617, 1483, 1763, 656, 183, 618, 1701, 1135]. Negroes [1226]. neige [1306]. Neil [1732, 1501]. nel [1555, 366]. nella [665]. Nelson [118, 323, 215]. Nemesis [319]. Nemore [1759, 336]. Neolithic [1406]. Neper [432]. nepreryvnosti [923]. Neptunian [1854]. Nervous [1052]. Nesta [371]. Netherlands [261]. Nettie [1514]. Neubauer [1192]. Neuchâtel [36]. Neue [1841, 213]. Neugebauer [1596]. Neumann [1093]. Neun [98]. neunziger [1617]. Neuroendocrinology [1511]. Neurology [469]. Neurosis [740]. Neuzeit [341]. Neuzeitlichen [341]. Nevill [1508]. Newcomb [1517]. Newcomen [1833]. News [14, 54, 87, 116, 175, 210, 239, 404, 547, 599, 645, 691, 785, 824, 863, 918, 970, 1000, 1040, 1133, 1224, 1280, 1390, 1586, 1636, 1725, 1873, 1811, 148, 294, 1682, 508, 733, 1084]. Newton [1189, 1415, 338, 461, 221, 526, 716, 1059, 1148, 1495, 1674, 427, 476, 1225, 232, 1083, 1220, 174, 688, 268, 501, 914, 1059, 47, 1277, 434, 35, 1740, 221, 369, 1908, 840, 1912]. Newtoni [1147]. Newtonian [272, 1417, 1805, 1078]. Newtoniana [1912, 313]. newtoniano [1495]. newtonismo [1497]. Newtons [157]. Ni [1106, 1106]. Niccolò [1651, 781, 1410]. Nice [1608]. Nicholas [414, 1458, 675, 611, 19, 1297, 886, 16]. Nichols [943]. Nico [1820]. Nicolai [568]. Nicolas [730, 700]. Nicolaus [1653, 1765, 303, 838]. Nicole [1433, 234, 472]. Nicollet [375, 1426]. Niderst [1416]. niederländischen [1780]. Niedersächsischen [1330]. Nielsen [1347]. Nierada [886]. Night [23]. Nightingale [1312]. Nikolai [460]. Nina [1749]. Nineteenth [1604, 159, 667, 1425, 1371, 1370, 487, 1566, 630, 904, 306, 1859, 1597, 480, 1925, 1920, 1788, 1369, 818, 1916, 577, 69, 1923, 288, 402, 1271, 777, 81, 3, 762, 579]. Nineteenth-Century [159, 1371, 630, 904, 1925, 1920, 1369, 1916, 577, 69, 288, 402, 1271, 777, 81, 3, 762, 579]. Ninety [138, 282, 496, 682, 851]. Ninety-Eighth [682]. Ninety-Fifth [138]. Ninety-Ninth [851]. Ninety-Seventh [496]. Ninety-Sixth [282]. Nino [1890]. Ninth [851, 716]. Nisbet [764, 270]. Nissen [1098, 517]. Nitrogen [504]. Nitske [765]. No [1466, 566, 118, 1353]. Nobel [637, 847, 1735, 847]. Nobis [1765]. Noble [1699]. Noel [1739, 973]. Nogar [866]. Non [1871]. Non-Ptolemaic [1871]. Nonconformity [1319]. Nonprofit [951]. Nontelescopic [1447]. Nootka [269]. Norbert [699]. Nordamerika [1015]. Norlind [262]. Norman [1144, 253, 1008, 306, 1754, 1911, 1605, 1089]. North [1069, 525, 1008, 807, 1055, 1478, 61, 1382, 1322, 1478]. Northern [155].

Northwest [1418]. Norwood [867, 412]. nos [561]. Nosema [1607]. Notation [1244]. Note [402, 544, 83, 1723, 85, 819, 1627]. Notes [375, 1426]. Noticias [269]. notion [1889, 1839]. Notvik [271]. Nousos [1607]. nouveau [775]. nouvelle [631]. nouvelles [892]. Novali [1192]. Novara [510]. November [1203, 1168]. Novembre [886]. Novembre-22 [886]. Nový [1094]. Nowe [660]. Nuclear [988, 1071, 583]. Nugent [712]. Numa [807]. Number [871, 60, 22, 60]. Numbering [302]. Numbers [731, 1568, 1862, 1895, 1505, 60]. Numerical [1823, 49]. numeris [336]. Numic [807]. Nunberg [1139]. Nunzio [842]. nuova [409]. Nursing [844]. Nussbaum [1897]. Nutka [269]. Nutrition [610]. Nuttall [525, 791]. Nutting [844]. nya [262]. Nye [533].

O [1596, 767, 1199, 1761, 1261, 353, 26, 1926, 97, 1089, 706, 1809]. O.M [160]. Oak [1778]. Oberschall [879]. Objectivity [1199]. Objects [929]. obra [1355, 884]. O’Brien [379]. Obscure [1872]. Observation [287]. Observations [1582, 1447, 1052, 1794, 308, 878]. Observatory [1281, 669, 1825, 1281, 1281]. Observer [1794]. Occasion [1375, 1010]. occidentales [1547]. Occult [711, 1242]. Ocean [251]. Oceanographers [997]. Oceanography [303, 251, 1342]. Ocherki [1883, 1813]. ocho [1407]. October [919]. Octonary [962]. Oesper [1680, 1397]. oeuvre [1930]. Oeuvres [1499]. Ogden [425, 258]. O’Gorman [28]. Ohio [552]. Ohm [916]. Ohrenkrank [564]. Ohrenkrank- [564]. Oil [1567]. Oklahoma [1728]. Oktoberrevolution [1617]. Okura [1567]. Olaf [1340, 1245]. Olagüe [1782, 1893]. Öland [1494]. Olby [1120]. Oldenburg [1654, 372, 372]. Oldest [1281]. Oleson [1286]. Oliphant [582]. Oliver [1117]. Olmi [1555]. Olson [514, 1178]. O’Malley [185, 82, 152]. Omnia [1490, 1348, 1170]. Omphalos [1217]. Omstreeks [362]. Oncology [305]. Onderdelen [244]. One [820, 1588, 1041, 1028, 1210, 1439, 1669, 1934, 1457, 26]. One-Hundredth [1028]. O’Neil [1398]. Ontogeny [590, 1523]. Open [1331]. Opera [1490, 1348, 1170]. Operations [1726]. Opere [745]. Ophthalmological [1347]. Oppenheimer [536, 1266, 75, 75]. Opper [1047]. Opposition [818]. opticae [1490]. Optical [1296, 3]. Optically [855]. Optics [1417, 1059, 361]. Optimus [473]. optique [1615]. Opuscula [983]. Orange [946]. orbe [563]. Order [1371, 590, 1445, 430, 119, 1818, 553]. Ordnance [634]. Ordway [94]. Oresme [472, 234]. Organic [120, 1375, 449, 1883, 303, 870]. organicheskikh [1883]. organicheskoi [1883, 1535]. Organicism [1373]. organisations [1563]. organisations- [1563]. organisés [1063]. Organisms [590]. Organization [1913]. Organizations [1379, 1818]. Organized [1710]. Organizer [195]. Organizing [1072]. Orientierung [1816]. Origin [712, 1854, 935, 1288, 381, 394, 541, 1882, 441, 728, 1193, 642, 29, 1747, 1322, 486]. Original [449, 801, 303, 337, 1550]. Originality [908]. Originalschriften [702]. origine [1839]. origines [1360]. Origini [313, 1690, 1878, 936]. Origins [670, 184, 468, 481, 503, 392, 1254, 1442, 37, 1018, 1549, 634, 1406, 231, 604, 1404, 1006, 1094, 1740, 320]. Ornithology [1381, 1180]. Oronce [886]. Orthoptic [419]. ortu [1476]. Osborne [322, 322]. Oscillating [1042]. Oskar [324]. Osler [1194]. Osservazioni [1846]. Ostwald [195, 85]. Oswald [1622]. Oswaldo [1619]. Other [1543, 661, 907, 1729, 298]. Otilie [245]. Otis [804, 421]. Otkrytie [1883]. Otology [1893]. ottica [1495]. Otto [1752, 1240, 63, 326, 420, 564, 928]. Our [1741, 416]. Outcasts [482]. Outline [1531, 356]. Outstanding [195]. Overskill [882]. Owen [1254, 1837, 1720, 1534, 411]. Owens [1372]. Owsei [1543, 469, 1049]. Oxford [297, 297].

P [184, 251, 1371, 1159, 1295, 1546, 558, 417, 1117, 1090, 1058, 1776, 1797, 516, 97, 101, 894, 130, 1453, 1711, 919, 155, 581, 23, 1155, 1558, 881]. P. [1375]. Pacey [1241]. Pach [1457]. Pacing [1840]. Pädagogik [621]. Padgitt [1156]. Padua [1250]. Pafnutii [1921]. Page [1451]. Pagel [1011]. Pain [308]. Painting [241]. Pais [75]. Pál [1457]. Palaeontology [875]. paléobotanique [1616]. Paleontology [1597, 1179]. Paley [401]. Palingénésie [200]. Palingenesis [544]. Palladius [1902]. Palm [1386]. Palter [526]. Paluello [983]. Pancaldi [1658]. Paola [895, 865, 1647]. Paolo [215, 313, 433, 713, 1301, 1875]. Paper [1129, 53]. Papers [767, 1203, 1746, 312, 1662, 907, 664, 70, 349, 1424, 1249, 386, 322, 1914, 221, 1688, 1781, 941, 1148, 919, 460, 1116, 812, 449, 1006]. Papyri [705]. Parabolic [963]. Paracelsian [1875]. Paracelsist [1875]. Paracelsus [621, 1875, 841, 1097, 621]. Paradigms [1243, 650]. Paradis [1856]. Paradoxes [650, 452]. Paralleles [61]. Parallelism [590]. Parallels [1336]. Parallelversion [1902]. Parasitology [1750]. Pardies [573]. Paris [1110, 405, 550, 1307]. parisien [885]. Parke [628]. Parker [705]. Parkinson [864]. Parry [226]. Part [1292, 1588, 1886, 70, 1424, 617, 1483, 559, 1763, 1744, 1054]. Particle [228]. Particularly [1323]. Particulars [1575]. Partington [559]. Partisanship [574]. partium [257]. Parts [257]. Pascal [1093, 78]. Pascual [570]. Pasoli [842]. Passage [1418]. Passion [725]. Passmore [1092]. Past [1458, 1228, 1095, 488]. Pastore [611]. Paszek [1172]. Pat [1818]. Patent [475]. Patents [1451]. Paterson [188]. Path [1120]. Pathological [380]. Pathology [380]. Patients [1784]. patologii [1564]. Patricia [1784]. Patrick [829, 250, 1530, 1400]. Patriot [627]. Patronage [1370, 1579]. Patrons [159, 943]. Patsy [929]. Pattaro [1555]. Pattern [871]. Patterns [1861, 1538, 711]. Paul [1851, 93, 1065, 769, 934, 28, 1140, 1480, 1473, 1503, 1556, 1560, 227, 1914, 1644, 1058, 1291, 377, 1103, 1794, 60, 1109, 1319, 84, 612, 170, 158]. Paul-Henri [934]. Paul-Marie [1109]. Pauline [305]. Pawel [886]. Payne [833]. Payne-Gallwey [833]. Pays [622]. Pays-Bas [622]. Peace [770]. Peano [181, 900]. Pearce [1859]. Pecham [708, 361]. Pécheux [346]. Pechey [761]. Pechey-Phipson [761]. Peder [1741]. Pedersen [1340, 1245]. Peel [766]. Peirce [64, 1588, 1919, 37, 1321, 1867]. Peking [1141]. Pelling [1554]. Pendulum [1218]. Pendulums [859]. Penicillin [1205]. Penniman [304]. Pennsylvania [1362]. Penrose [528]. pensée [227]. pensées [892]. People [807, 1018]. Perception [611, 1822, 1076, 308, 1674, 1632, 1637, 412]. Perceptual [1806]. Percussion [1002]. Perdu [5]. Perepiska [1783]. Peri [616, 257]. Periera [1647]. Perig [1355]. Period [805, 1805, 1876]. Periodic [203, 957, 196, 1041]. Periodica [1391]. Periodicals [1423]. Peripatos [255]. Permanent [1457]. Perplexities [1784]. Perrin [533]. Perseus [878]. Persia [520]. persischen [549]. Personal [327, 1630]. Personality [264]. Personnages [5]. Personnel [1560]. Persons [1145]. Perspectiva [361, 708, 1760]. Perspectivae [1760]. Perspective [986, 533, 662, 1253, 1594, 974]. Perspectives [1769, 1135, 1406, 1834, 603, 492, 1819]. Persuaded [1862]. pertinentes [1490]. Peter [1656, 536, 59, 1822, 1529, 381, 739, 1597, 1402, 1709, 1696, 1912, 1819, 79, 226]. Peterburga [1783]. Peterburgskaya [1783]. Petersburg [1783]. Petersburger [1558]. Peterson [1920, 1089]. Petrarch [1556]. Petroleum [1219]. Petry [378]. Peucer [1486]. Peumery [1360]. Pfeffer [1849]. Pfetsch [1068]. Pflanzenchemie [1792]. Pflanzengattung [858]. Pharaohs [647]. Pharmaceutical [406]. pharmaceutiques [1650]. pharmacie [1175, 190, 1175]. Pharmacology [1238, 1272]. Pharmacy [606, 657, 1054, 1541]. Pharmakologen [835]. Pharmakopöen [1066]. Pharmazie [1097, 1291, 1892]. Pharmaziegeschichte [1345]. Phases [28]. Phenomenology [243]. Phenomenon [1734]. Philadelphia [1197, 668, 1263]. Philip [388, 1734, 1217, 1453, 1715]. Philippa [1116]. Philippe [720]. Phillip [1533]. Phillips [1246]. Philo [1836]. Philology [1870]. philosophe [1705]. Philosopher [1116, 1358, 34, 1263, 479, 436, 629]. Philosophers [1412, 1240, 946]. Philosophiae [457]. Philosophical [650, 184, 1628, 17, 265, 1591, 1778, 1198, 1328, 1176, 1729, 946, 1711, 866, 1149, 1021]. Philosophicall [878]. Philosophie [1655, 170, 601, 65, 512, 1063, 341, 1590]. Philosophies [1719]. Philosophische [1708]. Philosophizing [688]. Philosophy [1773, 1088, 445, 890, 59, 1333, 1078, 192, 471, 1335, 1637, 119, 1143, 1822, 1178, 1461, 1529, 739, 768, 788, 735, 1303, 1870, 37, 1249, 580, 931, 1101, 1644, 270, 1346, 1530, 604, 118, 1138, 1192, 987, 1875, 378, 170, 633, 616, 345, 471, 790, 318, 1173, 926, 1049, 1086, 829, 976, 215]. Phineas [222]. Phipson [761]. Phlogiston [235]. Phoenix [1191]. Photograph [52]. Photographic [1310]. Photography [1500]. Phrenological [534]. Phylogeny [703, 1523]. physica [1490, 1081]. Physical [225, 1375, 1321, 1356, 690, 903, 1565, 121, 1117, 617, 1597, 393, 424]. Physician [370, 659, 1492, 375]. Physicians [518, 673, 1784, 1544, 1862, 1749]. physicien [573]. Physicist [1204, 479, 992, 989, 1193, 1715, 278, 1060, 158, 1201]. Physicists [540, 1663]. Physico [428]. Physico-Mechanical [428]. physicorum [1407]. Physics [1815, 1215, 857, 1221, 908, 1198, 154, 1178, 1461, 790, 458, 847, 1378, 1336, 1685, 1808, 617, 915, 1836, 1623, 143, 118, 1340, 722, 1583, 652, 515, 1265, 338, 329, 154, 557, 118, 182, 130, 1665, 1451, 350, 302, 678, 276, 1560, 987]. Physicus [473]. Physik [702, 1572, 74, 753]. Physikalische [1708]. physikalischen [874]. Physikbücher [1413]. Physiologe [274, 1849]. Physiological [1678, 1886, 380]. physiologie [1262, 1930, 1853]. Physiologist [370]. Physiology [1516, 889, 30, 1721, 1886, 876]. physiques [927]. Piaget [93]. Pickens [108]. Pickering [1165, 1443, 1312]. Picture [1741]. Pictures [537, 1693, 1228]. Pieces [721]. Pierce [64]. Pierre [888, 775, 1502, 1433, 754, 636, 1487, 190, 786, 583]. Pierson [919]. Pieter [1290]. Pihl [1340]. Piñero [740, 57, 57, 1045]. Ping [1406, 1483]. Ping-ti [1406]. Ping-yu [1483]. Pingree [749, 364, 749, 1352, 1471]. Pioneer [1664, 844, 447, 383]. Pioneering [761, 377]. Pioneers [1226, 90, 1511]. Pionier [1851]. pióra [838]. Pipping [1815]. Pirmin [260, 1299]. Pirquet [40]. Pis’ma [1783, 1713]. pitagorismo [1651]. Pitt [1769]. Pivotal [1453]. Pizor [298]. Place [1281, 1856, 1581, 207, 1051]. Plague [117, 1012]. Plains [1426]. plana [1759]. Planckschen [273]. planétaire [886]. planetarum [510]. Planetary [1323, 1165, 1270, 510]. Planktonforscher [274]. Plant [1599, 1767, 1722, 565]. Plants [525]. Plato [1086, 1882, 143]. Platonic [1810]. Plattmaker [1911]. Plessner [1761]. Plictho [66]. Pluralists [1346]. pluralité [1108, 775]. Plutarch [563]. Plutarchs [563]. Pluto [1165]. Pneumatic [1078, 289, 115]. Pneumatica [1836]. Pneumatics [307]. Poem [293]. Poetic [125]. Poetry [1258]. Poets [350]. Point [1218, 1665]. Points [704, 1751]. polarisation [1615]. Police [1656, 1184]. Policy [1278, 1571, 1212, 1619, 347]. Polio [846]. Poliomyelitis [612]. politica [665]. Political [1278, 1216, 1734, 144, 1212]. Politics [105, 1924, 1865, 1679, 1451, 1685, 583, 1928, 973, 246]. Polk [1798]. Pollock [1365]. Pologne [694]. polskiej [342, 1236]. Pomerans [276]. Poncé [1242]. Ponty [927]. Poole [1023]. Popper [1333]. Popple [756]. Popular [1258]. Population [468, 1745]. Porcelain [979]. Porep [274]. Pork [141]. Porter [1777]. Portfolio [177]. Portrait [714, 35, 1119, 864, 1486]. Portraits [1544, 864]. Portugal [1601]. Portuguese [1130]. Position [1083]. Positivism [59, 1368]. Positivist [131, 699]. Post [423, 1451]. Post-Aristotelian [423]. Posthumous [317]. Postscript [172]. Postwar [1278, 1212]. Poulle [1484]. pour [1770]. Powell [807, 73, 767]. Power [1833, 1022]. Powers [272]. Practical [644, 223, 1106]. Practice [610, 520, 1540, 352]. Praeface [1300]. Prag [1148]. Prager [311, 710, 1491, 1836]. Pragmatism [37]. Prague [1676, 245]. Prairies [1426]. praktischen [98]. Praxis [1291]. Pre [1805, 1470, 1295, 1467]. Pre-Columbian [1470]. Pre-Critical [1805]. Pre-Modern [1467]. Pre-Socratics [1295]. Precautionary [83]. Preceded [36]. Précédée [1616]. Precious [1183]. Precision [1218]. Prediction [46]. Predictions [1165]. predstavlenii [1813]. Prehistoric [1835]. Preiser [928, 1607]. Preliminary [32, 737]. Prelude [1882]. Premature [395]. premiers [1233]. preryvnosti [923]. Presence [1848]. Present [1458, 590, 488, 303, 1086, 1180, 1829, 56, 1095]. Presentation [501, 1274]. Presented [312, 625, 322, 460]. Preservation [306, 11]. Preserves [979]. President [1686, 1627]. Presidential [1119]. Presocratic [471, 1346]. Pressler [1098]. Pressure [1088]. Preston [303]. Presumed [590]. Prevention [760]. Prevost [1930]. Price [1247, 1573, 1454, 1594]. Prichard [903]. prima [367]. Primer [216, 570]. primera [1707]. Primitiven [218]. primo [179]. primtif [811]. Primus [1760]. Princeton [1203]. Principia [1147, 457, 174, 716, 461, 427]. Principle [1019, 1631, 780, 1475, 1573, 718]. Principles [748, 651, 1721, 1441, 2, 716, 1537, 28]. Printed [1343, 1693, 1230, 960]. Printing [187]. Prinzipien [1847]. Priorities [1582]. Prismata [1824]. Prismenfarben [157]. Private [951]. Prix [637, 847]. Prize [1395]. probabilité [1108]. Probability [1176, 1534, 975]. Problem [168, 1013, 923, 828, 554, 866, 841]. problema [936, 923]. Problemam [1783]. Problemata [1081]. Problematum [79]. Probleme [654, 1472]. Problemi [1890]. Problems [1638, 1092, 1198, 1328, 255, 1808, 81, 480, 1021, 1783, 633, 555, 1138]. Proceedings [1837, 1227, 944]. Process [653]. Processes [1136, 320]. Proclus [186]. Productivity [1560]. Products [1905]. Profession [1725, 1873, 1354, 1466, 1811, 1682, 1920]. Professional [1580]. Professionalization [1199]. Professions [96, 1334]. Professor [1622, 1486]. Profezia [1188]. Prognostics [473]. Program [634]. Progress [1597, 354, 220, 1638]. Progressive [1798, 1597, 1710]. Progressives [108]. Project [1802]. Projectile [833]. Projectile-Throwing [833]. Prolegomena [1703]. Prometheus [1022, 56]. promotion [432]. Propaganda [757]. Propagation [625]. Prophatius [145, 591]. Prophet [901, 719]. Prophetess [1505]. Prophets [943]. Proportionals [12]. Proprietary [1518]. prostranstva [923]. Protein [1883]. Protestantisme [730]. Protestants [1614]. Proteus [388]. Proust [1312]. prouvant [775]. Provenance [595]. Providence [1736]. Provinces [155]. Provincials [946]. Provine [468]. Prozesse [1563]. Prudnikov [1921]. Psalters [685]. Pseudo [616, 519, 1081, 1899]. Pseudo-Aristotele [519]. Pseudo-Aristotle [1081]. Pseudo-Hippocratic [616]. Pseudo-Thomas [1899]. Pseudoscience [1925]. PSI [1100]. psicologia [1890]. psicoterapia [740]. Psyche [1242]. psychiatrie [561, 1891]. psychiatriques [561]. Psychiatry [242, 1896, 149, 760]. Psychical [1076]. Psycho [1516]. Psycho-Physiology [1516]. Psychoanalyse [1853]. Psychoanalysis [1365]. Psychoanalytic [1139]. Psychological [1538, 1181]. Psychologie [1590]. Psychologist [719]. Psychology [93, 1100, 1237, 1590, 1748, 184, 446, 1503, 1052, 81, 1344, 878, 1695, 184, 1746, 977, 1392]. Psychometry [320]. Psychopathology [1538, 1052]. Psychophysical [940]. Psychotherapy [1463]. Ptolemaic [1871]. Ptolemy [1167]. Public [743, 1119, 223, 1012, 347, 1895, 1354, 757]. Publication [337]. Publications [132, 1385, 1171]. publicznej [660]. publiques [1650]. Published [1588, 440, 1326]. Publisher [937]. Pugwash [770]. Purcell [1643]. Purkynes [483]. Pursell [355]. Pursuit [1286, 1886].

Pyenson [1928]. Pythagorean [1882]. Pythagoreanism [834].

qarastun [1839]. Quantenhypothese [273]. Quantitative [969]. quantités [758]. Quantum [650, 1177, 1387, 492, 530, 771, 1812, 950, 1573, 415]. Quarta [1170]. Quarter [328]. Quarterly [1574]. quattro [842]. Quelques [227]. Quentin [1477]. Quer [570]. Querner [947]. Query [1681]. Quest [1313, 770]. Question [1582, 161]. Quff [564]. Quiet [1532]. Quine [1507]. quinquagesimum [532]. Quinquevalent [504]. Quotations [1532]. Qurra [1839]. Qusta [1396].

R [888, 1837, 805, 1687, 1069, 1618, 64, 1538, 1657, 1861, 1559, 1566, 1017, 934, 1769, 471, 1116, 1339, 1886, 625, 126, 985, 1561, 748, 768, 162, 1802, 703, 789, 1600, 191, 1206, 1257, 869, 1686, 1350, 489, 1772, 580, 1691, 1009, 1435, 559, 1346, 1348, 186, 979, 127, 462, 555, 873, 68, 1251, 612, 758, 836, 919, 1089, 363, 657, 1171, 1598, 123, 1073, 1068, 1349, 1051]. R. [1746]. Rabbi [49]. Rabbitt [73]. Rabelais [1335]. Rabi [75]. Rabinovitch [975]. Races [543]. Rachel [1462]. Racial [502, 482]. Racism [1887]. Radcliffe [1905]. Räderuhr [1698]. Radial [202]. Radiation [1166]. Radical [1069]. Radicalism [1271]. Radio [1926, 1404]. Radioactive [1036]. Radioactivity [578, 103, 1264]. Radiochemistry [578]. Radulphus [747]. ragione [1188]. Railway [485]. Rainald [1563]. Rainer [1816]. raison [1488]. Raisonnement [920]. Rakes [1742]. Ralph [833, 1397, 1680, 535]. Rama [709]. Ramelli [1496, 1496]. Ramsay [1261]. Rand [1746]. Randell [1006]. rang [432]. Ranger [1802]. Rankine [9]. Rao [709]. Rapaport [1234]. Rapp [1101]. Rapports [550]. Rashed [1396, 933, 1108, 1105]. Raspe [6, 271]. Rather [630]. Ratio [503]. Rational [1591, 1729, 878]. ratione [750, 619]. Ratzel [1785]. Raum [213]. Raven [531]. Ravetz [625, 555]. Ray [877, 1599, 907, 914, 1145, 765]. Rayhan [1649]. Rayleigh [321, 160]. Raymond [1204, 479, 1060, 584, 222, 866]. Rays [941, 1036]. razvitiia [1813]. Razzaz [1251]. Razzell [1696]. Re [722]. Re-Evaluation [722]. Reaction [1786, 1387, 948, 320]. Read [1781]. Reading [1131]. Readings [667, 1100, 1538, 1895, 516, 1099]. Real [1126]. Realism [1144, 828]. Realities [899]. Reality [1077, 533, 1329]. Realm [1359]. Reason [1592, 699, 577, 192, 430, 1257, 68]. Rebel [723]. Recapitulation [590]. Reception [1678, 1153, 1866, 642, 391]. Rechenbuch [98, 1900]. Recherche [5, 1770, 1416]. recherches [920]. Recht [1414]. rechte [750]. Rechung [801]. Reclus [1915]. Recollections [582]. Recorded [1327]. Recording [883]. Records [635, 1171]. recta [750]. Recurring [760]. Redaktion [1112]. Reddy [365]. Rede [959]. Rediscovery [1253]. Reduction [112, 1138, 554]. Reece [1177]. Reech [1858]. Reese [1310, 1341]. Referee [1129]. Reference [1589, 648]. Refining [1219]. Reflecting [955]. Reflections [1928, 1869, 1927, 150]. réflexion [1615]. Reform [1768, 1361]. Réforme [1488]. réfraction [1615]. Refuted [1507]. Regelungen [63]. Regiomontanus [1270]. Region [73]. Regional [552]. Register [1469, 1844]. Règles [1770]. Regna [1099]. Regnault [1338]. Regulative [1721]. Reich [769, 981, 1841, 1685]. Reid [1144, 162, 194]. Reign [1830]. Reihe [1330]. Reilly.J [88]. Reingold [919, 1203, 1660]. Reinhard [1892, 1545, 574, 1792, 1795]. Reinhold [1358]. Reinterpretation [722]. Reiser [1830]. Reiss [986]. Réjane [1233]. Related [303, 1138]. Relating [552]. Relation [857, 267, 1629, 655, 462]. Relations [241]. Relationship [232]. Relationships [1047]. Relativi [1846]. Relativity [780]. Relazione [842]. Relevance [789]. Religieux [1058, 1776]. Religion [455, 1917, 1901, 948, 1580, 1613, 379, 810, 407]. Religious [1005, 1614]. Remarkable [1311]. Remarks [1328, 781]. Remedies [1722]. Remes [1288]. Renaissance [1296, 711, 1904, 1473, 1556, 1253, 1011, 751, 620, 972, 499, 960, 1875, 1354, 1181, 1547, 1611, 409]. Renate [864]. Renaudot [757]. rendues [1108]. rendus [695]. Rene [1622, 1770, 524, 1820, 763]. René-Edouard [763]. Renne [1838]. Renzo [551]. Reorganization [144]. Repercussions [866]. Reply [965, 967]. Report [398, 883, 118, 442]. Repositories [118]. Representation [430]. Representations [1041]. Represented [864]. représenter [758]. Reprint [337]. Republic [1286]. Rescher [414]. Research [1278, 159, 306, 162, 825, 951, 1212, 949, 1619, 1076, 1456, 1443, 1880, 581, 1509, 1644, 1644]. Researcher [1117]. Researches [1435, 903, 406]. Resources [515]. Response [1618, 1720, 1569, 401]. Responses [1516]. Responsibility [1092]. Rest [286]. Restivo [1090]. Results [689]. Reti [779]. Retrospect [306, 1435]. Retrospective [1645]. Retrospectroscope [1602]. Reuchlin [93]. Revelations [1050]. Revell [419]. Reverdin [631, 631]. Review [1596, 1601, 225, 1604, 1853, 277, 1713, 614, 254, 264, 301, 888, 1330, 1411, 1490, 1824, 712, 1656, 1773, 105, 274, 414, 1428, 1434, 1492, 1574, 1664, 1924, 1661, 933, 1252, 1837, 1897, 1524, 368, 1057, 1458, 740, 805, 652, 1359, 1650, 1281, 871, 1832, 107, 226, 278, 515, 536, 767, 1001, 1071, 1088, 1266, 1372, 1571, 1850, 1928, 1469, 1733, 1687, 1292, 1108, 616, 745, 676, 1069, 1199, 799, 1814, 1618, 743, 1927, 159, 463, 1203, 1398, 1700, 40, 1310, 1739, 1753, 1706, 1190, 1731, 1752, 1815, 675, 1408, 339, 1884, 1595, 483, 532, 1848, 218, 1427]. Review [247, 1770, 1921, 246, 385, 1115, 667, 844, 734, 1890, 902, 345, 741, 407, 1611, 1856, 650, 701, 1425, 670, 891, 1086, 1144, 352, 635, 715, 1015, 1205, 1465, 1851, 315, 64, 388, 717, 1240, 1393, 445, 632, 1062, 1789, 117, 624, 986, 1354, 93, 1100, 1237, 1538, 1590, 1748, 1756, 15, 273, 1265, 1432, 1657, 63, 464, 1150, 973, 1202, 1454, 1734, 124, 185, 1466, 1543, 379, 1777, 364, 1638, 1431, 1766, 1151, 184, 242, 1196, 1235, 1746, 1861, 1602, 131, 251, 669, 692, 719, 932, 312, 316, 832, 1562, 1779, 1896, 1477, 1758, 1888, 413]. Review [423, 374, 890, 1170, 1413, 1692, 1092, 355, 17, 25, 649, 1065, 1559, 1793, 435, 519, 627, 253, 880, 1782, 375, 71, 1189, 1013, 1185, 262, 1286, 468, 870, 1371, 1184, 601, 744, 1172, 1887, 469, 1191, 490, 1147, 1197, 1542, 446, 527, 695, 809, 843, 59, 344, 384, 1374, 219, 883, 1554, 1876, 120, 895, 1375, 1110, 560, 1152, 1422, 443, 1892, 608, 1854, 405, 1370, 481, 487, 1641, 327, 535, 1662, 1204, 1566, 1858, 1017, 1067, 1502, 1885, 193, 370, 1228, 1830, 1737, 889, 769, 810, 265, 18, 1016, 1415, 1333, 1903, 934, 1587, 1771]. Review [1833, 755, 65, 188, 192, 224, 338, 842, 892, 1769, 1653, 1451, 798, 30, 1047, 471, 434, 708, 935, 971, 1296, 28, 1373, 1727, 100, 382, 426, 525, 877, 1008, 1048, 1157, 1290, 1343, 1399, 1494, 1537, 1570, 1599, 1694, 1745, 1588, 556, 1531, 479, 1800, 1116, 1335, 430, 714, 826, 1232, 1358, 1430, 1827, 804, 1180, 1262, 1889, 661, 1750, 444, 478, 1159, 1874, 103, 1099, 1545, 1847, 1533, 711, 275, 1506, 1620, 428, 1308, 1412, 1555, 1591, 1637, 1690, 1843, 1878, 1169, 362, 654, 1898, 43, 74, 119, 134, 241, 408, 907, 908]. Review [947, 992, 1640, 1684, 1714, 1801, 486, 419, 922, 1368, 671, 1283, 1356, 297, 518, 1360, 1557, 425, 978, 1834, 797, 720, 756, 603, 1140, 1786, 514, 1796, 1768, 827, 630, 1693, 243, 1426, 1019, 1523, 1007, 1470, 1835, 1244, 95, 522, 1553, 1765, 1825, 1904, 57, 109, 132, 570, 1045, 1433, 1658, 665, 319, 351, 1302, 812, 975, 1480, 1339, 440, 1778, 1738, 1111, 1395, 753, 1473, 55, 923, 1589, 1143, 256, 1198, 1239, 152, 903, 1423, 1568, 611, 1307, 1256, 1194, 1309, 329, 350, 1491, 149, 304, 807, 1328, 1102, 647, 1288, 904, 1917]. Review [1503, 1565, 1056, 875, 1334, 154, 255, 746, 1295, 272, 569, 637, 725, 1139, 150, 646, 656, 666, 833, 1053, 1134, 1254, 1355, 1610, 1757, 1886, 27, 92, 259, 309, 1742, 410, 625, 938, 1142, 1416, 1177, 472, 1300, 1547, 1901, 925, 1052, 982, 924, 62, 326, 1452, 910, 1775, 106, 674, 1822, 88, 126, 306, 467, 492, 302, 530, 557, 678, 771, 723, 752, 898, 985, 1003, 1230, 1361, 436, 989, 1178, 1461, 295, 1529, 1592, 387, 267, 381, 1153, 1561, 509, 1186, 748, 699, 739, 768, 788, 790, 828, 829, 830, 945, 1176, 811, 162]. Review [735, 1802, 342, 660, 694, 1479, 1005, 664, 161, 1303, 217, 1489, 1367, 920, 703, 458, 1060, 1429, 1823, 158, 365, 276, 1096, 584, 704, 1774, 1729, 789, 1826, 70, 183, 796, 847, 1600, 1798, 1504, 1187, 1482, 422, 1754, 653, 867, 1705, 626, 153, 574, 825, 1282, 1556, 1642, 1689, 1728, 1877, 99, 1500, 20, 102, 125, 222, 314, 1510, 1919, 121, 550, 750, 874, 1560, 1050, 1761, 1838, 35, 191, 244, 266, 268, 317, 318, 429, 447, 461, 488, 754, 787, 899, 1145, 1206, 1227, 1654, 1929, 24, 196, 299, 349, 449, 459, 1041, 1242]. Review [1261, 1336, 1397, 1532, 1112, 1304, 1085, 1195, 1257, 1511, 1643, 981, 1066, 721, 869, 951, 1686, 1546, 260, 310, 521, 524, 673, 1105, 1299, 1350, 578, 1264, 1785, 1915, 636, 1812, 1685, 764, 1014, 1859, 800, 37, 558, 1528, 1593, 600, 1044, 1509, 1119, 1200, 1284, 1534, 1882, 1911, 21, 441, 942, 1094, 1708, 533, 1424, 1621, 91, 489, 651, 1156, 1730, 759, 882, 1414, 1842, 360, 885, 1248, 1552, 1759, 1467, 1609, 513, 1098, 353, 417, 359, 1704, 346, 1249, 1818, 1117, 1772, 1712, 1468, 1018, 512, 553, 1258, 1732, 940, 868, 1922, 1464, 1181, 1784]. Review [929, 38, 1845, 1883, 128, 1289, 1608, 1173, 1487, 1655, 1683, 1820, 1930, 893, 31, 629, 580, 98, 617, 1616, 187, 240, 1253, 1478, 1740, 386, 1146, 32, 949, 1417, 609, 1474, 1549, 1174, 1055, 1881, 1305, 634, 742, 1597, 1508, 1114, 1460, 571, 1613, 1651, 1703, 1844, 1406, 1483, 1749, 554, 1569, 1879, 1497, 26, 298, 1619, 1603, 1710, 1893, 1755, 1183, 1536, 227, 211, 307, 322, 324, 354, 439, 475, 931, 980, 1101, 1698, 1799, 411, 1691, 1849, 1476, 1646, 424, 1836, 948, 321, 1023, 1072, 438, 480, 808, 896, 1914, 1463, 1061]. Review [1623, 258, 311, 747, 1351, 1551, 1009, 1435, 1925, 738, 1594, 1644, 221, 1909, 19, 250, 939, 724, 1743, 1021, 801, 1484, 270, 376, 897, 1364, 561, 1104, 529, 223, 1193, 1090, 1563, 1663, 707, 1136, 176, 213, 406, 559, 622, 668, 926, 1058, 1097, 1342, 1472, 1496, 1567, 1659, 1688, 1776, 1797, 1763, 1346, 1179, 58, 552, 737, 1913, 1255, 523, 872, 1241, 1314, 1605, 1926, 194, 465, 526, 566, 1247, 1614, 1855, 1419, 1420, 1348, 1530, 1622, 1024, 1456, 757, 437, 516, 1201, 1781, 1783, 1846, 573, 34, 303, 1564, 36, 1285, 383, 1238]. Review [1345, 1402, 1920, 1617, 1788, 1046, 1741, 1894, 108, 343, 950, 761, 1311, 418, 220, 575, 1063, 1338, 97, 186, 709, 1906, 1829, 214, 195, 1645, 905, 181, 369, 604, 1291, 1744, 979, 1421, 1507, 1182, 1548, 1607, 655, 623, 697, 420, 485, 927, 1020, 1918, 534, 760, 583, 887, 1535, 613, 1344, 1900, 101, 271, 127, 371, 367, 1767, 1821, 1880, 1404, 1751, 548, 41, 118, 263, 325, 462, 555, 930, 1154, 1394, 1526, 1699, 1817, 1697, 1012, 66, 1405, 482, 182, 1905, 377, 894, 562, 884, 1103, 1707, 1831, 793, 906, 1365, 1791]. Review [1787, 1341, 1369, 1709, 1792, 67, 710, 1485, 33, 873, 1306, 1459, 442, 1539, 1916, 749, 1352, 795, 1455, 1260, 610, 1226, 577, 766, 1234, 68, 151, 216, 296, 477, 864, 1011, 1312, 1337, 1544, 602, 1794, 976, 1138, 1525, 22, 1340, 1409, 1471, 1839, 252, 1233, 835, 1049, 1246, 1298, 1347, 1403, 1481, 839, 1902, 567, 1251, 1606, 130, 722, 1639, 612, 179, 215, 366, 409, 619, 865, 936, 983, 1188, 1652, 941, 1022, 29, 248, 572, 700, 716, 758, 1453, 42, 1004, 1735, 60, 1396, 943, 1790, 946, 372, 1059]. Review [1109, 1148, 1615, 1908, 1331, 836, 313, 433, 663, 713, 1113, 1301, 1495, 1910, 1711, 347, 421, 1118, 1229, 1780, 466, 189, 1313, 628, 919, 69, 618, 648, 1106, 1135, 1141, 1353, 1701, 61, 1363, 528, 1488, 1089, 257, 308, 363, 473, 520, 564, 1357, 1550, 1648, 96, 155, 1091, 878, 1158, 1243, 1392, 1501, 1575, 1695, 1263, 1702, 1462, 806, 1576, 412, 160, 762, 1499, 736, 156, 474, 565, 802, 1293, 1391, 1418, 1660, 1923, 658, 718, 901, 1064, 706, 1527, 633, 1862, 1895, 357, 606, 657, 1054, 1541, 269, 320, 470, 484, 531, 693, 803]. Review [879, 1860, 340, 341, 1612, 568, 157, 197, 228, 249, 415, 765, 1498, 615, 94, 579, 581, 909, 990, 549, 1736, 1297, 1475, 1899, 1236, 1457, 1841, 974, 1231, 1907, 546, 56, 89, 460, 984, 1171, 305, 328, 380, 448, 751, 763, 792, 1192, 1400, 1598, 1696, 1828, 1042, 659, 977, 1760, 23, 123, 245, 300, 323, 348, 356, 358, 838, 886, 1137, 1245, 677, 991, 1073, 1649, 582, 702, 952, 988, 1087, 1155, 1259, 1332, 1572, 1540, 1006, 1093, 432, 511, 1410, 1558, 576, 987, 1068, 1795, 900, 1294, 39, 517, 607]. Review [261, 1095, 1493, 1857, 133, 178, 876, 881, 1160, 1505, 1349, 551, 1175, 1107, 1362, 672, 840, 1813, 662, 361, 620, 972, 1250, 129, 212, 373, 416, 537, 791, 831, 937, 1043, 1287, 1366, 563, 621, 928, 1875, 190, 1891, 1816, 696, 837, 866, 1407, 1647, 104, 1573, 427, 431, 457, 476, 1726, 1912, 1010, 1486, 1665, 73, 122, 177, 944, 1070, 1852, 786, 1225, 378, 846, 510, 841, 770, 921, 1002, 1149, 16, 75, 1715, 845, 1747, 180, 337, 1819, 491, 1764, 90, 1401, 1762, 1840, 1120, 72, 698, 631, 605, 705, 794]. Review [1051, 834]. Revision [336]. revoliutsiia [1881]. Revolt [944]. Revolution [805, 1620, 1786, 1378, 1195, 1609, 1772, 1784, 1114, 1781, 873, 602, 1317, 190, 56, 264, 1257, 178]. Revolutionary [715, 101]. revolutionibus [1653, 1765]. Revolutions [1553, 1597, 1537, 886]. Reynolds [322]. Rezeptstruktur [1182]. Rg [1882]. Rh [1073]. Rheticu [84]. Rheticus [913, 367]. Rhetikus [33]. Rhine [1076].

Riaza [109]. Ricardo [570]. ricerche [409, 865]. Richard [1524, 107, 1071, 799, 40, 1562, 1758, 1737, 338, 514, 306, 1178, 1303, 626, 317, 1478, 1136, 906, 151, 477, 1089, 806, 909, 1294, 705, 1720, 505, 1502]. Richardson [1712, 153, 793]. Richter [1590, 653]. ricognizione [179]. Riddle [1718, 1838]. Rieber [1748]. Riemann [247]. Riese [1896, 1795]. Rieser [1088]. rifat [1251]. Righini [1257, 465, 1113, 1868, 1910]. Rigid [1185]. Rigorously [1858]. Riley [1600]. Rinascimento [409]. Rinpoche [801]. Ripley [1798]. Rise [443, 55, 1256, 439, 1699, 1049, 1331, 89, 1205]. Risse [876]. ritrovati [1485]. River [883, 73]. rivoluzione [433]. Road [872, 296]. Roads [460]. Roazen [1140]. Rob [319]. Robert [536, 345, 650, 1086, 1748, 71, 883, 1854, 487, 1333, 192, 1451, 479, 478, 907, 351, 350, 1742, 410, 1822, 302, 1529, 153, 1528, 1119, 1712, 475, 931, 1476, 321, 1179, 526, 1420, 1781, 575, 1338, 369, 1817, 1791, 421, 1135, 1363, 1695, 160, 901, 765, 677, 1095, 791, 431, 73, 75, 1120, 1276, 6, 1720, 1032, 1378, 1274, 1166, 447, 1772, 677]. Roberto [1407]. Roberts [75]. Robins [1007]. Robinson [1748, 1185, 769, 56]. Roche [885]. Rochot [1058]. Rocket [634]. Rocketry [94]. Rockets [634]. Rocks [1231, 1854]. Roderick [904]. Rodnyi [195]. Roe [1429]. Roebuck [96]. Roger [405, 1705, 1665]. Rogers [854, 310]. Rohault [318]. Roheim [769]. roku [1567, 838]. Role [688, 514, 777, 996, 742, 1341, 340]. Rolf [1115, 981, 1918, 928]. Rolle [1929]. Roller [1728, 603]. Rolling [234]. Rolt [1833]. Rom [651]. Roman [1700, 646, 155, 706, 1327, 980]. Romanes [996]. Romantic [50]. Romanticism [204]. Rome [566, 96]. Romer [578, 578]. Romuald [1564]. Ron [1759]. Ronald [490, 1568, 1842, 1022, 1862, 1895, 1505]. Ronan [300, 129, 831]. Ronchitti [1412]. Röntgen [765]. Root [7]. Roots [1100, 1694, 1533, 1244, 461, 1127]. Roper [1395]. Rose [1556]. Rosemary [743]. Rosen [1184, 367]. Rosenberg [1922, 1834, 797]. Rosenfield [527]. Rosenkrantz [1861]. Rosenthal [474]. Rosetti [66]. Roshdi [933, 1108, 1105, 1396]. Rosicrucian [1016]. Rösing [1594, 1640]. Rosminiani [1703]. Rosminiani-Celli [1703]. Ross [878, 719]. Rossi [215, 433, 1301]. Rossiter [1427]. Rösslin [1905]. Rotblat [770]. Roth [1466, 574, 1857]. Rothschuh [1831, 876, 1831]. Rothstein [673]. Rouse [1405]. Rousseau [870]. Rowan [1269]. Rowland [1888, 285, 916]. Roy [1777, 1116, 664, 1819]. Royal [225, 1281, 1815, 518, 1768, 1825, 529, 1913, 437, 1544, 1606, 598, 1129]. Royale [720]. Roysdon [1923]. Rozella [132]. Rubin [1351]. Rüdenberg [1067]. Rüdiger [274]. Rudolf [6, 30, 271, 308, 380, 1799]. Rudolph [1052, 1875]. Rudolphinae [522]. Rudolphinischen [522]. Rudwick [875]. Ruestow [987]. Rugged [42]. Rukeyser [368]. Rule [688, 503]. Rules [1215]. Rumford [223, 223]. Rupert [17, 1415, 1654, 1605, 372]. Rürup [1545]. Rus [1813]. Rush [715]. Rushton [1886]. Russel [728]. Russell [867, 121, 377, 412, 1753, 1750, 1796, 1005]. Russell-Gebbett [1796]. Russett [1569]. Russia [1313, 840, 1813]. Russian [642, 818]. russische [374]. Russo [58]. Ruth [1181, 1749, 383, 1311, 760, 1912]. Rutherford [952, 1155, 689, 1036, 103, 582, 988, 1259]. Ryal [71]. Rydén [1767].

S [1601, 1773, 767, 159, 1753, 352, 632, 1100, 316, 870, 1858, 1903, 1833, 338, 479, 1533, 907, 425, 1786, 1904, 807, 875, 1186, 704, 626, 1877, 1919, 521, 1915, 1528, 600, 1044, 1730, 882, 1467, 513, 868, 1881, 439, 1781, 220, 1020, 482, 864, 1409, 1298, 42, 836, 204, 320, 56, 1649, 1647, 1573, 73, 401, 510, 88]. S.J [573]. sa [1487]. Saadia [887]. Sabatier [1319]. Sabetai [1760]. Sabloff [1009]. Sabra [434]. Saburo [1141]. Sachs [830]. sächsische [1875]. sächsischer [802]. Sachwörter [702]. sacras [619]. Safavid [520]. Saffron [473]. Said [309]. Saidan [1903, 521]. Sailors [634]. Sajner [532]. Sal [1090]. Salah [933, 1105, 1480]. Salam [771]. Sale [1145]. Salerno [473]. Sales [1469]. Salinity [1342]. Sally [1367, 1828]. Salmane [1650]. Salmon [789]. Salomon [973]. Salonen [1293]. Salter [484]. Salts [504]. Saltzer [1824]. Salvator [215]. Samaw’al [933, 1105]. Sami [606, 1054, 546]. Sämtliche [1330]. Samuel [318, 1511, 1253, 224, 477]. San [1768]. Sanborn [1088, 1286, 223]. Sanders [1588, 37]. Sándor [1457]. Sandra [1555, 1749, 1507]. Sane [131]. Sang [1452]. Sang-Woon [1452]. sanguine [1360]. Sanskrit [709, 749, 1352, 749]. Santillana [122, 150, 1128]. sapiens [1356]. Saporta [1616]. Sarfatti [31, 887]. Sarton [403, 507, 823, 1388, 1450, 452]. Satellite [64]. satelliti [842]. Saul [1603]. Saunders [1007]. savant [1705, 315]. savantes [695]. Savin [717]. Sawers [306]. Saxon [1722]. Saxons [1722]. Scaglia [311, 710]. Scales [1531]. Scalpel [153]. Scattering [689]. Schädels [564]. Schafer [1840, 1816]. Schaffer [826]. Schaffner [579]. Schatzberg [1258]. Scheibe [950]. Scheiner [114]. Schepers [1414]. Scherz [435]. Scheubel [336]. Schiller [1262, 8, 1446, 1262, 1889]. Schiller-Nationalmuseum [8]. Schiøler [980]. Schipperges [947]. Schism [1465]. Schlumberger [1661]. Schmauderer [1096]. Schmitt [620]. Schmitz [1433, 1433]. Schmitz-Moormann [1433, 1433]. Schmorak [1309]. Schneer [62, 1536]. Schofield [192]. Scholarship [1125, 574]. School [1018, 949, 1871, 297]. Schopf [1019]. Schrift [359]. Schriften [1330, 1708, 1844, 1548, 619, 753]. Schrifttum [572]. Schrøder [1193]. Schrödinger’s [1387]. Schullian [1194]. Schultz [184]. Schüpfheimer [1408]. Schur [1235]. Schwartz [882, 477]. Schweigger [204]. Schwendinger [1069, 1069]. Schwind [1905]. Science [225, 1278, 1279, 288, 1928, 1199, 1927, 159, 1815, 1427, 1115, 345, 402, 1393, 379, 1777, 312, 1562, 1805, 649, 1037, 375, 1865, 71, 1679, 490, 398, 1370, 338, 1874, 908, 922, 603, 1786, 1275, 777, 1565, 1610, 306, 898, 1361, 1461, 589, 400, 829, 1367, 704, 626, 1282, 1689, 1728, 1510, 429, 447, 1242, 1304, 144, 1212, 1686, 1859, 1579, 1119, 1467, 1249, 1018, 1173, 1683, 187, 1740, 1879, 1619, 322, 424, 948, 1023, 1072, 1925, 1193, 1090, 395, 668, 138, 282, 496, 552, 682, 851, 1028, 1210, 1439, 1669, 1934, 1614, 1419, 1530, 757]. Science [516, 1125, 97, 195, 604, 147, 1526, 1699, 1817, 1103, 1365, 610, 1226, 864, 1633, 998, 1481, 409, 1331, 964, 189, 1313, 69, 618, 1135, 1353, 1089, 399, 1317, 89, 460, 1171, 987, 1580, 261, 361, 1444, 1875, 837, 1852, 180, 816, 834, 1472, 692, 1835, 56, 226, 246, 1115, 1371, 59, 1422, 1771, 478, 275, 1302, 1822, 1529, 739, 788, 830, 1303, 673, 1248, 893, 270, 1688, 872, 220, 1906, 1020, 1821, 462, 1369, 477, 976, 1022, 1118, 249, 1043, 696, 866, 809, 1458, 463, 339, 624, 986, 1432, 973, 1454, 1861, 1876, 1047]. Science [1335, 1637, 514, 1917, 1452, 735, 1005, 653, 222, 787, 1200, 1284, 1018, 580, 617, 1483, 1594, 1763, 343, 101, 1791, 1011, 1331, 347, 1091, 1660, 1736, 1297, 460, 1042, 886, 1137, 431, 16, 1764]. Sciences [225, 1815, 405, 1290, 711, 241, 243, 1565, 550, 1085, 1597, 1783, 749, 1352, 151, 1525, 1695, 1457, 1250, 1002, 1731, 601, 695, 694, 724, 58, 1748, 925, 121, 1528, 553, 1101, 943, 363, 1540, 346, 512, 1525, 190, 695]. scientiam [1490]. scientiarum [750, 1476]. Scientific [1281, 345, 717, 1638, 184, 413, 71, 1013, 1286, 405, 481, 826, 540, 392, 514, 1214, 812, 1589, 1271, 1256, 92, 982, 1003, 1178, 1592, 387, 828, 664, 161, 1367, 334, 1227, 349, 449, 1257, 951, 764, 1593, 205, 533, 651, 1730, 31, 1174, 1597, 1719, 211, 948, 896, 738, 939, 167, 1663, 1136, 1913, 1319, 1622, 1781, 1046, 575, 1443, 482, 793, 1004, 1576, 412, 689, 607, 178, 1813, 537, 791, 696, 1726, 1010, 1819, 414, 344, 907, 904, 1532, 1258, 1788, 623, 462, 555, 1394, 1735, 1287]. scientifica [1555, 1910, 433]. scientifique [1701, 1488, 1433]. scientifiques [1731]. Scientist [502, 312, 627, 375, 265, 88, 162, 584, 195, 776, 340, 1160, 1169, 1850, 714, 1117, 896]. Scientists [1927, 1145, 1686, 634, 1925, 642, 864, 1860, 725, 1397, 1425, 692, 869, 1685, 1818, 770]. Scienza [1555, 465, 1878, 179, 936, 1301, 551]. scienze [1843]. Scofield [1848]. Scope [93]. Scorpion [1547]. Scotland [1343]. Scott [925, 1003, 301, 657, 476]. Scottish [803, 1178, 1004]. Scriba [571, 22]. scritti [409, 713]. Sculp [756]. Sea [251, 692, 1013, 34]. Seaborg [75]. Seamanship [422]. Search [1756, 203, 869, 5, 491]. Searle [1924]. Sechzig [74]. secoli [179]. secolo [409, 366]. Second [349, 1439, 337]. Secondary [1588]. secondo [1555]. Secret [95]. Secrets [608, 1050]. Section [695]. Sects [673]. Secunda [1490]. Sedimentary [1231]. Seeger [1204, 479, 1060]. Seele [1403, 841]. Seeley [1737]. Seen [542]. Segrè [278]. Seicento [1188]. Sein [274, 295, 376]. seine [891, 1902, 858]. seiner [1572]. Seiyu [1567]. Sekundärliteratur [702]. Selectarum [568]. Selected [1746, 375, 1662, 1588, 1198, 678, 1461, 70, 314, 754, 1284, 1424, 1730, 386, 516, 303, 1453, 648, 357, 1006, 1401, 1100, 1157, 767, 1044, 900]. Selection [1005, 118, 1217, 576, 1706, 1793, 672]. Selections [1656, 343, 717]. Selective [883, 611]. Self [205, 776]. Self-Identity [205]. Self-Image [776]. selvaggi [663]. Seminar [1781]. Seminara [1552]. Seneta [1790]. Sense [1628, 1221, 308]. Senses [1837, 729]. senso [1843, 1497]. Sensory [1806, 1076]. Separate [1631, 475]. Sepp [621]. September [883, 961, 1098, 322, 1918, 1520, 961]. Serber [75]. seredina [1256]. Sergio [665, 663]. Series [1490, 1170, 349, 749, 1352]. Service [250]. Settlements [756]. Seven [207]. Seven-Place [207]. Seventeenth [624, 265, 338, 1631, 1911, 1114, 1255, 757, 1357, 987, 686, 1875, 431, 1322, 1216, 962, 333]. Seventeenth-Centry [1357]. Seventeenth-Century [1255, 1216, 962, 333].

Seventh [1837, 496]. Sex [1514, 1306, 1653]. sex-angulaire [1306]. Sexual [1466]. Seymour [413, 1007, 1369, 1573]. Seyyed [1610, 1481]. sfida [1410]. Shātir [732]. Shadow [104]. Shadows [1649]. Shakespeare [310]. Shakespearean [1767]. Shakti [565]. Shamin [1883]. Shankland [907]. shao [1107]. Shape [1793]. Shapere [1149]. Shapiro [375, 1196, 102]. Shapley [42]. Sharon [1700]. Shatir [1409]. Shaw [609]. Shea [965, 805, 1257]. Shea-Wolf [965]. Sheep [661]. Shepard [1116]. Sheperd [1193]. Sherover [464]. Shield [107]. Shiers [615]. Shifting [689]. Shigeru [1284]. Shih [1106]. Shih-ch’iu [1106]. Shine [1664]. Ships [422]. Shlomo [1761]. Short [1156]. Show [1784]. Showman [1559]. Shrock [1742]. Shryock [505]. Shu [1107, 98]. Shu-shu [1107]. Shujā [1522]. Shukhardin [1881]. Shumaker [711]. Sibinga [320]. Sibly [544]. Sickness [469, 1895]. Sidar [1428]. Side [1397]. Sidereo [842]. Sidney [829, 66]. Sidonius [1471]. siebzigsten [1098]. siècle [1616, 886, 1151, 885, 156]. siècles [1233]. Siegel [30, 1052, 308]. siglo [1613, 1707]. siglos [1551]. Sigmund [1312, 810, 906]. signe [1547]. Signers [1492]. Significance [689, 1294, 1288]. Significant [303, 1748]. Silk [66]. Silvester [837]. Silvio [1302, 523]. Simmons [485]. Simon [1132, 1517, 261]. Simple [819]. Simpson [910]. Sims [1620]. since [1454, 1044, 1623, 1894, 1526, 1331, 1660]. Sinclair [258, 1089, 1263]. Sine [207]. Singular [337]. sinibus [960]. Sir [1194, 1269, 526, 379, 440, 161, 1504, 1117, 1712, 1364]. Siraisi [1250]. sistema [1497]. Situation [1291]. Sivin [32, 1764]. Six [127]. Sixteenth [1479, 1875, 1486, 1386, 1583]. Sixteenth-Century [1386, 1583]. Sixth [282]. Size [1447]. Sketch [146]. Sketches [1727, 1503, 1883, 998, 529]. Skies [323]. Skinner [1575]. Skizzen [1096]. Skolnik [1341]. Skopová [245]. Sky [1185]. Slater [1576]. Small [1260]. Smell [308]. Smet [1484]. Smeur [362]. Smirnov [1170]. Smit [1290]. Smith [877, 1159, 132, 1053, 306, 99, 1055, 1828, 877, 132, 1778, 1429, 1754, 1605, 1382]. Smithsonian [737, 1220]. Smitt [1019]. Smyth [1825, 1825]. sobre [1045]. Social [1733, 1199, 1618, 502, 339, 352, 1734, 1184, 481, 1827, 1143, 1239, 461, 1528, 1772, 868, 949, 1925, 325, 555, 930, 1394, 943, 698, 1551, 1136, 1913, 1313]. Socialism [481]. sociedad [798, 57]. società [665]. société [1108]. sociétés [695]. Societies [1286, 1171]. Society [1279, 1458, 339, 1748, 17, 265, 1835, 1778, 995, 1886, 1442, 787, 1284, 1594, 529, 437, 697, 613, 1011, 946, 1313, 1089, 340, 431, 598, 398, 1129, 1223, 1449, 1562, 375, 1139, 1784, 1090, 1419, 1341, 1711, 474]. Sociological [1620, 1394]. Sociologist [629, 766]. Sociologists [1069]. Sociology [676, 1069, 1683, 879]. Socratics [1295]. Soddy [1264]. soedinenii [1883]. Software [1292]. sogenannte [213]. Sokolovskaia [1589]. Soldiers [635]. soleil [1552]. Solid [1576]. Solid-State [1576]. solide [1487]. Solitude [1420]. Solla [1454, 1594, 1247]. Solomon [1100, 757]. Solov’ev [195]. Solution [1887, 335, 633]. Solvay [1623]. Solvent [165]. Solving [1013]. Some [1518, 857, 1164, 1385, 1869, 979]. Somers [826]. Sometimes [661].

Sommerfeld [74]. Son [1217, 886]. Song [1882]. Sonne [981]. Sonnedecker [1541]. Sonnen [358]. Sonnen- [358]. Sorabji [799]. Soul [527, 1309]. Souls [1888]. Sound [1740, 269]. Source [1885, 942, 1248, 182, 998, 248]. Source-Book [182]. Sourcebook [1100, 1237, 1538]. Sources [777, 306, 1230, 303, 118, 1233, 1172, 1284, 308, 1282, 249]. Southwest [951]. Soviet [1687, 580, 1082, 1344]. Sowjetmacht [1929]. Sozialistischen [1617]. Space [1321, 923, 1631, 1328, 557, 1529, 393, 1021, 1639, 596, 94, 909, 690]. Spaceflight [1620]. Spagyrical [1483, 1763]. Spain [208, 1275]. Spallanzani [1846, 567]. Spanish [71, 756, 1768]. Spark [1404]. Spatial [1632]. Special [1574, 1686, 1597]. Species [381, 161, 642]. Spectra [762]. Spectroscopic [326]. Spectroscopy [762]. Speculations [957]. Speculum [1647]. Speech [878]. spektralen [273]. Spence [177, 944]. Spencer [1560, 766]. Spengler [1816]. spera [234, 1759]. Sperienze [1846]. Sperimentale [1846]. Sperry [628]. Spezialfächer [418]. Speziali [636]. Sphere [1007, 271]. Spheres [1553]. Spiegel [947, 1640, 1594]. Spies [564]. Spinal [152]. Spink [1298]. Spiral [202, 782]. Spirit [1524, 1164]. Spiritual [1479]. Spiritualism [728]. Spontaneous [674]. Spouting [1681]. Sprachgebiets [418]. Sprachraum [1698]. spravochnik [1589]. Spronsen [196]. Sputnik [1686]. Spy [1659]. Square [7]. Square-Root [7]. Staat [1709]. Staaten [1066]. Staats [566]. Staats- [566]. Stacey [370]. Stadt [1057]. Stadtbibliothek [566]. Staff [1742]. Standardisierung [1066]. Stanley [1372, 719, 1830, 1053, 1775, 1200, 1351, 872, 228, 1042]. Starfish [1371]. Stars [1447, 125, 42, 1840, 373]. State [253, 1035, 1034, 1633, 289, 1869, 1576]. Statei [1845]. States [107, 1172, 1334, 1405, 205, 749, 1735]. Statesman [370]. Statics [755]. Station [634]. Stationary [355]. Stationen [1017]. statique [1079, 1063, 1839]. statische [360]. Statistcs [1534]. Statistical [454, 100, 975, 1176, 1442, 789, 1790, 789]. Statistics [1216]. Status [1879, 819]. Staudinger [449, 449]. Stavenhagen [1050]. Steam [355, 1833, 475]. Stearn [426, 1494]. Stearns [222]. Steck [572]. Steele [524]. Steffens [1091, 1417]. Stehr [1820]. Stein [1504, 750]. Steinmetz [388]. Sten [1283, 1732]. Stendhal [809]. Steneck [1458, 1297]. Steno [435, 303]. Step [205]. Stepan [1619]. Stephanus [1902]. Stephen [652, 515, 734, 59, 804, 945, 153, 1539, 1498]. Stephens [904]. Stephenson [1595]. Stereochemistry [855]. Sterling [1157, 1570]. Stern [1930, 534]. Sterne [981]. Sternkatalog [1480]. Stevens [370, 743, 1514]. Stevin [261]. Stewart [1166]. Sticker [1521, 1057, 157, 1332]. Stieb [328]. Stiles [141]. Stillerman [306]. Stillman [264, 1412, 1726, 964]. Stillwell [187]. Stipisic [124]. Stirpium [568]. Stoa [915]. Stock [837]. Stockholm [1815]. Stocking [903]. Stoehr [1925]. Stoic [1549]. Stokes [1048]. Stone [1835]. Storia [822, 1658, 465, 1890, 179, 409]. storiografia [1690]. Story [1281, 880, 95, 1738, 1500, 951, 761, 1311, 1829, 536]. Stotz [796]. Strabismus. [419]. Strachan [528]. Strahlungsgesetzes [273]. Strata [1070]. Strategic [583]. Strategy [1721, 2]. Straton [255]. Strauss [790].

Stresemann [1180]. Strife [1216]. Strube [1590]. Structural [1604]. Structuralist [151]. Structure [1601, 1273, 1321, 504, 1593, 1101, 167, 1046, 1369, 1639, 1022, 952, 1536]. Struggle [392]. Struik [248]. Strutt [160, 160]. Struve [326]. Stuart [919]. Student [320]. Students [303]. Studi [1497, 409, 1188]. Studie [1655, 1489]. Studien [1824, 932, 1111, 1820, 341, 621, 571, 1097, 1182]. Studier [1356]. Studies [264, 986, 59, 1356, 1556, 121, 1320, 37, 1730, 1467, 1101, 311, 1104, 1090, 1744, 1453, 879, 1540, 620, 972, 1588, 907, 1005, 32, 471, 1822, 299, 1249, 1346, 36, 1138, 1243]. Studium [1250]. Study [956, 692, 355, 1793, 1679, 219, 265, 1047, 1874, 711, 241, 908, 392, 152, 1334, 306, 1178, 1176, 704, 501, 1579, 1772, 1379, 963, 1319, 1285, 462, 1012, 577, 803, 1294, 1505, 837, 866, 1762, 1432, 1620, 945, 1481, 1149, 340]. Stuewer [1665]. Stuttgart [400]. Styazhkin [181]. Style [28, 1178]. Styles [1719]. Suan [1762, 98, 396]. Subba [365]. Subject [187]. Subjectivity [935]. Subjects [428, 1689]. Sublime [1479]. Substance [1601]. Suchten [1875]. Sugar [254]. sui [842]. sujet [1891]. sul [1497]. sulla [865, 1188]. Sullivan [977]. Sulowski [747]. Suma [1407]. Summa [1407]. Sundials [1700, 180]. Sunspots [114]. Superbomb [1266]. Superficierum [595, 291]. Supernova [1327]. Supernovae [1595]. Suppe [1046]. Supper [1587]. Suppes [829, 1530]. Supplement [712, 944, 1825]. Supported [815]. Supporting [1742]. Surgeon [1712, 635]. Surgeons [1349]. Surgery [1768, 153, 1298, 520]. Surgical [635]. Survey [883, 625, 20, 609, 1483, 516, 657, 1171, 1245, 620, 64, 1573, 240]. Surveyor [1088]. Surveyors [706]. Survivals [1826]. Susan [1876]. Susanne [1640]. Susskind [1789]. Sutton [513]. Svedberg [1329, 1077]. Svedenstierna [1659, 1659]. Sven [671]. Sven-Eric [671]. Svensk [1283]. Sverige [1146]. sviluppo [1202]. Swain [1284, 1891]. Swazey [881]. Swedish [1815, 714, 348]. Swenson [632]. Swiagenin [1408]. Swieten [891]. Swift [536]. Switzerland [127]. Sykes [950]. Sylvia [1664, 387]. Symbol [1244, 1242]. Symbolism [1888]. Symbols [60]. Symmer [1032]. Symposium [1837, 1057, 312, 1375, 322, 952, 466]. Symptoms [1761]. Synposis [337]. Synthetic [1483]. Syntony [1404]. System [1656, 1393, 203, 957, 30, 1052, 268, 318, 196, 1041, 836, 50, 549, 1524, 21, 325]. Systematic [290]. Systematik [1831]. Systematiker [67]. Systems [1277]. Szabó [1847]. Szilard [812, 812]. Szymanowitz [584].

T [446, 1833, 426, 1494, 812, 304, 569, 447, 1511, 38, 996, 298, 1644, 221, 939, 897, 1659, 552, 1622, 1783, 1285, 195, 1584, 930, 1539, 716, 1148, 1363, 305, 988, 576, 1450, 75, 55]. T. [1856, 576]. Tūsī [234, 732]. Tabarroni [842]. Tabit [1839]. Tables [145, 959, 1350, 207, 1058, 749, 1484, 591]. Tabulae [522]. Taccola [710]. Taconic [1515]. Tactless [1358]. Taddeo [1213]. Tafeln [522]. Tahdid [1186]. Taisbak [800]. Taiwan [1291]. talets [890]. Tallmadge [257]. Talmud [1219]. Talwar [1797]. T’ang [1840]. Tangent [207]. Tannery [1058]. Tao [1336]. Taqi [1479]. Taqi-al-Din [1479]. Tarr [1645]. Tartaglia [1410]. Taste [308]. Tauer [1094]. Taviss [881]. Taxidermy [1381]. taxonomie [252]. Taylor [1925]. Teach [1496]. Teaches [66]. Teaching [652, 898, 466]. Technical [904, 1818, 245, 1813]. Technik [98, 1698, 1918]. Technikgeschichte [1545]. Technique [1661]. Techniques [1406, 223, 1525, 58, 1452, 419, 1053, 1832]. technischen [63]. Technischer [1330]. Technological [1687, 1922, 1101, 56]. Technology [1310, 1874, 1034, 1742, 1901, 1452, 306, 400, 1429, 704, 882, 1467, 617, 1483, 354, 1101, 311, 1594, 1644, 552, 1605, 97, 1645, 1699, 216, 215, 1331, 69, 1089, 1091, 56, 460, 431, 1764, 306, 91, 355, 1830, 1834, 603, 1754, 1206, 1686, 931, 1241, 420, 618, 614, 1734, 797, 930, 658, 698]. técnica [1045, 551]. Tectonics [1800]. Ted [415]. Teich [1135]. Teil [1187, 1617, 1548, 1558]. Teilhard [1433, 109]. tekhnicheskaia [1881]. tekhnicheskikh [1813]. Telescope [955, 686, 104, 1884]. Telle [1899]. Teller [1266, 1372]. Telliamed [34]. Telling [1281]. Temkin [1543, 469, 1049]. Temples [606]. Temporal [1445]. Temps [5, 1731]. Ten [399]. Tension [1730]. tentative [1615]. Tentoonstelling [263]. teoria [1890, 1782]. Terlecka [1236]. Terminology [31]. Terrell [1115]. Terrestrial [271]. Territory [1048]. Terry [943]. Tertia [1490]. Teske [1461]. Test [634, 593]. Testament [1050]. Testimonien [1548]. Tetty [1262]. Tetu [1324]. Texas [105, 1375, 1363]. Text [1298, 425]. Textbook [396, 1309]. Textes [550]. Textile [155]. Texts [888, 499, 305, 1538, 801]. Teyler [939]. Thackray [1817, 272, 893]. Thaddeus [1264]. Theaetetus [1628]. Their [1582, 1309, 869, 187, 609, 742, 19, 1606, 1639, 308, 633, 356]. Themes [1258]. Themon [885]. Theodora [327]. Theodore [265, 243, 703, 1614, 1349]. Théodoridè [809]. Théodoridès [133]. Theodosius [1138]. Theologian [455]. Theologians [1191]. Theologie [752, 752]. Theology [640, 401]. Théophraste [757]. Theophrastus [1474, 423]. Theorem [1229]. Theoretical [468, 158, 81, 559, 1198]. Theorica [1270]. Theoricae [510]. Theorie [157, 1615]. Theorien [273]. Theories [1323, 345, 514, 611, 557, 1507, 914, 927, 1046, 579, 434, 1460]. Theory [301, 290, 1318, 1273, 1638, 1793, 641, 1189, 590, 1854, 588, 9, 237, 1858, 1321, 503, 1810, 1637, 854, 287, 272, 1387, 1806, 492, 530, 334, 1033, 1812, 541, 441, 1181, 233, 1691, 1104, 1123, 1319, 1441, 575, 1338, 1906, 1344, 1217, 1790, 958, 204, 1576, 415, 1498, 1294, 913, 1632, 337, 1115, 1177, 1264, 558, 1394, 662, 510, 1115, 771, 1540]. Therapie [1182, 218]. thereto [756]. Thermodynamics [1858, 439]. Thesis [1387, 1507]. Theta [905]. Thies [259]. Thing [1879]. Things [553, 1127]. Thinkers [1302]. Thinking [1101, 651]. Third [688, 1685, 1669, 160]. Thirteenth [1107, 1762, 453, 1272, 396]. Thirteenth-Century [1762, 453]. Thistle [1314]. Thomas [712, 368, 1554, 1833, 525, 877, 1019, 1503, 1153, 310, 524, 441, 1730, 1305, 896, 1024, 101, 894, 1109, 1711, 189, 628, 528, 1899, 337, 1083, 1697, 1664, 1144, 1305, 438]. Thompson [1239, 223, 1311]. Thomson [1509, 1759, 717]. Thorkild [980]. Thorndike [131]. Thornton [258, 462]. Thought [701, 986, 1538, 147, 29, 1313, 1465, 1568, 267, 699, 1614, 793, 577, 991]. Thoughts [337]. Thousands [364, 491]. Three [1753, 1281, 367]. Thrower [1911, 655]. Throwing [833]. ti [1406]. Tibetan [801]. Tibor [1308]. tid [1146]. tiempo [798]. till [890]. Tilling [1415]. Tilman [1816]. Time [774, 923, 1328, 647, 1529, 303, 1639, 1163, 320, 969, 1326, 464, 1021, 1285, 122, 1398, 1104]. Times [1372, 1286, 1343, 1829, 1229, 701, 408]. Timofeevich [55]. Timothy [986, 1508]. Timpanaro [519]. Tinkers [1302]. Tissot [1909]. Tobey [490, 1527]. Tobias [938]. Today [322]. Toellner [806]. together [337]. Toledan [591]. Tome [695, 550, 1233, 432]. tomes [1058]. Tomkinson [1732]. Tommaso [1843]. Toms [756]. Tongues [162]. Tons [608]. Toomer [510, 1550]. Topics [303]. Topsell [894]. Tord [193]. Tore [890]. Tornado [817]. Torre [1690]. Totok [1414]. Touch [308]. Toulmin [734]. Tour [1659]. Tourmaline [1033]. Tours [695]. Town [99]. toxicologiques [920]. Trabulse [1613]. Traces [368]. Trackers [323]. Tract [616, 12]. Tractatus [472, 708]. Trade [462]. tradendarumque [750]. Tradition [264, 165, 1730, 1701, 764]. Traditional [1538, 1874, 1452, 1272]. traditionellen [1291]. Traditions [856, 1243, 565, 1547, 1092]. Traek [1741]. Trainor [347]. Traité [232, 1109]. Traitement [1615]. Traités [1552]. Trajectory [963]. Traktat [1403, 1899]. tramonto [665]. Transcribed [473]. Transformation [480]. Transformed [1926]. transfusion [1360]. Translated [473]. Translation [169, 1478, 1298, 1550, 1648, 1448]. Translations [1898, 243]. Transmutation [1719, 578]. Trattato [839]. Traumdeuter [1652]. travail [1839]. Travel [1659, 94]. Travels [177]. Treatise [1679, 1186, 268, 1836, 1539, 960, 1649, 887, 1762, 524]. Treatises [124, 367]. Treatment [760, 1309]. Tree [703, 968]. Trends [1082]. Trenn [1264]. Trento [909]. Trevirensis [1875]. Trevor [445, 1089, 1070]. Triads [1269]. Trial [846]. Triangulation [1167]. Tricarico [1890]. Trigg [1265, 329]. Trigon [46]. Trigonometric [960]. Triplets [1269]. Triumph [38, 1860]. Trochet [1262]. Trogir [124]. Tropical [1750]. Troy [883]. Truesdell [1858]. Truth [925]. Tsarist [1313]. Tudományos [1457]. Tugnoli [1555]. Tully [462]. Tunnel [168]. Tunzelmann [1833]. Turco [1497]. Turnbull [1822, 1529]. Turner [939, 1370, 948, 1024, 1287]. Turning [1665, 704, 1751]. Twelfth [473, 837, 1317]. Twelfth-Century [473, 1317]. Twentieth [1265, 1373, 152, 306, 678, 1085, 1621, 1238, 1378]. Twirled [815]. Two [775, 1032, 12, 1154, 1166, 1382, 1002]. Tycho [1679, 262]. Tyndall [1116]. Types [344]. Typologie [562].

U.R.S.S. [724]. U.S [64]. U.S.S.R. [460]. Überlieferung [358]. Übersetzung [1403]. uchenym [1783]. udviklingshistorie [1741]. uell [1355]. uhrung [1345]. Uitrusting [244]. Ulisse [1555, 1555]. Ullmann [1187, 1482]. Ulrich [1291, 1107]. umyslowej [342]. Unabridged [1106]. Unbound [56]. Uncertainty [1573]. Unconscious [242]. underlying [716]. Understanding [762, 712, 734]. Underwood [1779]. ungedruckte [572]. Unguru [1760]. Uniform [286]. Uniformly [286]. Union [826, 552, 580, 1227]. United [107, 1334, 1405, 749, 1735, 205, 1172]. Universalgelehrten [806]. universalis [366]. Universe [411, 300, 337, 337, 1904, 1042]. Universitat [1795, 1017]. Universitäten [418]. universitetsfilosofi [890]. University [890, 1728, 1227, 1732, 322, 864, 716, 943, 1059, 1331, 399, 952, 400, 167, 987]. Unnoticed [1385]. Unpublished [587, 938, 336, 1059, 1382]. Unsaturation [120]. Unschuld [1291]. unserer [928]. Unspecialized [1003]. Untersuchungen [563]. until [1343]. Unto [1288]. upon [1858, 1186, 337]. Uppsala [890, 1732]. Uqlidisi [1252, 1903, 1252, 1903]. Urbino [781]. Urdang [1541]. Urgeschichte [759]. Usage [1106]. Use [734, 1810, 1522]. Usefulness [257]. Uses [125, 1818, 1219, 608]. Ussery [659]. usu [257]. Uta [942]. utiles [1770].

V [1713, 1310, 64, 1170, 1415, 1800, 319, 365, 1883, 1881, 1483, 223, 1763, 1420, 1846, 34, 36, 271, 372, 1231, 1256, 98, 1783, 1813]. Vaccine [1696, 846]. Vagtborg [951]. Vaillancourt [305]. Val [1474]. Valencia [741]. Valency [459]. Valentin [1191, 840]. valeur [432]. Valle [254]. Valleys [883]. Vallicrosa [1275]. Values [1326, 216]. Vandermaelen [720]. Vanderpool [1090]. Vandervliet [674]. Vannevar [721]. Van’t [1261]. Varcoe [1072]. Variability [531]. Variants [1038]. Varieties [3, 1814]. Varignana [1213]. Various [428, 1496]. Vascular [836]. Vasoli [1188]. Vater [324]. vazhneishikh [1883]. Vector [21]. Vectorial [21]. Veda [1882]. Vegetti [745]. Vekenstyl [1484]. vel [472]. Velocities [202]. ventis [1474]. verantwortungsbewussten [621]. verdensbillede [1741]. Vereinigten [1066]. Verformungstendenzen [358]. Vergangenheit [1831]. Verhaltensforschung [324]. Verhandeling [362]. vérité [1770]. verk [262]. Vermeer [1899]. Vermischte [1708]. Vermorel [561]. Vern [178]. Vernon [1260]. Versammlungen [947]. Version [1081, 1836, 1059, 1836]. versus [1271, 474]. vertes [1537]. Vess [1195]. VI [372, 1148, 1493]. Viator. [972]. Vicinity [883]. Victor [274, 423, 1422, 678, 1474, 75, 1384, 1634]. Victorian [1562, 1876, 1152, 486, 1423, 1178, 899, 948, 1920, 1791, 1865, 1580, 306, 478]. vie [1930, 252, 1488, 448]. Vienna [1139]. Viertel [1408]. Viète [1841]. View [1062, 423, 113, 289]. Viewed [187]. Views [857, 1629]. VII [1845, 372, 1648]. VII-XII [1648]. VIII [1654]. Villanova [1348]. Vincenti [1922]. Vincenzo [1325]. vingt [1233]. Vinton [1740]. Virchow [380]. Virginia [253, 371]. Virology [1374]. Virus [1828]. Vis [172, 142]. Visible [337]. Visibles [1144]. Vision [233, 1460, 308, 1192]. Visual [1810, 611]. Vita [781, 1843, 1485, 1113]. Vital [1216, 1721]. Vite [1703, 1485]. Vitturi [124]. Viva [142, 172]. vivant [560]. Vivian [609]. Vladimirovich [1535]. Vogel [932, 883, 98, 1900, 1293]. Vogelfang [1293]. Voget [1464]. Void [1840]. Voisé [1525]. voix [1108]. Volker [522, 1112]. Volume [225, 1490, 136, 280, 494, 680, 849, 1026, 1208, 1437, 1667, 1932, 107, 1571, 1927, 1203, 734, 93, 1170, 1415, 471, 1898, 518, 425, 812, 1473, 1610, 1742, 126, 306, 985, 158, 1689, 1145, 1654, 1299, 1467, 617, 1483, 1698, 939, 223, 559, 668, 1058, 1776, 1763, 1346, 1605, 1348, 303, 127, 749, 1352, 941, 372, 1148, 919, 1527, 886, 1493, 972, 696, 476, 177, 944]. Volumen [1551]. Volumes [598, 1281, 820, 437]. vom [1918]. vor [1473]. Vorgeschichte [1684, 273]. vort [1741]. Vortex [815]. Vortex-Supported [815]. Vorträge [1332, 1918]. Vorzimmer [381]. Vossiani [1546]. vovich [1921]. Voyage [808]. Voyages [1232]. Vozniknovenie [1256]. vremeni [923]. Vrij [937]. vs [503]. Vucinich [1313]. vus [1731]. vv [1813].

W [888, 1824, 1524, 767, 1687, 463, 1398, 1427, 1748, 1431, 413, 1170, 355, 1147, 1152, 934, 1048, 444, 1159, 1835, 1111, 903, 904, 154, 1886, 467, 530, 898, 1592, 162, 1165, 191, 1528, 1911, 91, 1117, 1061, 1743, 668, 369, 655, 873, 68, 1525, 941, 706, 765, 328, 1095, 180, 342, 1564]. wa [1761]. wa-al-jawab [1761]. Waard [1058, 1058]. Wäckerlin [1408]. Waddington [1434]. Waerden [1137]. Wafā’ [778]. Wafa [521]. Wagley [535]. Wagner [40]. Wallace [387, 728, 1342, 696, 672]. Waller [317]. Wallingford [1478]. Wallis [571, 1912, 1912]. Wallwork [845]. Walser [357]. Walsh [878, 1466]. Walter [1490, 105, 1205, 382, 1423, 1258, 1922, 376, 1011, 928, 834, 1579, 397]. Walzer [1895]. Wandel [1096]. Wandruszka [891]. Wang [617]. Wansey [266]. Wanted [1466]. War [635, 540, 144, 1286]. Ward [254, 624]. Wardell [141]. Warnow [537, 118]. Wartime [1314]. Wartofsky [1528]. Warwick [1191]. Was [1222, 1132]. Washington [1421]. Wasserfuhr [1780]. Watches [1752]. Water [1164, 1754, 980, 979, 1681]. Water-Lifting [980]. Water-Spouting [1681]. Water-Works [979]. Waterston [9, 237]. Watson [1746, 1070, 1695]. Watt [439, 475]. Way [872]. Wayne [826, 711, 1462]. Ways [42]. Weapons [162]. Weart [1560]. Weather [250]. Weber [395, 574]. Webster [624, 1554, 224, 1361, 1114]. Webster-Ward [624]. Wechselbeziehungen [170]. Wechsler [1821]. Wednesday [1587]. Weg [483, 1799]. Wege [1017, 1795, 947]. Weights [1531, 235]. Weil [1057]. Weimar [540]. Weiner [537]. Weingart [1819]. Weinhandl [621]. Weismann [112]. Weiss [812]. Weisskopf [75, 678]. Welfare [757]. Well [66]. Wellcome [27, 864, 657]. Wellens [720]. Wellens-de [720]. Wellesley [1423]. Wells [861, 543]. Welt [1459]. Weltbild [256]. Weltgeltung [1017]. Werk [1853, 1785, 1799, 274, 376]. Werke [522, 1612, 511]. Werner [1892, 1714, 276, 70, 1572, 70, 504, 367, 70, 1508]. Wernerian [1854]. Wernher [94]. Wernicke [81]. Wesley [807, 789, 73]. West [627, 375, 589, 1179, 618, 1467]. Western [1092, 514, 807, 704, 1304, 1836, 485, 1526, 864, 69, 736, 56, 421, 90]. Westfall [338, 626]. Westman [1781]. Weyl [194]. Wheatstone [1364]. Wheeler [384]. Wheels [980]. Whewell [575]. Which [1793, 661, 66]. White [630, 1467, 1305, 742, 736, 1901, 829, 1743, 972, 1505]. Whitehead [1797]. Whiteside [221, 1148, 716, 1450]. Whitley [1136]. Whitmore [1305]. Whitney [416]. Whitrow [922, 1689]. Whitteridge [666]. Who [1205, 1222, 1738, 1055]. Wholes [1672]. Whyte [594]. wichtigsten [1847]. Wiege [932]. wieku [342, 1564]. Wiener [1453]. Wieslaw [1564]. Wigner [771]. Wikman [319]. Wilbert [698]. Wilbur [1294]. Wild [1738, 155]. Wilder [502]. Wilfrid [426]. Wilhelm [1330, 1477, 262, 769, 1684, 85, 1414, 1842, 1844, 531, 765, 841, 1849, 195]. Wilhelminischen [1684]. Wilkins [1196, 102]. Willard [1204, 867, 412]. Willem [196, 937]. Willey [1009, 1435]. William [805, 1927, 1062, 379, 880, 468, 426, 1494, 1180, 478, 1620, 1898, 756, 1194, 1610, 1257, 673, 37, 742, 1183, 480, 1342, 1829, 979, 347, 1229, 528, 160, 762, 658, 803, 39, 1362, 1366, 696, 1573, 1035, 1078, 1269, 959, 543, 1850, 384, 666, 575]. Williams [1818, 1859, 1362]. Williamsburg [253]. Willy [1824, 548, 403]. Wilson [301, 1428, 602, 1318, 1848, 725, 161, 1691, 269, 491]. Winau [928, 1918]. Winkelmann [928]. Winslow [486]. Winsor [1371]. Winter [693, 1558]. Winterl [50]. Wirkung [1414]. Wismann [746]. Wissenschaft [483, 1929, 1655, 341, 707]. Wissenschaften [1430, 1714, 1617, 1558, 1684, 750, 1929]. Wissenschaftliche [1558]. wissenschaftlichen [572, 1795, 1816]. wissenschaftlicher [1563]. Wissenschaftsentwicklung [1640]. Wissenschaftsforschung [1640]. wissenschaftsgeschichtliche [1489]. Wissenschaftsorganisatorische [1558]. Wissenschaftspolitik [1068, 1684]. Wissenschaftssoziologie. [1820]. wissenschaftssoziologische [1563]. Wissenschaftssteuerung [1640]. Witelo [1760, 453]. Witelonis [1760]. within [856, 716]. Wits [1181]. Wittenberg [1486, 913]. Wizards [1242]. Wlassics [1308]. Wolf [965, 1816, 1197, 905]. Wolfe [1798]. Wolff [899]. Wolffianism [890]. Wolffianismens [890]. Wolfgang [753, 492, 386, 633, 1816]. Wollheim [906]. Wolstenholme [1544]. Wolters [1755]. Woman [761, 1043]. Women [1466, 1749, 1444, 742, 1749]. Wood [125]. Woodcock [382]. Woodcroft [307]. Woodside [1749]. Woolen [66]. Woon [1452]. Word [1254]. Words [60]. Work [1554, 1152, 430, 238, 625, 126, 985, 945, 533, 81, 896, 939, 897, 1193, 194, 127, 1787, 1409, 96, 1519, 47, 1493, 937, 40, 1159, 321, 1726]. Works [1157, 1588, 812, 1561, 317, 754, 223, 979, 886, 900, 440, 648]. Worksheet [1723]. World [635, 1062, 1196, 251, 725, 1053, 1060, 1818, 1741, 1916, 197, 1277, 1010, 337, 540, 144, 614, 1205, 1542, 1757, 1775, 422, 268, 1051]. World-Edifice [1775]. Worlds [775, 188]. Worster [1694]. Worterbuch [1590]. Worth [1375]. Wortham [470]. Wound [1542]. Wright [537, 337, 1289]. Writers [542]. Writing [95]. Writings [1276, 375, 319, 152, 938, 703, 1478, 303, 1368, 1198, 576]. Written [1886, 864, 878]. Wrobel [1664]. Wunderlich [1842, 1842, 1799]. Wussing [106]. Wybourne [952]. Wyndham [1341]. Wynter [1287].

X [907, 522, 1654, 1058, 1900, 941, 765, 1813, 1081]. X-Ray [907]. X-Rays [941]. X-XVII [1813]. XI [1145]. XII [1058, 409, 1648]. XIIe [550]. XIII [1551, 1776]. XIV [179]. XIV-XVI [179]. XIVe [885]. XIX [1564, 1707]. XIXe [1151, 1616]. XVI [1551, 366, 409, 179]. XVIe [886]. XVII [1256, 1813, 1613]. XVIII [1256, 342]. XVIIIe [156]. XX [1490, 286]. XXII [587]. XXIV [1080].

Yabuuchi [1353]. Yabuuti [507]. Yagi [1284]. Yang [1762]. Yates [1016]. Yazid [1050]. Year [1326, 1573]. Years [536, 1069, 1393, 275, 304, 1742, 381, 1500, 26, 1855, 602, 399, 491, 919, 1203, 196, 1455, 1041, 693]. Yehuda [1173, 1817, 1064]. Yong [1762]. Yonge [1232]. York [883, 668, 1266]. Yorkshire [946]. Young [1793, 347, 421, 487, 1135, 1501]. Younger [1905]. Yourgrau [492, 633]. Yu [195, 1483]. yung [396]. Yutang [1106]. Yvette [1151, 1616].

Z [440, 1589, 1643, 702, 90, 660, 838]. Zaban [669]. Zacaire [1875]. Zacher [1061]. Zahlbegriffs [217]. Zahlensystems [1061]. Zahnarzt [1780]. Zambelli [895, 865, 1647]. Zamponi [1647]. Zane [375]. Zaslavsky [871]. Zassenhaus [1067]. Zauberpapyri [256]. Zaunick [1875]. Zdenek [245, 1095]. Zealand [952]. Zeit [1617, 928, 891, 1572]. Zeitalter [74]. Zeiten [1096]. Zeitmessers [1698]. Zentrums [1017]. Zhores [55]. Ziegler [1801]. Zier [1590]. Zierold [43]. Ziggelaar [573]. Zimmerman [1073]. Zimmermann [1892]. zodiacal [1547]. Zoltán [1826]. Zoological [1693, 1826, 1693]. zoologische [517]. Zoology [1154]. Zsigmond [1457]. Zuckerman [1817, 1735]. zum [218, 1892, 759, 1098, 1795, 928]. Zur [255, 752, 1698, 273, 1640, 1684, 1111, 1096, 1929, 1112, 1655, 571, 1061, 1097, 1617, 1182, 1548, 341, 1332, 63, 1480, 1755, 562, 1068]. Zusne [1237]. Zweite [1918]. zweiten [1408]. zwischen [1657, 1792, 170]. zyciorys [838]. Zygmunt [886].

References

Anonymous:1970:FMa

[1] Anonymous. Front matter. Isis, 61(1):1–3, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/229145.

Rudwick:1970:SLP

[2] Martin J. S. Rudwick. The strategy of Lyell’s principles of geology. Isis, 61(1):4–33, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

      6994 (electronic). URL http://www.jstor.org/stable/229146.

Spencer:1970:VNC

[3] J. Brookes Spencer. On the varieties of Nineteenth-Century magnetooptical discovery. Isis, 61(1):34–51, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

      stable/229147.

Fletcher:1970:ALC

[4] John E. Fletcher. Astronomy in the life and correspondence of athanasius Kircher. Isis, 61(1):52–67, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229148.

Lyon:1970:SFM

[5] John Lyon. The search for fossil man: Cinq personnages à la recherche du temps perdu. Isis, 61(1):68–84, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229149.

Carozzi:1970:RHR

[6] Albert V. Carozzi. Robert hooke, rudolf erich raspe, and the concept of “earthquakes”. Isis, 61(1):85–91, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229150.

Yong:1970:GBA

[7] Lam Lay Yong. The geometrical basis of the Ancient Chinese squareroot method. Isis, 61(1):92–102, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229151.

Kuhn:1970:CAI

[8] Dorothea Kuhn. Das Cotta-Archiv im Schiller-Nationalmuseum zu Marbach am Neckar. Korrespondenz mit deutschen Naturforschern im 19. Jahrhundert. (German) [The Cotta Archive in the National Schiller Museum in Marbach am Neckar. Correspondence with German naturalists in the 19th Century]. Isis, 61(1):103–104, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229152.

Daub:1970:WRC

[9] Edward E. Daub. Waterston, rankine, and Clausius on the kinetic theory of gases. Isis, 61(1):105–106, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229153.

White:1970:HGG

[10] Colin White and Robert J. Hardy. Huygens’ graph of Graunt’s data. Isis, 61(1):107–108, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229154.

Sciences:1970:APA

[11] History of the Conference Board of the Mathematical Sciences. An appeal for preservation of archival materials. Isis, 61(1):108–110, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229155.

Jensen:1970:ITA

[12] Claus Jensen. Identification of a tract in an Arabic manuscript: Eratosthenes on two mean proportionals. Isis, 61(1):111, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229156.

Theisen:1970:MMM

[13] Wilfred R. Theisen. The monastic manuscript microfilm library, collegeville, Minnesota. Isis, 61(1):111–112, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229157.

Anonymous:1970:Na

[14] Anonymous. News. Isis, 61(1):113–114, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229158.

Brush:1970:BRB

[15] Stephen G. Brush. Book review: Essays in the History of Mechanics. Isis, 61(1):115–118, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229159.

Wolf:1970:BRB

[16] Robert G. Wolf. Book review: Science: Men, Methods, Goals by Boruch Brody; Nicholas Capaldi. Isis, 61(1):118–119, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229160.

Cannon:1970:BRBa

[17] Walter F. Cannon. Book review: The Cambridge Philosophical Society: A History 1819–1969 by A. Rupert Hall. Isis, 61(1):119, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229161.

Dexter:1970:BRB

[18] Ralph W. Dexter. Book review: Bankers, Bones, and Beetles. The First Century of the America Museum of Natural History by Geoffrey Hellman. Isis, 61(1):119–120, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229162.

Middleton:1970:BRB

[19] W. E. Knowles Middleton. Book review: English Barometers 1680– 1860. A History of Domestic Barometers and Their Makers by Nicholas Goodison. Isis, 61(1):120–121, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229163.

Jones:1970:BRBa

[20] William Powell Jones. Book review: The High Firmament: A Survey of Astronomy in English Literature by A. J. Meadows. Isis, 61(1):121– 122, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229164.

Koppelman:1970:BRB

[21] Elaine H. Koppelman. Book review: A History of Vector Analysis. The Evolution of the Idea of a Vectorial System by Michael J. Crowe. Isis, 61(1):122–124, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229165.

Saidan:1970:BRB

[22] A. S. Saidan. Book review: The Concept of Number by Christoph J. Scriba. Isis, 61(1):124, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229166.

Thoren:1970:BRBa

[23] Victor E. Thoren. Book review: The Bowl of Night by F. P. Dickson. Isis, 61(1):125, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229167.

Kauffman:1970:BRB

[24] George B. Kauffman. Book review: The Evolution of Chemistry: A History of Its Ideas, Methods, and Materials by Eduard Farber. Isis, 61(1):126–127, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229168.

Cannon:1970:BRBb

[25] Walter F. Cannon. Book review: The Earth in Decay. A History of British Geomorphology, 1578–1878 by Gordon L. Davies. Isis, 61(1): 127–129, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229169.

Mark:1970:BRB

[26] Joan T. Mark. Book review: One Hundred Years of Anthropology by J. O. Brew. Isis, 61(1):129–130, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229170.

Hamarneh:1970:BRBa

[27] Sami Hamarneh. Book review: Medical Ceramics in the Wellcome Institute of the History of Medicine by John K. Crellin. Isis, 61(1):130–131, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229171.

Eaton:1970:BRB

[28] Leonard K. Eaton. Book review: Principles of Architectural History: The Four Phases of Architectural Style, 1420–1900 by Paul Frankl; James F. O’Gorman. Isis, 61(1):131, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229172.

Scriba:1970:BRB

[29] C. J. Scriba. Book review: Greek Mathematical Thought and the Origin of Algebra by Jacob Klein. Isis, 61(1):132–133, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229173.

Dyason:1970:BRB

[30] Diana Dyason. Book review: Galen’s System of Physiology and Medicine by Rudolf E. Siegel. Isis, 61(1):133–134, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229174.

Levey:1970:BRB

[31] Martin Levey. Book review: Mathematical Terminology in Hebrew Scientific Literature of the Middle Ages by Gad B. Sarfatti. Isis, 61(1): 135–136, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229175.

Lo:1970:BRB

[32] Karl Lo. Book review: Chinese Alchemy: Preliminary Studies by Nathan Sivin. Isis, 61(1):136–137, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229176.

Rosen:1970:BRB

[33] Edward Rosen. Book review: Georg Joachim Rhetikus, 1514–1574. Eine Bio-Bibliographie. Band III: Briefwechsel by Karl Heinz Burmeister. Isis, 61(1):137–139, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229177.

Ospovat:1970:BRB

[34] Alexander M. Ospovat. Book review: Telliamed or Conversations between an Indian Philosopher and a French Missionary on the Diminution of the Sea by Benoît de Maillet; Albert V. Carozzi. Isis, 61(1):139–141, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229178.

Kargon:1970:BRB

[35] Robert H. Kargon. Book review: A Portrait of Isaac Newton by Frank E. Manuel. Isis, 61(1):141–142, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229179.

Page:1970:BRB

[36] Leroy E. Page. Book review: Studies on Glaciers. Preceded by the Discourse of Neuchâtel by Louis Agassiz; Albert V. Carozzi. Isis, 61(1): 142–143, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229180.

Koehn:1970:BRB

[37] Donald R. Koehn. Book review: The Origins of Pragmatism: Studies in the Philosophy of Charles Sanders Peirce and William James by A. J. Ayer. Isis, 61(1):143–144, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229181.

Lehman:1970:BRB

[38] Hugh Lehman. Book review: The Triumph of the Darwinian Method by Michael T. Ghiselin. Isis, 61(1):144–145, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229182.

Uschmann:1970:BRB

[39] G. R. Uschmann. Book review: The Interpretation of Animal Form by William Coleman. Isis, 61(1):145–146, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229183.

Becker:1970:BRB

[40] Stanley L. Becker. Book review: Clemens von Pirquet. His Life and Work by Richard Wagner. Isis, 61(1):147, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229184.

Reingold:1970:BRBa

[41] Nathan Reingold. Book review: The Intellectual Migration, Europe and America, 1930–1960 by Donald Fleming; Bernard Bailyn. Isis, 61(1): 147–148, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229185.

Seeley:1970:BRB

[42] Daniel H. Seeley. Book review: Through Rugged Ways to the Stars by Harlow S. Shapley. Isis, 61(1):148–149, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229186.

Forman:1970:BRBb

[43] Paul Forman. Book review: Forschungsförderung in drei Epochen. Deutsche Forschungsgemeinschaft. Geschichte. Arbeitsweise. Kommentar by Kurt Zierold. Isis, 61(1):149–151, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229187.

Anonymous:1970:BMa

[44] Anonymous. Back matter. Isis, 61(1):152–154, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229188.

Anonymous:1970:FMb

[45] Anonymous. Front matter. Isis, 61(2):155–157, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229972.

Aston:1970:FTC

[46] Margaret Aston. The Fiery Trigon Conjunction: An Elizabethan astrological prediction. Isis, 61(2):158–187, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229973.

Stuewer:1970:CAN

[47] Roger H. Stuewer. A critical analysis of Newton’s work on diffraction. Isis, 61(2):188–205, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229974.

OBrien:1970:ECD

[48] Charles F. O’Brien. Eozoön canadense “the dawn animal of canada”. Isis, 61(2):206–223, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229975.

Rabinovitch:1970:RHC

[49] Nachum L. Rabinovitch. Rabbi Hasdai Crescas (1340–1410) on numerical infinities. Isis, 61(2):224–230, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229976.

Snelders:1970:IDS

[50] H. A. M. Snelders. The influence of the dualistic system of Jakob Joseph Winterl (1732–1809) on the German Romantic Era. Isis, 61(2):231–240, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229977.

Kauffman:1970:AM

[51] George B. Kauffman. Arturo Miolati (1869–1956). Isis, 61(2):241–253, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229978.

Jaffe:1970:MP

[52] Bernard Jaffe and Amabel N. Sollas. The Moseley photograph. Isis, 61 (2):254, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229979.

vonKlobusitzky:1970:CHP

[53] Dionysio von Klobusitzky. Contribution to the history of paper electrophoresis. Isis, 61(2):255–257, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229980.

Anonymous:1970:Nb

[54] Anonymous. News. Isis, 61(2):258–260, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229981.

Graham:1970:BRB

[55] Loren R. Graham. Book review: The Rise and Fall of T, by Zhores Aleksandrovich Medvedev; I. Michael Lerner; Genetika i dialektika by Ivan Timofeevich Frolov. D. Lysenko. Isis, 61(2):261–263, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229982.

Thackray:1970:BRBa

[56] Arnold Thackray. Book review: The Unbound Prometheus. Technological Change and Industrial Development in Western Europe from 1750 to the Present by David S. Landes; Science and Technology in the Industrial Revolution by A. E. Musson; Eric Robinson. Isis, 61(2):263–264, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229983.

Glick:1970:BRBa

[57] Thomas F. Glick. Book review: La introducción de la ciencia moderna en España by J. M. López Piñero; Medicina, historia, sociedad. Antología de clásicos médicos by J. M. López Piñero. Isis, 61(2):265–266, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229984.

Neu:1970:BRB

[58] John Neu. Book review: Éléments de bibliographie de l’histoire des sciences et des techniques by François Russo. Isis, 61(2):266–267, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229985.

Colodny:1970:BRB

[59] Robert G. Colodny. Book review: The Legacy of Logical Positivism. Studies in the Philosophy of Science by Peter Achinstein; Stephen F. Barker. Isis, 61(2):267–268, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229986.

Seidenberg:1970:BRB

[60] A. Seidenberg. Book review: Number Words and Number Symbols. A Cultural History of Numbers by Karl Menninger; Paul Broneer. Isis, 61 (2):268–269, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229987.

Sjostedt:1970:BRB

[61] C. E. Sjöstedt. Book review: Le Axiome de Paralleles de Euclides a Hilbert by J. D. North. Isis, 61(2):269–270, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229988.

Harrington:1970:BRB

[62] John W. Harrington. Book review: Toward a History of Geology by Cecil J. Schneer. Isis, 61(2):270–272, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229989.

Bryant:1970:BRB

[63] Lynwood Bryant. Book review: Zur Frühgeschichte der technischen Regelungen by Otto Mayr. Isis, 61(2):272–274, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229990.

Brittain:1970:BRB

[64] James E. Brittain. Book review: The Beginnings of Satellite Communications by J. R. Pierce; Early History of the Electron Microscope by L. Marton; Benjamin Peirce and the U.S. Coast Survey by V. F. Lenzen. Isis, 61(2):274–275, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229991.

Drake:1970:BRB

[65] Stillman Drake. Book review: La philosophie naturelle de Galilée by Maurice Clavelin. Isis, 61(2):275–277, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229992.

Reti:1970:BRB

[66] Ladislao Reti. Book review: The Plictho. Instructions in the Art of the Dyers Which Teaches the Dyeing of Woolen Cloths, Linens, Cottons, and Silk by the Great Art as Well as by the Common by Gioanventura Rosetti; Sidney M. Edelstein; Hector C. Borghetty. Isis, 61(2):277–278, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229993.

Roger:1970:BRB

[67] Jacques Roger. Book review: Carl von Linné, Arzt, Naturforscher, Systematiker by Heinz Goerke. Isis, 61(2):278–280, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229994.

Rousseau:1970:BRB

[68] G. S. Rousseau. Book review: Reason and Nature in the Eighteenth Century by R. W. Harris. Isis, 61(2):280, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229995.

Sivin:1970:BRB

[69] N. Sivin. Book review: John Fryer. The Introduction of Western Science and Technology into Nineteenth-Century China by Adrian Arthur Bennett. Isis, 61(2):280–282, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229996.

Ihde:1970:BRB

[70] Aaron J. Ihde. Book review: Alfred Werner, Founder of Coordination Chemistry by George B. Kauffman; Werner Centennial by George B. Kauffman; Classics in Coordination Chemistry, Part 1: The Selected Papers of Alfred Werner by George B. Kauffman. Isis, 61(2):282–283, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229997.

Chambers:1970:BRB

[71] David Wade Chambers. Book review: For Science and National Glory. The Spanish Scientific Expedition to America, 1862–1866 by Robert Ryal Miller. Isis, 61(2):283–284, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229998.

Zirkle:1970:BRB

[72] Conway Zirkle. Book review: Darwin and the Beagle by Alan Moorehead. Isis, 61(2):284–285, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229999.

White:1970:BRBa

[73] George W. White. Book review: The Colorado River Region and John Wesley Powell by Mary C. Rabbitt; Edwin D. McKee; John Wesley Powell and the Anthropology of the Canyon Country by Don D. Fowler; Robert C. Euler; Catherine S. Fowler. Isis, 61(2):285–287, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230000.

Forman:1970:BRBc

[74] Paul Forman. Book review: Briefwechsel. Sechzig Briefe aus dem goldenen Zeitalter der modernen Physik by Albert Einstein; Arnold Sommerfeld; Armin Hermann. Isis, 61(2):287–288, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230001.

Wolfenstein:1970:BRB

[75] Lincoln Wolfenstein. Book review: Oppenheimer by Abraham Pais; Glenn T. Seaborg; Robert Serber; Victor F. Weisskopf; I. I. Rabi; In the Matter of J. Robert Oppenheimer by Heinar Kipphardt; John Roberts. Isis, 61(2):288–289, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230002.

Anonymous:1970:BMb

[76] Anonymous. Back matter. Isis, 61(2):290–292, Summer 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230003.

Anonymous:1970:FMc

[77] Anonymous. Front matter. Isis, 61(3):293–295, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229683.

Baird:1970:PIN

[78] A. W. S. Baird. Pascal’s idea of nature. Isis, 61(3):296–320, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229684.

Siraisi:1970:BPA

[79] Nancy G. Siraisi. The Expositio Problematum Aristotelis of Peter of Abano. Isis, 61(3):321–339, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229685.

Burke:1970:KCE

[80] John G. Burke. Kirk and causality in edinburgh, 1805. Isis, 61(3):340– 354, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229686.

Marx:1970:NCM

[81] Otto M. Marx. Nineteenth-Century medical psychology: Theoretical problems in the work of Griesinger, Meynert, and Wernicke. Isis, 61 (3):355–370, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229687.

Burke:1970:ECD

[82] John G. Burke. Éloge: Charles Donald O’Malley 1907–1970. Isis, 61 (3):371–378, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229688.

Farber:1970:ECI

[83] Eduard Farber. Errors in chemical identification: A precautionary note to the history of chemistry. Isis, 61(3):379–383, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229689.

Rosen:1970:REE

[84] Edward Rosen and Karl Heinz Burmeister. Rheticus’ earliest extant letter to Paul eber. Isis, 61(3):384–386, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229690.

Holt:1970:NWO

[85] Niles R. Holt. A note on wilhelm Ostwald’s energism. Isis, 61(3):386– 389, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229691.

Victor:1970:JMA

[86] Stephen Victor. Johannes de Muris’ autograph of the De arte mensurandi. Isis, 61(3):389–395, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229692.

Anonymous:1970:Nc

[87] Anonymous. News. Isis, 61(3):396, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229693.

Heilbron:1970:BRBa

[88] John L. Heilbron. Book review: Francis Line, S, by Conor Reilly.J. An Exiled English Scientist 1595–1675. Isis, 61(3):397–398, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229694.

Thackray:1970:BRBb

[89] Arnold Thackray. Book review: The Rise of Modern Science: External or Internal Factors? by George Basalla. Isis, 61(3):398–399, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229695.

Yen:1970:BRB

[90] Florette Yen. Book review: Western Medical Pioneers in Feudal Japan by John Z. Bowers. Isis, 61(3):399–400, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229696.

Korber:1970:BRB

[91] Hans-Günther Körber. Book review: Invention of the Meteorological Instruments by W. E. Knowles Middleton; Catalog of Meteorological Instruments in the Museum of History and Technology. Isis, 61(3):400– 401, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229697.

Hamarneh:1970:BRBb

[92] Sami Hamarneh. Book review: The Scientific Enterprise and Christian Faith by Malcolm A. Jeeves. Isis, 61(3):401–402, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229698.

Brozek:1970:BRB

[93] Josef Brožek. Book review: Experimental Psychology: Its Scope and Method. Volume I: History and Method by Jean Piaget; Paul Fraisse; Maurice Reuchlin; Judith Chambers. Isis, 61(3):402–404, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229699.

Swenson:1970:BRB

[94] Loyd S. Swenson, Jr. Book review: History of Rocketry and Space Travel by Wernher von Braun; Frederick I. Ordway. Isis, 61(3):404–405, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229700.

Gingerich:1970:BRB

[95] Owen Gingerich. Book review: The Codebreakers. The Story of Secret Writing by David Kahn. Isis, 61(3):405–406, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229701.

Snyder:1970:BRB

[96] Walter F. Snyder. Book review: The Muses at Work. Arts, Crafts, and Professions in Ancient Greece and Rome by Carl Roebuck. Isis, 61(3): 406–407, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229702.

Pingree:1970:BRB

[97] David Pingree. Book review: History of Science and Technology in India by O. P. Jaggi. Isis, 61(3):407–408, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229703.

Libbrecht:1970:BRB

[98] Ulrich Libbrecht. Book review: Chiu chang suan shu. Neun Bücher arithmetischer Technik. Ein chinesisch Rechenbuch für den praktischen Gebrauch aus der frühen Hanzeit (202 v. Chr. bis 9 n. Chr.) by Kurt Vogel. Isis, 61(3):408–409, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229704.

Jeffries:1970:BRB

[99] Theodore W. Jeffries. Book review: Dr. Alexander Garden of Charles Town by Edmund Berkeley; Dorothy Smith Berkeley. Isis, 61(3):409– 410, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229705.

Egerton:1970:BRB

[100] Frank N. Egerton, III. Book review: Demography in Early America. Beginnings of the Statistical Mind, 1600–1800 by James H. Cassedy. Isis, 61(3):410–411, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229706.

Rappaport:1970:BRB

[101] Rhoda Rappaport. Book review: Science in France in the Revolutionary Era Described by Thomas Bugge by Maurice P. Crosland. Isis, 61(3):412– 413, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229707.

Jones:1970:BRBb

[102] William Powell Jones. Book review: John Wilkins 1614–1672: An Intellectual Biography by Barbara J. Shapiro. Isis, 61(3):413–414, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229708.

Feather:1970:BRB

[103] Norman Feather. Book review: Rutherford and Boltwood: Letters on Radioactivity by Lawrence Badash. Isis, 61(3):414–415, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229709.

Weitzel:1970:BRB

[104] Ronald L. Weitzel. Book review: The Shadow of the Telescope: A Biography of John Herschel by Günther Buttmann. Isis, 61(3):415–417, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229710.

Aldrich:1970:BRB

[105] Michele L. Aldrich. Book review: Geology and Politics in Frontier Texas 1845–1909 by Walter Keene Ferguson. Isis, 61(3):417–418, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229711.

Hawkins:1970:BRB

[106] Thomas Hawkins. Book review: Die Genesis des abstrakten Gruppenbegriffes by Hans Wussing. Isis, 61(3):418–419, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229712.

Badash:1970:BRB

[107] Lawrence Badash. Book review: Atomic Shield, 1947/1952. A History of the United States Atomic Energy Commission. Volume II by Richard G. Hewlett; Francis Duncan. Isis, 61(3):419–420, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229713.

Penick:1970:BRB

[108] James Penick. Book review: Eugenics and the Progressives by Donald K. Pickens. Isis, 61(3):420–421, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229714.

Glick:1970:BRBb

[109] Thomas F. Glick. Book review: Teilhard de Chardin y la evolución biológica by Fernando Riaza. Isis, 61(3):421–422, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229715.

Anonymous:1970:BMc

[110] Anonymous. Back matter. Isis, 61(3):423–424, Autumn 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229716.

Anonymous:1970:FMd

[111] Anonymous. Front matter. Isis, 61(4):425–458, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229457.

Churchill:1970:HWM

[112] Frederick B. Churchill. Hertwig, Weismann, and the meaning of reduction division circa 1890. Isis, 61(4):428–457, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229458.

McCormmach:1970:HLE

[113] Russell McCormmach. H. a. Lorentz and the electromagnetic view of nature. Isis, 61(4):459–497, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229459.

Shea:1970:GSI

[114] William R. Shea. Galileo, scheiner, and the interpretation of sunspots. Isis, 61(4):498–519, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229460.

Worthen:1970:PAK

[115] Thomas D. Worthen. Pneumatic action in the klepsydra and Empedocles’ account of breathing. Isis, 61(4):520–530, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229461.

Anonymous:1970:Nd

[116] Anonymous. News. Isis, 61(4):531–532, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229462.

Brown:1970:BRB

[117] Theodore M. Brown. Book review: A History of the Bubonic Plague in the British Isles. Isis, 61(4):533–534, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229463.

Reingold:1970:BRBb

[118] Nathan Reingold. Book review: A Selection of Manuscript Collections at American Repositories. National Catalog of Sources for History of Physics, Report No. 1, Center for the History and Philosophy of Physics of the American Institute of Physics by Joan Nelson Warnow. Isis, 61(4): 535, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229464.

Forman:1970:BRBd

[119] Paul Forman. Book review: Atomic Order. An Introduction to the Philosophy of Microphysics by Enrico Cantore. Isis, 61(4):535–536, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229465.

Costa:1970:BRB

[120] Albert B. Costa. Book review: Unsaturation in Organic Chemistry by A. Albert Baker. Isis, 61(4):536–537, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229466.

Kangro:1970:BRB

[121] Hans Kangro. Book review: Historical Studies in the Physical Sciences by Russell McCormmach. Isis, 61(4):537–540, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229467.

White:1970:BRBb

[122] Lynn White, Jr. Book review: Hamlet’s Mill. An Essay on Myth and the Frame of Time by Giorgio de Santillana; Hertha von Dechend. Isis, 61(4):540–541, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229468.

Thoren:1970:BRBb

[123] Victor E. Thoren. Book review: Early Greek Astronomy to Aristotle by D. R. Dicks. Isis, 61(4):541–542, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229469.

Bullough:1970:BRB

[124] Vern L. Bullough. Book review: The Medieval Treatises of Falconry and Hippiatry of Jacobellus Vitturi from Trogir by Mirko Drazen Grmek; Jakov Stipisic. Isis, 61(4):542–543, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229470.

Jones:1970:BRBc

[125] William Powell Jones. Book review: Chaucer and the Country of the Stars: Poetic Uses of Astrological Imagery by Chauncey Wood. Isis, 61 (4):543–544, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229471.

Heilbron:1970:BRBb

[126] John L. Heilbron. Book review: Martinus van Marum. Life and Work. Volume I by R. J. Forbes. Isis, 61(4):544–545, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229472.

Rauchenberg:1970:BRB

[127] Roy A. Rauchenberg. Book review: Martinus van Marum. Life and Work. Volume II: Journals of Six Journeys through Germany, France and Switzerland, 1782–1802 by R. J. Forbes. Isis, 61(4):545–546, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229473.

Leighly:1970:BRB

[128] John Leighly. Book review: Alexandre de Humboldt, historien et géographe de l’Amérique espagnole 1799/1804 by Charles Minguet. Isis, 61(4):546–547, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229474.

Warner:1970:BRB

[129] Deborah Jean Warner. Book review: Edmond Halley: Genius in Eclipse by Colin A. Ronan. Isis, 61(4):547–548, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229475.

Schlagel:1970:BRB

[130] Richard H. Schlagel. Book review: Bergson and the Evolution of Physics by P. A. Y. Gunter. Isis, 61(4):548–549, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229476.

Burstyn:1970:BRB

[131] Joan N. Burstyn. Book review: The Sane Positivist: A Biography of Edward L. Thorndike by Geraldine Joncich. Isis, 61(4):549–551, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229477.

Glick:1970:BRBc

[132] Thomas F. Glick. Book review: Early Foundations of Mexican Herpetology. An Annotated and Indexed Bibliography of the Herpetological Publications of Alfredo Dugès, 1826–1910 by Hobart M. Smith; Rozella B. Smith. Isis, 61(4):551–552, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229478.

Vandervliet:1970:BRB

[133] Glenn Vandervliet. Book review: Un grand médecin et biologiste Casimir-Joseph Davaine (1812–1882) by Jean Théodoridès. Isis, 61(4): 552–553, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229479.

Forman:1970:BRBe

[134] Paul Forman. Book review: Briefwechsel, 1916–1955, by Albert Einstein, Max Born, and Hedwig Born. Isis, 61(4):553–555, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229480;   http://www.jstor.
org/stable/pdfplus/229480.pdf.

Anonymous:1970:BMd

[135] Anonymous. Back matter. Isis, 61(4):556–566, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229481.

Anonymous:1970:VI

[136] Anonymous. Volume information. Isis, 61(4):558–563, Winter 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229456.

Anonymous:1970:FMe

[137] Anonymous. Front matter. Isis, 61(5):1–4, ???? 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229582.

Neu:1970:NFC

[138] John Neu. Ninety-fifth critical bibliography of the history of science and its cultural influences (to January 1970). Isis, 61(5):5–173+175–192, ???? 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229583.

Anonymous:1970:BMe

[139] Anonymous. Back matter. Isis, 61(5):??, ???? 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229584.

Anonymous:1971:FMa

[140] Anonymous. Front matter. Isis, 62(1):1–2, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228995.

Cassedy:1971:AMA

[141] James H. Cassedy. Applied microscopy and American pork diplomacy: Charles Wardell Stiles in Germany 1898–1899. Isis, 62(1):4–20, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228996.

Iltis:1971:LVV

[142] Carolyn Iltis. Leibniz and the vis viva controversy. Isis, 62(1):21–35, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228997.

Pohle:1971:MFP

[143] William Pohle. The mathematical foundations of Plato’s atomic physics. Isis, 62(1):36–46, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228998.

Kevles:1971:HPC

[144] Daniel J. Kevles. “Into hostile political camps”: The reorganization of international science in World War I. Isis, 62(1):47–60, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228999.

Harper:1971:PJM

[145] Richard I. Harper. Prophatius Judaeus and the Medieval astronomical tables. Isis, 62(1):61–68, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229000.

Hollerith:1971:BSH

[146] Virginia Hollerith and Herman Hollerith. Biographical sketch of herman Hollerith. Isis, 62(1):69–78, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229001.

Rechcigl:1971:IHC

[147] Miloslav Rechcigl, Jr. An introduction to the history of Czechoslovak thought and science: A critical bibliography. Isis, 62(1):79–95, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229002.

Greene:1971:N

[148] John C. Greene. News. Isis, 62(1):96–98, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229003.

Grossman:1971:BRB

[149] Gordon Grossman. Book review: Historic Derivations of Modern Psychiatry. Isis, 62(1):99–105, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229004.

Hall:1971:BRB

[150] A. Rupert Hall. Book review: Reflections on Men and Ideas by Giorgio de Santillana. Isis, 62(1):105–106, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229005.

Rousseau:1971:BRBa

[151] G. S. Rousseau. Book review: The Languages of Criticism and the Sciences of Man: The Structuralist Controversy by Richard Macksey; Eugenio Donato. Isis, 62(1):106–108, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229006.

Greenblatt:1971:BRB

[152] Samuel H. Greenblatt. Book review: The Human Brain and Spinal Cord. A Historical Study Illustrated by Writings from Antiquity to the Twentieth Century by Edwin Clarke; C. D. O’Malley. Isis, 62(1):109–110, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229007.

Jarcho:1971:BRB

[153] Saul Jarcho. Book review: The History of Cardiac Surgery, 1896–1955 by Stephen L. Johnson; The Scalpel and the Heart by Robert G. Richardson. Isis, 62(1):110–111, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229008.

Hahm:1971:BRBa

[154] David E. Hahm. Book review: Aristotle’s Physics by Hippocrates G. Apostle; Aristotle’s Physics, Books I and II by W. Charlton. Isis, 62 (1):111–113, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229009.

Snyder:1971:BRB

[155] Walter F. Snyder. Book review: Textile Manufacture in the Northern Roman Provinces by J. P. Wild. Isis, 62(1):113–114, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229010.

Stannard:1971:BRB

[156] Jerry Stannard. Book review: Agronomie et agronomes en France au XVIIIe siècle by André J. Bourde. Isis, 62(1):114–115, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229011.

Stuewer:1971:BRBa

[157] Roger H. Stuewer. Book review: Newtons Theorie der Prismenfarben by J. A. Lohne; Bernhard Sticker. Isis, 62(1):115, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229012.

Hoyer:1971:BRB

[158] Ulrich Hoyer. Book review: Paul Ehrenfest. Volume I: The Making of a Theoretical Physicist by Martin J. Klein. Isis, 62(1):115–116, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229013.

Beaver:1971:BRB

[159] Donald deB. Beaver. Book review: Dollars for Research. Science and Its Patrons in Nineteenth-Century America by Howard S. Miller. Isis, 62(1):116–117, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229014.

Spencer:1971:BRB

[160] J. Brookes Spencer. Book review: Life of John William Strutt Third Baron Rayleigh, O.M., F.R.S. by Robert John Strutt. Isis, 62(1):118, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229015.

Hodge:1971:BRB

[161] M. J. S. Hodge. Book review: Sir Charles Lyell’s Scientific Journals on the Species Question by Leonard G. Wilson. Isis, 62(1):119–120, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229016.

Hewlett:1971:BRB

[162] Richard G. Hewlett. Book review: Tongues of Conscience: Weapons Research and the Scientists’ Dilemma by R. W. Reid. Isis, 62(1):120– 121, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229017.

Anonymous:1971:BMa

[163] Anonymous. Back matter. Isis, 62(1):122–124, Spring 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229018.

Anonymous:1971:FMb

[164] Anonymous. Front matter. Isis, 62(2):125–127, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229236.

Holmes:1971:AFS

[165] Frederic L. Holmes. Analysis by fire and solvent extractions: The metamorphosis of a tradition. Isis, 62(2):128–148, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229237.

Heimann:1971:MHN

[166] P. M. Heimann. Maxwell, Hertz, and the nature of electricity. Isis, 62 (2):149–157, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229238.

Morrell:1971:UEL

[167] J. B. Morrell. The University of Edinburgh in the Late Eighteenth Century: Its scientific eminence and academic structure. Isis, 62(2):158–171, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229239.

Burns:1971:TET

[168] Alfred Burns. The tunnel of Eupalinus and the tunnel problem of Hero of Alexandria. Isis, 62(2):172–185, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229240.

Hewsen:1971:GPA

[169] Robert H. Hewsen and Pappus of Alexandria. The geography of Pappus of Alexandria: A translation of the Armenian fragments. Isis, 62(2):186– 207, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229241.

Thiele:1971:PCE

[170] Joachim Thiele, Paul Carus, and Ernst Mach. Paul Carus und Ernst Mach: Wechselbeziehungen zwischen deutscher und amerikanischer Philosophie um 1900. (German) [Paul Carus and Ernst Mach: interrelationships between German and American philosophy in 1900]. Isis, 62(2):208–219, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229242.

Multhauf:1971:EEF

[171] Robert P. Multhauf. Éloge: Eduard Farber, 1892–1969. Isis, 62(2):220– 224, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229243.

Heimann:1971:CVV

[172] P. M. Heimann and J. E. McGuire. Cavendish and the Vis Viva controversy: A Leibnizian postscript. Isis, 62(2):225–227, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229244.

McGraw:1971:CKK

[173] Donald J. McGraw. The class krassilnikoviae kriss: A new class of microorganisms? Isis, 62(2):227–230, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229245.

Fellmann:1971:LNP

[174] E. A. Fellmann. Leibniz and Newton’s Principia Mathematica. Isis, 62 (2):230, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229246.

Anonymous:1971:Na

[175] Anonymous. News. Isis, 62(2):231–232, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229247.

Multhauf:1971:BRBa

[176] Robert P. Multhauf. Book review: The Arts of the Alchemists. Isis, 62 (2):233–238, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229248.

White:1971:BRB

[177] George W. White. Book review: The Expeditions of John Charles Frémont, by John Charles Frémont; Donald Jackson; Mary Lee Spence. Volume I: Travels from 1838 to 1844, and Map Portfolio. Isis, 62(2):238– 244, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229249.

Vandervliet:1971:BRB

[178] Glenn Vandervliet. Book review: The Scientific Revolution by Vern L. Bullough. Isis, 62(2):245, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229250.

Schmitt:1971:BRBa

[179] Charles B. Schmitt. Book review: Atti del primo convegno internazionale di ricognizione delle fonti per la storia della scienza italiana: i secoli XIV-XVI by Carlo Maccagni. Isis, 62(2):245–246, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229251.

Woolf:1971:BRB

[180] Harry Woolf. Book review: Sundials: The Art and Science of Gnomonics by Frank W. Cousins. Isis, 62(2):247, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229252.

Polonoff:1971:BRB

[181] Irving Polonoff. Book review: History of Mathematical Logic from Leibniz to Peano by N. I. Styazhkin. Isis, 62(2):247–249, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229253.

Richter:1971:BRB

[182] Steffen Richter. Book review: A German Source-Book in Physics by K. B. Beaton; H. C. Bolton. Isis, 62(2):249–250, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229254.

Ihde:1971:BRB

[183] Aaron J. Ihde. Book review: The Chemistry of Life. Eight Lectures on the History of Biochemistry by Joseph Needham. Isis, 62(2):250–251, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229255.

Burnham:1971:BRBa

[184] John C. Burnham. Book review: A History of Scientific Psychology, Its Origins and Philosophical Backgrounds by D. B. Klein; A History of Modern Psychology by Duane P. Schultz. Isis, 62(2):251–252, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229256.

Bullough:1971:BRB

[185] Vern L. Bullough. Book review: The History of Medical Education by C. D. O’Malley. Isis, 62(2):252, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229257.

Pingree:1971:BRB

[186] David Pingree. Book review: A Commentary on the First Book of Euclid’s Elements by Proclus; Glenn R. Morrow. Isis, 62(2):252–253, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229258.

Lindberg:1971:BRBa

[187] David C. Lindberg. Book review: The Awakening Interest in Science during the First Century of Printing: An Annotated Checklist of First Editions Viewed from the Angle of Their Subject Content by Margaret Bingham Stillwell. Isis, 62(2):254, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229259.

Drake:1971:BRBa

[188] Stillman Drake. Book review: The Infinite Worlds of Giordano Bruno by Antoinette Mann Paterson. Isis, 62(2):254–255, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229260.

Silliman:1971:BRB

[189] Robert H. Silliman. Book review: Jean d’Alembert. Science and the Enlightenment by Thomas L. Hankins. Isis, 62(2):255–257, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229261.

Weiner:1971:BRB

[190] Dora B. Weiner. Book review: Sciences, médecine, pharmacie, de la révolution à l’empire (1789–1815) by Pierre Huard. Isis, 62(2):257–258, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229262.

Kargon:1971:BRBa

[191] Robert Kargon. Book review: The Life of the Honourable Robert Boyle, F.R.S. by R. E. W. Maddison. Isis, 62(2):258–259, Summer 1971. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229263.

Drake:1971:BRBb

[192] Stillman Drake. Book review: Mechanism and Materialism. British Natural Philosophy in an Age of Reason by Robert E. Schofield. Isis, 62 (2):259–260, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229264.

Davis:1971:BRB

[193] Miles Davis. Book review: Carl Friedrich Gauss. A Biography by Tord Hall; Albert Froderberg. Isis, 62(2):260–261, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229265.

North:1971:BRB

[194] J. D. North. Book review: Hilbert: With an Appreciation of His Mathematical Work by Hermann Weyl by Constance Reid. Isis, 62(2):261–262, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229266.

Plotkin:1971:BRB

[195] S. Ya. Plotkin. Book review: Wilhelm Ostwald. An Outstanding Scientist and Science Organizer by N. T. Rodnyi; Yu. I. Solov’ev. Isis, 62(2):262– 264, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229267.

Kauffman:1971:BRB

[196] George B. Kauffman. Book review: The Periodic System of Chemical Elements: A History of the First Hundred Years by Johannes Willem van Spronsen. Isis, 62(2):264–266, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229268.

Stuewer:1971:BRBb

[197] Roger H. Stuewer. Book review: My World Line: An Informal Autobiography by George Gamow. Isis, 62(2):266–267, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229269;   http://www.jstor.
org/stable/pdfplus/229269.pdf.

Anonymous:1971:BMb

[198] Anonymous. Back matter. Isis, 62(2):268–270, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229270.

Anonymous:1971:FMc

[199] Anonymous. Front matter. Isis, 62(3):271–272, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229941.

Marx:1971:AP

[200] Jacques Marx. Alchimie et Palingénésie. Isis, 62(3):274–289, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229942.

Kirby:1971:DA

[201] H. W. Kirby. The discovery of actinium. Isis, 62(3):290–308, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229943.

Hetherington:1971:MRV

[202] Norriss S. Hetherington. The measurement of radial velocities of spiral nebulae. Isis, 62(3):309–313, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229944.

Cassebaum:1971:PSC

[203] Heinz Cassebaum and George B. Kauffman. The periodic system of the chemical elements: The search for its discoverer. Isis, 62(3):314– 327, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229945.

Snelders:1971:JCS

[204] H. A. M. Snelders. J. S. C. Schweigger: His romanticism and his crystal electrical theory of matter. Isis, 62(3):328–338, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229946.

Kohlstedt:1971:STS

[205] Sally Kohlstedt. A step toward scientific self-identity in the United States: The failure of the National Institute, 1844. Isis, 62(3):339–362, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229947.

Grattan-Guinness:1971:MHM

[206] I. Grattan-Guinness and Magnus Gustav Mittag-Leffler. Materials for the history of mathematics in the institut Mittag-Leffler. Isis, 62(3):

363–374, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229948.

Mackay:1971:JSP

[207] Alan L. Mackay. Japanese seven-place Sine and tangent tables of 1856. Isis, 62(3):375–379, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229949.

Glick:1971:IKS

[208] Thomas F. Glick. On the influence of Kircher in Spain. Isis, 62(3):379– 381, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229950.

Mitchell:1971:EC

[209] M. E. Mitchell and Charles F. O’Brien. On Eozoön canadense. Isis, 62 (3):381–383, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229951.

Anonymous:1971:Nb

[210] Anonymous. News. Isis, 62(3):384–387, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229952.

Mayerhofer:1971:BRB

[211] Josef Mayerhöfer. Book review: Dictionary of Scientific Biography. Isis, 62(3):388–390, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229953.

Warner:1971:BRB

[212] Deborah Jean Warner. Book review: Libro dei Globi. Isis, 62(3):390– 394, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229954.

Multhauf:1971:BRBb

[213] Robert P. Multhauf. Book review: Neue (sogenannte) Magdeburger über den leeren Raum. Isis, 62(3):394–397, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229955.

Plessner:1971:BRB

[214] Martin Plessner. Book review: Astrologumena, die astrologische Literatur in der Antike und ihre Geschichte. Isis, 62(3):397–401, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229956.

Schmitt:1971:BRBb

[215] Charles B. Schmitt. Book review: Philosophy, Technology and the Arts in the Early Modern Era by Paolo Rossi; Salvator Attanasio; Benjamin Nelson. Isis, 62(3):401–402, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229957.

Rousseau:1971:BRBb

[216] G. S. Rousseau. Book review: The Children of Frankenstein: A Primer on Modern Technology and Human Values by Herbert J. Muller. Isis, 62 (3):402–404, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229958.

Hofmann:1971:BRB

[217] J. E. Hofmann. Book review: Geschichte des Zahlbegriffs by Helmuth Gericke. Isis, 62(3):404–405, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229959.

Berman:1971:BRB

[218] Alex Berman. Book review: Therapie von den Primitiven bis zum 20. Jahrhundert by Erwin H. Ackerknecht. Isis, 62(3):405–406, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229960.

Condit:1971:BRB

[219] Carl W. Condit. Book review: The Study of Architectural History by Bruce Alsopp. Isis, 62(3):406–407, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229961.

Pierson:1971:BRB

[220] Stuart Pierson. Book review: John Dalton and the Progress of Science by D. S. L. Cardwell. Isis, 62(3):407–409, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229962.

Merzbach:1971:BRB

[221] Uta Merzbach. Book review: The Mathematical Papers of Isaac Newton by Isaac Newton; D. T. Whiteside; M. A. Hoskin. Isis, 62(3):409–411, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229963.

Jones:1971:BRB

[222] William Powell Jones. Book review: Science in the British Colonies of America by Raymond Phineas Stearns. Isis, 62(3):411–412, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229964.

Morris:1971:BRB

[223] Robert J. Morris, Jr. Book review: The Collected Works of Count Rumford by Benjamin Thompson; Count Rumford; Sanborn C. Brown; The Collected Works of Count Rumford. Volume II: Practical Applications of Heat; The Collected Works of Count Rumford. Volume III: Devices and Techniques; The Collected Works of Count Rumford. Volume IV: Light and Armament; The Collected Works of Count Rumford. Volume V: Public Institutions. Isis, 62(3):412–413, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229965.

Drake:1971:BRBc

[224] Stillman Drake. Book review: Samuel Hartlib and the Advancement of Learning by Charles Webster. Isis, 62(3):413–414, Autumn 1971. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229966.

Abrams:1971:BRB

[225] John W. Abrams. Book review: The Royal Institution Library of Science. (Being the Friday Evening Discourses in Physical Sciences Held at the Royal Institution, 1851–1939.) Volume 1: Astronomy by Bernard Lovell. Isis, 62(3):414–415, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229967.

Badash:1971:BRBa

[226] Lawrence Badash. Book review: Peter Kapitsa on Life and Science by Albert Parry. Isis, 62(3):415, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229968.

May:1971:BRB

[227] Kenneth O. May. Book review: Quelques aspects de la pensée d’un mathématicien by Paul Levy. Isis, 62(3):415–416, Autumn 1971. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229969.

Stuewer:1971:BRBc

[228] Roger H. Stuewer. Book review: Particle Accelerators: A Brief History by M. Stanley Livingston. Isis, 62(3):416–417, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229970.

Anonymous:1971:BMc

[229] Anonymous. Back matter. Isis, 62(3):418–419, Autumn 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229971.

Anonymous:1971:FMd

[230] Anonymous. Front matter. Isis, 62(4):421–423, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229815.

Mayr:1971:MOC

[231] Otto Mayr. Maxwell and the origins of cybernetics. Isis, 62(4):424– 444, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229816.

Albury:1971:HTL

[232] William R. Albury. Halley and the Traité de la lumière of huygens: New light on Halley’s relationship with Newton. Isis, 62(4):445–468, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229817.

Lindberg:1971:ACE

[233] David C. Lindberg. Alkindi’s critique of Euclid’s theory of vision. Isis, 62(4):469–489, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229818.

Kren:1971:RDN

[234] Claudia Kren. The rolling device of Nas.ir al-Dīn al-T .ūsī in the de spera of Nicole oresme? Isis, 62(4):490–498, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229819.

Schufle:1971:EWB

[235] J. A. Schufle and George Thomas. Equivalent weights from Bergman’s data on phlogiston content of metals. Isis, 62(4):499–506, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229820.

Zeitlin:1971:EHM

[236] Jacob Israel Zeitlin and V. F. Lenzen. Éloge: Herbert M. Evans 1882– 1971. Isis, 62(4):507–511, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229821.

Daub:1971:WIK

[237] Edward E. Daub. Waterston’s influence on Krönig’s kinetic theory of gases. Isis, 62(4):512–515, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229822.

Grattan-Guinness:1971:MMC

[238] I. Grattan-Guinness. Missing materials concerning the life and work of Georg Cantor. Isis, 62(4):516–517, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229823.

Anonymous:1971:Nc

[239] Anonymous. News. Isis, 62(4):518–521, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229824.

Lindberg:1971:BRBb

[240] David C. Lindberg. Book review: The Nature of Light: An Historical Survey. Isis, 62(4):522–524, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229825.

Forman:1971:BRB

[241] Paul Forman. Book review: Behind Appearance: A Study of the Relations between Painting and the Natural Sciences in This Century. Isis, 62(4): 525–526, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229826.

Burnham:1971:BRBb

[242] John C. Burnham. Book review: The Discovery of the Unconscious, by Henri F. Ellenberger. The History and Evolution of Dynamic Psychiatry. Isis, 62(4):527–529, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229827.

Geran:1971:BRB

[243] Juliana Geran. Book review: Phenomenology and the Natural Sciences. Essays and Translations by Joseph J. Kockelmans; Theodore J. Kisiel. Isis, 62(4):529–530, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229828.

Kargon:1971:BRBb

[244] Robert H. Kargon. Book review: Henri van Heurck Museum. Geillustreerde Inventaris van de historische Microscopen Onderdelen en Uitrusting by Karel Frison. Isis, 62(4):530, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229829.

Thoren:1971:BRB

[245] Victor Thoren. Book review: Astronomy Gnomonics. A Catalogue of Instruments of the 15th to the 19th Centuries in the Collections of the National Technical Museum, Prague by Zdenek Horský; Otilie Skopová. Isis, 62(4):530, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229830.

Bieber:1971:BRB

[246] Jerome B. Bieber. Book review: Politics and the Community of Science by Joseph Haberer. Isis, 62(4):531–532, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229831.

Bernkopf:1971:BRB

[247] Michael Bernkopf. Book review: The Development of the Foundations of Mathematical Analysis from Euler to Riemann by Ivor Grattan-Guinness. Isis, 62(4):532–533, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229832.

Scriba:1971:BRB

[248] Christoph J. Scriba. Book review: A Source Book in Mathematics, 1200–1800 by D. J. Struik. Isis, 62(4):533–534, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229833.

Stuewer:1971:BRBd

[249] Roger H. Stuewer. Book review: Sources of Color Science by David L. MacAdam. Isis, 62(4):534–535, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229834.

Middleton:1971:BRB

[250] W. E. Knowles Middleton. Book review: A Century of Weather Service 1870–1970 by Patrick Hughes. Isis, 62(4):535, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229835.

Burstyn:1971:BRB

[251] Harold L. Burstyn. Book review: Exploring the Ocean World. A History of Oceanography by C. P. Idyll; Man and the Sea. Classic Accounts of Marine Explorations by Bernard L. Gordon. Isis, 62(4):535–537, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229836.

Sattler:1971:BRB

[252] Rolf Sattler. Book review: Le catalogue de la vie. Etude méthodologique sur la taxonomie by François Dagognet. Isis, 62(4):537, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229837.

Cassedy:1971:BRB

[253] James H. Cassedy. Book review: Disordered Minds. The First Century of Eastern State Hospital in Williamsburg, Virginia 1766–1866 by Norman Dain. Isis, 62(4):537–538, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229838.

Ahlborn:1971:BRB

[254] Richard Ahlborn. Book review: The Sugar Hacienda of the Marqueses del Valle by Ward J. Barrett. Isis, 62(4):538–539, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229839.

Hahm:1971:BRBb

[255] David E. Hahm. Book review: Die Naturphilosophie des Straton von Lampsakos: Zur Geschichte des Problems der Bewegung im Bereich des frühen Peripatos by Matthias Gatzemeier. Isis, 62(4):539–540, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229840.

Grant:1971:BRB

[256] Edward Grant. Book review: Weltbild und Astrologie in den griechischen Zauberpapyri by Hans Georg Gundel. Isis, 62(4):540, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229841.

Smith:1971:BRB

[257] Emilie Savage Smith. Book review: Galen on the Usefulness of the Parts of the Body (Peri Chreias Morion. De usu partium) by Margaret Tallmadge May. Isis, 62(4):540–542, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229842.

McVaugh:1971:BRB

[258] Michael McVaugh. Book review: The Liber de diversis medicinis in the Thornton Manuscript (MS Lincoln Cathedral A.5.2) by Margaret Sinclair Ogden. Isis, 62(4):542, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229843.

Hamarneh:1971:BRB

[259] Sami Hamarneh. Book review: Der Diabetestraktat ’Abd al-Lat.if al-Bagdadi’s by Hans–Jürgen Thies. Isis, 62(4):542–543, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229844.

King:1971:BRB

[260] Lester S. King. Book review: Achilles Pirmin Gasser, 1505–1577: Arzt und Naturforscher, Historiker und Humanist by Karl Heinz Burmeister. Isis, 62(4):543–544, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229845.

VanHelden:1971:BRB

[261] A. Van Helden. Book review: Simon Stevin: Science in the Netherlands around 1600 by E. J. Dijksterhuis. Isis, 62(4):544–545, Winter 1971.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229846.

Christianson:1971:BRB

[262] J. R. Christianson. Book review: Tycho Brahe. En levnadsteckning med nya bidrag belysande hans liv och verk by Wilhelm Norlind. Isis, 62(4): 545–546, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229847.

Reingold:1971:BRB

[263] Nathan Reingold. Book review: Herman Boerhaave 1668–1968. Tentoonstelling. Isis, 62(4):546, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229848.

Aiton:1971:BRB

[264] E. J. Aiton. Book review: Galileo Studies: Personality, Tradition and Revolution by Stillman Drake. Isis, 62(4):546–547, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229849.

Dewhurst:1971:BRB

[265] Kenneth Dewhurst. Book review: The Common Scientist in the Seventeenth Century. A Study of the Dublin Philosophical Society 1683–1708 by K. Theodore Hoppen. Isis, 62(4):547–549, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229850.

Kargon:1971:BRBc

[266] Robert H. Kargon. Book review: Henry Wansey and His American Journal, 1794 by David Jeremy. Isis, 62(4):549–550, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229851.

Herbert:1971:BRB

[267] Sandra Herbert. Book review: Kant’s Concept of Geography and Its Relation to Recent Geographical Thought by J. A. May. Isis, 62(4):550– 551, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229852.

Kargon:1971:BRBd

[268] Robert H. Kargon. Book review: A Treatise of the System of the World by Issac Newton. Isis, 62(4):551, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229854.

Stocking:1971:BRB

[269] George W. Stocking, Jr. Book review: Noticias de Nutka: An Account of Nootka Sound in 1792 by José Mariano Moziño; Iris Higbie Wilson. Isis, 62(4):551, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229853.

Molella:1971:BRB

[270] Arthur P. Molella. Book review: Herder and the Philosophy and History of Science by H. B. Nisbet. Isis, 62(4):551–552, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229855.

Rappaport:1971:BRB

[271] Rhoda Rappaport. Book review: An Introduction to the Natural History of the Terrestrial Sphere by Rudolf Erich Raspe; Audrey Notvik Iversen; Albert V. Carozzi. Isis, 62(4):552–554, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229856.

Hahn:1971:BRB

[272] Roger Hahn. Book review: Atoms and Powers. An Essay on Newtonian Matter-Theory and the Development of Chemistry by Arnold Thackray. Isis, 62(4):554–555, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229857.

Brush:1971:BRB

[273] Stephen Brush. Book review: Vorgeschichte des Planckschen Strahlungsgesetzes. Messungen und Theorien der spektralen Energieverteilung bis zur Begründung der Quantenhypothese by Hans Kangro. Isis, 62(4):555– 556, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229858.

Alexander:1971:BRB

[274] Theodore Alexander. Book review: Der Physiologe und Planktonforscher Victor Hensen (1835–1924). Sein Leben und sein Werk by Rüdiger Porep. Isis, 62(4):556–557, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229859.

Filner:1971:BRB

[275] Robert E. Filner. Book review: Fifty Years with Science by J. G. Crowther. Isis, 62(4):557–558, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229860.

Hufbauer:1971:BRB

[276] Karl Hufbauer. Book review: Physics and Beyond. Encounters and Conversations by Werner Heisenberg; Arnold J. Pomerans. Isis, 62(4):558– 560, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229861.

Adams:1971:BRB

[277] Mark B. Adams. Book review: The Lysenko Affair by David Joravsky. Isis, 62(4):560–561, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229862.

Badash:1971:BRBb

[278] Lawrence Badash. Book review: Enrico Fermi, Physicist by Emilio Segrè. Isis, 62(4):561–562, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229863.

Anonymous:1971:BMd

[279] Anonymous. Back matter. Isis, 62(4):563–565, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229864.

Anonymous:1971:VI

[280] Anonymous. Volume information. Isis, 62(4):566–572, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229814.

Anonymous:1971:FMe

[281] Anonymous. Front matter. Isis, 62(5):1–4, ???? 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228835.

Neu:1971:NSC

[282] John Neu. Ninety-sixth critical bibliography of the history of science and its cultural influences (to January 1971). Isis, 62(5):5–11+13–173, ????

1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228836.

Anonymous:1971:BMe

[283] Anonymous. Back matter. Isis, 62(5):??, ???? 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228837.

Anonymous:1972:FMa

[284] Anonymous. Front matter. Isis, 63(1):1–3, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229189.

Miller:1972:RNE

[285] John David Miller. Rowland and the nature of electric currents. Isis, 63 (1):4–27, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229190.

Drake:1972:UME

[286] Stillman Drake. The uniform motion equivalent to a uniformly accelerated motion from rest (galileo gleanings xx). Isis, 63(1):28–38, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229191.

Goldstein:1972:TOM

[287] Bernard R. Goldstein. Theory and observation in Medieval astronomy. Isis, 63(1):39–47, March 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229192.

Badash:1972:CNC

[288] Lawrence Badash. The completeness of Nineteenth-Century science. Isis, 63(1):48–58, March 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229193.

Siegfried:1972:LVG

[289] Robert Siegfried. Lavoisier’s view of the gaseous state and its early application to pneumatic chemistry. Isis, 63(1):59–78, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229194.

Ariotti:1972:BCE

[290] Piero Ariotti and Benedetto Castelli. Benedetto castelli: Early systematic experiments and theory of the differential absorption of heat by colors. Isis, 63(1):79–87, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229195.

Rose:1972:CJD

[291] Paul Lawrence Rose. Commandino, John Dee, and the De superficierum Divisionibus of Machometus Bagdedinus. Isis, 63(1):88–93, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229196. See [595].

Jarrell:1972:LDC

[292] Richard A. Jarrell. The latest date of composition of Gilbert’s de mundo. Isis, 63(1):94–95, March 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229197.

Schmitt:1972:PMG

[293] Charles B. Schmitt. On a poem misattributed to Galileo. Isis, 63(1): 95–97, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229198.

Greene:1972:N

[294] John C. Greene. News. Isis, 63(1):98–100, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229199.

Hellman:1972:BRB

[295] C. Doris Hellman. Book review: Johann Kepler: Sein Leben in Bildern und eigenen Berichten. Isis, 63(1):101–103, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229200.

Rousseau:1972:BRBa

[296] G. S. Rousseau. Book review: The Road to Medical Enlightenment 1650– 1695. Isis, 63(1):103–106, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229201.

Frank:1972:BRB

[297] Robert G. Frank, Jr. Book review: Oxford Medicine: Essays on the Evolution of the Oxford Clinical School by Kenneth Dewhurst. Isis, 63 (1):106–107, March 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229202.

Marsak:1972:BRB

[298] Ann Marsak. Book review: The Man in the Moone and Other Lunar Fantasies by Faith K. Pizor; T. Allan Comp. Isis, 63(1):108, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229203.

Kauffman:1972:BRBa

[299] George B. Kauffman. Book review: Studies in the History of Chemistry by Harold Hartley. Isis, 63(1):109–110, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229204.

Thoren:1972:BRBa

[300] Victor E. Thoren. Book review: Discovering the Universe. A History of Astronomy by Colin A. Ronan. Isis, 63(1):110, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229205.

Aiton:1972:BRC

[301] E. J. Aiton. Book review: The Conflict between Atomism and Conservation Theory 1644 to 1860 by Wilson L. Scott. Isis, 63(1):110–111, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229206.

Heilbron:1972:BRP

[302] John L. Heilbron. Book review: Physics in My Generation by Max Born; Bahnbrecher des Atomzeitalters. Grosse Naturforscher von Maxwell bis Heisenberg by Friedrich Herneck; Moseley and the Numbering of the Elements by Bernard Jaffe; The Big Machine by Robert Jungk. Isis, 63 (1):111–112, March 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229207;
 http://www.jstor.org/stable/pdfplus/229207.pdf.

Ospovat:1972:BRB

[303] Alexander M. Ospovat. Book review: Adventures in Earth History, Being a Volume of Significant Writings from Original Sources, on Cosmology, Geology, Climatology, Oceanography, Organic Evolution, and Related Topics of Interests to Students of Earth History, from the Time of Nicolaus Steno to the Present. Selected, Edited, and with Introductions… by Preston Cloud. Isis, 63(1):112–113, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229208.

Gruber:1972:BRB

[304] Jacob W. Gruber. Book review: A Hundred Years of Anthropology by T. K. Penniman. Isis, 63(1):113–114, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229209.

Theodorides:1972:BRBa

[305] J. Théodoridès. Book review: A Medical Bibliography (Garrison and Morton). An Annotated Checklist of Texts Illustrating the History of Medicine by Leslie T. Morton; Bibliographic Control of the Literature of Oncology, 1800–1960 by Pauline M. Vaillancourt. Isis, 63(1):114–115, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229210.

Heilbron:1972:BRBc

[306] John L. Heilbron. Book review: The Sources of Invention. A Study of the Causes and Consequences of Industrial Innovation through the Inventions of the Nineteenth and Twentieth Centuries by John Jewkes; David Sawers; Richard Stillerman; Victorian Technology and Its Preservation in Modern Britain by Norman A. F. Smith; Technology in Retrospect and Critical Events in Science. Volume I by Illinois Institute of Technology Research Institute. Isis, 63(1):115, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229211.

Mayr:1972:BRBa

[307] Otto Mayr. Book review: The Pneumatics of Hero of Alexandria: A Facsimile of the 1851 Woodcroft Edition. Isis, 63(1):116, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229212.

Smith:1972:BRBa

[308] Emilie Savage Smith. Book review: Galen on Sense Perception, His Doctrines, Observations and Experiments on Vision, Hearing, Smell, Taste, Touch and Pain, and Their Historical Sources by Rudolf E. Siegel. Isis, 63(1):116–118, March 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229213.

Hamarneh:1972:BRB

[309] Sami Hamarneh. Book review: Ibn al-Haitham by Hakim Mohammed Said. Isis, 63(1):118–119, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229214.

King:1972:BRB

[310] Lester S. King. Book review: Chronicle from Aldgate: Life and Death in Shakespeare’s London by Thomas Rogers Forbes. Isis, 63(1):119– 120, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229215.

McVaugh:1972:BRB

[311] Michael McVaugh. Book review: Brunelleschi. Studies of His Technology and Inventions by Frank D. Prager; Gustina Scaglia. Isis, 63(1):120– 121, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229216.

Bylebyl:1972:BRBa

[312] Jerome J. Bylebyl. Book review: Captain Cook, Navigator and Scientist. Papers Presented at the Cook Bicentenary Symposium, Australian Academy of Science, Canberra, 1 May 1969 by G. M. Badger. Isis, 63(1): 121, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229217.

Shea:1972:BRBa

[313] William R. Shea. Book review: L’universo-macchina. Origini della filosofia newtoniana by Paolo Casini. Isis, 63(1):121–122, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229218.

Jones:1972:BRB

[314] William Powell Jones. Book review: Chain of Friendship: Selected Letters of Dr. John Fothergill, 1735–1780 by Betsy C. Corner; Christopher C. Booth. Isis, 63(1):122–123, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229219.

Briggs:1972:BRB

[315] J. Morton Briggs, Jr. Book review: Lazare Carnot, Savant by Charles C. Gillispie. Isis, 63(1):123–124, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229220.

Bylebyl:1972:BRBb

[316] Jerome J. Bylebyl. Book review: Fundamenta medicinae by Friedrich Hoffmann; Lester S. King. Isis, 63(1):124, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229221.

Kargon:1972:BRBa

[317] Robert H. Kargon. Book review: The Posthumous Works of Robert Hooke by Robert Hooke; Richard Waller. Isis, 63(1):124, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229222.

Kargon:1972:BRBb

[318] Robert H. Kargon. Book review: A System of Natural Philosophy by Jacques Rohault; John Clarke; Samuel Clarke. Isis, 63(1):124–125, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229223.

Goerke:1972:BRB

[319] H. Goerke. Book review: Lachesis and Nemesis. Four Chapters on the Human Condition in the Writings of Carl Linnaeus by K. Rob. V. Wikman. Isis, 63(1):125, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229224.

Stocking:1972:BRBa

[320] George W. Stocking, Jr. Book review: Origins of Psychometry: Johan Jacob De Jaager, Student of F. C. Donders, on Reaction Time and Mental Processes (1865) by Josef Brožek; Maarten S. Sibinga. Isis, 63(1): 125, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229225.

McGucken:1972:BRB

[321] William McGucken. Book review: Lord Rayleigh: The Man and His Work by Robert Bruce Lindsay. Isis, 63(1):126, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229226.

Mayr:1972:BRBb

[322] Otto Mayr. Book review: Osborne Reynolds and Engineering Science Today: Papers Presented at the Osborne Reynolds Centenary Symposium, University of Manchester, September, 1968 by D. M. McDowell; J. D. Jackson. Isis, 63(1):126–127, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229227.

Thoren:1972:BRBb

[323] Victor E. Thoren. Book review: Trackers of the Skies by E. Nelson Hayes. Isis, 63(1):127, March 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229228.

Mayr:1972:BRBc

[324] Ernst Mayr. Book review: Oskar Heinroth, Vater der Verhaltensforschung 1871–1945 by Katharina Heinroth. Isis, 63(1):127–128, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229229.

Reingold:1972:BRBa

[325] Nathan Reingold. Book review: L. J. Henderson on the Social System by Bernard Barber. Isis, 63(1):128–129, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229230.

Hart:1972:BRB

[326] Richard Hart. Book review: Spectroscopic Astrophysics. An Assessment of the Contributions of Otto Struve by G. H. Herbig. Isis, 63(1):129, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229231.

Darnell:1972:BRB

[327] Regna Darnell. Book review: Alfred Kroeber: A Personal Configuration by Theodora Kroeber. Isis, 63(1):130–131, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229232.

Theodorides:1972:BRBb

[328] J. Théodoridès. Book review: History of the American College of Apothecaries. The First Quarter Century, 1940–1965 by Ernst W. Stieb. Isis, 63(1):131, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229233.

Gross:1972:BRBa

[329] Walter E. Gross. Book review: Crucial Experiments in Modern Physics by George L. Trigg. Isis, 63(1):131–132, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229234.

Anonymous:1972:BMa

[330] Anonymous. Back matter. Isis, 63(1):133–138, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229235.

Anonymous:1972:FMb

[331] Anonymous. Front matter. Isis, 63(2):139–141, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229042.

Salaff:1972:BHK

[332] Stephen Salaff. A biography of hua lo-keng. Isis, 63(2):142–183, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229043.

Westfall:1972:CMS

[333] Richard S. Westfall. Circular motion in Seventeenth-Century mechanics. Isis, 63(2):184–189, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229044.

Home:1972:ABE

[334] R. W. Home. Aepinus and the British electricians: The dissemination of a scientific theory. Isis, 63(2):190–204, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229045.

Hahm:1972:CSD

[335] David E. Hahm. Chrysippus’ solution to the Democritean dilemma of the cone. Isis, 63(2):205–220, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229046.

Scheubel:1972:JSR

[336] Johann Scheubel and Barnabas B. Hughes. Johann Scheubel’s revision of Jordanus de Nemore’s De numeris datis: An analysis of an unpublished manuscript. Isis, 63(2):221–234, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229047.

Woolf:1972:BRB

[337] Harry Woolf. Book review: An Original Theory or New Hypothesis of the Universe (1750), by Thomas Wright; Michael A. Hoskin; Second or Singular Thoughts upon the Theory of the Universe by Thomas Wright; Michael A. Hoskin; Clavis Coelestis. Being the Explication of a Diagram Entitled A Synposis of the Universe or, the Visible World Epitomised by Thomas Wright (1742) by Thomas Wright. A Facsimile Reprint together with the First Publication of a Theory of the Universe (1734). Isis, 63(2): 235–241, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229048.

Drake:1972:BRB

[338] Stillman Drake. Book review: Force in Newton’s Physics, by Richard S. Westfall. The Science of Dynamics in the Seventeenth Century. Isis, 63(2):242–244, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229049.

Benison:1972:BRB

[339] Saul Benison. Book review: Science in American Society: A Social History. Isis, 63(2):245–249, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229050.

Storer:1972:BRB

[340] Norman W. Storer. Book review: The Scientist’s Role in Society: A Comparative Study. Isis, 63(2):249–251, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229051.

Stricklen:1972:BRB

[341] Charles G. Stricklen, Jr. Book review: Neuzeit und Aufklärung, by Jürgen Mittelstrass. Studien zur Entstehung der Neuzeitlichen Wissenschaft und Philosophie. Isis, 63(2):251–253, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229052.

Hilfstein:1972:BRB

[342] Erna Hilfstein and Edward Rosen. Book review: Kopernik i heliocentryzm w polskiej kulturze umyslowej do końca XVIII wieku by Barbara Bieńkowska. Isis, 63(2):254–255, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229053.

Penick:1972:BRB

[343] James Penick, Jr. Book review: Science in America: Historical Selections by John C. Burnham. Isis, 63(2):255, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229054.

Colodny:1972:BRBa

[344] Robert G. Colodny. Book review: Scientific Types by J. G. Crowther. Isis, 63(2):255–256, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229055.

Bolzan:1972:BRB

[345] J. E. Bolzan. Book review: The Nature and Function of Scientific Theories. Essays in Contemporary Science and Philosophy by Robert G. Colodny. Isis, 63(2):256–257, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229056.

LaPorte:1972:BRB

[346] Deirdre La Porte. Book review: Sur l’histoire des sciences by Michel Pécheux; Michel Fichant. Isis, 63(2):257–258, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229057.

Sherwood:1972:BRBa

[347] Morgan Sherwood. Book review: Science and Public Policy by Louise B. Young; William J. Trainor. Isis, 63(2):258–259, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229058.

Thoren:1972:BRBc

[348] Victor E. Thoren. Book review: Swedish Astronomers 1477–1900 by Per Collinder. Isis, 63(2):259, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229059.

Kauffman:1972:BRBb

[349] George B. Kauffman. Book review: Classical Scientific Papers– Chemistry, Second Series. Papers on the Nature and Arrangement of the Chemical Elements. Isis, 63(2):259–261, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229060.

Gross:1972:BRBb

[350] Walter E. Gross. Book review: Physics for Poets by Robert H. March. Isis, 63(2):261–262, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229061.

Goetzmann:1972:BRB

[351] William H. Goetzmann. Book review: The Makers of Modern Geography by Robert E. Dickinson. Isis, 63(2):262–263, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229062.

Brieger:1972:BRB

[352] Gert H. Brieger. Book review: Mainstreams of Medicine. Essays on the Social and Intellectual Context of Medical Practice by Lester S. King. Isis, 63(2):263–264, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229063.

Kruta:1972:BRBa

[353] Vladislav Kruta. Book review: The History of Coronary Heart Disease by J. O. Leibowitz. Isis, 63(2):264–265, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229064.

Mayr:1972:BRBd

[354] Otto Mayr. Book review: A History of Technology and Invention: Progress through the Ages by Maurice Daumas; Eileen B. Hennessy. Isis, 63(2):265–267, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229065.

Calvert:1972:BRB

[355] Monte A. Calvert. Book review: Early Stationary Steam Engines in America: A Study in the Migration of a Technology by Carroll W. Pursell. Isis, 63(2):267–268, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229066.

Thoren:1972:BRBd

[356] Victor E. Thoren. Book review: The Astrologers and Their Creed: An Historical Outline by Christopher McIntosh. Isis, 63(2):268, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229067.

Stewart:1972:BRB

[357] T. D. Stewart. Book review: Medicine and Ethnology. Selected Essays by Erwin H. Ackerknecht; H. H. Walser; H. M. Koelbing. Isis, 63(2):268– 269, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229068.

Thoren:1972:BRBe

[358] Victor E. Thoren. Book review: Verformungstendenzen in der Überlieferung antiker Sonnen- und Mondfinsternisse by Alexander Demandt. Isis, 63(2):269, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229069.

Kudlien:1972:BRB

[359] F. Kudlien. Book review: Die hippokratische Schrift “Über die heilige Krankheit” by Hermann Grensemann. Isis, 63(2):269–270, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229070.

Kren:1972:BRB

[360] Claudia Kren. Book review: Dynamische und statische Betrachtungsweise in der antiken Mechanik by Fritz Krafft. Isis, 63(2):270–271, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229071.

Wallace:1972:BRB

[361] William A. Wallace. Book review: John Pecham and the Science of Optics. Perspectiva Communis by David C. Lindberg. Isis, 63(2):271– 272, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229072.

Folkerts:1972:BRB

[362] Menso Folkerts. Book review: De Verhandeling over de Cirkelkwadratuur van Franco van Luik van Omstreeks 1050 by A. J. E. M. Smeur. Isis, 63(2):272–274, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229073.

Smith:1972:BRBb

[363] Emilie Savage Smith. Book review: Bibliography of Mediaeval Arabic and Jewish Medicine and Allied Sciences by R. Y. Ebied. Isis, 63(2): 274–275, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229074.

Burckhardt:1972:BRB

[364] J. J. Burckhardt. Book review: The Thousands of Abu Ma’shar by David Pingree. Isis, 63(2):275–276, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229075.

Huard:1972:BRB

[365] P. Huard. Book review: Glimpses of Health and Medicine in Mauryan Empire by D. V. Subba Reddy. Isis, 63(2):276–277, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229076.

Schmitt:1972:BRBa

[366] Charles B. Schmitt. Book review: Mathesis universalis: genesi di una idea nel XVI secolo by Giovanni Crapulli. Isis, 63(2):277–278, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229077.

Ravetz:1972:BRB

[367] Jerome R. Ravetz. Book review: Three Copernican Treatises: The “Commentariolus” of Copernicus, the “Letter against Werner,” the “Narratio prima” of Rheticus by Edward Rosen. Isis, 63(2):278, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229078.

Applebaum:1972:BRB

[368] Wilbur Applebaum. Book review: The Traces of Thomas Hariot by Muriel Rukeyser. Isis, 63(2):278–280, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229079.

Polonoff:1972:BRB

[369] Irving Polonoff. Book review: The Methodological Heritage of Newton by Robert E. Butts; John W. Davis. Isis, 63(2):280–282, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229080.

Davis:1972:BRB

[370] Audrey Davis. Book review: Edward Stevens: Gastric Physiologist, Physician and American Statesman by Stacey B. Day. Isis, 63(2):282, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229081.

Rauschenberg:1972:BRB

[371] Roy A. Rauschenberg. Book review: John Banister and His Natural History of Virginia, 1678–1692 by Joseph Ewan; Nesta Ewan. Isis, 63 (2):282–283, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229082.

Shapiro:1972:BRB

[372] Barbara Shapiro. Book review: The Correspondence of Henry Oldenburg by Henry Oldenburg; A. Rupert Hall; Marie Boas Hall; The Correspondence of Henry Oldenburg. Volume V: 1668–1669; The Correspondence of Henry Oldenburg. Volume VI: 1669–1670; The Correspondence of Henry Oldenburg. Volume VII: 1670–1671. Isis, 63(2):283–284, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229083.

Warner:1972:BRBa

[373] Deborah Jean Warner. Book review: Johannes Hevelius and His Catalog of Stars. Isis, 63(2):284–285, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229084.

Calinger:1972:BRB

[374] Ronald Calinger. Book review: Der russische Militärarzt im 18. Jahrhundert by Heinz Müller–Dietz. Isis, 63(2):285–286, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229085.

Caswell:1972:BRB

[375] John E. Caswell. Book review: Physician to the West: Selected Writings of Daniel Drake on Science and Society by Henry D. Shapiro; Zane L. Miller; The Journals of Joseph N. Nicollet: A Scientist on the Mississippi Headwaters with Notes on Indian Life, 1836–37 by Martha Coleman Bray; Andre Fertey. Isis, 63(2):286–288, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229086.

Molella:1972:BRB

[376] Arthur P. Molella. Book review: Dietrich Georg Kieser, Sein Leben und Werk by Walter Brednow. Isis, 63(2):288–289, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229087.

Rinker:1972:BRB

[377] Harry L. Rinker. Book review: Lewis and Clark: Pioneering Naturalists by Paul Russell Cutright. Isis, 63(2):289–290, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229088.

Williams:1972:BRB

[378] L. Pearce Williams. Book review: Hegel’s Philosophy of Nature by Michael John Petry. Isis, 63(2):290–291, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229089.

Burchfield:1972:BRB

[379] Joe D. Burchfield. Book review: Sir William Dawson: A Life in Science and Religion by Charles F. O’Brien. Isis, 63(2):291–292, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229090.

Theodorides:1972:BRBc

[380] J. Théodoridès. Book review: Cellular Pathology, as Based on Physiological and Pathological Histology by Rudolf Virchow; Frank Chance. Isis, 63(2):292, June 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229091.

Herbert:1972:BRB

[381] Sandra Herbert. Book review: Charles Darwin: The Years of Controversy. The Origin of Species and Its Critics 1859–1882 by Peter Vorzimmer. Isis, 63(2):292–294, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229092.

Egerton:1972:BRBa

[382] Frank N. Egerton, III. Book review: Henry Walter Bates, Naturalist of the Amazons by George Woodcock. Isis, 63(2):295–296, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229093.

Parascandola:1972:BRB

[383] John Parascandola. Book review: Pioneer in Modern Medicine: David Linn Edsall of Harvard by Joseph C. Aub; Ruth K. Hapgood. Isis, 63 (2):296, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229094.

Colodny:1972:BRBb

[384] Robert G. Colodny. Book review: William Morton Wheeler, Biologist by Mary Alice Evans; Howard Ensign Evans. Isis, 63(2):297, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229095.

Bieber:1972:BRB

[385] Jerome B. Bieber. Book review: Cortisone. Memoirs of a Hormone Hunter by Edward C. Kendall. Isis, 63(2):297–298, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229096.

Lindenfeld:1972:BRB

[386] David Lindenfeld. Book review: The Selected Papers of Wolfgang Köhler by Wolfgang Köhler; Mary Henle. Isis, 63(2):298, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229097.

Hendrickson:1972:BRB

[387] Walter B. Hendrickson. Book review: Charles Kenneth Leith, Scientific Adviser by Sylvia Wallace McGrath. Isis, 63(2):298–299, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229098.

Brittain:1972:BRB

[388] James E. Brittain. Book review: Lectures on Electrical Engineering by Charles Proteus Steinmetz; Philip L. Alger. Isis, 63(2):299–300, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229099.

Anonymous:1972:BMb

[389] Anonymous. Back matter. Isis, 63(2):301–303, June 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229100.

Anonymous:1972:FMc

[390] Anonymous. Front matter. Isis, 63(3):305–307, September 1972. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229271.

Shimao:1972:RLC

[391] Eikoh Shimao. The reception of Lavoisier’s chemistry in Japan. Isis, 63 (3):308–320, September 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229272.

Gale:1972:DCS

[392] Barry G. Gale. Darwin and the concept of a struggle for existence: A study in the extrascientific origins of scientific ideas. Isis, 63(3):321–344, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229273.

Miller:1972:MGE

[393] Arthur I. Miller. The myth of Gauss’ experiment on the Euclidean nature of physical space. Isis, 63(3):345–348, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229274. See comments [690].

Kohler:1972:OLF

[394] Robert E. Kohler, Jr. The origin of Lavoisier’s first experiments on combustion. Isis, 63(3):349–355, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229275.

Multhauf:1972:PSA

[395] Robert P. Multhauf. A premature science advisor: Jacob A. Weber (1737–1792). Isis, 63(3):356–369, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229276.

Yong:1972:BYS

[396] Lam Lay Yong. The Jih yung suan fa: An elementary arithmetic textbook of the Thirteenth Century. Isis, 63(3):370–383, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229277.

Scriba:1972:EWB

[397] Christoph J. Scriba and Bernhard Sticker. Éloge: Walter Baron, 1904– 1971. Isis, 63(3):384–387, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229278.

Cohen:1972:CCB

[398] I. Bernard Cohen. A cumulative critical bibliography of the history of science: A report to the History of Science Society. Isis, 63(3):388–392, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229279.

Sticker:1972:IHS

[399] Bernhard Sticker. The Institute for the History of Science at the University of Hamburg: The first ten years, 1960–1970. Isis, 63(3):393–394, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229280.

Hermann:1972:HST

[400] Armin Hermann. History of science and technology at the University of stuttgart. Isis, 63(3):394–395, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229281.

Wilson:1972:PNT

[401] Leonard G. Wilson. Paley and natural theology: A response to m. j. s. Hodge. Isis, 63(3):396, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229282.

Bozeman:1972:SNC

[402] Theodore Dwight Bozeman and George H. Daniels. Science and Nineteenth-Century American culture: A note on George H. Daniels’ science in the age of Jackson. Isis, 63(3):397–402, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229283.

Anonymous:1972:GSM

[403] Anonymous. The George Sarton Medal awarded to Willy Hartner August 1971. Isis, 63(3):403, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229284.

Anonymous:1972:N

[404] Anonymous. News. Isis, 63(3):404, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229285.

Crosland:1972:BRB

[405] Maurice Crosland. Book review: The Anatomy of a Scientific Institution, by Roger Hahn. The Paris Academy of Sciences, 1666–1803. Isis, 63(3): 405–407, September 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229286.

Multhauf:1972:BRB

[406] Robert P. Multhauf. Book review: Researches in the History of Pharmaceutical Chemistry. Isis, 63(3):408–412, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229287.

Bowden:1972:BRB

[407] Mary Ellen Bowden. Book review: Religion and the Decline of Magic. Isis, 63(3):412–416, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229288.

Forman:1972:BRB

[408] Paul Forman. Book review: Einstein: The Life and Times. Isis, 63 (3):417–418, September 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229289.

Schmitt:1972:BRBb

[409] Charles B. Schmitt. Book review: Bibliografia degli scritti di Eugenio Garin; L’età nuova: ricerche di storia della cultura dal XII al XVI secolo; Dal Rinascimento all’illuminismo: studi e ricerche; Science and Civic Life in the Italian Renaissance. Isis, 63(3):419–422, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229290.

Hankins:1972:BRB

[410] Thomas L. Hankins. Book review: Esquisse d’une histoire de l’idée de Nature by Robert Lenoble. Isis, 63(3):422–423, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229291.

Maziarz:1972:BRB

[411] Edward A. Maziarz. Book review: The Universe of the Mind by George E. Owen. Isis, 63(3):423–424, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229292.

Spector:1972:BRB

[412] Marshall Spector. Book review: Perception and Discovery. An Introduction to Scientific Inquiry by Norwood Russell Hanson; Willard C. Humphreys. Isis, 63(3):424–425, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229293.

Cahnman:1972:BRB

[413] Werner J. Cahnman. Book review: Ernst Cassirer: Scientific Knowledge and the Concept of Man by Seymour W. Itzkoff. Isis, 63(3):425–427, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229294.

Alexander:1972:BRB

[414] Peter Alexander. Book review: Scientific Explanation by Nicholas Rescher. Isis, 63(3):427–428, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229295.

Stuewer:1972:BRB

[415] Roger H. Stuewer. Book review: Quantum Theory and Beyond. Essays and Discussions Arising from a Colloquium by Ted Bastin. Isis, 63(3): 428–429, September 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229296.

Warner:1972:BRBb

[416] Deborah Jean Warner. Book review: The Discovery of Our Galaxy by Charles A. Whitney. Isis, 63(3):429, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229297.

Kruta:1972:BRBb

[417] Vladislav Kruta. Book review: The Battle against Heart Disease by P. E. Baldry. Isis, 63(3):429–430, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229298.

Pfetsch:1972:BRB

[418] Frank Pfetsch. Book review: Die Entwicklung der medizinischen Spezialfächer an den Universitäten des deutschen Sprachgebiets by Hans-Heinz Eulner. Isis, 63(3):430–431, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229299.

Fox:1972:BRB

[419] Samuel L. Fox. Book review: Strabismus. A History of Orthoptic Techniques by M. J. Revell. Isis, 63(3):431–432, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229300.

Pursell:1972:BRBa

[420] Carroll Pursell. Book review: Feedback Mechanisms in the Historical Collections of the National Museum of History and Technology by Otto Mayr. Isis, 63(3):433, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229301.

Sherwood:1972:BRBb

[421] Morgan Sherwood. Book review: Western Mining by Otis E. Young,; Robert Lenon. Isis, 63(3):433–434, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229302.

Jackson:1972:BRB

[422] Melvin H. Jackson. Book review: Ships and Seamanship in the Ancient World by Lionel Casson. Isis, 63(3):434–435, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229303.

Caley:1972:BRB

[423] Earle R. Caley. Book review: Theophrastus De igne. A Post-Aristotelian View of the Nature of Fire by Victor Coutant. Isis, 63(3):435–436, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229304.

McCullough:1972:BRB

[424] E. J. McCullough. Book review: Physical Science in the Middle Ages by Edward Grant. Isis, 63(3):436–437, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229305.

Freeman:1972:BRB

[425] Eric J. Freeman. Book review: The Cyrurgie of Guy de Chauliac. Volume I: Text by Margaret S. Ogden. Isis, 63(3):437–438, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229306.

Egerton:1972:BRBb

[426] Frank N. Egerton. Book review: The Compleat Naturalist: A Life of Linnaeus by Wilfrid Blunt; William T. Stearn. Isis, 63(3):438–439, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229307.

Westfall:1972:BRBa

[427] Richard S. Westfall. Book review: Introduction to Newton’s ’Principia’ by I. Bernard Cohen. Isis, 63(3):439–440, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229308.

Finn:1972:BRB

[428] Bernard S. Finn. Book review: Physico-Mechanical Experiments on Various Subjects by Francis Hauksbee. Isis, 63(3):440, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229309.

Kargon:1972:BRBc

[429] Robert H. Kargon. Book review: Galileo, Science and the Church by Jerome J. Langford. Isis, 63(3):440–441, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229310.

Ewan:1972:BRB

[430] Joseph Ewan. Book review: Reason and Experience. The Representation of Natural Order in the Work of Carl von Linné by James L. Larson. Isis, 63(3):441–442, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229311.

Westfall:1972:BRBb

[431] Richard S. Westfall. Book review: Science, Technology and Society in Seventeenth Century England by Robert K. Merton. Isis, 63(3):442–443, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229312.

Truesdell:1972:BRB

[432] C. Truesdell. Book review: Histoire des logarithmes de Neper à Euler. Tome II: La promotion des logarithmes au rang de valeur analytique by Charles Naux. Isis, 63(3):443–444, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229313.

Shea:1972:BRBb

[433] William R. Shea. Book review: Aspetti della rivoluzione scientifica by Paolo Rossi. Isis, 63(3):444–445, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229314.

Eastwood:1972:BRB

[434] Bruce S. Eastwood. Book review: Theories of Light from Descartes to Newton by A. I. Sabra. Isis, 63(3):445–446, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229315.

Carozzi:1972:BRB

[435] Albert V. Carozzi. Book review: Dissertations on Steno as Geologist by Gustav Scherz. Isis, 63(3):446–448, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229316.

Heimann:1972:BRB

[436] P. M. Heimann. Book review: Faraday as a Natural Philosopher by Joseph Agassi. Isis, 63(3):448–449, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229317.

Olson:1972:BRB

[437] Richard Olson. Book review: Archives of the Royal Society of Great Britain in Facsimile. Minutes of Manager’s Meetings, 1799–1900. Volumes I-II: 1799–1802. Isis, 63(3):449–450, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229318.

McKinney:1972:BRBa

[438] H. Lewis McKinney. Book review: Thomas Henry Huxley by Albert Ashforth. Isis, 63(3):450, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229319.

Mayr:1972:BRBe

[439] Otto Mayr. Book review: From Watt to Clausius: The Rise of Thermodynamics in the Early Industrial Age by D. S. L. Cardwell. Isis, 63 (3):451–452, September 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229320.

Goodstein:1972:BRB

[440] Judith R. Goodstein. Book review: Sir Humphry Davy’s Published Works by June Z. Fullmer. Isis, 63(3):452–453, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229321.

Koppelman:1972:BRB

[441] Elaine Koppelman. Book review: Lebesgue’s Theory of Integration. Its Origin and Development by Thomas Hawkins. Isis, 63(3):454–455, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229322.

Rosenberg:1972:BRB

[442] Charles Rosenberg. Book review: Insanity and Idiocy in Massachusetts: Report of the Commission on Lunacy, 1855 by Edward Jarvis. Isis, 63 (3):455–456, September 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229323.

Cowen:1972:BRB

[443] David L. Cowen. Book review: Homeopathy in America. The Rise and Fall of a Medical Heresy by Martin Kaufman. Isis, 63(3):456–457, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229324.

Farrall:1972:BRBa

[444] Lyndsay Farrall. Book review: The Letters of F. W. Ludwig Leichhardt by M. Aurousseau. Isis, 63(3):457–458, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229325.

Brock:1972:BRB

[445] W. H. Brock. Book review: Affinity and Matter. Elements of Chemical Philosophy 1800–1865 by Trevor H. Levere. Isis, 63(3):458–459, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229326.

Cole:1972:BRB

[446] L. E. Cole. Book review: The Psychology of Maine de Biran by F. C. T. Moore. Isis, 63(3):459–460, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229327.

Kargon:1972:BRBd

[447] Robert H. Kargon. Book review: Robert T. Gunther: A Pioneer in the History of Science, 1869–1940 by A. E. Gunther. Isis, 63(3):460, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229328.

Theodorides:1972:BRBd

[448] J. Théodoridès. Book review: Of Microbes and Life. Les microbes et la vie by Jacques Monod; Ernest Borek. Isis, 63(3):461, September 1972.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229329.

Kauffman:1972:BRBc

[449] George B. Kauffman. Book review: From Organic Chemistry to Macromolecules: A Scientific Autobiography Based on My Original Papers by Hermann Staudinger; Magda Staudinger. Isis, 63(3):461–463, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229330.

Anonymous:1972:BMc

[450] Anonymous. Back matter. Isis, 63(3):464–468, September 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229331.

Anonymous:1972:FMd

[451] Anonymous. Front matter. Isis, 63(4):469–471, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229771.

Thackray:1972:DBP

[452] Arnold Thackray and Robert K. Merton. On discipline building: The paradoxes of George Sarton. Isis, 63(4):472–495, December 1972. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229772.

Unguru:1972:WTC

[453] Sabetai Unguru. Witelo and Thirteenth-Century mathematics: An assessment of his contributions. Isis, 63(4):496–508, December 1972. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229773.

Cowan:1972:FGS

[454] Ruth Schwartz Cowan. Francis Galton’s statistical ideas: The influence of eugenics. Isis, 63(4):509–528, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229774.

Guralnick:1972:GRB

[455] Stanley M. Guralnick. Geology and religion before darwin: The case of edward hitchcock, theologian and geologist (1793–1864). Isis, 63(4): 529–543, December 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229775.

Palter:1972:EMM

[456] Robert Palter and James Hynd. Early measurements of magnetic force. Isis, 63(4):544–558, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229776.

Westfall:1972:BRBc

[457] Richard S. Westfall. Book review: Philosophiae naturalis principia mathematica. Isis, 63(4):559–560, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229777.

Home:1972:BRB

[458] R. W. Home. Book review: Coulomb and the Evolution of Physics and Engineering in Eighteenth-Century France. Isis, 63(4):560–562, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229778.

Kauffman:1972:BRBd

[459] George B. Kauffman. Book review: The History of Valency. Isis, 63 (4):563–566, December 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229779.

Thackray:1972:BRB

[460] Arnold Thackray. Book review: Science at the Cross Roads. Papers Presented to the International Congress of the History of Science and Technology Held in London from June 29 to July 3, 1931, by Delegates of the U.S.S.R. by Nikolai I. Bukharin. Isis, 63(4):566, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229780.

Kargon:1972:BRBe

[461] Robert H. Kargon. Book review: The Social and Economic Roots of Newton’s ’Principia’ by Boris Hessen. Isis, 63(4):567, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229781.

Reingold:1972:BRBb

[462] Nathan Reingold. Book review: Scientific Books, Libraries, and Collectors. A Study of Bibliography and the Book Trade in Relation to Science by John L. Thornton; R. I. J. Tully. Isis, 63(4):567–568, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229782.

Beaver:1972:BRB

[463] Donald deB. Beaver. Book review: Science: Growth and Change by Henry W. Menard. Isis, 63(4):568–569, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229783.

Buchdahl:1972:BRB

[464] Gerd Buchdahl. Book review: Heidegger, Kant and Time by Charles M. Sherover. Isis, 63(4):569–570, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229784.

North:1972:BRB

[465] J. D. North. Book review: Il Museo di Storia della Scienza a Firenze by Maria Luisa Righini Bonelli. Isis, 63(4):570–571, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229785.

Siegfried:1972:BRB

[466] Robert Siegfried. Book review: Teaching the History of Chemistry, A Symposium by George B. Kauffman. Isis, 63(4):571–572, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229786.

Heilbron:1972:BRBd

[467] J. L. Heilbron. Book review: A History of Electricity and Magnetism by Herbert W. Meyer. Isis, 63(4):572, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229787.

Churchill:1972:BRB

[468] Frederick B. Churchill. Book review: The Origins of Theoretical Population Genetics by William B. Provine. Isis, 63(4):572–574, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229788.

Clarke:1972:BRB

[469] Edwin Clarke. Book review: The Falling Sickness. A History of Epilepsy from the Greeks to the Beginnings of Modern Neurology by Owsei Temkin. Isis, 63(4):574–575, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229789.

Stocking:1972:BRBb

[470] George W. Stocking, Jr. Book review: The Genesis of British Egyptology: 1549–1906 by John David Wortham. Isis, 63(4):575–576, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229790.

Easterling:1972:BRB

[471] H. J. Easterling. Book review: Studies in Presocratic Philosophy. Volume I. The Beginnings of Philosophy by D. J. Furley; R. E. Allen. Isis, 63 (4):576–577, December 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229791.

Hansen:1972:BRB

[472] Bert Hansen. Book review: Nicole Oresme and the Kinematics of Circular Motion: “Tractatus de commensurabilitate vel incommensurabilitate motuum celi” by Edward Grant. Isis, 63(4):577–579, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229792.

Smith:1972:BRBc

[473] Emilie Savage Smith. Book review: Maurus of Salerno, Twelfth-Century “Optimus Physicus” with His Commentary on the Prognostics of Hippocrates, Now First Transcribed from Manuscript and Translated into English by Morris Harold Saffron. Isis, 63(4):579–580, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229793.

Stannard:1972:BRB

[474] Jerry Stannard. Book review: The Herb. Hashish versus Medieval Muslim Society by Franz Rosenthal. Isis, 63(4):580–581, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229794.

Mayr:1972:BRBf

[475] Otto Mayr. Book review: Bicentenary of the James Watt Patent for a Separate Condenser for the Steam Engine by Robert Donaldson. Isis, 63(4):581, December 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229795.

Westfall:1972:BRBd

[476] Richard S. Westfall. Book review: The Correspondence of Isaac Newton. Volume IV: 1694–1704 by J. F. Scott. Isis, 63(4):581–582, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229796.

Rousseau:1972:BRBb

[477] G. S. Rousseau. Book review: Samuel Johnson and the New Science by Richard B. Schwartz. Isis, 63(4):582–584, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229797.

Farrall:1972:BRBb

[478] Lyndsay Farrall. Book review: Victorian Science by George Basalla; William Coleman; Robert H. Kargon. Isis, 63(4):584–585, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229798.

Elkana:1972:BRB

[479] Yehuda Elkana. Book review: Ernst Mach. Physicist and Philosopher by Robert S. Cohen; Raymond J. Seeger. Isis, 63(4):585–586, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229799.

McKinney:1972:BRBb

[480] H. Lewis McKinney. Book review: Biology in the Nineteenth Century: Problems of Form, Function, and Transformation by William Coleman. Isis, 63(4):586–587, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229800.

Culotta:1972:BRB

[481] C. A. Culotta. Book review: The Scientific Origins of National Socialism. Social Darwinism in Ernst Haeckel and the German Monist League by Daniel Gasman. Isis, 63(4):587–588, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229801.

Richardson:1972:BRB

[482] R. Alan Richardson. Book review: Outcasts from Evolution: Scientific Attitudes of Racial Inferiority, 1859–1900 by John S. Haller. Isis, 63 (4):588–589, December 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229802.

Berg:1972:BRB

[483] John R. Berg. Book review: Purkynes Weg: Wissenschaft, Bildung und Nation by Erna Lesky. Isis, 63(4):589–590, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229803.

Stocking:1972:BRBc

[484] George W. Stocking, Jr. Book review: Daisy Bates by Elizabeth Salter. Isis, 63(4):590, December 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229804.

Pursell:1972:BRBb

[485] Carroll Pursell. Book review: The Birth of the Great Western Railway: Extracts from the Diary and Correspondence of George Henry Gibbs by Jack Simmons. Isis, 63(4):590–591, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229805.

Foster:1972:BRB

[486] W. D. Foster. Book review: Darwin’s Victorian Malady Evidence for Its Medically Induced Origin by John H. Winslow. Isis, 63(4):591–592, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229806.

Dain:1972:BRB

[487] Norman Dain. Book review: Mind, Brain and Adaptation in the Nineteenth Century: Cerebral Localization and Its Biological Context from Gall to Ferrier by Robert M. Young. Isis, 63(4):592–593, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229807.

Kargon:1972:BRBf

[488] Robert H. Kargon. Book review: Past and Present Causes in Geology by Lucien Cayeux; A. Carozzi. Isis, 63(4):593–594, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229808.

Korner:1972:BRB

[489] Stephan Körner. Book review: Development of Mathematical Logic by R. L. Goodstein. Isis, 63(4):594, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229809.

Coben:1972:BRB

[490] Stanley Coben. Book review: The American Ideology of National Science 1919–1930 by Ronald C. Tobey. Isis, 63(4):594–595, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229810.

Wortham:1972:BRB

[491] John D. Wortham. Book review: Thousands of Years: An Archaeologist’s Search for Ancient Egypt by John A. Wilson. Isis, 63(4):595–596, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229811.

Heilbron:1972:BRBe

[492] J. L. Heilbron. Book review: Perspectives in Quantum Theory. Essays in Honor of Alfred Lande by Wolfgang Yourgrau; Alwayn van der Merwe. Isis, 63(4):596, December 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229812.

Anonymous:1972:BMd

[493] Anonymous. Back matter. Isis, 63(4):597–619, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229813.

Anonymous:1972:VI

[494] Anonymous. Volume information. Isis, 63(4):600–618, December 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229770.

Anonymous:1972:FMe

[495] Anonymous. Front matter. Isis, 63(5):1–4, ???? 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229585.

Neu:1972:NSC

[496] John Neu. Ninety-seventh critical bibliography of the history of science and its cultural influences (to January 1972). Isis, 63(5):5–11+13– 205+207–233, ???? 1972. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229586.

Anonymous:1972:BMe

[497] Anonymous. Back matter. Isis, 63(5):234, ???? 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229587.

Anonymous:1973:FMa

[498] Anonymous. Front matter. Isis, 64(1):1–3, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229865.

Debus:1973:MCT

[499] Allen G. Debus. Motion in the chemical texts of the Renaissance. Isis, 64(1):4–17, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229866.

Hine:1973:MC

[500] William L. Hine. Mersenne and copernicanism. Isis, 64(1):18–32, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229867.

Kitcher:1973:FLI

[501] Philip Kitcher. Fluxions, limits, and infinite littlenesse. a study of Newton’s presentation of the calculus. Isis, 64(1):33–49, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229868.

Beardsley:1973:ASS

[502] Edward H. Beardsley. The American scientist as social activist: Franz Boas, Burt G. Wilder, and the cause of racial justice, 1900–1915. Isis, 64(1):50–66, March 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229869.

Drake:1973:MRT

[503] Stillman Drake. Medieval ratio theory vs compound medicines in the origins of Bradwardine’s Rule. Isis, 64(1):67–77, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229870.

Kauffman:1973:QNS

[504] George B. Kauffman. “Quinquevalent” nitrogen and the structure of ammonium salts: Contributions of Alfred Werner and others. Isis, 64 (1):78–95, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229871.

Reingold:1973:ERH

[505] Nathan Reingold. Éloge: Richard Harrison Shryock, 1895–1972. Isis, 64(1):96–100, March 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229872.

Rudolph:1973:FCD

[506] Richard C. Rudolph. The first carbon-14 dating in China. Isis, 64(1): 101–102, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229873.

Anonymous:1973:GSM

[507] Anonymous. The 1972 George Sarton Medal awarded to Kiyosi Yabuuti. Isis, 64(1):103–104, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229874.

Stuewer:1973:N

[508] Roger H. Stuewer. News. Isis, 64(1):105–107, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229875.

Hermelink:1973:BRB

[509] Heinrich Hermelink. Book review: The Astrological History of Masha’allah. Isis, 64(1):108–110, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229876.

Wilson:1973:BRB

[510] Curtis Wilson. Book review: Campanus of Novara and Medieval Planetary Theory, by Francis S. Benjamin,; G. J. Toomer. Theoricae planetarum. Isis, 64(1):110–112, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229877.

Truesdell:1973:BRB

[511] C. Truesdell. Book review: Die Werke von Jakob Bernoulli, by Jakob Bernoulli. Band 1. Isis, 64(1):112–114, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229878.

Larson:1973:BRBa

[512] James L. Larson. Book review: Études d’histoire et de philosophie des sciences by Georges Canguilhem. Isis, 64(1):115–116, March 1973. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229879.

Krikorian:1973:BRB

[513] A. D. Krikorian. Book review: The Arnold Arboretum: The First Century by S. B. Sutton. Isis, 64(1):116, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229880.

Gale:1973:BRB

[514] Barry G. Gale. Book review: Science as Metaphor: The Historical Role of Scientific Theories in Forming Western Culture by Richard Olson. Isis, 64(1):117–118, March 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229881.

Badash:1973:BRBa

[515] Lawrence Badash. Book review: Resources for the History of Physics by Stephen G. Brush. Isis, 64(1):118, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229882.

Olson:1973:BRB

[516] Richard Olson. Book review: The Science of Matter, A Historical Survey. Selected Readings by Maurice P. Crosland. Isis, 64(1):118–119, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229883.

Uschmann:1973:BRB

[517] G. R. Uschmann. Book review: Die zoologische Buchillustration. Band I: Bibliographie by Claus Nissen. Isis, 64(1):119–120, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229884.

Frank:1973:BRB

[518] Robert G. Frank, Jr. Book review: A History of the Royal College of Physicians of London. Volume III by A. M. Cooke. Isis, 64(1):120– 121, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229885.

Carruccio:1973:BRB

[519] Ettore Carruccio. Book review: De lineis insecabilibua by Pseudo-Aristotele; Maria Timpanaro Cardini. Isis, 64(1):121–122, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229886.

Smith:1973:BRBa

[520] Emilie Savage Smith. Book review: Safavid Medical Practice, or The Practice of Medicine, Surgery and Gynaecology in Persia between 1500 A.D. and 1750 A.D. by Cyril Elgood. Isis, 64(1):122–123, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229887.

King:1973:BRBa

[521] David A. King. Book review: Arabic Arithmetic. The Arithmetic of Abu al-Wafa^{⊃} al-Buzajani (10th Century) MSS. Or. 103 Leiden & 42 M Cairo and The Arithmetic of al-Karaji (11th Century), MS 855 Istanbul, edited by A. S. Saidan. Isis, 64(1):123–125, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229888.

Gingerich:1973:BRB

[522] Owen Gingerich. Book review: Johannes Kepler Gesammelte Werke. Band X: Tabulae Rudolphinae by Franz Hammer; Die Rudolphinischen Tafeln von Johannes Kepler by Volker Bialas. Isis, 64(1):125–126, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229889.

Norberg:1973:BRB

[523] Arthur L. Norberg. Book review: The Life of Benjamin Banneker by Silvio A. Bedini. Isis, 64(1):126–127, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229890.

King:1973:BRBb

[524] Lester S. King. Book review: Treatise of Man by René Descartes; Thomas Steele Hall. Isis, 64(1):127–128, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229891.

Egerton:1973:BRB

[525] Frank N. Egerton. Book review: The Genera of North American Plants by Thomas Nuttall. Isis, 64(1):128–129, March 1973. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229892.

North:1973:BRBa

[526] J. D. North. Book review: The Annus Mirabilis of Sir Isaac Newton 1666–1966 by Robert Palter. Isis, 64(1):129–130, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229893.

Coleman:1973:BRB

[527] William Coleman. Book review: From Beast-Machine to Man-Machine. Animal Soul in French Letters from Descartes to La Mettrie by Leonora Cohen Rosenfield. Isis, 64(1):130, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229894.

Skov:1973:BRB

[528] Niels A. Skov. Book review: The East India Company Journals of Captain William Keeling and Master Thomas Bonner, 1615–1617 by Michael Strachan; Boies Penrose. Isis, 64(1):130–131, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229895.

Morrell:1973:BRB

[529] J. B. Morrell. Book review: Sketches of the Royal Society and Royal Society Club by John Barrow. Isis, 64(1):132, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229896.

Heilbron:1973:BRBa

[530] J. L. Heilbron. Book review: The Genesis of Quantum Theory (1899– 1913) by Armin Hermann; Claude W. Nash. Isis, 64(1):132–133, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229897.

Stocking:1973:BRB

[531] George W. Stocking, Jr. Book review: Linguistic Variability and Intellectual Development by Wilhelm von Humboldt; George C. Buck; Frithjof A. Raven. Isis, 64(1):133–134, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229898.

Berg:1973:BRB

[532] John R. Berg. Book review: Autobiographia iuvenulia. Ad centesimum quinquagesimum J. G. Mendel natalem by Gregorii Mendel; Josef Sajner. Isis, 64(1):134, March 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229899.

Kopperl:1973:BRB

[533] Sheldon J. Kopperl. Book review: Molecular Reality. A Perspective on the Scientific Work of Jean Perrin by Mary Jo Nye. Isis, 64(1):135– 136, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229900.

Quen:1973:BRB

[534] Jacques M. Quen. Book review: Heads and Headlines. The Phrenological Fowlers by Madeleine B. Stern. Isis, 64(1):136–137, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229901.

Darnell:1973:BRB

[535] Regna Darnell. Book review: Ralph Linton by Adelin Linton; Charles Wagley. Isis, 64(1):137–138, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229902.

Badash:1973:BRBb

[536] Lawrence Badash. Book review: The Swift Years. The Robert Oppenheimer Story by Peter Michelmore. Isis, 64(1):138, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229903.

Warner:1973:BRB

[537] Deborah Jean Warner. Book review: The Legacy of George Ellery Hale. Evolution of Astronomy and Scientific Institutions in Pictures and Documents by Helen Wright; Joan N. Warnow; Charles Weiner. Isis, 64(1): 138–139, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229904.

Anonymous:1973:BMa

[538] Anonymous. Back matter. Isis, 64(1):140–146, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229905.

Anonymous:1973:FMb

[539] Anonymous. Front matter. Isis, 64(2):147–149, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229594.

Forman:1973:SIW

[540] Paul Forman. Scientific internationalism and the Weimar physicists: The ideology and its manipulation in Germany after World War I. Isis, 64(2): 150–180, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229595.

Kohler:1973:ETO

[541] Robert E. Kohler, Jr. The enzyme theory and the origin of biochemistry. Isis, 64(2):181–196, June 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229596.

White:1973:HGM

[542] George W. White. The history of geology and mineralogy as seen by American writers, 1803–1835: A bibliographic essay. Isis, 64(2):197– 214, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229597.

Wells:1973:WCW

[543] Kentwood D. Wells. William Charles Wells and the races of Man. Isis, 64(2):215–225, June 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229598.

Debus:1973:FNP

[544] Allen G. Debus. A further note on palingenesis: The account of ebenezer sibly in the illustration of astrology (1792). Isis, 64(2):226–230, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229599.

Smeaton:1973:BEC

[545] W. A. Smeaton and J. A. Schufle. Bergman’s “equivalents”: A correction. Isis, 64(2):231–232, June 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229600.

Tee:1973:SHR

[546] Garry J. Tee. On sami Hamarneh’s review of der diabetestraktat ’abd al-Lat.if al-Baġdādī’s. Isis, 64(2):232, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229601.

Anonymous:1973:Na

[547] Anonymous. News. Isis, 64(2):233–235, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229602.

Reiche:1973:BRB

[548] Harald A. T. Reiche. Book review: Die Goldhörner von Gallehus, by Willy Hartner. Die Inschriften, die ikonographischen und literarischen Beziehungen, das Entstehungsdatum. Isis, 64(2):236–239, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229603.

Swerdlow:1973:BRB

[549] N. Swerdlow. Book review: Das heliozentrische System in der griechischen, persischen und indischen Astronomie. Isis, 64(2):239–243, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229604.

Kangro:1973:BRB

[550] Hans Kangro. Book review: Actes du XIIe Congrès International d’Histoire des Sciences, Paris, 1968, Tome I (A): Colloques, Textes des Rapports. Isis, 64(2):244–245, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229605.

Vernet:1973:BRB

[551] J. Vernet. Book review: Scienza e tecnica by Renzo Elisa Gavioli; Maria Elisa Gavioli. Isis, 64(2):245, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229606.

Neu:1973:BRB

[552] John Neu. Book review: A Regional Union Catalog of Manuscripts Relating to the History of Science and Technology Located in Indiana, Michigan, and Ohio by Edwin T. Layton, Jr. Isis, 64(2):245–246, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229607.

Larson:1973:BRBb

[553] James L. Larson. Book review: The Order of Things. An Archaeology of the Human Sciences by Michel Foucault. Isis, 64(2):246–247, June 1973.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229608.

Manier:1973:BRB

[554] Edward Manier. Book review: Interpretations of Life and Mind. Essays around the Problem of Reduction by Marjorie Grene. Isis, 64(2):248– 249, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229609.

Reingold:1973:BRB

[555] Nathan Reingold. Book review: Scientific Knowledge and Its Social Problems by Jerome R. Ravetz. Isis, 64(2):249, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229610.

Eklund:1973:BRB

[556] Jon Eklund. Book review: History of Hydrology by Asit K. Biswas. Isis, 64(2):250–251, June 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229611.

Heilbron:1973:BRBb

[557] J. L. Heilbron. Book review: Concepts of Space. The History of Theories of Space in Physics by Max Jammer. Isis, 64(2):251, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229612.

Koertge:1973:BRB

[558] Noretta Koertge. Book review: The Evolution of the Atomic Theory by David P. Mellor. Isis, 64(2):251–252, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229613.

Multhauf:1973:BRBa

[559] Robert P. Multhauf. Book review: A History of Chemistry. Volume I, Part 1: Theoretical Background by J. R. Partington. Isis, 64(2):252– 253, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229614.

Cowan:1973:BRB

[560] Ruth Schwartz Cowan. Book review: La logique du vivant. Une histoire de l’hérédité by François Jacob. Isis, 64(2):253–254, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229615.

Mora:1973:BRB

[561] George Mora. Book review: Cent ans de psychiatrie. Essai sur l’histoire des institutions psychiatriques en France de 1870 à nos jours by Henri Vermorel; André Meylan. Isis, 64(2):255–256, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229616.

Risse:1973:BRB

[562] Guenter B. Risse. Book review: Zur Typologie des Irrenhauses in Frankreich und Deutschland (1780–1840) by Dieter Jetter. Isis, 64(2): 256–257, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229617.

Wasserstein:1973:BRB

[563] A. Wasserstein. Book review: Das Mondgesicht. (De facie in orbe lunae) by Plutarch; Herwig Görgemanns; Untersuchungen zu Plutarchs Dialog De facie in orbe lunae by Herwig Gorgemanns. Isis, 64(2):257–258, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229618.

Smith:1973:BRBb

[564] Emilie Savage Smith. Book review: Anatomie und Chirurgie des Schädels, insbesondere der Hals-, Nasen- und Ohrenkrank- heiten nach Ibn al-Quff by Otto Spies; Horst Müller–Bütow. Isis, 64(2):258–259, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229619.

Stannard:1973:BRB

[565] Jerry Stannard. Book review: Plant Myths and Traditions in India by Shakti M. Gupta. Isis, 64(2):259–260, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229620.

North:1973:BRBb

[566] J. D. North. Book review: The Almanus Manuscript. Staats- und Stadtbibliothek Augsburg, Codex in 2 degrees No. 209. Rome circa 1475– circa 1485 by J. H. Leopold. Isis, 64(2):260–261, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229621.

Schiller:1973:BRB

[567] Joseph Schiller. Book review: Lettres à M. l’Abbé Spallanzani by Charles Bonnet; Carlo Castellani. Isis, 64(2):261–262, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229622.

Stuckey:1973:BRB

[568] Ronald L. Stuckey. Book review: Selectarum Stirpium Americanarum Historia by Nicolai Josephi Jacquin. Isis, 64(2):263, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229623.

Hahn:1973:BRBa

[569] Roger Hahn. Book review: The Modernity of the Eighteenth Century by Louis T. Milic. Isis, 64(2):264–265, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229624.

Glick:1973:BRB

[570] Thomas F. Glick. Book review: El botánico José Quer (1695–1764), primer apologista de la ciencia española by Ricardo Pascual. Isis, 64(2): 265, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229625.

Mahoney:1973:BRB

[571] Michael S. Mahoney. Book review: Studien zur Mathematik des John Wallis (1616–1703) by Christoph J. Scriba. Isis, 64(2):265–266, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229626.

Scriba:1973:BRB

[572] Christoph J. Scriba. Book review: Bibliographia Lambertiana. Ein Führer durch das gedruckte und ungedruckte Schrifttum und den wissenschaftlichen Briefwechsel von Johann Heinrich Lambert, 1728–1777 by Max Steck. Isis, 64(2):266–268, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229627.

Orvell:1973:BRB

[573] Betty Orvell. Book review: Le physicien Ignace Gaston Pardies S.J. (1636–1673) by August Ziggelaar. Isis, 64(2):268–269, June 1973. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229628.

Jay:1973:BRB

[574] Martin Jay. Book review: Scholarship and Partisanship: Essays on Max Weber by Reinhard Bendix; Guenther Roth. Isis, 64(2):269–271, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229629.

Pierson:1973:BRB

[575] Stuart Pierson. Book review: William Whewell’s Theory of Scientific Method by Robert E. Butts. Isis, 64(2):271–272, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229630.

Turner:1973:BRB

[576] R. Steven Turner. Book review: T. H. Huxley on Education. A Selection from His Writings by T. H. Huxley; Cyril Bibby. Isis, 64(2):272, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229631.

Rothblatt:1973:BRB

[577] Sheldon Rothblatt. Book review: History, Man, and Reason: A Study in Nineteenth-Century Thought by Maurice Mandelbaum. Isis, 64(2):273– 274, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229632.

Kirby:1973:BRB

[578] H. W. Kirby. Book review: The Discovery of Radioactivity and Transmutation by Alfred Romer; Radiochemistry and the Discovery of Isotopes by Alfred Romer. Isis, 64(2):274–275, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229633.

Swenson:1973:BRBa

[579] Loyd S. Swenson, Jr. Book review: Nineteenth-Century Aether Theories by Kenneth F. Schaffner. Isis, 64(2):275–276, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229634.

Levy:1973:BRB

[580] Edwin Levy. Book review: Science and Philosophy in the Soviet Union by Loren R. Graham. Isis, 64(2):276–278, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229635.

Swenson:1973:BRBb

[581] Loyd S. Swenson, Jr. Book review: Adventures in Research: A History of Ames Research Center, 1940–1965 by Edwin P. Hartman. Isis, 64(2): 278, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229636.

Trenn:1973:BRB

[582] Thaddeus J. Trenn. Book review: Rutherford: Recollections of the Cambridge Days by Mark Oliphant. Isis, 64(2):278–279, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229637.

Quester:1973:BRB

[583] George H. Quester. Book review: Nuclear Politics: The British Experience with an Independent Strategic Force 1939–1970 by Andrew J. Pierre. Isis, 64(2):280, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229638.

Hughes:1973:BRB

[584] Thomas Parke Hughes. Book review: Edward Goodrich Acheson: Inventor, Scientist, Industrialist by Raymond Szymanowitz. Isis, 64(2): 281–282, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229639.

Anonymous:1973:BMb

[585] Anonymous. Back matter. Isis, 64(2):283–286, June 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229640.

Anonymous:1973:FMc

[586] Anonymous. Front matter. Isis, 64(3):287–289, September 1973. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229717.

Drake:1973:GEC

[587] Stillman Drake. Galileo’s experimental confirmation of horizontal inertia: Unpublished manuscripts (galileo gleanings xxii). Isis, 64(3):290–305, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229718.

Crosland:1973:LTA

[588] Maurice Crosland. Lavoisier’s theory of acidity. Isis, 64(3):306–325, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229719.

Hendrickson:1973:SCA

[589] Walter B. Hendrickson. Science and culture in the American middle west. Isis, 64(3):326–340, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229720.

Coleman:1973:LRT

[590] William Coleman. Limits of the recapitulation theory: Carl Friedrich Kielmeyer’s critique of the presumed parallelism of Earth history, ontogeny, and the present order of organisms. Isis, 64(3):341–350, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229721.

Toomer:1973:PJT

[591] G. J. Toomer. Prophatius Judaeus and the Toledan Tables. Isis, 64 (3):351–355, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229722.

Iltis:1973:DCM

[592] Carolyn Iltis. The decline of Cartesianism in mechanics: The Leibnizian– Cartesian debates. Isis, 64(3):356–373, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229723.

MacLachlan:1973:TIE

[593] James MacLachlan. A test of an “imaginary” experiment of Galileo’s. Isis, 64(3):374–379, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229724.

Drabkin:1973:ELL

[594] Miriam Drabkin. Éloge: Lancelot Law Whyte, 1896–1972. Isis, 64(3): 380–381, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229725.

Watson:1973:PJD

[595] Andrew G. Watson. The provenance of John Dee’s manuscript of the De Superficierum Divisionibus of Machometus Bagdedinus. Isis, 64(3):

382–383, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229726.

See [291].

Schwarz:1973:AES

[596] Michael Schwarz. The affirmation of empty space by an Eleventh-Century Muctazilite. Isis, 64(3):384–385, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229727.

Parascandola:1973:LE

[597] John Parascandola, Chauncey D. Leake, Hans Georg Gundel, Martin Plessner, and W. D. M. Paton. Letters to the Editor. Isis, 64(3):385–386, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229728.

Anonymous:1973:CAR

[598] Anonymous. Correction: Archives of the Royal Society of great Britain in facsimile. minutes of Manager’s meetings, 1799–1900. volumes i-ii: 1799– 1802. Isis, 64(3):386, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229729.

Anonymous:1973:Nb

[599] Anonymous. News. Isis, 64(3):387–388, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229730.

Kohler:1973:BRB

[600] Robert E. Kohler, Jr. Book review: Molecules and Life, by Joseph S. Fruton. Historical Essays on the Interplay of Chemistry and Biology. Isis, 64(3):389–390, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229731.

Clagett:1973:BRB

[601] Marshall Clagett. Book review: Études d’histoire des sciences et de la philosophie du Moyen Age. Isis, 64(3):390–393, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229732.

Rudwick:1973:BRB

[602] Martin J. S. Rudwick. Book review: Charles Lyell, by Leonard G. Wilson. The Years to 1841: The Revolution in Geology. Isis, 64(3):394–395, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229733.

Fruchtbaum:1973:BRB

[603] Harold Fruchtbaum. Book review: Perspectives in the History of Science and Technology by Duane H. D. Roller. Isis, 64(3):395–398, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229734.

Polonoff:1973:BRB

[604] Irving Polonoff. Book review: Metaphysics and the Philosophy of Science. The Classical Origins: Descartes to Kant by Gerd Buchdahl. Isis, 64(3): 398–399, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229735.

delCarril:1973:BRB

[605] Mario F. del Carril. Book review: Criticism and the Growth of Knowledge by Imre Lakatos; Alan Musgrave. Isis, 64(3):399–400, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229736.

Stieb:1973:BRBa

[606] Ernst W. Stieb. Book review: Temples of the Muses and a History of Pharmacy Museums by Sami K. Hamarneh. Isis, 64(3):401, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229737.

Vachon:1973:BRB

[607] P. David Vachon. Book review: Invisible Colleges: Diffusion of Knowledge in Scientific Communities by Diana Crane. Isis, 64(3):401–402, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229738.

Coxeter:1973:BRB

[608] H. S. M. Coxeter. Book review: The Secrets of Ancient Geometry — And Its Uses by Tons Brunés; Charles M. Napier. Isis, 64(3):402–404, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229739.

Longcore:1973:BRB

[609] Randall Longcore. Book review: A Survey of Mathematics, Elementary Concepts and Their Historical Development by Vivian Shaw Groza.

Isis, 64(3):404–405, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229740.

Rossiter:1973:BRB

[610] Margaret Rossiter. Book review: Crop Nutrition: Science and Practice before Liebig by G. E. Fussell. Isis, 64(3):405–406, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229741.

Gregory:1973:BRB

[611] Richard L. Gregory. Book review: Selective History of Theories of Visual Perception 1650–1950 by Nicholas Pastore. Isis, 64(3):406–408, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229742.

Schlesinger:1973:BRB

[612] R. Walter Schlesinger. Book review: A History of Poliomyelitis by John R. Paul. Isis, 64(3):408–409, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229743.

Rae:1973:BRB

[613] John B. Rae. Book review: A Biographical Dictionary of American Civil Engineers by American Society of Civil Engineers. Isis, 64(3):409–410, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229744.

Africa:1973:BRB

[614] Thomas W. Africa. Book review: Technology in the Ancient World by Henry Hodges. Isis, 64(3):410–411, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229745.

Susskind:1973:BRB

[615] Charles Süsskind. Book review: Bibliography of the History of Electronics by George Shiers. Isis, 64(3):411–412, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229746.

Balme:1973:BRB

[616] D. M. Balme. Book review: The Pseudo-Hippocratic Tract Peri Hebdomadon Chapters 1–11 and Greek Philosophy by J. Mansfeld. Isis, 64 (3):412–413, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229747.

Libbrecht:1973:BRB

[617] U. Libbrecht. Book review: Science and Civilisation in China. Volume IV: Physics and Physical Technology. Part III: Civil Engineering and Nautics by Joseph Needham; Wang Ling; Lu Gwei-dien. Isis, 64(3): 413–416, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229748.

Sivin:1973:BRBa

[618] N. Sivin. Book review: Clerks and Craftsmen in China and the West. Lectures and Addresses on the History of Science and Technology by Joseph Needham. Isis, 64(3):417–418, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229749.

Schmitt:1973:BRB

[619] Charles B. Schmitt. Book review: De ratione cognoscendi sacras literas. Über den Erkenntnisgrund der heiligen Schriften by Matthias Flacius Illvricus; Lutz Geldsetzer. Isis, 64(3):419, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229750.

Wallace:1973:BRB

[620] William A. Wallace. Book review: A Critical Survey and Bibliography of Studies on Renaissance Aristotelianism 1958–1969 by Charles B. Schmitt. Isis, 64(3):419–420, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229751.

Webster:1973:BRB

[621] C. Webster. Book review: Paracelsus Studien by Ferdinand Weinhandl; Sepp Domandl; Erziehung und Menschenbild bei Paracelsus. Anfänge einer verantwortungsbewussten Pädagogik by Sepp Domandl. Isis, 64 (3):420, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229752.

Multhauf:1973:BRBb

[622] Robert P. Multhauf. Book review: Les débuts de l’industrie chimique dans les Pays-Bas autrichiens by Annette André-Félix. Isis, 64(3):420– 421, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229753.

Price:1973:BRB

[623] Derek de Solla Price. Book review: Scientific Instruments of the 17th and 18th Centuries by Maurice Daumas; Mary Holbrook. Isis, 64(3): 421–422, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229754.

Brown:1973:BRB

[624] Theodore M. Brown. Book review: Science and Education in the Seventeenth Century: The Webster-Ward Debate by Allen G. Debus. Isis, 64 (3):422–424, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229755.

Hankins:1973:BRB

[625] Thomas L. Hankins. Book review: Joseph Fourier 1768–1830. A Survey of His Life and Work, Based on a Critical Edition of His Monograph on the Propagation of Heat, Presented to the Institut de France in 1807 by I. Grattan-Guinness; J. R. Ravetz. Isis, 64(3):424–425, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229756.

Jaouiche:1973:BRB

[626] K. Jaouiche. Book review: The Construction of Modern Science. Mechanisms and Mechanics by Richard S. Westfall. Isis, 64(3):425–426, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229757.

Casada:1973:BRB

[627] James A. Casada. Book review: Africanus Horton, 1835–1883: West African Scientist and Patriot by Christopher Fyfe. Isis, 64(3):426–427, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229758.

Sinclair:1973:BRB

[628] Bruce Sinclair. Book review: Elmer Sperry: Inventor and Engineer by Thomas Parke Hughes. Isis, 64(3):427, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229759.

Levine:1973:BRB

[629] Donald N. Levine. Book review: Emile Durkheim: Sociologist and Philosopher by Dominick LaCapra. Isis, 64(3):427–429, September 1973.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229760.

Geison:1973:BRB

[630] Gerald L. Geison. Book review: Addison and the White Corpuscles: An Aspect of Nineteenth-Century Biology by L. J. Rather. Isis, 64(3): 429–430, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229761.

deMoulin:1973:BRB

[631] D. de Moulin. Book review: Jaques-Louis Reverdin 1842–1929. Un chirurgien à l’aube d’une ère nouvelle by Henri Reverdin. Isis, 64(3): 430–431, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229762.

Bromberg:1973:BRB

[632] Joan Bromberg. Book review: The Ethereal Aether. A History of the Michelson–Morley-Miller Aether-Drift Experiments, 1880–1930 by Loyd S. Swenson. Isis, 64(3):431–432, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229763.

Stent:1973:BRB

[633] Gunther S. Stent. Book review: Challenging Biological Problems: Directions toward Their Solution by John A. Behnke; Biology, History, and Natural Philosophy by Allen D. Breck; Wolfgang Yourgrau. Isis, 64(3): 432–433, September 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229764.

Lundeberg:1973:BRB

[634] Philip K. Lundeberg. Book review: Sailors, Scientists and Rockets: Origins of the Navy Rocket Program and of the Naval Ordnance Test Station, Inyokern by Albert B. Christman. Isis, 64(3):433–435, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229765.

Brieger:1973:BRB

[635] Gert H. Brieger. Book review: Surgeon to Soldiers. Diary and Records of the Surgical Consultant Allied Force Headquarters, World War II by Edward D. Churchill. Isis, 64(3):435, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229766.

Klein:1973:BRB

[636] Martin J. Klein. Book review: Correspondance 1903–1955 by Albert Einstein; Michele Besso; Pierre Speziali. Isis, 64(3):435–436, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229767.

Hahn:1973:BRBb

[637] Roger Hahn. Book review: Les Prix Nobel en 1970 by The Nobel Foundation. Isis, 64(3):436–437, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229768.

Anonymous:1973:BMc

[638] Anonymous. Back matter. Isis, 64(3):438–440, September 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229769.

Anonymous:1973:FMd

[639] Anonymous. Front matter. Isis, 64(4):441–443, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229642.

Bynum:1973:AMN

[640] William F. Bynum. The anatomical method, natural theology, and the functions of the brain. Isis, 64(4):444–468, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229643.

Chalmers:1973:LME

[641] A. F. Chalmers. The limitations of Maxwell’s electromagnetic theory. Isis, 64(4):469–483, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229644.

Rogers:1973:RDO

[642] James Allen Rogers. The reception of Darwin’s origin of species by Russian scientists. Isis, 64(4):484–503, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229645.

Osler:1973:GME

[643] Margaret J. Osler. Galileo, motion, and essences. Isis, 64(4):504–509, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229646.

Jayawardene:1973:IPA

[644] S. A. Jayawardene and Di Rafael Bombelli. The influence of practical arithmetics on the algebra of Rafael Bombelli. Isis, 64(4):510–523, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229647.

Anonymous:1973:Nc

[645] Anonymous. News. Isis, 64(4):524–526, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229648.

Hall:1973:BRBa

[646] Bert S. Hall. Book review: Greek and Roman Artillery: Historical Development. Isis, 64(4):527–533, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229649.

Guggenbuhl:1973:BRB

[647] Laura Guggenbuhl. Book review: Mathematics in the Time of the Pharaohs. Isis, 64(4):533–534, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229650.

Sivin:1973:BRBb

[648] N. Sivin. Book review: An Annotated Bibliography of Selected Chinese Reference Works. Isis, 64(4):534–539, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229651.

Cannon:1973:BRB

[649] W. F. Cannon. Book review: Natural Science Books in English 1600– 1900 by David M. Knight. Isis, 64(4):540, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229652.

Bowman:1973:BRB

[650] Peter A. Bowman. Book review: Paradigms and Paradoxes. The Philosophical Challenge of the Quantum Domain by Robert G. Colodny. Isis, 64(4):540–541, December 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229653.

Koslow:1973:BRB

[651] Arnold Koslow. Book review: The Principles of Scientific Thinking by Rom Harré. Isis, 64(4):541–542, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229654.

Arons:1973:BRB

[652] Arnold Arons. Book review: History in the Teaching of Physics by Stephen G. Brush; Allen L. King. Isis, 64(4):542–543, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229655.

Jaki:1973:BRB

[653] Stanley L. Jaki. Book review: Science as a Cultural Process by Maurice N. Richter. Isis, 64(4):544, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229656.

Folkerts:1973:BRB

[654] Menso Folkerts. Book review: Geschichte der Mathematik. Probleme und Gestalten by Gerhard Kropp. Isis, 64(4):545–546, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229657.

Pounds:1973:BRB

[655] Norman J. G. Pounds. Book review: Maps and Man. An Examination of Cartography in Relation to Culture and Civilization by Norman J. W. Thrower. Isis, 64(4):546–547, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229658.

Hall:1973:BRBb

[656] Diana Long Hall. Book review: Machina Carnis. The Biochemistry of Muscular Contraction in Its Historical Development by Dorothy M. Needham. Isis, 64(4):547–548, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229659.

Stieb:1973:BRBb

[657] Ernst W. Stieb. Book review: Glass and British Pharmacy, 1699–1900; A Survey, and Guide to the Wellcome Collection of British Glass by J.

K. Crellin; J. R. Scott. Isis, 64(4):548–549, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229660.

Stearns:1973:BRB

[658] Peter N. Stearns. Book review: Technology and Culture: An Anthology by Melvin Kranzberg; William H. Davenport. Isis, 64(4):549–550, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229661.

Thomas:1973:BRB

[659] Phillip Drennon Thomas. Book review: Chaucer’s Physician. Medicine and Literature in Fourteenth Century England by Huling E. Ussery. Isis, 64(4):550–551, December 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229662.

Hilfstein:1973:BRB

[660] Erna Hilfstein and Edward Rosen. Book review: Nowe materialy do dzialalno’sci publicznej Mikolaja Kopernika z lat 1512–1537 by Marian Biskup. Isis, 64(4):551, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229663.

Farley:1973:BRB

[661] John Farley. Book review: Letter from G. Bidloo to Anthony van Leeuwenhoek about the Animals Which are Sometimes Found in the Liver of Sheep and Other Beasts by J. Jansen. Isis, 64(4):552, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229664.

Waff:1973:BRB

[662] Craig B. Waff. Book review: The Euler-Mayer Correspondence (1751– 1755): A New Perspective on Eighteenth Century Advances in the Lunar Theory by Eric G. Forbes. Isis, 64(4):552–554, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229665.

Shea:1973:BRB

[663] William R. Shea. Book review: I filosofi e i selvaggi 1580–1780 by Sergio Landucci. Isis, 64(4):554–555, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229666.

Hindle:1973:BRB

[664] Brooke Hindle. Book review: The Scientific Papers of James Logan by Roy N. Lokken. Isis, 64(4):555–556, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229667.

Godechot:1973:BRB

[665] Jacques Godechot. Book review: Il tramonto dell’illuminismo. Filosofia e politica nella società francese (1770–1810) by Sergio Moravia. Isis, 64 (4):556–558, December 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229668.

Hall:1973:BRBc

[666] Thomas S. Hall. Book review: William Harvey and the Circulation of the Blood by Gweneth Whitteridge. Isis, 64(4):558–559, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229669.

Blake:1973:BRB

[667] John B. Blake. Book review: Medical America in the Nineteenth Century. Readings from the Literature by Gert H. Brieger. Isis, 64(4):559–560, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229670.

Multhauf:1973:BRBc

[668] Robert P. Multhauf. Book review: The American Mineralogical Journal. Volume 1, New York, 1810–1814 by Archibald Bruce; The Monthly American Journal of Geology and Natural Science. Volume 1, Philadelphia, 1831–1832 by G. W. Featherstonhaugh. Isis, 64(4):560, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229671.

Burstyn:1973:BRB

[669] Harold L. Burstyn. Book review: The Harvard College Observatory: The First Four Directorships, 1839–1919 by Bessie Zaban Jones; Lyle Gifford Boyd; A History of Nautical Astronomy by Charles H. Cotter. Isis, 64 (4):560–563, December 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229672.

Brazier:1973:BRB

[670] Mary A. B. Brazier. Book review: John Hunter, A. A. Berthold and the Origins of Endocrinology by C. Barker Jørgensen. Isis, 64(4):563–565, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229673.

Frangsmyr:1973:BRB

[671] Tore Frängsmyr. Book review: Israel Hwasser by Sven-Eric Liedman. Isis, 64(4):565, December 1973. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229674.

Vorzimmer:1973:BRB

[672] Peter J. Vorzimmer. Book review: Wallace and Natural Selection by H. Lewis McKinney. Isis, 64(4):565–566, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229675.

King:1973:BRBc

[673] Lester S. King. Book review: American Physicians in the 19th Century: From Sects to Science by William G. Rothstein. Isis, 64(4):567–568, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229676.

Haynes:1973:BRB

[674] R. H. Haynes. Book review: Microbiology and the Spontaneous Generation Debate during the 1870’s by Glenn Vandervliet. Isis, 64(4):568–569, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229677.

Beltz:1973:BRB

[675] John Stuart Beltz. Book review: The Kremlin and the Cosmos by Nicholas Daniloff. Isis, 64(4):569–570, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229678.

Bannister:1973:BRB

[676] Robert C. Bannister. Book review: Chicago Sociology, 1920–1932 by Robert E. L. Faris. Isis, 64(4):570–571, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229679.

Thoresen:1973:BRB

[677] Timothy H. H. Thoresen. Book review: Robert H. Lowie by Robert F. Murphy. Isis, 64(4):571–572, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229680.

Heilbron:1973:BRBc

[678] J. L. Heilbron. Book review: Physics in the Twentieth Century. Selected Essays by Victor F. Weisskopf. Isis, 64(4):572–573, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229681.

Anonymous:1973:BMd

[679] Anonymous. Back matter. Isis, 64(4):574–594, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229682.

Anonymous:1973:VI

[680] Anonymous. Volume information. Isis, 64(4):577–593, December 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229641.

Anonymous:1973:FMe

[681] Anonymous. Front matter. Isis, 64(5):1–4, ???? 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229938.

Neu:1973:NEC

[682] John Neu. Ninety-eighth critical bibliography of the history of science and its cultural influences (to January 1973). Isis, 64(5):5–173+175– 196, ???? 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229939.

Anonymous:1973:BMe

[683] Anonymous. Back matter. Isis, 64(5):??, ???? 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229940.

Anonymous:1974:FMa

[684] Anonymous. Front matter. Isis, 65(1):1–3, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228878.

Hutchinson:1974:ATN

[685] G. Evelyn Hutchinson. Attitudes toward nature in Medieval England: The Alphonso and Bird psalters. Isis, 65(1):5–37, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228879.

VanHelden:1974:TSC

[686] Albert Van Helden. The telescope in the seventeenth century. Isis, 65(1): 38–58, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228880.

LeGrand:1974:DCF

[687] H. E. Le Grand. Determination of the composition of the fixed alkalis 1789–1810. Isis, 65(1):59–65, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228881.

Finocchiaro:1974:NTR

[688] Maurice A. Finocchiaro. Newton’s third rule of philosophizing: A role for logic in historiography. Isis, 65(1):66–73, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228882.

Trenn:1974:GMS

[689] Thaddeus J. Trenn, Hans Geiger, Ernest Marsden, and E. Rutherford. The Geiger–Marsden scattering results and Rutherford’s atom, July 1912 to July 1913: The shifting significance of scientific evidence. Isis, 65(1): 74–82, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228883.

Goe:1974:CMB

[690] George Goe, B. L. van der Waerden, and Arthur I. Miller. Comments on Miller’s “The Myth of Gauss’ Experiment on the Euclidean Nature of Physical Space”. Isis, 65(1):83–87, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228884. See [393].

Anonymous:1974:Na

[691] Anonymous. News. Isis, 65(1):88–91, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228885.

Burstyn:1974:BRBa

[692] Harold L. Burstyn. Book review: Scientists and the Sea 1650–1900, by Margaret Deacon. A Study of Marine Science. Isis, 65(1):92–95, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228886.

Stocking:1974:BRBa

[693] George W. Stocking, Jr. Book review: Blackberry Winter: My Earlier Years. Isis, 65(1):95–97, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228887.

Hilfstein:1974:BRB

[694] Erna Hilfstein and Edward Rosen. Book review: Etudes d’histoire des sciences en Pologne by Aleksander Birkenmajer. Isis, 65(1):98–99, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228888.

Coleman:1974:BRBa

[695] William Coleman. Book review: Comptes rendus du 93e Congrès national des sociétés savantes. Tours, 1968. Section des sciences. Tome II: Histoire des sciences. Isis, 65(1):99, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228889.

Weisheipl:1974:BRBa

[696] James A. Weisheipl. Book review: Causality and Scientific Explanation. Volume I: Medieval and Early Classical Science by William A. Wallace. Isis, 65(1):99–100, March 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228890.

Provine:1974:BRB

[697] William Provine. Book review: Genetics and American Society: A Historical Appraisal by Kenneth M. Ludmerer. Isis, 65(1):101–102, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228891.

deCamp:1974:BRB

[698] L. Sprague de Camp. Book review: Technology and Social Change by Wilbert E. Moore. Isis, 65(1):102–103, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228892.

Hesse:1974:BRBa

[699] Mary Hesse. Book review: The Alienation of Reason. A History of Positivist Thought by Leszek Kolakowski; Norbert Guterman. Isis, 65(1): 103–104, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228893.

Scriba:1974:BRBa

[700] Christoph J. Scriba. Book review: Éléments d’histoire des mathématiques by Nicolas Bourbaki. Isis, 65(1):104, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228894.

Boyer:1974:BRB

[701] Carl B. Boyer. Book review: Mathematical Thought from Ancient to Modern Times by Morris Kline. Isis, 65(1):104–106, March 1974. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228895.

Trenn:1974:BRB

[702] Thaddeus J. Trenn. Book review: Lexikon Geschichte der Physik A-Z. Biographien, Sachwörter, Originalschriften und Sekundärliteratur by Armin Hermann. Isis, 65(1):106–107, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228896.

Holtzman:1974:BRB

[703] Stephen F. Holtzman. Book review: Climbing Man’s Family Tree. A Collection of Major Writings on Human Phylogeny, 1699 to 1971 by Theodore D. McCown; Kenneth A. R. Kennedy. Isis, 65(1):107–108, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228897.

Hughes:1974:BRB

[704] Thomas Parke Hughes. Book review: Turning Points in Western Technology: A Study of Technology, Science, and History by D. S. L. Cardwell. Isis, 65(1):108–110, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228898.

vanderWaerden:1974:BRBa

[705] B. L. van der Waerden. Book review: Demotic Mathematical Papyri by Richard A. Parker. Isis, 65(1):110–111, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228899.

Steiner:1974:BRB

[706] Grundy Steiner. Book review: The Roman Land Surveyors: An Introduction to the Agrimensores by O. A. W. Dilke. Isis, 65(1):111–112, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228900.

Mueller:1974:BRB

[707] Ian Mueller. Book review: Grundprobleme der Geschichte der antiken Wissenschaft by Kurt von Fritz. Isis, 65(1):112–114, March 1974. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228901.

Eastwood:1974:BRB

[708] Bruce S. Eastwood. Book review: Tractatus de perspectiva by John Pecham; David C. Lindberg. Isis, 65(1):114–115, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228902.

Pingree:1974:BRB

[709] David Pingree. Book review: A Check-list of Sanskrit Medical Manuscripts in India by B. Rama Rao. Isis, 65(1):115–116, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228903.

Rose:1974:BRB

[710] Paul Lawrence Rose. Book review: Mariano Taccola and His Book “De Ingeneis” by Frank D. Prager; Gustina Scaglia. Isis, 65(1):116, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228904.

Figala:1974:BRB

[711] Karin Figala and Joachim O. Fleckenstein. Book review: The Occult Sciences in the Renaissance. A Study in Intellectual Patterns by Wayne Shumaker. Isis, 65(1):116–118, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228905.

Albury:1974:BRB

[712] William R. Albury. Book review: An Essay on the Origin of Human Knowledge, Being a Supplement to Mr. Locke’s Essay on the Human Understanding by Etienne Bonnot de Condillac; Thomas Nugent. Isis, 65(1):118–119, March 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228906.

Shea:1974:BRB

[713] William R. Shea. Book review: Scritti di idraulica fluviale e di canalizzazione by Paolo Frisi; Cristiana Fischer. Isis, 65(1):119, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228907.

Ewan:1974:BRBa

[714] Joseph Ewan. Book review: Linnaeus, A Modern Portrait of the Great Swedish Scientist by Heinz Goerke. Isis, 65(1):119–120, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228908.

Brieger:1974:BRB

[715] Gert H. Brieger. Book review: Benjamin Rush: Revolutionary Gadfly by David Freeman Hawke. Isis, 65(1):120, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228909.

Scriba:1974:BRBb

[716] Christoph J. Scriba. Book review: The Mathematical Principles underlying Newton’s ‘Principia Mathematica,’ Being the Ninth Gibson Lecture in the History of Mathematics Delivered within the University of Glasgow by D. T. Whiteside. Isis, 65(1):121, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228910.

Brittain:1974:BRB

[717] James Brittain. Book review: Selections from the Scientific Correspondence of Elihu Thomson by Harold J. Abrahams; Marion B. Savin. Isis, 65(1):121–122, March 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228911.

Steffens:1974:BRB

[718] Henry John Steffens. Book review: Ludvig Colding and the Conservation of Energy Principle by Per F. Dahl. Isis, 65(1):122–124, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228912.

Burstyn:1974:BRBb

[719] Joan N. Burstyn. Book review: G. Stanley Hall: The Psychologist as Prophet by Dorothy Ross. Isis, 65(1):124–125, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228913.

Friis:1974:BRBa

[720] Herman R. Friis. Book review: Inventaire du fonds Philippe Vandermaelen conservé à la Bibliothèque Royale Albert Ier by Liliane Wellens-de Donder. Isis, 65(1):125, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228914.

Kevles:1974:BRB

[721] Daniel J. Kevles. Book review: Pieces of the Action by Vannevar Bush. Isis, 65(1):125–126, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228915.

Schlagel:1974:BRB

[722] Richard H. Schlagel. Book review: Bergson and Modern Physics: A Reinterpretation and Re-Evaluation by Milic Capek. Isis, 65(1):127– 128, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228916.

Heilbron:1974:BRBa

[723] J. L. Heilbron. Book review: Albert Einstein Creator and Rebel by Banesh Hoffmann; Helen Dukas. Isis, 65(1):128–129, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228917.

Mikulak:1974:BRB

[724] Maxim W. Mikulak. Book review: Logique et méthodologie des sciences en U.R.S.S. Isis, 65(1):129–130, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228918.

Hahn:1974:BRB

[725] Roger Hahn. Book review: Passion to Know. The World’s Scientists by Mitchell Wilson. Isis, 65(1):130–131, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228919.

Anonymous:1974:BMa

[726] Anonymous. Back matter. Isis, 65(1):133–140, March 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228920.

Anonymous:1974:FMb

[727] Anonymous. Front matter. Isis, 65(2):141–143, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229368.

Kottler:1974:ARW

[728] Malcolm Jay Kottler. Alfred russel wallace, the origin of man, and spiritualism. Isis, 65(2):144–192, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229369.

Steneck:1974:AGC

[729] Nicholas H. Steneck. Albert the great on the classification and localization of the internal senses. Isis, 65(2):193–211, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229370.

Guedon:1974:PCM

[730] Jean-Claude Guedon. Protestantisme et chimie: Le milieu intellectuel de nicolas Lémery. Isis, 65(2):212–228, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229371.

Rabinovitch:1974:FCL

[731] Nachum L. Rabinovitch. A Fifteenth-Century law of large numbers. Isis, 65(2):229–238, June 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229372.

Rosinska:1974:NAD

[732] Grazyna Rosińska. Nas.īr al-Dīn al-T .ūsī and ibn al-Shāt.ir in cracow? Isis, 65(2):239–243, June 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229373.

Stuewer:1974:N

[733] Roger H. Stuewer. News. Isis, 65(2):244–248, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229374.

Bloor:1974:BRB

[734] David Bloor. Book review: Human Understanding, by Stephen Toulmin. Volume I: The Collective Use and Evolution of Concepts. Isis, 65(2):

249–253, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229375.

Hiebert:1974:BRB

[735] Erwin N. Hiebert. Book review: Mach’s Philosophy of Science. Isis, 65(2):253–255, June 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229376.

Spring:1974:BRB

[736] David Spring. Book review: Machina ex Deo: Essays in the Dynamism of Western Culture by Lynn White, Jr. Isis, 65(2):255–256, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229377.

Neu:1974:BRB

[737] John Neu. Book review: Preliminary Guide to the Smithsonian Archives. 125th Anniversary of the Smithsonian Institution by Smithsonian Institution. Isis, 65(2):256, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229378.

Menkes:1974:BRB

[738] Diana D. Menkes. Book review: Dictionary of Scientific Biography. Isis, 65(2):256–259, June 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229379.

Hesse:1974:BRBb

[739] Mary Hesse. Book review: Law and Explanation. An Essay in the Philosophy of Science by Peter Achinstein. Isis, 65(2):259–260, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229380.

Ardila:1974:BRB

[740] Ruben Ardila. Book review: Neurosis y psicoterapia, un estudio histórico by José M. López Piñero; José M. Morales Meseguer. Isis, 65(2):260, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229381.

Bolzan:1974:BRB

[741] J. E. Bolzan. Book review: Actas del III Congreso Nacional de Historia de la Medicina, Valencia (España), 10–12 de abril de 1969. Isis, 65(2): 260–261, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229382.

Lutzker:1974:BRB

[742] Edythe Lutzker. Book review: Women in White. Their Role as Doctors through the Ages by Geoffrey Marks; William K. Beatty. Isis, 65(2):261– 262, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229383.

Bean:1974:BRB

[743] William B. Bean. Book review: American Medicine and the Public Interest by Rosemary Stevens. Isis, 65(2):262–264, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229384.

Clagett:1974:BRB

[744] Marshall Clagett. Book review: Archimède by Charles Mugler. Isis, 65(2):264–267, June 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229385.

Balme:1974:BRB

[745] D. M. Balme. Book review: Opere biologiche by Aristotele; Diego Lanza; Mario Vegetti. Isis, 65(2):267–268, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229386.

Hahm:1974:BRB

[746] David E. Hahm. Book review: La lettre d’Épicure by Jean Bollack; Mayotte Bollack; Heinz Wismann. Isis, 65(2):268–269, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229387.

McVaugh:1974:BRB

[747] Michael R. McVaugh. Book review: In Anticlaudianum Alani commentum by Radulphus de Longo Campo; Jan Sulowski. Isis, 65(2):269, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229388.

Hermeunk:1974:BRB

[748] Heinrich Hermeunk. Book review: Al-Bit.ruji: On the Principles of Astronomy by Bernard R. Goldstein. Isis, 65(2):269–271, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229389.

Rosenfeld:1974:BRB

[749] B. Rosenfeld. Book review: Sanskrit Astronomical Tables in the United States by David Pingree; Census of the Exact Sciences in Sanskrit. Series A, Volume I by David Pingree; Census of the Exact Sciences in Sanskrit. Series A, Volume II by David Pingree. Isis, 65(2):271–272, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229390.

Kangro:1974:BRB

[750] Hans Kangro. Book review: De Methodo, hoc est de recta investigandarum tradendarumque artium ac scientiarum ratione. Über die Methode, d. h. über die rechte Forschung und Lehre in den Künsten und Wissenschaften by Jacobus Acontius; Alois von der Stein. Isis, 65(2):273– 274, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229391.

Theodorides:1974:BRBa

[751] Jean Théodoridès. Book review: African Animals in Renaissance Literature and Art by Joan E. Barclay Lloyd. Isis, 65(2):274–275, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229392.

Heilbron:1974:BRBb

[752] J. L. Heilbron. Book review: Theologie der Elektrizität. Zur Begegnung und Auseinandersetzung von Theologie und Naturwissenschaft im 17. und 18. Jahrundert by Ernst Benz. Isis, 65(2):275, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229393.

Grabiner:1974:BRB

[753] Judith V. Grabiner. Book review: Schriften über die Grundlagen der Mathematik und Physik by George Berkeley; Wolfgang Breidert. Isis, 65(2):275–277, June 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229394.

Kargon:1974:BRBa

[754] Robert Kargon. Book review: The Selected Works of Pierre Gassendi by Craig B. Brush. Isis, 65(2):277, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229395.

Dorn:1974:BRB

[755] Harold Dorn. Book review: Coulomb’s Memoir on Statics. An Essay in the History of Civil Engineering by Jacques Heyman. Isis, 65(2):277– 279, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229396.

Friis:1974:BRBb

[756] Herman R. Friis. Book review: A Map of the British Empire in America with the French and Spanish Settlements Adjacent thereto…, C. Lempriere & del., B. Baron Sculp., London Engrav’d by William Henry Toms, 1733 by Henry Popple. Isis, 65(2):279, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229397.

Olmsted:1974:BRB

[757] John W. Olmsted. Book review: Public Welfare, Science, and Propaganda in Seventeenth Century France: The Innovations of Théophraste Renaudot by Howard M. Solomon. Isis, 65(2):279–280, June 1974. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229398.

Scriba:1974:BRBc

[758] Christoph J. Scriba. Book review: Essai sur une manière de représenter les quantités imaginaires dans les constructions géométriques by R. Argand. Isis, 65(2):281, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229399.

Krader:1974:BRB

[759] Lawrence Krader. Book review: Berliner Gesellschaft für Anthropologie, Ethnologie und Urgeschichte. Festschrift zum Hundertjährigen Bestehen der Berliner Gesellschaft für Anthropologie, Ethnologie und Urgeschichte, 1869–1969. Isis, 65(2):281–282, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229400.

Quen:1974:BRB

[760] Jacques M. Quen. Book review: Psychiatry and the Community in 19th-Century America. The Recurring Concern with Environment in the Prevention and Treatment of Mental Disorders by Ruth B. Caplan; Gerald Caplan. Isis, 65(2):282, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229401.

Peterson:1974:BRB

[761] M. Jeanne Peterson. Book review: Edith Pechey-Phipson, M.D.: The Story of England’s Foremost Pioneering Woman Doctor by Edythe Lutzker. Isis, 65(2):282–284, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229402.

Spencer:1974:BRB

[762] J. Brookes Spencer. Book review: Nineteenth-Century Spectroscopy. Development of the Understanding of Spectra 1802–1897 by William McGucken. Isis, 65(2):284–285, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229403.

Theodorides:1974:BRBb

[763] Jean Théodoridès. Book review: Lettres de René-Edouard Claparède (1832–1871) by G. de Morsier. Isis, 65(2):285–286, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229404.

Knight:1974:BRB

[764] David Knight. Book review: Goethe and the Scientific Tradition by H. B. Nisbet. Isis, 65(2):286–287, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229405.

Stuewer:1974:BRB

[765] Roger H. Stuewer. Book review: The Life of Wilhelm Conrad Röntgen, Discoverer of the X Ray by W. Robert Nitske. Isis, 65(2):287, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229406.

Rothblatt:1974:BRB

[766] Sheldon Rothblatt. Book review: Herbert Spencer: The Evolution of a Sociologist by J. D. Y. Peel. Isis, 65(2):287–289, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229407.

Badash:1974:BRB

[767] Lawrence Badash. Book review: Selected Papers of Cecil Frank Powell by E. H. S. Burhop; W. O. Lock; M. G. K. Menon. Isis, 65(2):289–290, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229408.

Hesse:1974:BRBc

[768] Mary Hesse. Book review: Critical Essays on the Philosophy of R. G. Collingwood by Michael Krausz. Isis, 65(2):290, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229409.

Decker:1974:BRBa

[769] Hannah S. Decker. Book review: The Freudian Left: Wilhelm Reich, Geza Roheim, Herbert Marcuse by Paul A. Robinson. Isis, 65(2):290– 292, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229410.

Winnik:1974:BRB

[770] Herbert Winnik. Book review: Scientists in the Quest for Peace: A History of the Pugwash Conferences by J. Rotblat. Isis, 65(2):292, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229411.

Heilbron:1974:BRA

[771] J. L. Heilbron. Book review: Abdus Salam and Eugene Wigner, Aspects of Quantum Theory. Isis, 65(2):292–293, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229412.

Anonymous:1974:BMb

[772] Anonymous. Back matter. Isis, 65(2):294–296, June 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229413.

Anonymous:1974:FMc

[773] Anonymous. Front matter. Isis, 65(3):297–299, September 1974. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228954.

Burchfield:1974:DDG

[774] Joe D. Burchfield. Darwin and the dilemma of geological time. Isis, 65 (3):300–321, September 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228955.

Carre:1974:MBT

[775] Marie-Rose Carré. A man between two worlds: Pierre Borel and his discours nouveau prouvant la pluralité des mondes of 1657. Isis, 65(3):

322–335, September 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228956.

Sonntag:1974:MSS

[776] Otto Sonntag. The motivations of the scientist: The self-image of Albrecht von Haller. Isis, 65(3):336–351, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228957.

Guralnick:1974:SMR

[777] Stanley M. Guralnick. Sources of misconception on the role of science in the Nineteenth-Century American College. Isis, 65(3):352–366, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228958.

Saidan:1974:AAW

[778] A. S. Saidan, Abu’l-Wafa’, Muhammad ibn Muhammad ibn Yahya, and al Buzajani. The arithmetic of Abū’l–Wafā’. Isis, 65(3):367–375, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228959.

Dibner:1974:ELR

[779] Bern Dibner. Éloge: Ladislao Reti 1901–1973. Isis, 65(3):376–378, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228960.

Newburgh:1974:FDP

[780] Ronald Newburgh. Fresnel drag and the principle of relativity. Isis, 65 (3):379–386, September 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228961.

Rose:1974:CUR

[781] Paul Lawrence Rose. Copernicus and urbino: Remarks on bernardino Baldi’s vita di Niccolò copernico (1588). Isis, 65(3):387–389, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228962.

Hetherington:1974:EHA

[782] Norriss S. Hetherington. Edwin Hubble on Adriaan van Maanen’s internal motions in spiral nebulae. Isis, 65(3):390–393, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228963.

Smith:1974:BEF

[783] Cyril Stanley Smith. Bergman’s “equivalents”: A further comment. Isis, 65(3):393–394, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228964.

Ronchi:1974:LE

[784] Vasco Ronchi. Letter to the Editor. Isis, 65(3):394–395, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228965.

Anonymous:1974:Nb

[785] Anonymous. News. Isis, 65(3):396–397, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228966.

Whiteside:1974:BRB

[786] D. T. Whiteside. Book review: The Mathematical Career of Pierre de Fermat (1601–1665). Isis, 65(3):398–400, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228967.

Kargon:1974:BRBb

[787] Robert Kargon. Book review: Science and Society 1600–1900. Isis, 65 (3):400–402, September 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228968.

Hesse:1974:BRBd

[788] Mary Hesse. Book review: A Historical Introduction to the Philosophy of Science by John Losee. Isis, 65(3):402–403, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228969.

Hunt:1974:BRB

[789] G. M. K. Hunt. Book review: Statistical Explanation and Statistical Relevance by Wesley C. Salmon; R. C. Jeffrey; J. G. Greeno. Isis, 65 (3):403–404, September 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228970.

Hesse:1974:BRBe

[790] Mary Hesse. Book review: Modern Physics and Its Philosophy by Martin Strauss. Isis, 65(3):404, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228971.

Warner:1974:BRBa

[791] Deborah Jean Warner. Book review: Early Scientific Instruments. The Arthur Frank Loan Collection by Robert H. Nuttall. Isis, 65(3):404–405, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228972.

Theodorides:1974:BRBc

[792] Jean Théodoridès. Book review: Animals in Archaeology by A. Houghton Brodrick. Isis, 65(3):405–406, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228973.

Ritterbush:1974:BRB

[793] Philip C. Ritterbush. Book review: Modern Art and Scientific Thought by John Adkins Richardson. Isis, 65(3):406, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228974.

vanderWaerden:1974:BRBb

[794] B. L. van der Waerden. Book review: New and Full Moons, 1001 B.C. to A.D. 1561 by Herman H. Goldstine. Isis, 65(3):407, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228975.

Rosenkrantz:1974:BRB

[795] Barbara G. Rosenkrantz. Book review: Men, Beasts, and Gods: A History of Cruelty and Kindness to Animals by Gerald Carson. Isis, 65 (3):407–408, September 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228976.

Ihde:1974:BRBa

[796] Aaron J. Ihde. Book review: A History of Biochemistry by Marcel Florkin; Elmer H. Stotz. Isis, 65(3):409, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228977.

Fries:1974:BRB

[797] Russell I. Fries. Book review: Technology and American Economic Growth by Nathan Rosenberg. Isis, 65(3):409–410, September 1974.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228978.

Durling:1974:BRB

[798] Richard J. Durling. Book review: Galeno en la sociedad y en la ciencia de su tiempo (c. 130–c. 200 d. de C.) by Luis García Ballester. Isis, 65 (3):410–411, September 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228979.

Barnes:1974:BRB

[799] Jonathan Barnes. Book review: Aristotle on Memory by Richard Sorabji. Isis, 65(3):412–413, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228980.

Knorr:1974:BRB

[800] Wilbur Knorr. Book review: Division and Logos by Christian Marinus Taisbak. Isis, 65(3):413–414, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228981.

Miyasita:1974:BRB

[801] Saburô Miyasita. Book review: Tibetan Medicine, Illustrated in Original Texts by Rechung Rinpoche. Isis, 65(3):414–415, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228982.

Stannard:1974:BRB

[802] Jerry Stannard. Book review: Johannes Kentmann 1518–1574. Ein sächsischer Arzt und Naturforscher by Johannes Helm. Isis, 65(3):415– 416, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228983.

Stocking:1974:BRBb

[803] George W. Stocking, Jr. Book review: James Burnett, Lord Monboddo by E. L. Cloyd; Henry Home, Lord Kames, and the Scottish Enlightenment: A Study in National Character and in the History of Ideas by William C. Lehmann. Isis, 65(3):416–418, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228984.

Farber:1974:BRB

[804] Paul Lawrence Farber. Book review: Buffon by Otis E. Fellows; Stephen F. Milliken. Isis, 65(3):418–419, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228985.

Ariotti:1974:BRB

[805] Piero E. Ariotti. Book review: Galileo’s Intellectual Revolution, Middle Period, 1610–1632 by William R. Shea. Isis, 65(3):419–420, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228986.

Sonntag:1974:BRB

[806] Otto Sonntag. Book review: Albrecht von Haller. Über die Einheit im Denken des letzten Universalgelehrten by Richard Toellner. Isis, 65(3): 421, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228987.

Gruber:1974:BRB

[807] Jacob W. Gruber. Book review: Anthropology of the Numa; John Wesley Powell’s Manuscripts on the Numic People of Western North America, 1868–1880 by Don D. Fowler; Catherine S. Fowler. Isis, 65(3):421–422, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228988.

McKinney:1974:BRB

[808] H. Lewis McKinney. Book review: The Voyage of the Challenger by Eric Linklater. Isis, 65(3):422–423, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228989.

Coleman:1974:BRBb

[809] William Coleman. Book review: Stendhal du côté de la science by Jean Théodoridè. Isis, 65(3):423, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228990.

Decker:1974:BRBb

[810] Hannah S. Decker. Book review: Sigmund Freud und die Religion by Kasimir Birk. Isis, 65(3):424–425, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228991.

Hetherington:1974:BRB

[811] Norriss S. Hetherington. Book review: L’hypothèse de l’atome primtif by Georges Lemaître. Isis, 65(3):425, September 1974. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228992.

Goldberg:1974:BRB

[812] Stanley Goldberg. Book review: The Collected Works of Leo Szilard. Volume I: Scientific Papers by Bernard T. Feld; Gertrud Weiss Szilard. Isis, 65(3):425–426, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228993.

Anonymous:1974:BMc

[813] Anonymous. Back matter. Isis, 65(3):427–428, September 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228994.

Anonymous:1974:FMd

[814] Anonymous. Front matter. Isis, 65(4):429–430, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229333.

Tigner:1974:ETL

[815] Steven S. Tigner. Empedocles’ twirled ladle and the vortex-supported Earth. Isis, 65(4):432–447, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229334.

Zochert:1974:SCM

[816] Donald Zochert. Science and the common man in ante-bellum America. Isis, 65(4):448–473, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229335.

Leighly:1974:EDD

[817] John Leighly. An early drawing and description of a Tornado. Isis, 65 (4):474–486, December 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229336.

Rogers:1974:ROD

[818] James Allen Rogers. Russian opposition to Darwinism in the nineteenth century. Isis, 65(4):487–505, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229337.

Oldroyd:1974:NSF

[819] D. R. Oldroyd. A note on the status of a. f. Cronstedt’s simple Earths and his analytical methods. Isis, 65(4):506–512, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229338.

Barr:1974:BMF

[820] E. Scott Barr. Biographical material in the first one hundred volumes of the astrophysical journal. Isis, 65(4):513–515, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229339.

DiCurcio:1974:LE

[821] Robert A. DiCurcio. Letter to the Editor. Isis, 65(4):516, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229340.

Finocchiaro:1974:CHA

[822] Maurice A. Finocchiaro. A curious history of astronomy: Leopardi’s storia dell’astronomia. Isis, 65(4):517–519, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229341.

Anonymous:1974:GSM

[823] Anonymous. The 1973 George Sarton Medal awarded to Henry Guerlac. Isis, 65(4):520–521, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229342.

Anonymous:1974:Nc

[824] Anonymous. News. Isis, 65(4):522–523, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229343.

Jayawardene:1974:BRB

[825] S. A. Jayawardene. Book review: Bibliography and Research Manual of the History of Mathematics. Isis, 65(4):524–526, December 1974. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229344.

Ewan:1974:BRBb

[826] Joseph Ewan. Book review: Early Scientific Books in the Schaffer Library, Union College by Wayne Somers. Isis, 65(4):526–527, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229345.

Gauthier:1974:BRB

[827] Yvon Gauthier. Book review: La mathématisation des doctrines informes by Georges Canguilhem. Isis, 65(4):527–528, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229346.

Hesse:1974:BRBf

[828] Mary Hesse. Book review: The Problem of Scientific Realism by Edward A. MacKinnon. Isis, 65(4):528, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229347.

Hesse:1974:BRBg

[829] Mary Hesse. Book review: Philosophy, Science, and Method. Essays in Honor of Ernest Nagel by Sidney Morgenbesser; Patrick Suppes; Morton White. Isis, 65(4):528–529, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229348.

Hesse:1974:BRBh

[830] Mary Hesse. Book review: The Field Concept in Contemporary Science by Mendel Sachs. Isis, 65(4):529, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229350.

Warner:1974:BRBb

[831] Deborah Jean Warner. Book review: Astronomy by Colin A. Ronan. Isis, 65(4):529, December 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229349.

Bynum:1974:BRB

[832] W. F. Bynum. Book review: An Illustrated History of Brain Function by Edwin Clarke; Kenneth Dewhurst. Isis, 65(4):530, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229351.

Hall:1974:BRB

[833] Bert S. Hall. Book review: The Projectile-Throwing Engines of the Ancients by Ralph Payne-Gallwey. Isis, 65(4):530–531, December 1974.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229352.

vonFritz:1974:BRB

[834] Kurt von Fritz. Book review: Lore and Science in Ancient Pythagoreanism by Walter Burkert; Edwin L. Minar. Isis, 65(4):531, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229353.

Savage-Smith:1974:BRB

[835] Emilie Savage-Smith. Book review: Galens Exzerpte aus älteren Pharmakologen by Cajus Fabricius. Isis, 65(4):531–532, December 1974. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229354.

Shaw:1974:BRB

[836] James Rochester Shaw. Book review: The Heart and the Vascular System in Ancient Greek Medicine: From Alcmaeon to Galen by C. R. S. Harris. Isis, 65(4):533–534, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229355.

Weisheipl:1974:BRBb

[837] James A. Weisheipl. Book review: Myth and Science in the Twelfth Century: A Study of Bernard Silvester by Brian Stock. Isis, 65(4):534– 535, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229356.

Thoren:1974:BRB

[838] Victor E. Thoren. Book review: Najstarszy zyciorys Mikolaja Kopernika z roku 1588 pióra Bernardina Baldiego. [The Earliest Biography of Nicolaus Copernicus Dating from 1588, by Bernardino Baldi.] by Bronislaw Biliński. Isis, 65(4):535, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229357.

Scaglia:1974:BRB

[839] Gustina Scaglia. Book review: Trattato di architettura by Antonio Averlino detto Il Filarete; Anna Maria Finoli; Liliana Grassi. Isis, 65(4): 536–537, December 1974. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229358.

Vucinich:1974:BRB

[840] Alexander Vucinich. Book review: Newton and Russia: The Early Influence, 1698–1796 by Valentin Boss. Isis, 65(4):537–538, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229359.

Winder:1974:BRB

[841] Marianne Winder. Book review: Das Leib-Seele-Geist-Problem bei Paracelsus und einigen Autoren des 17. Jahrhunderts by Ernst Wilhelm Kämmerer. Isis, 65(4):538–539, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229360.

Drake:1974:BRB

[842] Stillman Drake. Book review: Discussione col Nunzio Sidereo e Relazione sui quattro satelliti di Giove by Johann Kepler; Elio Pasoli; Giorgio Tabarroni. Isis, 65(4):539–540, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229361.

Coleman:1974:BRBc

[843] William Coleman. Book review: L’équilibre de la nature by C. Linné; Bernard Jasmin. Isis, 65(4):540–541, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229362.

Blake:1974:BRB

[844] John B. Blake. Book review: Mary Adelaide Nutting, Pioneer of Modern Nursing by Helen E. Marshall. Isis, 65(4):541–542, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229363.

Wolff:1974:BRB

[845] Kurt H. Wolff. Book review: Durkheim: Morality and Milieu by Ernest Wallwork. Isis, 65(4):542–543, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229364.

Williams:1974:BRB

[846] L. Pearce Williams. Book review: Trial by Fury: The Polio Vaccine Controversy by Aaron E. Klein. Isis, 65(4):543–544, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229365.

Ihde:1974:BRBb

[847] Aaron J. Ihde. Book review: Les Prix Nobel en 1971 by Institutions Nobel; Nobel Lectures, Chemistry, 1963–1970 by Nobel Foundation; Nobel Lectures, Physics, 1963–1970 by Nobel Foundation. Isis, 65(4):544, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229366.

Anonymous:1974:BMd

[848] Anonymous. Back matter. Isis, 65(4):545–546, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229367.

Anonymous:1974:VI

[849] Anonymous. Volume information. Isis, 65(4):547–561, December 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229332.

Anonymous:1974:FMe

[850] Anonymous. Front matter. Isis, 65(5):1–4, ???? 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229520.

Neu:1974:NNC

[851] John Neu. Ninety-ninth critical bibliography of the history of science and its cultural influences (to January 1974). Isis, 65(5):5–165+167–206, ???? 1974. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229521.

Anonymous:1975:FMa

[852] Anonymous. Front matter. Isis, 66(1):1–3, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229522.

Delair:1975:EDD

[853] Justin B. Delair and William A. S. Sarjeant. The earliest discoveries of dinosaurs. Isis, 66(1):4–25, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229523.

Gerstner:1975:DTM

[854] Patsy A. Gerstner. A dynamic theory of mountain building: Henry Darwin rogers, 1842. Isis, 66(1):26–37, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229524.

Kauffman:1975:DOA

[855] George B. Kauffman. The discovery of optically active coordination compounds: A milestone in stereochemistry. Isis, 66(1):38–62, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229525.

Farber:1975:BDD

[856] Paul Lawrence Farber. Buffon and daubenton: Divergent traditions within the histoire naturelle. Isis, 66(1):63–74, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229526.

Daniels:1975:LSA

[857] Norman Daniels. Lobachevsky: Some anticipations of later views on the relation between geometry and physics. Isis, 66(1):75–85, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229527.

Schlunegger:1975:FLA

[858] E. Schlunegger, K. F. Meyer, and P. C. Boeren. Frédéric-Louis Allamand, seine Familie und die nach ihm benannte Pflanzengattung. (German) [Frédéric-Louis Allamand, his family, and the genus named after him]. Isis, 66(1):86–91, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229528.

Drake:1975:NLG

[859] Stillman Drake. New light on a Galilean claim about pendulums. Isis, 66(1):92–95, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229529.

Steinsaltz:1975:LMJ

[860] Adin Steinsaltz. The loadstone in Medieval Jewish literature. Isis, 66(1): 95–96, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229530.

Erickson:1975:AEC

[861] Paul A. Erickson. Anthropology and evolution: A comment on Wells. Isis, 66(1):96–97, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229531.

Schwarz:1975:LE

[862] Michael Schwarz, Robert H. Hewsen, and Rom Harré. Letters to the Editor. Isis, 66(1):98–101, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229532.

Anonymous:1975:Na

[863] Anonymous. News. Isis, 66(1):102–104, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229533.

Rousseau:1975:BRBa

[864] G. S. Rousseau. Book review: A Catalogue of Western Manuscripts on Medicine and Science in the Wellcome Historical Medical Library, by S. A. J. Moorat; Portraits of Doctors and Scientists in the Wellcome Institute of the History of Medicine. A Catalogue by Renate Burgess; Catalogue of Medical Books in Manchester University Library 1480–1700 by Ethel M. Parkinson; Audrey E. Lumb; Biographical Dictionary of Botanists Represented in the Hunt Institute Portrait Collection by Hunt Botanical Library. II: Manuscripts Written after A.D. 1650. Isis, 66(1): 105–108, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229534.

Schmitt:1975:BRBa

[865] Charles B. Schmitt. Book review: Ricerche sulla cultura dell’Italia moderna by Paola Zambelli. Isis, 66(1):108–109, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229535.

Weisheipl:1975:BRB

[866] James A. Weisheipl. Book review: The Problem of Evolution: A Study of the Philosophical Repercussions of Evolutionary Science by John N. Deely; Raymond J. Nogar. Isis, 66(1):109–110, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229536.

Jaki:1975:BRB

[867] Stanley L. Jaki. Book review: Constellations and Conjectures by Norwood Russell Hanson; Willard C. Humphreys. Isis, 66(1):110–112, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229537.

Layton:1975:BRB

[868] Edwin T. Layton, Jr. Book review: Engineering Education: A Social History by George S. Emmerson. Isis, 66(1):112, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229538.

Kevles:1975:BRBa

[869] Daniel J. Kevles. Book review: Scientists in Search of Their Conscience by Anthony R. Michaelis; Hugh Harvey. Isis, 66(1):112–113, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229539.

Churchill:1975:BRB

[870] Frederick B. Churchill. Book review: Organic Form, The Life of an Idea by G. S. Rousseau. Isis, 66(1):113–114, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229540.

Aull:1975:BRB

[871] Charlotte H. Aull. Book review: Africa Counts: Number and Pattern in African Culture by Claudia Zaslavsky. Isis, 66(1):114–115, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229541.

Norberg:1975:BRB

[872] Arthur L. Norberg. Book review: The Milky Way: An Elusive Road for Science by Stanley L. Jaki. Isis, 66(1):115–116, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229542.

Rosen:1975:BRB

[873] Edward Rosen. Book review: The Astronomical Revolution: Copernicus– Kepler–Borelli by Alexandre Koyré; R. E. W. Maddison. Isis, 66(1): 116, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229543.

Kangro:1975:BRB

[874] Hans Kangro. Book review: Von Galilei bis Einstein. Entwicklung der physikalischen Ideen by B. G. Kuznecov. Isis, 66(1):116–118, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229544.

Haber:1975:BRB

[875] Francis C. Haber. Book review: The Meaning of Fossils: Episodes in the History of Palaeontology by Martin J. S. Rudwick. Isis, 66(1):118– 119, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229545.

Vandervliet:1975:BRBa

[876] Glenn Vandervliet. Book review: History of Physiology by Karl E. Rothschuh; Guenter B. Risse. Isis, 66(1):119–120, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229546.

Egerton:1975:BRBa

[877] Frank N. Egerton, III. Book review: History of Entomology by Ray F. Smith; Thomas E. Mittler; Carroll N. Smith. Isis, 66(1):120, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229547.

Sokal:1975:BRB

[878] Michael M. Sokal. Book review: A Philosophicall Discourse concerning Speech (1668) and A Discourse Written to a Learned Frier (1670) by Géraud de Cordemoy; Barbara Ross; Observations on Mental Derangement by Andrew Combe; Anthony A. Walsh; Rational Psychology (1849) by Laurens Perseus Hickok; Ernest Harms. Isis, 66(1):120–122, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229548.

Stocking:1975:BRB

[879] George W. Stocking, Jr. Book review: The Establishment of Empirical Sociology: Studies in Continuity, Discontinuity, and Institutionalization by Anthony Oberschall. Isis, 66(1):122, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229549.

Cassedy:1975:BRB

[880] James H. Cassedy. Book review: The Story of Medicine in America by Geoffrey Marks; William K. Beatty. Isis, 66(1):122–123, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229550.

Vandervliet:1975:BRBb

[881] Glenn Vandervliet. Book review: Human Aspects of Biomedical Innovation by Everett Mendelsohn; Judith P. Swazey; Irene Taviss. Isis, 66(1):123–124, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229551.

Kranzberg:1975:BRB

[882] Melvin Kranzberg. Book review: Overskill: The Decline of Technology in Modern Civilization by Eugene S. Schwartz. Isis, 66(1):124–125, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229552.

Condit:1975:BRB

[883] Carl W. Condit. Book review: A Report of the Mohawk-Hudson Area Survey: A Selective Recording Survey of the Industrial Archeology of the Mohawk and Hudson River Valleys in the Vicinity of Troy, New York, June-September 1969 by Robert M. Vogel. Isis, 66(1):125–126, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229553.

Risse:1975:BRB

[884] Guenter B. Risse. Book review: Alma y enfermedad en la obra de Galeno by Luis Garcia Ballester. Isis, 66(1):126–127, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229554.

Kren:1975:BRB

[885] Claudia Kren. Book review: L’oeuvre astronomique de Themon Juif, maître parisien du XIVe siècle by Henri Hugonnard-Roche. Isis, 66(1): 127–128, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229555.

Thoren:1975:BRB

[886] Victor E. Thoren. Book review: Science et astrologie au XVIe siècle. Oronce Fine et son horloge planétaire, 22 Novembre-22 Décembre. Catalogue by Madeleine Boy; Complete Works. Volume I: The Manuscript of Nicholas Copernicus’ “On the Revolutions” Facsimile by Nicholas Copernicus; Pawel Czartoryski; Zygmunt Nierada; Erna Hilfstein. Isis, 66(1): 128, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229556.

Rabinovitch:1975:BRB

[887] N. L. Rabinovitch. Book review: Mone Mispar. An Arithmetical Treatise by Saadia Chouraqui; Gad B. Sarfatti. Isis, 66(1):129, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229557.

Aiton:1975:BRB

[888] E. J. Aiton. Book review: Leibniz and Dynamics. The Texts of 1692 by Pierre Costabel; R. E. W. Maddison. Isis, 66(1):129–130, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229558.

Debus:1975:BRB

[889] Allen G. Debus. Book review: Circulation Physiology and Medical Chemistry in England 1650–1680 by Audrey B. Davis. Isis, 66(1):130–131, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229559.

Calinger:1975:BRB

[890] Ronald Calinger. Book review: Wolffianismens genombrott i Uppsala: Frihetstida universitetsfilosofi till 1700-talets mitt (The Emergence of Wolffianism at Uppsala: University Philosophy from 1720 up to the Middle of the Eighteenth Century) by Tore Frängsmyr. Isis, 66(1):132–133, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229560.

Brechka:1975:BRB

[891] Frank T. Brechka. Book review: Gerard van Swieten und seine Zeit by Erna Lesky; Adam Wandruszka. Isis, 66(1):134–135, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229561.

Drake:1975:BRB

[892] Stillman Drake. Book review: Les nouvelles pensées de Galilée, mathématicien et ingénieure du Duc de Florence by Marin Mersenne. Isis, 66(1):135–136, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229562.

Levere:1975:BRB

[893] Trevor H. Levere. Book review: John Dalton. Critical Assessments of His Life and Science by Arnold Thackray. Isis, 66(1):136–137, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229563.

Ripley:1975:BRB

[894] S. Dillon Ripley. Book review: The Fowles of Heauen or History of Birdes by Edward Topsell; Thomas P. Harrison; F. David Hoeniger. Isis, 66(1):137–138, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229564.

Costa:1975:BRB

[895] Gustavo Costa. Book review: La formazione filosofica di Antonio Genovesi by Paola Zambelli. Isis, 66(1):138–139, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229565.

McKinney:1975:BRB

[896] H. Lewis McKinney. Book review: Scientist Extraordinary. The Life and Scientific Work of Thomas Henry Huxley, 1825–1895 by Cyril Bibby. Isis, 66(1):139–140, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229566.

Molella:1975:BRB

[897] Arthur P. Molella. Book review: Ernst Mach, His Work, Life, and Influence by John T. Blackmore. Isis, 66(1):140–142, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229567.

Heilbron:1975:BRBa

[898] J. L. Heilbron. Book review: H. E. Armstrong and the Teaching of Science, 1880–1930 by W. H. Brock. Isis, 66(1):142, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229568.

Kargon:1975:BRB

[899] Robert H. Kargon. Book review: The Victorian City: Images and Realities by H. J. Dyos; Michael Wolff. Isis, 66(1):142–144, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229569.

Umezawa:1975:BRB

[900] Toshio Umezawa. Book review: Selected Works of Giuseppe Peano by Hubert C. Kennedy. Isis, 66(1):144, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229570.

Steffens:1975:BRB

[901] Henry John Steffens. Book review: Julius Robert Mayer, Prophet of Energy by Robert Bruce Lindsay. Isis, 66(1):144–146, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229571.

Boley:1975:BRB

[902] Bruno A. Boley. Book review: Memorie by Luigi Federico Menabrea; Letterio Briguglio; Luigi Bulferetti. Isis, 66(1):146–147, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229572.

Greene:1975:BRB

[903] John C. Greene. Book review: Researches into the Physical History of Man by James Cowles Prichard; George W. Stocking. Isis, 66(1):147– 148, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229573.

Guralnick:1975:BRB

[904] Stanley M. Guralnick. Book review: Scientific and Technical Education in Nineteenth-Century England by Gordon W. Roderick; Michael D. Stephens. Isis, 66(1):148–149, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229574.

Pollack:1975:BRB

[905] Robert H. Pollack. Book review: Alfred Binet by Theta H. Wolf. Isis, 66(1):149–150, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229575.

Ritvo:1975:BRB

[906] Lucille B. Ritvo. Book review: Sigmund Freud by Richard Wollheim. Isis, 66(1):150–152, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229576.

Forman:1975:BRBa

[907] Paul Forman. Book review: Scientific Papers of Arthur Holly Compton: X-Ray and Other Studies by Arthur Holly Compton; Robert S. Shankland. Isis, 66(1):152, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229577.

Forman:1975:BRBb

[908] Paul Forman. Book review: Originality and Competition in Science. A Study of the British High Energy Physics Community by Jerry Gaston.

Isis, 66(1):152, March 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229578.

Swenson:1975:BRBa

[909] Loyd D. Swenson, Jr. Book review: The National Aeronautics and Space Administration by Richard Hirsch; Joseph John Trento. Isis, 66(1):152– 153, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229579.

Hatch:1975:BRB

[910] Elvin Hatch. Book review: Melville J. Herskovits by George Eaton Simpson. Isis, 66(1):153, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

229580.

Anonymous:1975:BMa

[911] Anonymous. Back matter. Isis, 66(1):154–160, March 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229581.

Anonymous:1975:FMb

[912] Anonymous. Front matter. Isis, 66(2):161–163, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229414.

Westman:1975:MCR

[913] Robert S. Westman. The Melanchthon circle, Rheticus, and the Wittenberg interpretation of the Copernican theory. Isis, 66(2):164–193, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229415.

Shapiro:1975:NDL

[914] Alan E. Shapiro. Newton’s definition of a light ray and the diffusion theories of chromatic dispersion. Isis, 66(2):194–210, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229416.

Longrigg:1975:EPL

[915] James Longrigg. Elementary physics in the lyceum and stoa. Isis, 66(2): 211–229, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229417.

Miller:1975:RMA

[916] John David Miller. Rowland’s magnetic analogy to Ohm’s law. Isis, 66(2):230–241, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229418.

Mazumdar:1975:JMB

[917] Pauline M. H. Mazumdar. Johannes Müller on the blood, the lymph, and the chyle. Isis, 66(2):242–253, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229419.

Anonymous:1975:Nb

[918] Anonymous. News. Isis, 66(2):254–255, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229420.

Sinclair:1975:BRB

[919] Bruce Sinclair. Book review: The Papers of Joseph Henry, by Nathan Reingold; Stuart Pierson; Arthur P. Molella; James M. Hobbins; John R. Kerwood. Volume I: December 1797–October 1832: The Albany Years. Isis, 66(2):256–258, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229421.

Holmes:1975:BRB

[920] Frederic L. Holmes. Book review: Raisonnement expérimental et recherches toxicologiques chez Claude Bernard. Isis, 66(2):259–261, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229422.

Wisan:1975:BRBa

[921] Winifred Wisan. Book review: Florence in the Forgotten Centuries: 1527–1800: A History of Florence and the Florentines in the Age of the Grand Dukes by Eric Cochrane. Isis, 66(2):261–262, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229423.

Fox:1975:BRB

[922] Robert Fox. Book review: Isis Cumulative Bibliography. A Bibliography of the History of Science formed from Isis Critical Bibliographies 1–90, 1913–65 by Magda Whitrow. Isis, 66(2):262–263, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229424.

Graham:1975:BRB

[923] Loren Graham. Book review: Problema preryvnosti i nepreryvnosti prostranstva i vremeni (The Problem of Discontinuity and Continuity in Space and Time) by M. D. Akhundov. Isis, 66(2):264, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229425.

Harre:1975:BRM

[924] R. Harré. Book review: Method, Model and Matter by Mario Bunge. Isis, 66(2):264–266, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229426.

Harder:1975:BRB

[925] Allen J. Harder. Book review: Truth in the Sciences by Scott Buchanan. Isis, 66(2):266–267, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229427.

Multhauf:1975:BRB

[926] Robert P. Multhauf. Book review: A History of the Bibliography of Philosophy by Michael Jasenas. Isis, 66(2):267–268, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229428.

Pyenson:1975:BRBa

[927] Lewis Pyenson. Book review: Leçons sur la genèse des théories physiques. Galilée, Amprère, Einstein by Jacques Merleau-Ponty. Isis, 66(2):268– 270, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229429.

Webster:1975:BRB

[928] Charles Webster. Book review: Medizingeschichte in unserer Zeit. Festgabe für Edith Heischkel-Artelt und Walter Artelt zum 65. Geburtstag by Hans-Heinz Eulner; Gunter Mann; Gert Preiser; Rolf Winau; Otto Winkelmann. Isis, 66(2):270–271, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229430.

Leckie:1975:BRB

[929] Doris Leckie. Book review: The Care and Exhibition of Medical History Museum Objects by Patsy A. Gerstner. Isis, 66(2):271, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229431.

Reingold:1975:BRB

[930] Nathan Reingold. Book review: Technology and Social Change in America by Edwin T. Layton. Isis, 66(2):271, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229432.

Mayr:1975:BRBa

[931] Otto Mayr. Book review: Bibliography of the Philosophy of Technology by Carl Mitcham; Robert Mackey. Isis, 66(2):271–273, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229433.

Busard:1975:BRB

[932] H. L. L. Busard. Book review: Der Donauraum die Wiege Mathematischer Studien in Deutschland by Kurt Vogel. Isis, 66(2):273–274, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229434.

Anbouba:1975:BRB

[933] Adel Anbouba. Book review: Al-Bahir en algèbre by As-Samaw’al; Salah Ahmad; Roshdi Rashed. Isis, 66(2):274–275, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229435.

Donahue:1975:BRB

[934] W. H. Donahue. Book review: The Cosmology of Giordano Bruno by Paul-Henri Michel; R. E. W. Maddison. Isis, 66(2):276, June 1975. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229436.

Eastwood:1975:BRBa

[935] Bruce S. Eastwood. Book review: The Origin of Subjectivity. An Essay on Descartes by Hiram Caton. Isis, 66(2):276–278, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229437.

Schmitt:1975:BRBb

[936] C. B. Schmitt. Book review: Il problema metodologico alle origini della scienza moderna by Angelo Crescini. Isis, 66(2):278, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229438.

Warner:1975:BRB

[937] Deborah Jean Warner. Book review: Willem Jansz. Blaeu. A Biography and History of His Work as a Cartographer and Publisher by J. Keuning; Marijke Donkersloot-de Vrij. Isis, 66(2):279, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229439.

Hankins:1975:BRB

[938] Thomas L. Hankins. Book review: The Unpublished Writings of Tobias Mayer by Eric G. Forbes. Isis, 66(2):279–280, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229440.

Middleton:1975:BRB

[939] W. E. Knowles Middleton. Book review: Van Marum’s Scientific Instruments in Teyler’s Museum. (Martinus van Marum, Life and Work, Volume IV.) by G. L’E. Turner; T. H. Levere. Isis, 66(2):280–281, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229441.

Laverty:1975:BRB

[940] Carroll D. Laverty. Book review: Charles Henry and the Formation of a Psychophysical Aesthetic by José A. Argüelles. Isis, 66(2):281–282, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229442.

Schneer:1975:BRB

[941] Cecil J. Schneer. Book review: Early Papers on Diffraction of X-Rays by Crystals. Volume II by J. M. Bijvoet; W. G. Burgers; G. Hägg. Isis, 66(2):282–284, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229443.

Koppelman:1975:BRB

[942] Elaine Koppelman. Book review: A Source Book in Classical Analysis by Garrett Birkhoff; Uta Merzbach. Isis, 66(2):284–285, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229444.

Sevier:1975:BRB

[943] John Sevier. Book review: Prophets and Patrons: The French University and the Emergence of the Social Sciences by Terry Nichols Clark. Isis, 66(2):285–287, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229445.

White:1975:BRB

[944] George W. White. Book review: The Expeditions of John Charles Frémont. Volume II: The Bear Flag Revolt and the Court Martial, and Volume II Supplement: Proceedings of the Court Martial by John Charles Frémont; Mary Lee Spence; Donald Jackson. Isis, 66(2):287–288, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229446.

Hesse:1975:BRB

[945] Mary Hesse. Book review: Émile Durkheim, His Life and Work: A Historical and Critical Study by Stephen Lukes. Isis, 66(2):288, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/229447.

Shapin:1975:BRB

[946] Steven Shapin. Book review: Philosophers and Provincials: The Yorkshire Philosophical Society from 1822 to 1844 by A.D. Orange. Isis, 66 (2):288–289, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/229448.

Forman:1975:BRBc

[947] Paul Forman. Book review: Wege der Naturforschung 1822–1972 im Spiegel der Versammlungen Deutscher Naturforscher und ärzte by Hans Querner; Heinrich Schipperges. Isis, 66(2):289–290, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229449.

McGhee:1975:BRB

[948] Richard D. McGhee. Book review: Between Science and Religion: The Reaction to Scientific Naturalism in Late Victorian England by Frank Miller Turner. Isis, 66(2):290–291, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229450.

Loewenberg:1975:BRB

[949] Peter Loewenberg. Book review: The Dialectical Imagination: A History of the Frankfurt School and the Institute of Social Research, 1923–1950 by Martin Jay. Isis, 66(2):291–293, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229451.

Petersen:1975:BRB

[950] Aage Petersen. Book review: The Logical Analysis of Quantum Mechanics by Erhard Scheibe; J. B. Sykes. Isis, 66(2):293–294, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229452.

Kevles:1975:BRBb

[951] Daniel J. Kevles. Book review: The Story of Southwest Research Center, A Private, Nonprofit, Scientific Research Adventure by Harold Vagtborg. Isis, 66(2):294–295, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229453.

Trenn:1975:BRBa

[952] Thaddeus J. Trenn. Book review: The Structure of Matter: Rutherford Centennial Symposium, Held at the University of Canterbury, Christchurch, New Zealand, 7–9 July 1971 by B. G. Wybourne. Isis, 66(2):295–296, June 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/229454.

Anonymous:1975:BMb

[953] Anonymous. Back matter. Isis, 66(2):297–298, June 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/229455.

Anonymous:1975:FMc

[954] Anonymous. Front matter. Isis, 66(3):299–301, September 1975. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228838.

Ariotti:1975:BCM

[955] Piero E. Ariotti. Bonaventura cavalieri, Marin mersenne, and the reflecting telescope. Isis, 66(3):302–321, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228839.

Allen:1975:IDS

[956] Garland E. Allen. The introduction of Drosophila into the study of heredity and evolution: 1900–1910. Isis, 66(3):322–333, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228840.

Cole:1975:EAS

[957] Theron M. Cole, Jr. Early atomic speculations of Marc Antoine Gaudin: Avogadro’s hypothesis and the periodic system. Isis, 66(3):334–360, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228841.

Siegel:1975:CGM

[958] Daniel M. Siegel. Completeness as a goal in Maxwell’s electromagnetic theory. Isis, 66(3):361–368, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228842.

Harper:1975:ATW

[959] Richard Harper. The astronomical tables of William rede. Isis, 66(3): 369–378, September 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228843.

Ross:1975:OFS

[960] Richard P. Ross and Oronce Fine. Oronce Fine’s de sinibus libri ii: The first printed trigonometric treatise of the French Renaissance. Isis, 66 (3):379–386, September 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228844.

Dyason:1975:EEB

[961] Diana Dyason. Éloge: Elizabeth Belmont Gasking, September 2, 1914– September 10, 1973. Isis, 66(3):387–389, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228845.

Smith:1975:SCO

[962] Cyril Stanley Smith. A Seventeenth-Century octonary arithmetic. Isis, 66(3):390–394, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228846.

Naylor:1975:AGS

[963] Ronald Naylor. An aspect of Galileo’s study of the parabolic trajectory. Isis, 66(3):394–396, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228847.

Shea:1975:SDA

[964] William R. Shea and Neil S. Wolf. Stillman Drake and the Archimedean grandfather of experimental science. Isis, 66(3):397–400, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228848.

Drake:1975:RSW

[965] Stillman Drake and James MacLachlan. Reply to the shea-wolf critique. Isis, 66(3):400–403, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228849.

Anonymous:1975:EBE

[966] Anonymous. Erratum: Bergman’s “equivalents”: A further comment. Isis, 66(3):404, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228851.

Schufle:1975:FRB

[967] J. A. Schufle. Further reply on Bergman’s “equivalents”. Isis, 66(3):404, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228850.

Kennedy:1975:CMF

[968] Kenneth A. R. Kennedy and Stephen F. Holtzman. Climbing Man’s family tree. Isis, 66(3):404–406, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228852.

Tasch:1975:QEG

[969] Paul Tasch. A quantitative estimate of geological time by lyell. Isis, 66(3):406, September 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228853.

Anonymous:1975:Nc

[970] Anonymous. News. Isis, 66(3):407–408, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228854.

Eastwood:1975:BRBb

[971] Bruce S. Eastwood. Book review: The Ascent of Man. Isis, 66(3): 409–411, September 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228855.

Wallace:1975:BRB

[972] William A. Wallace. Book review: Viator. Medieval and Renaissance Studies, 1972, Volume 3 by Lynn White. Isis, 66(3):411–412, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228856.

Buck:1975:BRB

[973] Peter Buck. Book review: Science and Politics by Jean-Jacques Salomon; Noël Lindsay. Isis, 66(3):412–413, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228857.

Taylor:1975:BRB

[974] Angus E. Taylor. Book review: Functional Analysis in Historical Perspective by A. F. Monna. Isis, 66(3):413–414, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228858.

Goldstein:1975:BRB

[975] Bernard R. Goldstein. Book review: Probability and Statistical Inference in Ancient and Medieval Jewish Literature by Nachum L. Rabinovitch. Isis, 66(3):414–415, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228859.

Ruse:1975:BRB

[976] Michael Ruse. Book review: Philosophy of Biological Science by David Hull. Isis, 66(3):416–417, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228860.

Thomas:1975:BRB

[977] Richard J. Thomas and Josef Brožek. Book review: Historical Conceptions of Psychology by Mary Henle; Julian Jaynes; John J. Sullivan. Isis, 66(3):417–418, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

228861.

Freeman:1975:BRB

[978] Eric J. Freeman. Book review: Bibliography of the History of Medicine 1964–1969 by National Library of Medicine. Isis, 66(3):418–419, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228862.

Post:1975:BRB

[979] Robert C. Post. Book review: Construction of Mill Dams by James Leffel; Some Details of Water-Works Construction by William R. Billings; The Manufacture of Liquors and Preserves by J. De Brevans; The Manufacture of Porcelain and Glass by Dionysius Lardner. Isis, 66(3):419, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228863.

Mayr:1975:BRBb

[980] Otto Mayr. Book review: Roman and Islamic Water-Lifting Wheels by Thorkild Schiøler. Isis, 66(3):419–420, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228864.

Kelley:1975:BRB

[981] David H. Kelley. Book review: Sonne, Mond und Sterne über dem Reich der Inka by Rolf Müller. Isis, 66(3):420–421, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228865.

Harper:1975:BRB

[982] Richard Harper. Book review: The Scientific Achievement of the Middle Ages by Richard C. Dales. Isis, 66(3):421–422, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228866.

Schmitt:1975:BRBc

[983] Charles B. Schmitt. Book review: Opuscula: The Latin Aristotle by Lorenzo Minio-Paluello; Medieval Latin Aristotle Commentaries by Charles H. Lohr; A Bibliography of Aristotle Editions, 1501–1600 by F. Edward Cranz. Isis, 66(3):422–425, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228867.

Thackray:1975:BRB

[984] Arnold Thackray. Book review: The Life of Captain James Cook by J. C. Beaglehole. Isis, 66(3):425–426, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228868.

Heilbron:1975:BRBb

[985] J. L. Heilbron. Book review: Martinus Van Marum, Life and Work. Volume III by R. J. Forbes. Isis, 66(3):427, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228869.

Brown:1975:BRB

[986] Harcourt Brown. Book review: Science, Language, and the Perspective Mind: Studies in Literature and Thought from Campanella to Bayle by Timothy J. Reiss. Isis, 66(3):427–429, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228870.

Turner:1975:BRB

[987] G. L’E. Turner. Book review: Physics at Seventeenth and Eighteenth-Century Leiden: Philosophy and the New Science in the University by Edward G. Ruestow. Isis, 66(3):429–430, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228871.

Trenn:1975:BRBb

[988] Thaddeus J. Trenn. Book review: Rutherford: The Father of Nuclear Energy by T. E. Allibone. Isis, 66(3):430–431, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228872.

Heimann:1975:BRB

[989] P. M. Heimann. Book review: H. G. J. Moseley. The Life and Letters of an English Physicist, 1887–1915 by J. L. Heilbron. Isis, 66(3):431–432, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228873.

Swenson:1975:BRBb

[990] Loyd S. Swenson, Jr. Book review: The Master of Light: A Biography of Albert A. Michelson by Dorothy Michelson Livingston. Isis, 66(3): 432–434, September 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228874.

Tobey:1975:BRB

[991] Ronald Tobey. Book review: Ellsworth Huntington. His Life and Thought by Geoffrey J. Martin. Isis, 66(3):434–435, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228875.

Forman:1975:BRBd

[992] Paul Forman. Book review: The Physicist’s Conception of Nature by Jagdish Mehra. Isis, 66(3):435–436, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/228876.

Anonymous:1975:BMc

[993] Anonymous. Back matter. Isis, 66(3):437–438, September 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228877.

Anonymous:1975:FMd

[994] Anonymous. Front matter. Isis, 66(4):439–441, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228922.

Greene:1975:FAC

[995] John C. Greene, Robert P. Multhauf, Arnold Thackray, George Basalla, Derek de Solla Price, Charles C. Gillispie, and Erwin N. Hiebert. Fiftieth anniversary celebration of the society. Isis, 66(4):442–482, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228923.

Lesch:1975:RIE

[996] John E. Lesch. The role of isolation in evolution: George J. Romanes and John T. Gulick. Isis, 66(4):483–503, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228924.

Rehbock:1975:HHO

[997] Philip F. Rehbock. Huxley, haeckel, and the oceanographers: The case of bathybius haeckelii. Isis, 66(4):504–533, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228925.

Rudwick:1975:CSH

[998] Martin J. S. Rudwick. Caricature as a source for the history of science: De la Beche’s anti-lyellian sketches of 1831. Isis, 66(4):534–560, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228926.

Rosen:1975:ECD

[999] Edward Rosen. Éloge: C. Doris Hellman, August 28, 1910–March 28, 1973. Isis, 66(4):561–562, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228927.

Anonymous:1975:Nd

[1000] Anonymous. News. Isis, 66(4):563–565, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228928.

Badash:1975:BRB

[1001] Lawrence Badash. Book review: Marie Curie. Isis, 66(4):566–568, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228929.

Wisan:1975:BRBb

[1002] Winifred L. Wisan. Book review: Two New Sciences, Including Centers of Gravity and Force of Percussion. Isis, 66(4):568–569, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228930.

Heilbron:1975:BRBc

[1003] J. L. Heilbron. Book review: An Index to Biographical Fragments in Unspecialized Scientific Journals by E. Scott Barr. Isis, 66(4):569–570, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228931.

Seidel:1975:BRB

[1004] Robert W. Seidel. Book review: Scottish Scientific Instrument-Makers, 1600–1900 by D. J. Bryden. Isis, 66(4):570, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228932.

Himrod:1975:BRB

[1005] David K. Himrod. Book review: Science and Religious Belief: A Selection of Recent Historical Studies by C. A. Russell. Isis, 66(4):570–572, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228933.

Tropp:1975:BRB

[1006] Henry S. Tropp. Book review: The Origins of Digital Computers: Selected Papers by Brian Randell. Isis, 66(4):572–573, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228934.

Gibbs:1975:BRB

[1007] Sharon Gibbs. Book review: The Armillary Sphere by Seymour Robins; The Astrolabe by Harold Saunders. Isis, 66(4):573–574, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228935.

Egerton:1975:BRBb

[1008] Frank N. Egerton, III. Book review: A Century of Fisheries in North America by Norman G. Benson. Isis, 66(4):574–575, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228936.

Meggers:1975:BRB

[1009] Betty J. Meggers. Book review: A History of American Archaeology by Gordon R. Willey; Jeremy A. Sabloff. Isis, 66(4):575–576, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/228937.

Westman:1975:BRB

[1010] Robert S. Westman. Book review: The Scientific World of Copernicus. On the Occasion of the 500th Anniversay of His Birth, 1473–1973 by Barbara Bieńkowska; Christina Cenkalska. Isis, 66(4):576–577, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228938.

Rousseau:1975:BRBb

[1011] G. S. Rousseau. Book review: Science, Medicine and Society in the Renaissance: Essays to Honor Walter Pagel by Allen G. Debus. Isis, 66 (4):577–579, December 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228939.

Renaldo:1975:BRB

[1012] John J. Renaldo. Book review: Cristofano and the Plague: A Study in the History of Public Health in the Age of Galileo by Carlo M. Cipolla. Isis, 66(4):579, December 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228940.

Chapin:1975:BRB

[1013] Seymour L. Chapin. Book review: The Birth of Scientific Navigation. The Solving in the 18th Century of the Problem of Finding Longitude at Sea by Eric G. Forbes. Isis, 66(4):579–580, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228941.

Knight:1975:BRB

[1014] David M. Knight. Book review: Erasmus Darwin by Donald M. Hassler. Isis, 66(4):580, December 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228942.

Brieger:1975:BRB

[1015] Gert H. Brieger. Book review: Geschichte des Hospitals. Band III: Nordamerika: Kolonialzeit (1600–1776) by Dieter Jetter. Isis, 66(4):580– 581, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/228943.

Dobbs:1975:BRB

[1016] Betty Jo Teeter Dobbs. Book review: The Rosicrucian Enlightenment by Frances A. Yates. Isis, 66(4):581–582, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/228944.

Dauben:1975:BRB

[1017] Joseph W. Dauben. Book review: Die Mathematik und ihre Dozenten an der Berliner Universität 1810–1920. Stationen auf dem Wege eines mathematischen Zentrums von Weltgeltung by Kurt-R. Biermann. Isis, 66(4):582–585, December 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228945.

Laqueur:1975:BRB

[1018] Thomas W. Laqueur. Book review: Science for the People: The Origins of the School Science Curriculum in England by David Layton. Isis, 66 (4):585–586, December 1975. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/228946.

Ghiselin:1975:BRB

[1019] Michael T. Ghiselin. Book review: F. A. Smitt, Marine Bryozoa, and an Application of the Darwinian Principle of Descent in 1868 by Thomas J. M. Schopf; Edward L. Bassett. Isis, 66(4):586, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228947.

Pyenson:1975:BRBb

[1020] Lewis Pyenson. Book review: Einstein and the Generations of Science by Lewis S. Feuer. Isis, 66(4):586–590, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228948.

Miller:1975:BRB

[1021] Arthur I. Miller. Book review: Philosophical Problems of Space and Time by Adolf Grünbaum. Isis, 66(4):590–594, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228949.

Schroeder-Gudehus:1975:BRB

[1022] Brigitte Schroeder-Gudehus. Book review: The Chaining of Prometheus. Evolution of a Power Structure for Canadian Science by F. Ronald Hayes. Isis, 66(4):594–595, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/228950.

McGucken:1975:BRB

[1023] William McGucken. Book review: The Government of Science in Britain by J. B. Poole; Kay Andrews. Isis, 66(4):595–596, December 1975. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228951.

Olch:1975:BRB

[1024] Peter D. Olch. Book review: Heritage of Excellence: The Johns Hopkins Medical Institutions, 1914–1947 by Thomas B. Turner. Isis, 66(4):596– 597, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/228952.

Anonymous:1975:BMd

[1025] Anonymous. Back matter. Isis, 66(4):598–599, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/228953.

Anonymous:1975:VI

[1026] Anonymous. Volume information. Isis, 66(4):600–615, December 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/228921.

Anonymous:1975:FMe

[1027] Anonymous. Front matter. Isis, 66(5):1–4, ???? 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/229591.

Neu:1975:OHC

[1028] John Neu. One-hundredth critical bibliography of the history of science and its cultural influences (to January 1975). Isis, 66(5):5–199+201– 223, ???? 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229592.

Anonymous:1975:BMe

[1029] Anonymous. Back matter. Isis, 66(5):??, ???? 1975. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/229593.

Anonymous:1976:FMa

[1030] Anonymous. Front matter. Isis, 67(1):1–3, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231128.

Multhauf:1976:DBD

[1031] Robert P. Multhauf. Dedication to bern dibner. Isis, 67(1):4–6, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231129.

Heilbron:1976:RST

[1032] J. L. Heilbron. Robert symmer and the two electricities. Isis, 67(1): 7–20, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231130.

Home:1976:ATC

[1033] R. W. Home. Aepinus, the tourmaline crystal, and the theory of electricity and magnetism. Isis, 67(1):21–30, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231131.

Finn:1976:HET

[1034] Bernard S. Finn. History of electrical technology: The state of the art. Isis, 67(1):31–35, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231132.

DeKosky:1976:WCF

[1035] Robert K. DeKosky. William Crookes and the fourth state of matter. Isis, 67(1):36–60, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231133.

Trenn:1976:RAB

[1036] Thaddeus J. Trenn. Rutherford on the alpha–beta–gamma classification of radioactive rays. Isis, 67(1):61–75, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231134.

Cassedy:1976:MAM

[1037] James H. Cassedy. The microscope in American medical science, 1840– 1860. Isis, 67(1):76–97, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231135.

Hine:1976:MV

[1038] William L. Hine. Mersenne variants. Isis, 67(1):98–103, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231136.

Reid:1976:LE

[1039] Robert Reid, Georg Nádor, and N. L. Rabinovitch. Letters to the Editor. Isis, 67(1):103–105, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231137.

Anonymous:1976:Na

[1040] Anonymous. News. Isis, 67(1):106–108, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231138.

Kauffman:1976:BRB

[1041] George B. Kauffman. Book review: Graphic Representations of the Periodic System during One Hundred Years. Isis, 67(1):109–111, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231139.

Thibodeau:1976:BRB

[1042] Kenneth F. Thibodeau. Book review: Science and Creation: From Eternal Cycles to an Oscillating Universe by Stanley L. Jaki. Isis, 67(1):

112, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231140.

Warner:1976:BRB

[1043] Deborah Jean Warner. Book review: Woman in Science by H. J. Mozans. Isis, 67(1):112–113, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231141.

Kohler:1976:BRB

[1044] Robert E. Kohler. Book review: Selected Bibliography of Biographical Data for the History of Biochemistry since 1800 by Joseph S. Fruton. Isis, 67(1):113, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231142.

Glick:1976:BRB

[1045] Thomas F. Glick. Book review: Bibliografía histórica sobre la ciencia y la técnica en España by J. M. López Piñero. Isis, 67(1):114, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231143.

Pecchioni:1976:BRB

[1046] Frank Pecchioni. Book review: The Structure of Scientific Theories by Frederick Suppe. Isis, 67(1):114–116, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231144.

Dyck:1976:BRB

[1047] Martin Dyck. Book review: Science and the Arts: A Study in Relationships from 1600–1900 by Jacob Opper. Isis, 67(1):116–117, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231145.

Egerton:1976:BRBa

[1048] Frank N. Egerton, III. Book review: Territory by Allen W. Stokes. Isis, 67(1):117–118, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231146.

Savage-Smith:1976:BRB

[1049] Emilie Savage-Smith. Book review: Galenism. Rise and Decline of a Medical Philosophy by Owsei Temkin. Isis, 67(1):118–119, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231147.

Kaplan:1976:BRB

[1050] Barbara Beigun Kaplan. Book review: A Testament of Alchemy, Being the Revelations of Morienus, Ancient Adept and Hermit of Jerusalem to Khalid Ibn Yazid ibn Mu’awiyya, King of the Arabs, of the Divine Secrets of the Magisterium and Accomplishment of the Alchemical Art by Lee Stavenhagen. Isis, 67(1):119–121, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231148.

vanderWaerden:1976:BRB

[1051] B. L. van der Waerden. Book review: The Place of Astronomy in the Ancient World by F. R. Hodson. Isis, 67(1):121, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231149.

Harkins:1976:BRB

[1052] Michael J. Harkins. Book review: Galen on Psychology, Psychopathology, and Functions and Diseases of the Nervous System. An Analysis of His Doctrines, Observations and Experiments by Rudolph E. Siegel. Isis, 67(1):121–123, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231150.

Hall:1976:BRBa

[1053] Bert S. Hall. Book review: Mappae clavicula: A Little Key to the World of Medieval Techniques by Cyril Stanley Smith; John G. Hawthorne. Isis, 67(1):123–124, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231151.

Stieb:1976:BRB

[1054] Ernst W. Stieb. Book review: Al-Biruni’s Book on Pharmacy and Materia Medica: Introduction, Commentary and Evaluation. Part II by Sami K. Hamarneh. Isis, 67(1):124–125, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231152.

Love:1976:BRB

[1055] J. Barry Love. Book review: Doctor John Mitchell: The Man Who Made the Map of North America by Edmund Berkeley; Dorothy Smith Berkeley. Isis, 67(1):125–126, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231153.

Haase:1976:BRB

[1056] Rudolf Haase. Book review: Der Musiktheoretiker Johannes Kepler by Michael Dickreiter. Isis, 67(1):126–127, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231154.

Applebaum:1976:BRB

[1057] Wilbur Applebaum. Book review: Internationales Kepler-Symposium, Weil der Stadt 1971 by Fritz Krafft; Karl Meyer; Bernhard Sticker. Isis, 67(1):127–128, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231155.

Multhauf:1976:BRBa

[1058] Robert P. Multhauf. Book review: Correspondance du P. Marin Mersenne, Religieux Minime. Volume XII: 1643 by Marin Mersenne; Paul Tannery; Cornélis De Waard; Bernard Rochot; Tables et index cumulatif des tomes I à X (années 1617–1641) by Cornélis De Waard. Isis, 67(1):128–129, March 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231156.

Shapiro:1976:BRBa

[1059] Alan E. Shapiro. Book review: The Unpublished First Version of Isaac Newton’s Cambridge Lectures on Optics 1670–72: A Facsimile of the Autograph, Now Cambridge University Library MS. Add. 4002 by Isaac Newton. Isis, 67(1):129–130, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231157.

Home:1976:BRB

[1060] R. W. Home. Book review: Benjamin Franklin: New World Physicist by Raymond J. Seeger. Isis, 67(1):130–131, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231158.

McMillan:1976:BRB

[1061] Cynthia Anne McMillan. Book review: Die Hauptschriften zur Dyadik von G. W. Leibniz. Ein Beitrag zur Geschichte des binären Zahlensystems by Hans J. Zacher. Isis, 67(1):131–132, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231159.

Bromberg:1976:BRB

[1062] Joan Bromberg. Book review: Fields of Force. The Development of a World View from Faraday to Einstein by William Berkson. Isis, 67(1): 132–134, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231160.

Pierson:1976:BRB

[1063] Stuart Pierson. Book review: Leçons sur la philosophie chimique and Essai de statique chimique des êtres organisés by Jean-Baptiste Dumas; Albert Bruylants. Isis, 67(1):134–137, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231161.

Steffens:1976:BRB

[1064] Henry John Steffens. Book review: The Discovery of the Conservation of Energy by Yehuda Elkana. Isis, 67(1):137–139, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231162.

Cannon:1976:BRB

[1065] W. F. Cannon. Book review: Darwin on Man by Howard E. Gruber; Paul H. Barrett; Howard Gruber; Paul Barrett. Isis, 67(1):139–141, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231163.

Kempler:1976:BRB

[1066] K. Kempler and F. Szabadváry. Book review: Arzneimittel-Standardisierung im 19. Jahrhundert in den Pharmakopöen Deutschlands, Frankreichs, Grossbritanniens und der Vereinigten Staaten von Amerika by Erika Hickel. Isis, 67(1):141, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231164.

Dauben:1976:BRB

[1067] Joseph W. Dauben. Book review: Briefe an David Hilbert by Hermann Minkowski; L. Rüdenberg; H. Zassenhaus. Isis, 67(1):142–143, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231165.

Turner:1976:BRB

[1068] R. Steven Turner. Book review: Zur Entwicklung der Wissenschaftspolitik in Deutschland, 1750–1914 by Frank R. Pfetsch. Isis, 67(1):143–145, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231166.

Bannister:1976:BRBa

[1069] Robert C. Bannister. Book review: The Sociologists of the Chair: A Radical Analysis of the Formative Years of North American Sociology 1883–1922 by Herman Schwendinger; Julia R. Schwendinger. Isis, 67(1): 145–146, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231167.

White:1976:BRB

[1070] George W. White. Book review: The Strata of Derbyshire by White Watson; Trevor D. Ford. Isis, 67(1):146–147, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231168.

Badash:1976:BRBa

[1071] Lawrence Badash. Book review: Nuclear Navy, 1946–1962 by Richard G. Hewlett; Francis Duncan. Isis, 67(1):147–148, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231169.

McGucken:1976:BRB

[1072] William McGucken. Book review: Organizing for Science in Britain by Ian Varcoe. Isis, 67(1):148–149, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231170.

Tobey:1976:BRB

[1073] Ronald Tobey. Book review: Rh: The Intimate History of a Disease and Its Conquest by David R. Zimmerman. Isis, 67(1):149–150, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231171.

Anonymous:1976:BMa

[1074] Anonymous. Back matter. Isis, 67(1):151–156, March 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231172.

Anonymous:1976:FMb

[1075] Anonymous. Front matter. Isis, 67(2):157–159, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230920.

McVaugh:1976:JBR

[1076] Michael McVaugh and Seymour H. Mauskopf. J. b. Rhine’s extra-sensory perception and its background in psychical research. Isis, 67(2):160– 189, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230921.

Kerker:1976:SMR

[1077] Milton Kerker. The Svedberg and molecular reality. Isis, 67(2):190–216, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230922. See comments

[1329].

Donovan:1976:PCN

[1078] Arthur Donovan. Pneumatic chemistry and Newtonian natural philosophy in the Eighteenth Century: William Cullen and Joseph Black. Isis, 67(2):217–228, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230923.

LeGrand:1976:BES

[1079] H. E. Le Grand. Berthollet’s essai de statique chimique and acidity. Isis, 67(2):229–238, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230924.

Drake:1976:EMG

[1080] Stillman Drake. The evolution of de motu (galileo gleanings xxiv). Isis, 67(2):239–250, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230925.

Kruk:1976:PAA

[1081] Remke Kruk. Pseudo-Aristotle: An Arabic version of Problemata physica X. Isis, 67(2):251–256, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230926.

Rabkin:1976:TFS

[1082] Yakov M. Rabkin. Trends and forces in the Soviet history of chemistry. Isis, 67(2):257–273, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230927.

Cane:1976:JAB

[1083] Edric Cane and Thomas L. Hankins. Jean d’Alembert between Descartes and Newton: A critique of Thomas L. Hankins’ position. Isis, 67(2):274– 278, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230928.

Stuewer:1976:N

[1084] Roger H. Stuewer. News. Isis, 67(2):279–283, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230929.

Keiner:1976:BRBa

[1085] Melvyn Keiner. Book review: Life Sciences in the Twentieth Century. Isis, 67(2):284–286, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230930.

Breidert:1976:BRBa

[1086] Wolfgang Breidert. Book review: Philosophy and Mathematics. From Plato to the Present by Robert J. Baum. Isis, 67(2):286–287, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230931.

Trenn:1976:BRBa

[1087] Thaddeus J. Trenn. Book review: The Cavendish Laboratory 1874–1974 by J. G. Crowther. Isis, 67(2):287–289, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230932.

Badash:1976:BRBb

[1088] Lawrence Badash. Book review: Natural Philosophy at Dartmouth. From Surveyors’ Chains to the Pressure of Light by Sanborn C. Brown; Leonard M. Rieser. Isis, 67(2):289–290, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230933.

Smart:1976:BRB

[1089] Christopher C. Smart. Book review: Let Us Be Honest and Modest: Technology and Society in Canadian History by Bruce Sinclair; Norman R. Ball; James O. Peterson; A Curious Field Book: Science and Society in Canadian History by Trevor H. Levere; Richard A. Jarrell. Isis, 67(2): 290–292, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230934.

Mortimer:1976:BRB

[1090] Joanne S. Mortimer. Book review: Comparative Studies in Science and Society by Sal P. Restivo; Christopher K. Vanderpool. Isis, 67(2):292– 293, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230935.

Snyder:1976:BRB

[1091] Walter F. Snyder. Book review: Science, Technology, and Culture by Henry John Steffens; H. N. Muller. Isis, 67(2):293–294, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230936.

Callicott:1976:BRB

[1092] John Baird Callicott. Book review: Man’s Responsibility for Nature. Ecological Problems and Western Traditions by John Passmore. Isis, 67(2):294–295, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230937.

Tropp:1976:BRB

[1093] Henry S. Tropp. Book review: The Computer from Pascal to von Neumann by Herman H. Goldstine. Isis, 67(2):295–297, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230938.

Koppelman:1976:BRB

[1094] Elaine Koppelman. Book review: Origins of Modern Algebra by Lubos Nový; Jaroslav Tauer. Isis, 67(2):297, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230939.

VanHelden:1976:BRB

[1095] Albert Van Helden. Book review: Mapping of the Moon: Past and Present by Zdenek Kopal; Robert W. Carder. Isis, 67(2):297–298, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230940.

Hufbauer:1976:BRB

[1096] Karl Hufbauer. Book review: Der Chemiker im Wandel der Zeiten. Skizzen zur geschichtlichen Entwicklung des Berufsbildes by Eberhard Schmauderer. Isis, 67(2):298–299, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230941.

Multhauf:1976:BRBb

[1097] Robert P. Multhauf. Book review: Studien zur Geschichte der Mineralnamen in Pharmazie, Chemie und Medizin von den Anfängen bis Paracelsus by Dietlinde Goltz. Isis, 67(2):299–300, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230942.

Krikorian:1976:BRB

[1098] A. D. Krikorian. Book review: Festschrift für Claus Nissen zum siebzigsten Geburtstag, 2 September 1971 by Elisabeth Geck; Guido Pressler. Isis, 67(2):300, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230943.

Fee:1976:BRB

[1099] Elizabeth Fee. Book review: Readings in the History of Anthropology by Regna Darnell. Isis, 67(2):300–301, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230944.

Brozek:1976:BRB

[1100] Josef Brožek. Book review: The Roots of Psychology: A Sourcebook in the History of Ideas by Solomon Diamond; Selected Readings in the History of Psychology: A PSI Companion by Fred S. Keller. Isis, 67(2): 301–303, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230945.

Mayr:1976:BRB

[1101] Otto Mayr. Book review: Contributions to a Philosophy of Technology: Studies in the Structure of Thinking in the Technological Sciences by Friedrich Rapp. Isis, 67(2):303–304, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230946.

Guedon:1976:BRB

[1102] Jean-Claude Guédon. Book review: Histoire de l’enseignement en France by Antoine Léon. Isis, 67(2):304, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230947.

Risse:1976:BRB

[1103] Guenter B. Risse. Book review: The House of Life: Magical and Medical Science in Ancient Egypt by Paul Ghalioungui. Isis, 67(2):304–306, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230948.

Moravcsik:1976:BRB

[1104] J. M. Moravcsik. Book review: Time and Necessity. Studies in Aristotle’s Theory of Modality by Jaakko Hintikka. Isis, 67(2):306–307, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230949.

King:1976:BRB

[1105] David A. King. Book review: Al-Bahir en Algèbre d’As-Samaw’al by Salah Ahmad; Roshdi Rashed. Isis, 67(2):307–308, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230950.

Sivin:1976:BRBa

[1106] N. Sivin. Book review: Chu-Ni dai jiten (Unabridged Chinese–Japanese Dictionary) by Aichi Daigaku Chu-Ni dai jiten Hensansho; A New Practical Chinese–English Dictionary by Liang Shih-ch’iu; Lin Yutang’s Chinese–English Dictionary of Modern Usage. Isis, 67(2):308–309, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230951.

Vogel:1976:BRB

[1107] Kurt Vogel. Book review: Chinese Mathematics in the Thirteenth Century. The Shu-shu chiu-chang of Ch’in Chiu-shao by Ulrich Libbrecht. Isis, 67(2):309–311, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230952.

Baker:1976:BRB

[1108] Keith Michael Baker. Book review: Essai sur l’application de l’analyse à la probabilité des décisions rendues à la pluralité des voix by M. le Marquis de Condorcet; Condorcet. Mathématique et société by Roshdi Rashed. Isis, 67(2):311–312, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230953.

Shapiro:1976:BRBb

[1109] Alan E. Shapiro. Book review: De homine: Traité de l’homme by Thomas Hobbes; Paul-Marie Maurin. Isis, 67(2):312–313, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230954.

Costabel:1976:BRB

[1110] Pierre Costabel. Book review: Leibniz in Paris (1672–1676). His Growth to Mathematical Maturity by Joseph Ehrenfried Hofmann. Isis, 67(2):

313–315, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230955.

Gould:1976:BRB

[1111] Henry W. Gould. Book review: Die mathematischen Studien von G. W. Leibniz zur Kombinatorik by Eberhard Knobloch. Isis, 67(2):315– 316, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230956.

Kaunzner:1976:BRB

[1112] Wolfgang Kaunzner. Book review: Die Coss von Jost Bürgi in der Redaktion von Johannes Kepler, Ein Beitrag zur frühen Algebra by Martha List; Volker Bialas. Isis, 67(2):316–317, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230957.

Shea:1976:BRB

[1113] William R. Shea. Book review: Vita di Galileo by Maria Luisa Righini Bonelli. Isis, 67(2):318, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230958.

MacLachlan:1976:BRB

[1114] James MacLachlan. Book review: The Intellectual Revolution of the Seventeenth Century by Charles Webster. Isis, 67(2):318–319, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230959.

Blackmore:1976:BRB

[1115] John T. Blackmore. Book review: Theory of Science by Bernard Bolzano; Rolf George; Theory of Science by Bernard Bolzano; Jan Berg; Burnham Terrell. Isis, 67(2):319–321, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230960.

Elliott:1976:BRB

[1116] Clark A. Elliott. Book review: John Tyndall, Natural Philosopher, 1820– 1893. Catalogue of Correspondence, Journals and Collected Papers by James R. Friday; Roy M. MacLeod; Philippa Shepard. Isis, 67(2):321, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230961.

Landefeld:1976:BRB

[1117] Seth Landefeld. Book review: Sir Oliver Lodge: Physical Researcher and Scientist by W. P. Jolly. Isis, 67(2):321–323, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230962.

Sherwood:1976:BRB

[1118] Morgan Sherwood. Book review: Nature and the American Mind: Louis Agassiz and the Culture of Science by Edward Lurie. Isis, 67(2):323, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230963.

Kohlstedt:1976:BRBa

[1119] Sally Gregory Kohlstedt. Book review: The Maturing of American Science: A Portrait of Science in Public Life Drawn from the Presidential Addresses of the American Association for the Advancement of Science 1920–1970 by Robert A. Kargon. Isis, 67(2):323–325, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230964.

Yoxen:1976:BRB

[1120] E. J. Yoxen. Book review: The Path to the Double Helix by Robert Olby. Isis, 67(2):325–326, June 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230965.

Anonymous:1976:BMb

[1121] Anonymous. Back matter. Isis, 67(2):327–330, June 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230966.

Anonymous:1976:FMc

[1122] Anonymous. Front matter. Isis, 67(3):331–332, September 1976. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230677.

Nye:1976:HMF

[1123] Robert A. Nye. Heredity or milieu: The foundations of modern European criminological theory. Isis, 67(3):334–355, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230678.

Sloan:1976:BLC

[1124] Phillip R. Sloan. The buffon-linnaeus controversy. Isis, 67(3):356–375, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230679.

Paul:1976:SIC

[1125] Harry W. Paul. Scholarship and ideology: The Chair of the General History of Science at the Collège de France, 1892–1913. Isis, 67(3): 376–397, September 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230680.

Naylor:1976:GRE

[1126] Ronald Naylor. Galileo: Real experiment and didactic demonstration. Isis, 67(3):398–419, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230681.

Longrigg:1976:RAT

[1127] James Longrigg. The “roots of all things”. Isis, 67(3):420–438, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230682.

Sivin:1976:EGD

[1128] N. Sivin. Éloge: Giorgio Diaz de Santillana, 1902–1974. Isis, 67(3): 439–443, September 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230683.

Smith:1976:FRJ

[1129] Crosbie W. Smith. Faraday as referee of Joule’s Royal Society paper “On the Mechanical Equivalent of Heat”. Isis, 67(3):444–449, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230684.

Kapil:1976:PCI

[1130] R. N. Kapil and A. K. Bhatnagar. Portuguese contributions to Indian botany. Isis, 67(3):449–452, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230685.

Egerton:1976:DER

[1131] Frank N. Egerton, III. Darwin’s early reading of Lamarck. Isis, 67(3): 452–456, September 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230686.

Drake:1976:WSM

[1132] Stillman Drake. Was Simon Mayr Galileo’s “ancient adversary” in 1607? Isis, 67(3):456–459, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230687.

Anonymous:1976:Nb

[1133] Anonymous. News. Isis, 67(3):460–462, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230688.

Hall:1976:BRBb

[1134] Bert S. Hall. Book review: The Madrid Codices. Isis, 67(3):463–475, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230689.

Sivin:1976:BRBb

[1135] N. Sivin. Book review: Changing Perspectives in the History of Science. Essays in Honour of Joseph Needham by Mikulas Teich; Robert Young. Isis, 67(3):475–477, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230690.

Mullins:1976:BRB

[1136] N. Mullins. Book review: Social Processes of Scientific Development by Richard Whitley. Isis, 67(3):477–478, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230691.

Thoren:1976:BRB

[1137] Victor E. Thoren. Book review: Science Awakening II: The Birth of Astronomy by B. L. van der Waerden. Isis, 67(3):478–479, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230692.

Ruse:1976:BRB

[1138] Michael Ruse. Book review: Studies in the Philosophy of Biology. Reduction and Related Problems by Franciso José Ayala; Theodosius Dobzhansky. Isis, 67(3):479–481, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230693.

Hale:1976:BRB

[1139] Nathan G. Hale, Jr. Book review: Minutes of the Vienna Psychoanalytic Society by Herman Nunberg; Ernst Federn. Isis, 67(3):481–482, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230694.

Galatzer-Levy:1976:BRB

[1140] Robert Galatzer-Levy. Book review: Freud and His Followers by Paul Roazen. Isis, 67(3):482–483, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230695.

Sivin:1976:BRBc

[1141] N. Sivin. Book review: Chugoku igakushi kogi (Lectures on the History of Chinese Medicine) by Peking College of Chinese Medicine; Natsu Saburo. Isis, 67(3):483–484, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230696.

Hankins:1976:BRB

[1142] Thomas L. Hankins. Book review: Actes de la journée Maupertuis (Créteil, 1er décembre 1973). Isis, 67(3):484, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230697.

Granger:1976:BRB

[1143] Gilles G. Granger. Book review: Condorcet: From Natural Philosophy to Social Mathematics by Keith Baker. Isis, 67(3):484–485, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230698.

Breidert:1976:BRBb

[1144] Wolfgang Breidert. Book review: Thomas Reid’s Inquiry–The Geometry of Visibles and the Case for Realism by Norman Daniels. Isis, 67(3): 485–486, September 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230699.

Kargon:1976:BRBa

[1145] Robert H. Kargon. Book review: Sale Catalogues of Libraries of Eminent Persons. Volume XI: Scientists–Elias Ashmole, Edmond Halley, Robert Hooke, John Ray by H. A. Feisenberger. Isis, 67(3):486–487, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230700.

Lindskog:1976:BRB

[1146] Bengt I. Lindskog. Book review: Magnetisörernas tid. Den animala magnetismen i Sverige by Karin Johannisson. Isis, 67(3):487–488, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230701.

Cohen:1976:BRBa

[1147] I. Bernard Cohen. Book review: Marginalia in Newtoni Principia Mathematica 1687 by G. W. Leibniz; E. A. Fellmann. Isis, 67(3):488–489, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230702.

Shapiro:1976:BRBc

[1148] Alan E. Shapiro. Book review: The Mathematical Papers of Isaac Newton. Volume VI: 1684–1691 by Isaac Newton; D. T. Whiteside; M. A. Hoskin; A. Prag. Isis, 67(3):489–491, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230703.

Wisan:1976:BRB

[1149] Winifred Lovell Wisan. Book review: Galileo: A Philosophical Study by Dudley Shapere. Isis, 67(3):491–492, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230704.

Buchwald:1976:BRB

[1150] Jed Z. Buchwald. Book review: Lord Kelvin and the Age of the Earth by Joe D. Burchfield. Isis, 67(3):492–494, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230705.

Burkhardt:1976:BRB

[1151] Richard W. Burkhardt, Jr. Book review: L’introduction du Darwinisme en France au XIXe siècle by Yvette Conry. Isis, 67(3):494–496, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230706.

Cowan:1976:BRB

[1152] Ruth Schwartz Cowan. Book review: Francis Galton: The Life and Work of a Victorian Genius by D. W. Forrest. Isis, 67(3):496–497, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230707.

Herbert:1976:BRB

[1153] Sandra Herbert. Book review: The Comparative Reception of Darwinism by Thomas F. Glick. Isis, 67(3):497–499, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230708.

Reingold:1976:BRB

[1154] Nathan Reingold. Book review: A Century of Zoology at the British Museum through the Lives of Two Keepers, 1815–1914 by Albert E. Gunther. Isis, 67(3):499–500, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230709.

Trenn:1976:BRBb

[1155] Thaddeus J. Trenn. Book review: Ernest Rutherford and the Atom by P. B. Moon. Isis, 67(3):500, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230710.

Kottler:1976:BRB

[1156] Malcolm J. Kottler. Book review: A Short History of the Early American Microscopes by Donald L. Padgitt. Isis, 67(3):500–501, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230711.

Egerton:1976:BRBb

[1157] Frank N. Egerton, III. Book review: Last of the Naturalists: The Career of C. Hart Merriam by Keir B. Sterling; Selected Works by Clinton Hart Merriam. Isis, 67(3):502–503, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230712.

Sokal:1976:BRB

[1158] Michael M. Sokal. Book review: The Life and Mind of John Dewey by George Dykhuizen; Jo Ann Boydston. Isis, 67(3):503–505, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230713.

Farrall:1976:BRB

[1159] Lyndsay A. Farrall. Book review: Grafton Elliot Smith: The Man and His Work by A. P. Elkin; N. W. G. Macintosh. Isis, 67(3):505–506, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230714.

Vandervliet:1976:BRB

[1160] Glenn Vandervliet. Book review: Edwin Broun Fred: Scientist, Administrator, Gentleman by Diane Johnson. Isis, 67(3):506, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230715.

Anonymous:1976:BMc

[1161] Anonymous. Back matter. Isis, 67(3):507–508, September 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230716.

Anonymous:1976:FMd

[1162] Anonymous. Front matter. Isis, 67(4):509–511, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230558.

Sivin:1976:CAM

[1163] N. Sivin. Chinese alchemy and the manipulation of time. Isis, 67(4): 512–526, December 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230559.

Eklund:1976:SWS

[1164] Jon Eklund. Of a spirit in the water: Some early ideas on the aerial dimension. Isis, 67(4):527–550, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230560.

Hoyt:1976:WHP

[1165] William Graves Hoyt. W. h. Pickering’s planetary predictions and the discovery of Pluto. Isis, 67(4):551–564, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230561.

Siegel:1976:BSG

[1166] Daniel M. Siegel. Balfour Stewart and Gustav Robert Kirchhoff: Two independent approaches to “Kirchhoff’s Radiation Law”. Isis, 67(4):565– 600, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230562.

Carmody:1976:PTE

[1167] Francis J. Carmody. Ptolemy’s triangulation of the Eastern Mediterranean. Isis, 67(4):601–609, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230563.

Gillispie:1976:ECB

[1168] Charles C. Gillispie. Éloge: Carl B. Boyer, November 3, 1906–April 26, 1976. Isis, 67(4):610–614, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230564.

Florkin:1976:BRB

[1169] Marcel Florkin. Book review: Claude Bernard and Animal Chemistry, by Frederic L. Holmes. The Emergence of a Scientist. Isis, 67(4):615–617, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230565.

Calinger:1976:BRB

[1170] Ronald Calinger. Book review: Opera omnia, by Leonhard Euler; A. Juskevic; V. Smirnov; W. Habicht. Briefwechsel. Series Quarta A: Commercium Epistolicum. Volume I. Isis, 67(4):617–619, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230566.

Thackray:1976:BRB

[1171] Arnold Thackray. Book review: The Corresponding Societies of the British Association for the Advancement of Science. 1883–1929. A Survey of Historical Records, Archives and Publications by R. M. MacLeod; James R. Friday; C. Gregor. Isis, 67(4):619–620, December 1976. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230567.

Clark:1976:BRB

[1172] Paul W. Clark. Book review: United States Air Force History: A Guide to Documentary Sources by Lawrence J. Paszek. Isis, 67(4):620–621, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230568.

Lenoir:1976:BRB

[1173] Timothy Lenoir. Book review: The Interaction between Science and Philosophy by Yehuda Elkana. Isis, 67(4):621–622, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230569.

Losee:1976:BRB

[1174] John P. Losee. Book review: The Justification of Scientific Change by Carl Kordig. Isis, 67(4):622–624, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230570.

Virieux-Reymond:1976:BRB

[1175] Antoinette Virieux-Reymond. Book review: L’École de pharmacie de l’Université de Lausanne by Jean Hugli; Centenaire de l’École de pharmacie. Isis, 67(4):624, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230571.

Hesse:1976:BRB

[1176] Mary Hesse. Book review: The Emergence of Probability. A Philosophical Study of Early Ideas about Probability, Induction and Statistical Inference by Ian Hacking. Isis, 67(4):624–625, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230572.

Hanle:1976:BRB

[1177] Paul Hanle. Book review: The History of Quantum Theory by Friedrich Hund; Gordon Reece. Isis, 67(4):625–626, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230573.

Heimann:1976:BRB

[1178] P. M. Heimann. Book review: Scottish Philosophy and British Physics 1750–1880. A Study in the Foundations of the Victorian Scientific Style by Richard Olson. Isis, 67(4):626–628, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230574.

Nelson:1976:BRB

[1179] Clifford M. Nelson. Book review: The Dawnseekers. The First History of American Paleontology by Robert West Howard. Isis, 67(4):628–629, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230575.

Farber:1976:BRB

[1180] Paul Lawrence Farber. Book review: Ornithology from Aristotle to the Present by Erwin Stresemann; Cathleen Epstein; G. William Cottrell; Hans J. Epstein. Isis, 67(4):629–630, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230576.

Leary:1976:BRB

[1181] David E. Leary. Book review: The Inward Wits: Psychological Theory in the Middle Ages and the Renaissance by E. Ruth Harvey. Isis, 67(4): 630, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230577.

Potter:1976:BRB

[1182] Paul Potter. Book review: Studien zur altorientalischen und griechischen Heilkunde. Therapie–Arzneibereitung–Rezeptstruktur by Dietlinde Goltz. Isis, 67(4):630–631, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230578.

Mauskopf:1976:BRB

[1183] Seymour H. Mauskopf. Book review: The Precious Metals of Medicine by Geoffrey Marks; William K. Beatty. Isis, 67(4):631–632, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230579.

Cipolla:1976:BRB

[1184] Carlo M. Cipolla. Book review: From Medical Police to Social Medicine: Essays on the History of Health Care by George Rosen. Isis, 67(4):632– 633, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230580.

Chapin:1976:BRB

[1185] Seymour L. Chapin. Book review: Giants in the Sky. A History of the Rigid Airship by Douglas H. Robinson. Isis, 67(4):633–634, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230581.

Hermelink:1976:BRB

[1186] Heinrich Hermelink. Book review: A Commentary upon Biruni’s Kitab Tah.did al-Amakin, an 11th Century Treatise on Mathematical Geography by E. S. Kennedy. Isis, 67(4):634–636, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230582.

Iskandar:1976:BRB

[1187] Albert Z. Iskandar. Book review: Katalog der arabischen alchemistischen Handschriften der Chester Beatty Library. Teil I: Beschreibung der Handschriften by Manfred Ullmann. Isis, 67(4):636–637, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230583.

Schmitt:1976:BRB

[1188] Charles B. Schmitt. Book review: Profezia e ragione: studi sulla cultura del Cinquecento e del Seicento by Cesare Vasoli. Isis, 67(4):637–638, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230584.

Chandler:1976:BRB

[1189] Philip Chandler. Book review: Isaac Newton’s Theory of the Moon’s Motion (1702) by I. Bernard Cohen. Isis, 67(4):638–639, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230585.

Bedini:1976:BRB

[1190] Silvio A. Bedini. Book review: The Art of Breguet by George Daniels. Isis, 67(4):639–640, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230586.

Clulee:1976:BRB

[1191] Nicholas H. Clulee. Book review: Cross and Crucible: Johann Valentin Andreae (1586–1654), Phoenix of the Theologians by John Warwick Montgomery. Isis, 67(4):640–641, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230587.

Theodorides:1976:BRB

[1192] Jean Theodorides. Book review: Bifocal Vision. Novalis’ Philosophy of Nature and Disease by John Neubauer. Isis, 67(4):641–642, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230588.

Morris:1976:BRB

[1193] Robert J. Morris, Jr. Book review: The Argand Burner, Its Origin and Development in France and England, 1780–1800: An Epoch in the History of Science Illustrated by the Life and Work of the Physicist Ami Argand (1750–1803) by Michael Schrøder; Hugh Sheperd. Isis, 67(4): 642–643, December 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230589.

Grmek:1976:BRB

[1194] M. D. Grmek. Book review: The Baglivi Correspondence from the Library of Sir William Osler by Dorothy M. Schullian. Isis, 67(4):643–644, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230590.

Keiner:1976:BRBb

[1195] Melvyn Keiner. Book review: Medical Revolution in France, 1789– 1796 by David M. Vess. Isis, 67(4):644–645, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230591.

Burnham:1976:BRB

[1196] Frederic B. Burnham. Book review: The Discovery of a World in the Moone by John Wilkins; Barbara Shapiro. Isis, 67(4):645–646, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230592.

Cohen:1976:BRBb

[1197] I. Bernard Cohen. Book review: The Library of James Logan of Philadelphia 1674–1751 by Edwin Wolf. Isis, 67(4):646–648, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230593.

Grant:1976:BRB

[1198] Charles T. Grant. Book review: Theoretical Physics and Philosophical Problems. Selected Writings by Ludwig Boltzmann; Brian McGuinness. Isis, 67(4):648–649, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230594.

Bannister:1976:BRBb

[1199] Robert C. Bannister. Book review: Advocacy and Objectivity: A Crisis in the Professionalization of American Social Science, 1865–1905 by Mary O. Furner. Isis, 67(4):649–650, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230595.

Kohlstedt:1976:BRBb

[1200] Sally Gregory Kohlstedt. Book review: Science and the Ante-Bellum American College by Stanley M. Guralnick. Isis, 67(4):650–651, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230596.

Olson:1976:BRB

[1201] Richard Olson. Book review: Joseph Fourier: The Man and the Physicist by John Herivel. Isis, 67(4):651–652, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230597.

Buck:1976:BRB

[1202] Barbara Reeves Buck. Book review: Istruzione e sviluppo industriale in Italia 1859–1914 by Carlo G. Lacaita. Isis, 67(4):652–654, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230598.

Beaver:1976:BRB

[1203] Donald deB. Beaver. Book review: The Papers of Joseph Henry. Volume II: November 1832–December 1835. The Princeton Years by Nathan Reingold. Isis, 67(4):654–655, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230599.

Daub:1976:BRB

[1204] Edward E. Daub. Book review: Josiah Willard Gibbs, American Mathematical Physicist Par Excellence by Raymond J. Seeger. Isis, 67(4): 655–657, December 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230600.

Brieger:1976:BRB

[1205] Gert H. Brieger. Book review: Rise up to Life. A Biography of Howard Walter Florey, Who Gave Penicillin to the World by Lennard Bickel. Isis, 67(4):657, December 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230601.

Kargon:1976:BRBb

[1206] Robert H. Kargon. Book review: Guide to the Robert Andrews Millikan Collection at the California Institute of Technology by Albert F. Gunns; Judith R. Goodstein. Isis, 67(4):657–658, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230602.

Anonymous:1976:BMd

[1207] Anonymous. Back matter. Isis, 67(4):659–661, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230603.

Anonymous:1976:VI

[1208] Anonymous. Volume information. Isis, 67(4):662–677, December 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230557.

Anonymous:1976:FMe

[1209] Anonymous. Front matter. Isis, 67(5):1–4, ???? 1976. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230367.

Neu:1976:OHF

[1210] John Neu. One hundred first critical bibliography of the history of science and its cultural influences (to January 1976). Isis, 67(5):5–15+17– 179+181–200, ???? 1976. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230368.

Anonymous:1977:FMa

[1211] Anonymous. Front matter. Isis, 68(1):1–3, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230369.

Kevles:1977:NSF

[1212] Daniel J. Kevles. The National Science Foundation and the debate over postwar research policy, 1942–1945: A political interpretation of science — the endless frontier. Isis, 68(1):4–26, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230370.

Siraisi:1977:TAB

[1213] Nancy G. Siraisi. Taddeo Alderotti and Bartolomeo da Varignana on the nature of medical learning. Isis, 68(1):27–39, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230371.

Gillespie:1977:DAE

[1214] Neal C. Gillespie. The Duke of Argyll, evolutionary anthropology, and the art of scientific controversy. Isis, 68(1):40–54, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230372.

Clarke:1977:IRD

[1215] Desmond M. Clarke. The impact rules of Descartes’ physics. Isis, 68(1): 55–66, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230373.

Buck:1977:SCP

[1216] Peter Buck. Seventeenth-Century political arithmetic: Civil strife and vital statistics. Isis, 68(1):67–84, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230374.

Ross:1977:PGO

[1217] Frederic R. Ross. Philip Gosse’s omphalos, edmund Gosse’s father and son, and Darwin’s theory of natural selection. Isis, 68(1):85–96, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230375.

Naylor:1977:GNP

[1218] Ronald Naylor. Galileo’s need for precision: The “point” of the fourth day pendulum experiment. Isis, 68(1):97–103, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230376.

Steinsaltz:1977:URP

[1219] Adin Steinsaltz. The uses and the refining of petroleum as mentioned in the talmud. Isis, 68(1):104–105, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230377.

Dobbs:1977:NMS

[1220] B. J. T. Dobbs. Newton manuscripts at the Smithsonian institution. Isis, 68(1):105–107, March 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230378.

Drake:1977:GSE

[1221] Stillman Drake. Galileo on sense experience and foundations of physics. Isis, 68(1):108–110, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230379.

Crowe:1977:WWH

[1222] Michael J. Crowe. Who was H. J. Mozans? Isis, 68(1):111, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230380.

Stuewer:1977:AMS

[1223] Roger H. Stuewer. Annual meeting of the Society, 1976. Isis, 68(1):112– 113, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230381.

Anonymous:1977:Na

[1224] Anonymous. News. Isis, 68(1):114–115, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230382.

Whiteside:1977:BRB

[1225] Derek T. Whiteside. Book review: The Foundations of Newton’s Alchemy; Or “The Hunting of the Greene Lyon”. Isis, 68(1):116–121, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230383.

Rossiter:1977:BRB

[1226] Margaret W. Rossiter. Book review: Black Pioneers of Science and Invention by Louis Haber; Negroes in Science: Natural Science Doctorates, 1876–1969 by James M. Jay. Isis, 68(1):121–122, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230384.

Kargon:1977:BRB

[1227] Robert H. Kargon. Book review: Union Catalogue of Scientific Libraries in the University of Cambridge. Scientific Conference Proceedings 1644– 1972. Isis, 68(1):122–123, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230385.

Davis:1977:BRB

[1228] Audrey B. Davis. Book review: Pictures from the Past of the Healing Arts by Jozsef Antall. Isis, 68(1):123–124, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230386.

Sheynin:1977:BRB

[1229] Oscar B. Sheynin. Book review: The Life and Times of the Central Limit Theorem by William J. Adams. Isis, 68(1):124, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230387.

Heilbron:1977:BRBa

[1230] J. L. Heilbron. Book review: Electricity, Magnetism and Animal Magnetism. A Checklist of Printed Sources, 1600–1850 by Ellen G. Gartrell. Isis, 68(1):124–125, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230388.

Taylor:1977:BRB

[1231] Kenneth L. Taylor. Book review: Sedimentary Rocks: Concepts and History by Albert V. Carozzi. Isis, 68(1):125, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230389.

Ewan:1977:BRBa

[1232] Joseph Ewan. Book review: Voyages to Hawaii before 1860 by Bernice Judd; Helen Yonge Lind. Isis, 68(1):126, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230390.

Sattler:1977:BRB

[1233] Rolf Sattler. Book review: Aux sources de la biologie. Tome 1: Les vingt premiers siècles. La classification by Réjane Bernier. Isis, 68(1):126– 127, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230391.

Rothblatt:1977:BRB

[1234] Sheldon Rothblatt. Book review: The History of the Concept of Association of Ideas by David Rapaport. Isis, 68(1):127–128, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230392.

Burnham:1977:BRB

[1235] John C. Burnham. Book review: Freud: Living and Dying by Max Schur. Isis, 68(1):128–129, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230393.

Symmons-Symonolewicz:1977:BRB

[1236] Konstantin Symmons-Symonolewicz. Book review: Historia Etnografii Polskiej by Malgorzata Terlecka. Isis, 68(1):129–130, March 1977. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230394.

Brozek:1977:BRB

[1237] Josef Brožek. Book review: Names in the History of Psychology: A Biographical Sourcebook by Leonard Zusne. Isis, 68(1):130–131, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230395.

Parascandola:1977:BRBa

[1238] John Parascandola. Book review: An Historical Account of Pharmacology to the Twentieth Century by Chauncey Leake. Isis, 68(1):131–133, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230396.

Greenbaum:1977:BRB

[1239] Louis S. Greenbaum. Book review: The Hospital: A Social and Architectural History by John D. Thompson; Grace Goldin. Isis, 68(1):133–134, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230397.

Brittain:1977:BRBa

[1240] James E. Brittain. Book review: Philosophers and Machines by Otto Mayr. Isis, 68(1):134, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230398.

Norberg:1977:BRBa

[1241] Arthur L. Norberg. Book review: The Maze of Ingenuity: Ideas and Idealism in the Development of Technology by Arnold Pacey. Isis, 68(1): 135, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230399.

Kauffman:1977:BRBa

[1242] George B. Kauffman. Book review: The Game of Wizards: Psyche, Science and Symbol in the Occult by Charles Poncé. Isis, 68(1):135– 136, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230400.

Sokal:1977:BRBa

[1243] Michael M. Sokal. Book review: Studies in the History of Linguistics: Traditions and Paradigms by Dell Hymes. Isis, 68(1):136–137, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230401.

Gillings:1977:BRB

[1244] R. J. Gillings. Book review: The Roots of Civilization. The Cognitive Beginnings of Man’s First Art, Symbol and Notation by Alexander Marshack. Isis, 68(1):137–139, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230402.

Thoren:1977:BRB

[1245] Victor E. Thoren. Book review: A Survey of the Almagest by Olaf Pedersen. Isis, 68(1):139–141, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230403.

Savage-Smith:1977:BRBa

[1246] Emilie Savage-Smith. Book review: Aspects of Greek Medicine by E. D. Phillips. Isis, 68(1):141–142, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230404.

North:1977:BRD

[1247] J. D. North. Book review: Derek de Solla Price: Gears from the Greeks: The Antikythera Mechanism. A Calendar Computer from ca. 80 B.C. Isis, 68(1):142–143, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230405.

Kren:1977:BRB

[1248] Claudia Kren. Book review: A Source Book in Medieval Science by Edward Grant. Isis, 68(1):143–144, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230406.

LaPorte:1977:BRB

[1249] Deirdre La Porte. Book review: Studies in Medieval Philosophy, Science, and Logic. Collected Papers, 1933–1969 by Ernest A. Moody. Isis, 68 (1):144–145, March 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230407.

Wallace:1977:BRB

[1250] William A. Wallace. Book review: Arts and Sciences at Padua. The Studium of Padua before 1350 by Nancy G. Siraisi. Isis, 68(1):146– 147, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230408.

Schioler:1977:BRB

[1251] Thorkild Schiøler. Book review: The Book of Knowledge of Ingenious Mechanical Devices (Kitab fi ma ‘rifat al-h.iyal al-handasiyya) by Ibn al-Razzaz al-Jazari; Donald R. Hill. Isis, 68(1):147–148, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230409.

Anbouba:1977:BRB

[1252] Adel Anbouba. Book review: Al-Fus.ul fil-H .isab al-Hindi. (The Arithmetic of al-Uqlidisi.) by Al-Uqlidisi. Isis, 68(1):148–150, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230410.

Lindberg:1977:BRB

[1253] David C. Lindberg. Book review: The Renaissance Rediscovery of Linear Perspective by Samuel Y. Edgerton. Isis, 68(1):150–152, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230411.

Hall:1977:BRBa

[1254] Marie Boas Hall. Book review: The Chemists and the Word: The Didactic Origins of Chemistry by Owen Hannaway. Isis, 68(1):152–153, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230412.

Nidditch:1977:BRB

[1255] P. H. Nidditch. Book review: Change and Continuity in Seventeenth-Century England by Christopher Hill. Isis, 68(1):153–154, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230413.

Griffiths:1977:BRB

[1256] David M. Griffiths. Book review: Vozniknovenie nauchnykh akademii (seredina XVII–seredina XVIII v.) (The Rise of Scientific Academies [From the Middle of the 17th to the Middle of the 18th Century]) by Iudif’ Khaimovna Kopelevich. Isis, 68(1):154–156, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230414.

Keiner:1977:BRB

[1257] Melvyn Keiner. Book review: Reason, Experiment, and Mysticism in the Scientific Revolution by Maria Luisa Righini Bonelli; William R. Shea. Isis, 68(1):156–158, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230415.

Larson:1977:BRB

[1258] James L. Larson. Book review: Scientific Themes in the Popular Literature and the Poetry of the German Enlightenment, 1720–1760 by Walter Schatzberg. Isis, 68(1):158, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230416.

Trenn:1977:BRBa

[1259] Thaddeus J. Trenn. Book review: Rutherford Correspondence Catalog by Lawrence Badash. Isis, 68(1):158–159, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230417.

Ross:1977:BRB

[1260] Dorothy Ross. Book review: The Legacy of Albion Small by Vernon K. Dibble. Isis, 68(1):159, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230418.

Kauffman:1977:BRBb

[1261] George B. Kauffman. Book review: Van’t Hoff-Le Bel Centennial by O. Bertrand Ramsay. Isis, 68(1):159–161, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230419.

Farber:1977:BRB

[1262] Paul Lawrence Farber. Book review: Henri Dutrochet (Henri du Trochet 1776–1847). Le matérialisme mécaniste et la physiologie générale by Joseph Schiller; Tetty Schiller. Isis, 68(1):161–162, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230420.

Solberg:1977:BRB

[1263] Winton U. Solberg. Book review: Philadelphia’s Philosopher Mechanics: A History of the Franklin Institute, 1824–1865 by Bruce Sinclair. Isis, 68(1):162–163, March 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230421.

Kirby:1977:BRB

[1264] H. W. Kirby. Book review: Radioactivity and Atomic Theory by Thaddeus J. Trenn; Frederick Soddy. Isis, 68(1):163–165, March 1977. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230422.

Brush:1977:BRBa

[1265] Stephen G. Brush. Book review: Landmark Experiments in Twentieth Century Physics by George L. Trigg. Isis, 68(1):165, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230423.

Badash:1977:BRBa

[1266] Lawrence Badash. Book review: The Advisors. Oppenheimer, Teller, and the Superbomb by Herbert York. Isis, 68(1):165–167, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230424.

Anonymous:1977:BMa

[1267] Anonymous. Back matter. Isis, 68(1):168–170, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230425.

Anonymous:1977:FMb

[1268] Anonymous. Front matter. Isis, 68(2):171–173, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230068.

Hankins:1977:TTS

[1269] Thomas L. Hankins. Triplets and triads: Sir William rowan Hamilton on the metaphysics of mathematics. Isis, 68(2):174–193, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230069.

Kren:1977:PLT

[1270] Claudia Kren. Planetary latitudes, the Theorica Gerardi, and Regiomontanus. Isis, 68(2):194–205, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230070.

Gregory:1977:SVD

[1271] Frederick Gregory. Scientific versus dialectical materialism: A clash of ideologies in Nineteenth-Century German radicalism. Isis, 68(2):206– 223, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230071.

Unschuld:1977:TCP

[1272] Ulrike Unschuld. Traditional Chinese pharmacology: An analysis of its development in the Thirteenth Century. Isis, 68(2):224–248, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230072.

Bowler:1977:EDC

[1273] Peter J. Bowler. Edward drinker cope and the changing structure of evolutionary theory. Isis, 68(2):249–265, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230073.

Moyer:1977:RHA

[1274] Albert E. Moyer. Robert Hooke’s ambiguous presentation of “Hooke’s Law”. Isis, 68(2):266–275, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230074.

Glick:1977:EJM

[1275] Thomas F. Glick. Éloge: José María Millás Vallicrosa (1897–1970) and the founding of the history of science in Spain. Isis, 68(2):276–283, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230075.

Agassi:1977:RBA

[1276] Joseph Agassi. Robert Boyle’s anonymous writings. Isis, 68(2):284– 287, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230076.

Tamny:1977:NGD

[1277] Martin Tamny. Newton and Galileo’s dialogue on the great world systems. Isis, 68(2):288–289, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230077.

Anonymous:1977:CNS

[1278] Anonymous. Correction: The national science foundation and the debate over postwar research policy, 1942–1945: A political interpretation of science–the endless frontier. Isis, 68(2):289, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230078.

Anonymous:1977:HSS

[1279] Anonymous. History of science society endowment fund: Contributors through January 31, 1977. Isis, 68(2):290–292, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230079.

Anonymous:1977:Nb

[1280] Anonymous. News. Isis, 68(2):293–295, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230080.

Atkinson:1977:BRB

[1281] R. d’E. Atkinson. Book review: Greenwich Observatory, by Eric G. Forbes; A. J. Meadows; Derek Howse; Francis Place and the Early History of the Greenwich Observatory by Derek Howse. Three Volumes Telling the Story of Britain’s Oldest Scientific Institution, The Royal Observatory at Greenwich and Herstmonceux 1675–1975. Isis, 68(2):296– 299, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230081.

Jayawardene:1977:BRB

[1282] S. A. Jayawardene. Book review: Sources for the History of Science, 1660–1914 by David Knight. Isis, 68(2):299–302, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230082.

Frangsmyr:1977:BRBa

[1283] Tore Frängsmyr. Book review: Svensk lärdomshistoria by Sten Lindroth. Isis, 68(2):302–303, June 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230083.

Koizumi:1977:BRB

[1284] Kenkichiro Koizumi. Book review: Science and Society in Modern Japan: Selected Historical Sources by Shigeru Nakayama; David L. Swain; Eri Yagi. Isis, 68(2):303–305, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230084.

Palter:1977:BRB

[1285] Robert Palter. Book review: The Study of Time by J. T. Fraser; F. C. Haber; G. H. Müller. Isis, 68(2):305–306, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230085.

Christianson:1977:BRB

[1286] Eric H. Christianson. Book review: The Pursuit of Knowledge in the Early American Republic: American Scientific and Learned Societies from Colonial Times to the Civil War by Alexandra Oleson; Sanborn C. Brown. Isis, 68(2):306–308, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230086.

Warner:1977:BRBa

[1287] Deborah Jean Warner. Book review: Scientific Instruments by Harriet Wynter; Anthony Turner. Isis, 68(2):308, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230087.

Guggenbuhl:1977:BRB

[1288] Laura Guggenbuhl. Book review: The Method of Analysis. Its Geometrical Origin and Its General Significance by Jaako Hintikka; Unto Remes. Isis, 68(2):308–309, June 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230088.

Leighly:1977:BRB

[1289] John Leighly. Book review: Geographies of the Mind: Essays in Historical Geosophy in Honor of John Kirtland Wright by David Lowenthal; Martyn J. Bowden. Isis, 68(2):309–310, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230089.

Egerton:1977:BRBa

[1290] Frank N. Egerton, III. Book review: History of the Life Sciences. An Annotated Bibliography by Pieter Smit. Isis, 68(2):310–311, June 1977.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230090.

Porkert:1977:BRB

[1291] Manfred Porkert. Book review: Die Praxis des traditionellen Heilsystems. Unter Einschluss der Pharmazie dargestellt an der heutigen Situation auf Taiwan by Paul Ulrich Unschuld. Isis, 68(2):311–312, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230091.

Baity:1977:BRB

[1292] Elizabeth Chesley Baity. Book review: Megalithic Software. Part 1: England by L. B. Borst; B. M. Borst. Isis, 68(2):312–314, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230092.

Stannard:1977:BRBa

[1293] Jerry Stannard. Book review: Vögel und Vogelfang im alten Mesopotamien by Armas Salonen. Isis, 68(2):314, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230093.

Unguru:1977:BRB

[1294] Sabetai Unguru. Book review: The Evolution of the Euclidean Elements: A Study of the Theory of Incommensurable Magnitudes and Its Significance for Early Greek Geometry by Wilbur Richard Knorr. Isis, 68(2): 314–316, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230094.

Hahm:1977:BRB

[1295] David E. Hahm. Book review: The Pre-Socratics. A Collection of Critical Essays by Alexander P. D. Mourelatos. Isis, 68(2):316–317, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230095.

Eastwood:1977:BRB

[1296] Bruce Eastwood. Book review: A Catalogue of Medieval and Renaissance Optical Manuscripts by David C. Lindberg. Isis, 68(2):317–318, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230096.

Sylla:1977:BRB

[1297] Edith Sylla. Book review: Science and Creation in the Middle Ages. Henry of Langenstein (d. 1397) on Genesis by Nicholas H. Steneck. Isis, 68(2):318–319, June 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230097.

Savage-Smith:1977:BRBb

[1298] Emilie Savage-Smith. Book review: Albucasis on Surgery and Instruments. A Definitive Edition of the Arabic Text with English Translation and Commentary by M. S. Spink; G. L. Lewis. Isis, 68(2):319–321, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230098.

King:1977:BRBa

[1299] Lester S. King. Book review: Achilles Pirmin Gasser 1505–1577. Arzt und Naturforscher, Historiker und Humanist. Volume III: Briefwechsel by Karl Heinz Burmeister. Isis, 68(2):321, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230099.

Hansen:1977:BRB

[1300] Bert Hansen. Book review: The Mathematicall Praeface to the Elements of Geometrie of Euclid of Megara (1570) by John Dee. Isis, 68(2):321– 323, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230100.

Shea:1977:BRB

[1301] William R. Shea. Book review: Francesco Bacone, Dalla Magia alla Scienza by Paolo Rossi. Isis, 68(2):323–324, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230101.

Goetzmann:1977:BRB

[1302] William H. Goetzmann. Book review: Thinkers and Tinkers: Early American Men of Science by Silvio A. Bedini. Isis, 68(2):324–325, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230102.

Hodge:1977:BRB

[1303] M. J. S. Hodge. Book review: Berkeley’s Philosophy of Science by Richard J. Brook. Isis, 68(2):325–326, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230103.

Kearney:1977:BRB

[1304] Hugh F. Kearney. Book review: The Emergence of Science in Western Europe by Maurice Crosland. Isis, 68(2):327, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230104.

Lukens:1977:BRB

[1305] David C. Lukens. Book review: Thomas White’s De Mundo Examined by Thomas Hobbes; Harold Whitmore. Isis, 68(2):328–329, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230105.

Rosen:1977:BRB

[1306] Edward Rosen. Book review: L’Étrenne ou la neige sex-angulaire by Johann Kepler. Isis, 68(2):329–331, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230106.

Gregory:1977:BRB

[1307] Frederick Gregory. Book review: D’Holbach’s Coterie: An Enlightenment in Paris by Alan C. Kors. Isis, 68(2):331–332, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230107.

Finocchiaro:1977:BRBa

[1308] Maurice A. Finocchiaro. Book review: Galilei critico letterario by Tibor Wlassics. Isis, 68(2):332–333, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230108.

Grob:1977:BRB

[1309] Gerald N. Grob. Book review: Textbook of Disturbances of Mental Life or Disturbances of the Soul and Their Treatment by Johann Christian Heinroth; J. Schmorak. Isis, 68(2):333–334, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230109.

Becker:1977:BRB

[1310] Stanley L. Becker. Book review: Images and Enterprise: Technology and the American Photographic Industry, 1839 to 1925 by Reese V. Jenkins.

Isis, 68(2):334–335, June 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230110.

Peterson:1977:BRB

[1311] M. Jeanne Peterson. Book review: The Remarkable Gamgees. A Story of Achievement by Ruth D’Arcy Thompson. Isis, 68(2):335–336, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230111.

Rousseau:1977:BRBa

[1312] G. S. Rousseau. Book review: Creative Malady. Illness in the Lives and Minds of Charles Darwin, Florence Nightingale, Mary Baker Eddy, Sigmund Freud, Marcel Proust, Elizabeth Barrett Browning by George Pickering. Isis, 68(2):336–337, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230112.

Simirenko:1977:BRB

[1313] Alex Simirenko. Book review: Social Thought in Tsarist Russia: The Quest for a General Science of Society, 1861–1917 by Alexander Vucinich. Isis, 68(2):337–338, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230113.

Norberg:1977:BRBb

[1314] Arthur L. Norberg. Book review: The Mackenzie-McNaughton Wartime Letters by Mel Thistle. Isis, 68(2):338–339, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230114.

Anonymous:1977:BMb

[1315] Anonymous. Back matter. Isis, 68(2):340–342, June 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230115.

Anonymous:1977:FMc

[1316] Anonymous. Front matter. Isis, 68(3):343–345, September 1977. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231311.

Stiefel:1977:HST

[1317] Tina Stiefel. The heresy of science: A Twelfth-Century conceptual revolution. Isis, 68(3):346–362, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231312.

Baxter:1977:BWD

[1318] Alice Levine Baxter. E. B. Wilson’s “destruction” of the germ-layer theory. Isis, 68(3):363–374, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231313.

Nye:1977:NCS

[1319] Mary Jo Nye. Nonconformity and creativity: A study of Paul sabatier, chemical theory, and the French scientific community. Isis, 68(3):375– 391, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231314.

Kauffman:1977:EES

[1320] George B. Kauffman. Early experimental studies of cobalt-ammines. Isis, 68(3):392–403, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231315.

Dipert:1977:PTG

[1321] Randall R. Dipert. Peirce’s theory of the geometrical structure of physical space. Isis, 68(3):404–413, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231316.

Yeomans:1977:ONA

[1322] Donald K. Yeomans. The origin of North American astronomy — seventeenth century. Isis, 68(3):414–425, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231317.

Babb:1977:APT

[1323] Stanley E. Babb, Jr. Accuracy of planetary theories, particularly for Mars. Isis, 68(3):426–434, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231318.

Nisio:1977:ETH

[1324] Sigeko Nisio. Éloge: Tetu Hirosige, August 28, 1928–January 7, 1975. Isis, 68(3):435–436, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231319.

Eade:1977:AVC

[1325] J. C. Eade. The accuracy of vincenzo Coronelli’s celestial globe. Isis, 68 (3):437–440, September 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231320.

Tasch:1977:LGM

[1326] Paul Tasch. Lyell’s geochronological model: Published year values for geological time. Isis, 68(3):440–442, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231321.

Stothers:1977:SDR

[1327] Richard Stothers. Is the supernova of A.D. 185 recorded in Ancient Roman literature? Isis, 68(3):443–447, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231322.

Grunbaum:1977:RMR

[1328] Adolf Grunbaum and Arthur I. Miller. Remarks on Miller’s review of philosophical problems of space and time. Isis, 68(3):447–450, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231323.

Schufle:1977:CSM

[1329] J. A. Schufle. Comments on “The Svedberg and Molecular Reality”. Isis, 68(3):450–451, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 231324. See [1077].

Aiton:1977:BRBa

[1330] E. J. Aiton. Book review: Sämtliche Schriften und Briefe, by Gottfried Wilhelm Leibniz; Leibniz-Archiv der Niedersächsischen Landesbibliothek. Dritte Reihe: Mathematischer, Naturwissenschaftlicher und Technischer Briefwechsel. Isis, 68(3):452–453, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231325.

Sharlin:1977:BRB

[1331] Harold Issadore Sharlin. Book review: Science and the Rise of Technology since 1800 by The Open University; Science and Belief: From Copernicus to Darwin. Isis, 68(3):453–456, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231326.

Trenn:1977:BRBb

[1332] Thaddeus J. Trenn. Book review: Erfahrung und Erkenntnis: Vorträge und Aufsätze zur Geschichte der naturwissenschaftlichen Denkweisen 1943–1973 by Bernhard Sticker. Isis, 68(3):456, September 1977. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231327.

Dolby:1977:BRB

[1333] R. G. A. Dolby. Book review: The Philosophy of Karl Popper by Robert John Ackermann. Isis, 68(3):456–457, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231328.

Haber:1977:BRB

[1334] Samuel Haber. Book review: Professions and Monopoly: A Study of Medicine in the United States and Great Britain by Jeffrey L. Berlant. Isis, 68(3):457–458, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231329.

Ellrich:1977:BRB

[1335] Robert J. Ellrich. Book review: Science and the Human Comedy. Natural Philosophy in French Literature from Rabelais to Maupertuis by Harcourt Brown. Isis, 68(3):458–460, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231330.

Kauffman:1977:BRBc

[1336] George B. Kauffman. Book review: The Tao of Physics: An Exploration of the Parallels between Modern Physics and Eastern Mysticism by Fritjof Capra. Isis, 68(3):460–461, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231331.

Rousseau:1977:BRBb

[1337] G. S. Rousseau. Book review: The Electrical Imagination: Electricity in Literature and Music by Brian Coleman. Isis, 68(3):461–462, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231332.

Pierson:1977:BRB

[1338] Stuart Pierson. Book review: The Caloric Theory of Gases from Lavoisier to Regnault by Robert Fox. Isis, 68(3):462–464, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231333.

Gooding:1977:BRB

[1339] David C. Gooding. Book review: Energy: Historical Development of the Concept by R. Bruce Lindsay. Isis, 68(3):464–465, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231334.

Saliba:1977:BRBa

[1340] George Saliba. Book review: Early Physics and Astronomy: A Historical Introduction by Olaf Pedersen; Mogens Pihl. Isis, 68(3):465–466, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231335.

Rocke:1977:BRBa

[1341] Alan J. Rocke. Book review: A Century of Chemistry: The Role of Chemists and the American Chemical Society by Herman Skolnik; Kenneth M. Reese; American Chemists and Chemical Engineers by Wyndham D. Miles. Isis, 68(3):466–467, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231336.

Multhauf:1977:BRB

[1342] Robert P. Multhauf. Book review: The Development of the Chlorinity/ Salinity Concept in Oceanography by William J. Wallace. Isis, 68(3):467, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231337.

Egerton:1977:BRBb

[1343] Frank N. Egerton, III. Book review: British Botanical and Horticultural Literature before 1800, Comprising a History and Bibliography of Botanical and Horticultural Books Printed in England, Scotland, and Ireland from the Earliest Times until 1800 by Blanche Henrey. Isis, 68(3): 467–469, September 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231338.

Rahmani:1977:BRB

[1344] Levy Rahmani. Book review: Soviet Psychology. History, Theory, Content by John McLeish. Isis, 68(3):469–470, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231339.

Parascandola:1977:BRBb

[1345] John Parascandola. Book review: Einf“uhrung in die Pharmaziegeschichte by Georg Edmund Dann. Isis, 68(3):470, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231340.

Nehamas:1977:BRB

[1346] Alexander Nehamas. Book review: Studies in Presocratic Philosophy. Volume II: Eleatics and Pluralists by R. E. Allen; David J. Furley. Isis, 68 (3):470–471, September 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231341.

Savage-Smith:1977:BRBc

[1347] Emilie Savage-Smith. Book review: Ancient Ophthalmological Agents by Harald Nielsen. Isis, 68(3):471–474, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231342.

ONeill:1977:BRB

[1348] Ynez Violé O’Neill. Book review: Arnaldi de Villanova Opera medica omnia. Volume II: Aphorismi de gradibus by Michael R. McVaugh. Isis, 68 (3):474–475, September 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231343.

Veith:1977:BRB

[1349] Ilza Veith. Book review: The Cutting Edge: Early History of the Surgeons of London by R. Theodore Beck. Isis, 68(3):475–476, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231344.

King:1977:BRBb

[1350] David A. King. Book review: The Astronomical Tables of Levi Ben Gerson by Bernard R. Goldstein. Isis, 68(3):476–477, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231345.

McVaugh:1977:BRB

[1351] Michael McVaugh. Book review: Medieval English Medicine by Stanley Rubin. Isis, 68(3):477–478, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231346.

Rosenfeld:1977:BRB

[1352] B. Rosenfeld. Book review: Census of the Exact Sciences in Sanskrit. Series A, Volume III by David Pingree. Isis, 68(3):478, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231347.

Sivin:1977:BRB

[1353] N. Sivin. Book review: Chûgoku no kagaku (Chinese Science) by Yabuuchi Kiyoshi. Isis, 68(3):478–479, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231348.

Brown:1977:BRB

[1354] Theodore M. Brown. Book review: Public Health and the Medical Profession in the Renaissance by Carlo M. Cipolla. Isis, 68(3):479–480, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231349.

Hall:1977:BRBb

[1355] Thomas S. Hall. Book review: La introducción del modelo físicomatemático en la medicina moderna: Análisis de la obra de G. A. Borelli (1608–1679), De motu animalium by Emilio Balaguer Perig“uell. Isis, 68 (3):480–481, September 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231350.

Frangsmyr:1977:BRBb

[1356] Tore Frangsmyr. Book review: Homo sapiens L. Studier i Carl von Linnés naturuppfattning och människolära. (Homo sapiens L. Studies in Linnaeus’ Concept of Nature and His Physical Anthropology) by Gunnar Broberg. Isis, 68(3):481, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231351.

Smith:1977:BRB

[1357] A. D. Smith. Book review: Medicine in Seventeenth-Centry England by Allen G. Debus. Isis, 68(3):482, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231352.

Ewan:1977:BRBb

[1358] Joseph Ewan. Book review: The Tactless Philosopher. Johann Reinhold Forster (1729–1798) by Michael E. Hoare. Isis, 68(3):482–483, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231353.

Ashworth:1977:BRB

[1359] William B. Ashworth, Jr. Book review: John Aubrey and the Realm of Learning by Michael Hunter. Isis, 68(3):483–484, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231354.

Frank:1977:BRB

[1360] Robert G. Frank, Jr. Book review: Les origines de la transfusion sanguine by Jean-Jacques Peumery. Isis, 68(3):484–485, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231355.

Heilbron:1977:BRBb

[1361] J. L. Heilbron. Book review: The Great Instauration. Science, Medicine and Reform, 1626–1660 by Charles Webster. Isis, 68(3):485–487, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231356.

Vogel:1977:BRB

[1362] Morris J. Vogel. Book review: America’s First Hospital: The Pennsylvania Hospital, 1751–1841 by William H. Williams. Isis, 68(3):487–488, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231357.

Skinner:1977:BRB

[1363] Hubert C. Skinner. Book review: Father of Texas Geology: Robert T. Hill by Nancy Alexander. Isis, 68(3):488–489, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231358.

Molella:1977:BRB

[1364] Arthur P. Molella. Book review: Sir Charles Wheatstone, F.R.S., 1802– 1875 by Brian Bowers. Isis, 68(3):489–490, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231359.

Roazen:1977:BRB

[1365] Paul Roazen. Book review: Freud: The Fusion of Science and Humanism. The Intellectual History of Psychoanalysis by John E. Gedo; George H. Pollock. Isis, 68(3):490–491, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231360.

Warner:1977:BRBb

[1366] Deborah Jean Warner. Book review: Lowell and Mars by William Graves Hoyt. Isis, 68(3):491–492, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231361.

Hollinger:1977:BRB

[1367] David A. Hollinger. Book review: The Formation of the American Scientific Community: The American Association for the Advancement of Science, 1848–1860 by Sally Gregory Kohlstedt. Isis, 68(3):492–493, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231362.

Fox:1977:BRB

[1368] Robert Fox. Book review: Auguste Comte and Positivism. The Essential Writings by Gertrud Lenzer. Isis, 68(3):493, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231363.

Rocke:1977:BRBb

[1369] Alan J. Rocke. Book review: Crystals and Compounds: Molecular Structure and Composition in Nineteenth-Century French Science by Seymour H. Mauskopf. Isis, 68(3):493–494, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231364.

Crosland:1977:BRB

[1370] Maurice Crosland. Book review: The Patronage of Science in the Nineteenth Century by G. L’E. Turner. Isis, 68(3):494–496, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231365.

Churchill:1977:BRB

[1371] Frederick B. Churchill. Book review: Starfish, Jellyfish, and the Order of Life. Issues in Nineteenth-Century Science by Mary P. Winsor. Isis, 68 (3):496–497, September 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231366.

Badash:1977:BRBb

[1372] Lawrence Badash. Book review: Energy and Conflict. The Life and Times of Edward Teller by Stanley A. Blumberg; Gwinn Owens. Isis, 68 (3):497–498, September 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231367.

Ebert:1977:BRB

[1373] James D. Ebert. Book review: Crystals, Fabrics and Fields. Metaphors of Organicism in Twentieth–Century Developmental Biology by Donna Jeanne Haraway. Isis, 68(3):498–500, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231368.

Colodny:1977:BRB

[1374] Robert G. Colodny. Book review: Animal Cell Culture and Virology by Robert J. Kuchler. Isis, 68(3):500, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231369.

Costa:1977:BRB

[1375] Albert B. Costa. Book review: P. D. and the Bartlett Group at Harvard 1934–1974: A Group Autobiography on the Occasion of the Bartlett Symposium on Physical Organic Chemistry, Fort Worth, Texas, August 14– 16, 1975. Isis, 68(3):500–501, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231370.

Anonymous:1977:BMc

[1376] Anonymous. Back matter. Isis, 68(3):502–504, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231371.

Anonymous:1977:FMd

[1377] Anonymous. Front matter. Isis, 68(4):505–507, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230005.

Kargon:1977:CMR

[1378] Robert H. Kargon. The conservative mode: Robert a. Millikan and the Twentieth–Century revolution in physics. Isis, 68(4):508–526, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230006.

Laudan:1977:IOB

[1379] Rachel Laudan. Ideas and organizations in British geology: A case study in institutional history. Isis, 68(4):527–538, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230007.

Heilbron:1977:FHF

[1380] J. L. Heilbron. Franklin, haller, and franklinist history. Isis, 68(4):539– 549, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230008.

Farber:1977:DTH

[1381] Paul Lawrence Farber. The development of taxidermy and the history of ornithology. Isis, 68(4):550–566, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230009.

Spencer:1977:TUE

[1382] Frank Spencer and Benjamin Smith Barton. Two unpublished essays on the anthropology of North America by Benjamin Smith Barton. Isis, 68 (4):567–573, December 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230010.

Blackwell:1977:CHB

[1383] Richard J. Blackwell and Christiaan Huygen. Christiaan Huygens’ The Motion of Colliding Bodies. Isis, 68(4):574–597, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230011.

Heilbron:1977:EVF

[1384] J. L. Heilbron. Éloge: Victor Fritz Lenzen, 1890–1975. Isis, 68(4):598– 600, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230012.

Klein:1977:SUP

[1385] Martin J. Klein and Allan Needell. Some unnoticed publications by Einstein. Isis, 68(4):601–604, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230013.

Harries:1977:SCD

[1386] H. C. Harries. Sixteenth-Century dissemination of the coconut palm. Isis, 68(4):605–606, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230014.

Hanle:1977:ESR

[1387] Paul A. Hanle. Erwin Schrödinger’s reaction to Louis de Broglie’s thesis on the quantum theory. Isis, 68(4):606–609, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230015.

Cohen:1977:ASM

[1388] I. Bernard Cohen. Award of the 1976 Sarton Medal to Bern Dibner. Isis, 68(4):610–615, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230016.

Harris:1977:E

[1389] John N. Harris. To the Editor. Isis, 68(4):616–617, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230017.

Anonymous:1977:Nc

[1390] Anonymous. News. Isis, 68(4):618, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230018.

Stannard:1977:BRBb

[1391] Jerry Stannard and Robert P. Multhauf. Book review: Indices naturwissenschaftlich-medizinischer Periodica bis 1850. Isis, 68(4):619– 620, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230019.

Sokal:1977:BRBb

[1392] Michael M. Sokal. Book review: Eminent Contributors to Psychology. Isis, 68(4):621–623, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230020.

Brittain:1977:BRBb

[1393] James E. Brittain. Book review: A History of Engineering and Science in the Bell System: The Early Years (1875–1925) by M. D. Fagen. Isis, 68(4):624, December 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230021.

Reingold:1977:BRB

[1394] Nathan Reingold. Book review: Scientific Knowledge and Sociological Theory by Barry Barnes; Knowledge and Social Imagery by David Bloor. Isis, 68(4):625–626, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230022.

Gould:1977:BRB

[1395] Stephen Jay Gould. Book review: Ancient Eugenics. The Arnold Prize Essay for 1913 by Allen G. Roper. Isis, 68(4):626–627, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230023.

Sesiano:1977:BRB

[1396] Jacques Sesiano. Book review: L’art de l’algèbre de Diophante by Roshdi Rashed; Qust.a ibn Luqa. Isis, 68(4):627–630, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230024.

Kauffman:1977:BRBd

[1397] George B. Kauffman. Book review: The Human Side of Scientists by Ralph E. Oesper. Isis, 68(4):630–631, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230025.

Beaver:1977:BRB

[1398] Donald deB. Beaver. Book review: Time and the Calendars by W. M. O’Neil. Isis, 68(4):632–633, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230026.

Egerton:1977:BRBc

[1399] Frank N. Egerton, III. Book review: Introduction to the History of Mycology by Geoffrey C. Ainsworth. Isis, 68(4):633, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230027.

Theodorides:1977:BRB

[1400] J. Théodoridès. Book review: The Development of Microbiology by Patrick Collard. Isis, 68(4):633–634, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230028.

Yochelson:1977:BRB

[1401] Ellis L. Yochelson. Book review: Evolution and the Diversity of Life. Selected Essays by Ernst Mayr. Isis, 68(4):634–635, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230029.

Parascandola:1977:BRBc

[1402] John Parascandola. Book review: The Battle against Bacteria: A Fresh Look by Peter Baldry. Isis, 68(4):635, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230030.

Savage-Smith:1977:BRBd

[1403] Emilie Savage-Smith. Book review: Galens Traktat “Dass die Kräfte der Seele den Mischungen des Körpers folgen” in arabischer Übersetzung by Hans Heinrich Biesterfeldt. Isis, 68(4):635–636, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230031.

Reich:1977:BRB

[1404] Leonard S. Reich. Book review: Syntony and Spark. The Origins of Radio by Hugh G. J. Aitken. Isis, 68(4):636–638, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230032.

Reynolds:1977:BRB

[1405] Terry S. Reynolds. Book review: Hydraulics in the United States, 1776– 1976 by Hunter Rouse. Isis, 68(4):638–639, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230033.

Major:1977:BRB

[1406] John S. Major. Book review: The Cradle of the East. An Inquiry into the Indigenous Origins of Techniques and Ideas of Neolithic and Early Historic China 5000–1000 B.C. by Ping-ti Ho; Early Chinese Civilization. Anthropological Perspectives by K. C. Chang. Isis, 68(4):639–640, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230034.

Weisheipl:1977:BRB

[1407] James A. Weisheipl. Book review: Suma de los ocho libros de la “Fisica” de Aristóteles (Summa physicorum) by Roberto Grosseteste; J. E. Bolzán; Celina Lértora Mendoza. Isis, 68(4):641, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230035.

Benedek:1977:BRB

[1408] Thomas G. Benedek. Book review: Der “Schüpfheimer Codex,” ein Medizinalbuch aus dem zweiten Viertel des 15. Jahrhunderts by Katharina Wäckerlin–Swiagenin. Isis, 68(4):642, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230036.

Saliba:1977:BRBb

[1409] George A. Saliba. Book review: The Life and Work of Ibn al-Shatir an Arab Astronomer of the 14th Century by E. S. Kennedy; Imad Ghanem. Isis, 68(4):642–643, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230037.

Truesdell:1977:BRB

[1410] C. Truesdell. Book review: Cartelli di sfida matematica by Lodovico Ferrari; Niccolò Tartaglia; Arnaldo Masotti. Isis, 68(4):643–644, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230038.

Aiton:1977:BRBb

[1411] E. J. Aiton. Book review: The Gresham Lectures of John Flamsteed by Eric G. Forbes. Isis, 68(4):644–645, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230039.

Finocchiaro:1977:BRBb

[1412] Maurice A. Finocchiaro. Book review: Galileo against the Philosophers in his Dialogue of Cecco di Ronchitti (1605) and Considerations of Alimberto Mauri (1606) by Stillman Drake. Isis, 68(4):645–647, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230040.

Calinger:1977:BRB

[1413] Ronald Calinger. Book review: Die allgemeinverständlichen Physikbücher der französischen Aufklärung by Andreas Kleinert. Isis, 68(4):647–649, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230041.

Kreiling:1977:BRB

[1414] Frederick C. Kreiling. Book review: Akten des II. Internationalen Leibniz-Kongresses, Hannover, 17–22 Juli 1972. Band I: Begrüssungsansprachen, Gesamtinterpretationen, Geschichte, Recht, Gesellschaftstheorie, Historische Wirkung, Berichte by Kurt Müller; Heinrich Schepers; Wilhelm Totok. Isis, 68(4):649–650, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230042.

Dobbs:1977:BRB

[1415] B. J. T. Dobbs. Book review: The Correspondence of Isaac Newton. Volume V: 1709–1713 by A. Rupert Hall; Laura Tilling. Isis, 68(4): 650–651, December 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230043.

Hankins:1977:BRB

[1416] Thomas L. Hankins. Book review: Fontenelle à la recherche de lui-même (1657–1702) by Alain Niderst. Isis, 68(4):651–652, December 1977. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230044.

Lohne:1977:BRB

[1417] J. A. Lohne. Book review: The Development of Newtonian Optics in England by Henry John Steffens. Isis, 68(4):652, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230045.

Stanton:1977:BRB

[1418] William Stanton. Book review: Passage through the Garden: Lewis and Clark and the Image of the American Northwest by John Logan Allen. Isis, 68(4):652–653, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230046.

Nye:1977:BRB

[1419] Mary Jo Nye. Book review: American and French Culture, 1800–1900. Interchanges in Art, Science, Literature, and Society by Henry Blumenthal. Isis, 68(4):653–654, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230047.

ODowd:1977:BRB

[1420] Patrick O’Dowd. Book review: Alexander Graham Bell and the Conquest of Solitude by Robert V. Bruce. Isis, 68(4):654–655, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230048.

Post:1977:BRB

[1421] Robert C. Post. Book review: Desideratum in Washington: The Intellectual Community in the Capital City, 1870–1900 by J. Kirkpatrick Flack. Isis, 68(4):655–657, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230049.

Cowan:1977:BRB

[1422] Ruth Schwartz Cowan. Book review: A Guide to Francis Galton’s English Men of Science by Victor L. Hilts. Isis, 68(4):657–658, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230050.

Greene:1977:BRB

[1423] John C. Greene. Book review: The Wellesley Index to Victorian Periodicals 1824–1900 by Walter E. Houghton. Isis, 68(4):658–659, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230051.

Kopperl:1977:BRB

[1424] Sheldon J. Kopperl. Book review: Classics in Coordination Chemistry. Part II: Selected Papers (1798–1899) by George B. Kauffman. Isis, 68 (4):659, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230052.

Branagan:1977:BRB

[1425] David Branagan. Book review: Scientists in Nineteenth Century Australia: A Documentary History by Ann Mozley Moyal. Isis, 68(4):659– 661, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230053.

Gerstner:1977:BRB

[1426] Patsy A. Gerstner. Book review: Joseph N. Nicollet on the Plains and Prairies. The Expeditions of 1838–39 with Journals, Letters, and Notes on the Dakota Indians by Edmund C. Bray; Martha Coleman Bray. Isis, 68(4):661, December 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230054.

Berman:1977:BRB

[1427] Morris Berman. Book review: The Emergence of Agricultural Science: Justus Liebig and the Americans, 1840–1880 by Margaret W. Rossiter. Isis, 68(4):661–663, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230055.

Alexander:1977:BRB

[1428] Nancy Alexander. Book review: George Hammell Cook: A Life in Agriculture and Geology, 1818–1889 by Jean Wilson Sidar. Isis, 68(4):663– 664, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230056.

Hounshell:1977:BRB

[1429] David A. Hounshell. Book review: Harpers Ferry Armory and the New Technology: The Challenge of Change by Merritt Roe Smith. Isis, 68 (4):664–665, December 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230057.

Eyde:1977:BRB

[1430] Richard H. Eyde. Book review: Akademie der Wissenschaften und der Literatur Mainz 1949–1974. Isis, 68(4):665, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230058.

Burke:1977:BRB

[1431] John G. Burke. Book review: Chicago 1910–1929 by Carl W. Condit; Chicago 1930–1970 by Carl W. Condit. Isis, 68(4):666–667, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230059.

Brush:1977:BRBb

[1432] Stephen G. Brush. Book review: Science Development: An Evaluation Study by David E. Drew. Isis, 68(4):667–668, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230060.

Glick:1977:BRB

[1433] Thomas F. Glick. Book review: L’oeuvre scientifique by Pierre Teilhard de Chardin; Karl Schmitz-Moormann; Nicole Schmitz-Moormann. Isis, 68(4):668–669, December 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230061.

Allen:1977:BRB

[1434] Garland E. Allen. Book review: The Evolution of an Evolutionist by C. H. Waddington. Isis, 68(4):669–670, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230062.

Meggers:1977:BRB

[1435] Betty J. Meggers. Book review: Archaeological Researches in Retrospect by Gordon R. Willey. Isis, 68(4):670–671, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230063.

Anonymous:1977:BMd

[1436] Anonymous. Back matter. Isis, 68(4):672–674, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230064.

Anonymous:1977:VI

[1437] Anonymous. Volume information. Isis, 68(4):675–692, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230004.

Anonymous:1977:FMe

[1438] Anonymous. Front matter. Isis, 68(5):1–4, ???? 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230555.

Neu:1977:OHS

[1439] John Neu. One hundred second critical bibliography of the history of science and its cultural influences (to January 1977). Isis, 68(5):5–181+183– 202, ???? 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230556.

Anonymous:1978:FMa

[1440] Anonymous. Front matter. Isis, 69(1):1–3, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230604.

ORourke:1978:CJH

[1441] J. E. O’Rourke. A comparison of James Hutton’s principles of knowledge and theory of the Earth. Isis, 69(1):4–20, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230605.

Hilts:1978:AEO

[1442] Victor L. Hilts. Aliis exterendum, or, the origins of the statistical society of London. Isis, 69(1):21–43, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230606.

Plotkin:1978:ECP

[1443] Howard Plotkin. Edward C. Pickering and the endowment of scientific research in America, 1877–1918. Isis, 69(1):44–57, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230607.

Warner:1978:SEW

[1444] Deborah Jean Warner. Science education for women in antebellum America. Isis, 69(1):58–67, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230608.

Corish:1978:ATO

[1445] Denis Corish. Aristotle on temporal order: “Now,” “Before,” and “After”. Isis, 69(1):68–74, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230609.

Burchfield:1978:EJS

[1446] Joe D. Burchfield, Paul L. Farber, and Richard S. Westfall. Éloge: Joseph Schiller, 1906–1977. Isis, 69(1):75–76, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230610.

Frankel:1978:IGN

[1447] Henry R. Frankel. The importance of Galileo’s nontelescopic observations concerning the size of the fixed stars. Isis, 69(1):77–82, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230611.

vanderPas:1978:LTB

[1448] Peter W. van der Pas. The Latin translation of Benjamin Franklin’s letters on electricity. Isis, 69(1):82–85, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230612.

Stuewer:1978:AMS

[1449] Roger H. Stuewer. Annual meeting of the Society, 1977. Isis, 69(1): 85–86, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230613.

Westfall:1978:ASM

[1450] Richard S. Westfall. Award of the 1977 Sarton Medal to D. T. Whiteside. Isis, 69(1):86–87, March 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230614.

Dupree:1978:BRB

[1451] A. Hunter Dupree. Book review: Physics, Patents, and Politics: A Biography of Charles Grafton Page, by Robert Charles Post. 227. Isis, 69(1):88–89, March 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230615.

Hartner:1978:BRB

[1452] Willy Hartner. Book review: Science and Technology in Korea, by Sang-Woon Jeon. Traditional Instruments and Techniques. Isis, 69(1):89– 92, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230616.

Scriba:1978:BRB

[1453] Christoph J. Scriba. Book review: Dictionary of the History of Ideas. Studies of Selected Pivotal Ideas by Philip P. Wiener. Isis, 69(1):92– 93, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230617.

Buck:1978:BRB

[1454] Peter Buck. Book review: Science since Babylon by Derek de Solla Price. Isis, 69(1):93–94, March 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230618.

Rosenkrantz:1978:BRB

[1455] Barbara Gutmann Rosenkrantz. Book review: Guy’s Hospital-250 Years by Clive E. Handler. Isis, 69(1):94–95, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230619.

Olch:1978:BRB

[1456] Peter D. Olch. Book review: Adventures in Medical Research: A Century of Discovery at Johns Hopkins by A. McGehee Harvey. Isis, 69(1): 95, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230620.

Szabadvary:1978:BRB

[1457] Ferenc Szabadváry. Book review: A Magyar Tudományos Akadémia másfél évszázada 1825–1975 (One and a Half Centuries of the Hungarian Academy of Sciences 1825–1975) by Kónya Sándor; Zsigmond Pál Pach; A Magyar Tudományos Akadémia állandó bizottságai 1854–1949 (Permanent Committees of the Hungarian Academy of Sciences 1854–1949) by Jánosné Fráter. Isis, 69(1):96–97, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230621.

Applebaum:1978:BRB

[1458] Wilbur Applebaum. Book review: Science and Society: Past, Present, and Future by Nicholas H. Steneck. Isis, 69(1):97–98, March 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230622.

Rosen:1978:BRB

[1459] Edward Rosen. Book review: Die Genesis der kopernikanischen Welt by Hans Blumenberg. Isis, 69(1):98–99, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230623.

Machamer:1978:BRB

[1460] Peter Machamer. Book review: Theories of Vision from Al-Kindi to Kepler by David C. Lindberg. Isis, 69(1):99–100, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230624.

Heimann:1978:BRB

[1461] P. M. Heimann. Book review: The History of Physics and the Philosophy of Science: Selected Essays by Armin Teske. Isis, 69(1):100, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230625.

Sommer:1978:BRB

[1462] Margaret F. Sommer. Book review: Natural History in America, from Mark Catesby to Rachel Carson by Wayne Hanley. Isis, 69(1):100–101, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230626.

McMahon:1978:BRB

[1463] Carol E. McMahon. Book review: The History of Psychotherapy: From Healing Magic to Encounter by Jan Ehrenwald. Isis, 69(1):101–102, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230627.

Leaf:1978:BRB

[1464] Murray J. Leaf. Book review: A History of Ethnology by Fred Voget. Isis, 69(1):102–103, March 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230628.

Brieger:1978:BRB

[1465] Gert H. Brieger. Book review: Divided Legacy: A History of the Schism in Medical Thought by Harris L. Coulter. Isis, 69(1):103–105, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230629.

Bullough:1978:BRBa

[1466] Vern L. Bullough. Book review: “Doctors Wanted: No Women Need Apply”: Sexual Barriers in the Medical Profession, 1835–1975 by Mary Roth Walsh. Isis, 69(1):105, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230630.

Kreutz:1978:BRBa

[1467] Barbara M. Kreutz. Book review: On Pre-Modern Technology and Science: A Volume of Studies in Honor of Lynn White, jr. by Bert S. Hall; Delno C. West. Isis, 69(1):105–107, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230631.

Lankton:1978:BRB

[1468] Larry D. Lankton. Book review: The Archaeology of Industry by Kenneth Hudson. Isis, 69(1):107–108, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230632.

Baesel:1978:BRB

[1469] Don R. Baesel. Book review: Natural History Auctions 1700–1972. A Register of Sales in the British Isles by J. M. Chalmers-Hunt. Isis, 69(1): 108, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230633.

Gibbs:1978:BRB

[1470] Sharon Gibbs. Book review: Archaeoastronomy in Pre-Columbian America by Anthony F. Aveni. Isis, 69(1):108–109, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230634.

Saliba:1978:BRB

[1471] George Saliba. Book review: Carmen astrologicum by Dorotheus Sidonius; David Pingree. Isis, 69(1):109–110, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230635.

Multhauf:1978:BRBa

[1472] Robert P. Multhauf. Book review: Le problème des métaux dans la science antique by Robert Halleux. Isis, 69(1):110–111, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230636.

Graeser:1978:BRB

[1473] Andreas Graeser. Book review: Der Aristotelismus bei den Griechen, von Andronikos bis Alexander von Aphrodisias. Volume I: Die Renaissance des Aristotelismus im l. Jahrhundert vor Chr. by Paul Moraux. Isis, 69(1):111–112, March 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230637.

Longrigg:1978:BRBa

[1474] James Longrigg. Book review: De ventis by Theophrastus; Victor Coutant; Val L. Eichenlaub. Isis, 69(1):112–113, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230638.

Sylla:1978:BRB

[1475] Edith Sylla. Book review: The Principle of Inertia in the Middle Ages by Allan Franklin. Isis, 69(1):113–114, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230639.

McCluskey:1978:BRBa

[1476] Stephen C. McCluskey. Book review: De ortu scientiarum by Robert Kilwardby; Albert G. Judy. Isis, 69(1):114–115, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230640.

Cadden:1978:BRB

[1477] Joan Cadden. Book review: Naturkenntnisse und Naturanschauungen bei Wilhelm von Auvergne by Albrecht Quentin. Isis, 69(1):115–116, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230641.

Lindberg:1978:BRB

[1478] David C. Lindberg. Book review: An Edition of His Writings with Introductions, English Translation and Commentary by J. D. North by Richard of Wallingford; J. D. North. Isis, 69(1):116–117, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230642.

Hill:1978:BRB

[1479] Donald R. Hill. Book review: Taqi-al-Din and Arabic Mechanical Engineering. With the Sublime Methods of Spiritual Machines. An Arabic Manuscript of the Sixteenth Century by Ahmad Y. Hassan. Isis, 69(1): 117–118, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230643.

Goldstein:1978:BRB

[1480] Bernard R. Goldstein. Book review: Zur Kritik der Koordinatenüberlieferung im Sternkatalog des Almagest by Ibn as-S.alah.; Paul Kunitzsch. Isis, 69(1):118–120, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230644.

Savage-Smith:1978:BRB

[1481] Emilie Savage-Smith. Book review: Islamic Science. An Illustrated Study by Seyyed Hossein Nasr. Isis, 69(1):120–121, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230645.

Iskandar:1978:BRB

[1482] Albert Z. Iskandar. Book review: Katalog der arabischen alchemistischen Handschriften der Chester Beatty Library by Manfred Ullmann. Isis, 69 (1):121, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230646.

Major:1978:BRB

[1483] John S. Major. Book review: Science and Civilisation in China. Volume V: Chemistry and Chemical Technology. Part 2: Spagyrical Discovery and Invention: Historical Survey, from Cinnabar Elixirs to Synthetic Insulin by Joseph Needham; Ho Ping-yu; Lu Gwei-djen. Isis, 69(1):121– 122, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230647.

Moesgaard:1978:BRB

[1484] Kristian Peder Moesgaard. Book review: Les tables astronomiques de Louvain de 1528 by Henri Baers ou Vekenstyl; Emmanuel Poulle; Antoine De Smet. Isis, 69(1):122–124, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230648.

Rose:1978:BRB

[1485] Paul Lawrence Rose. Book review: La vita di Copernico di Bernardino Baldi dell’anno 1588 alla luce dei ritrovati manoscritti delle Vite dei matematici by Bronislaw Biliński. Isis, 69(1):124–125, March 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230649.

Westman:1978:BRB

[1486] Robert S. Westman. Book review: Caspar Peucer’s Library: Portrait of a Wittenberg Professor of the Mid-Sixteenth Century by Robert Kolb. Isis, 69(1):125–126, March 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230650.

Lenoir:1978:BRBa

[1487] Timothy Lenoir. Book review: Doctrine de l’angle solide. Inventaire de sa bibliothèque by Florimond de Beaune; Pierre Costabel. Isis, 69(1):126– 127, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230651.

Sloan:1978:BRB

[1488] Phillip R. Sloan. Book review: La raison et la vie: Idéal scientifique et idéologie en Allemagne de la Réforme jusqu’à Kant by Francis Courtès. Isis, 69(1):127–128, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230652.

Holder:1978:BRB

[1489] H. Hölder. Book review: Die Geographie in der “Encyclopédie.” Eine wissenschaftsgeschichtliche Studie by Johannes Dörflinger. Isis, 69(1): 128–129, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230653.

Aiton:1978:BRB

[1490] E. J. Aiton. Book review: Leonhardi Euleri Opera Omnia. Series Secunda: Opera Mechanica. Volume XX: Commentationes mechanicae et astronomicae ad scientiam navalem pertinentes by Walter Habicht; Leonhardi Euleri Opera Omnia. Series Tertia: Opera physica miscellanea epistolae. Volume IX: Commentationes opticae by Walter Habicht; Emil Alfred Fellmann. Isis, 69(1):129–130, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230654.

Gross:1978:BRB

[1491] Walter E. Gross. Book review: The Autobiography of John Fitch by Frank D. Prager. Isis, 69(1):130–131, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230655.

Allen:1978:BRBa

[1492] Virginia E. Allen. Book review: Physician Signers of the Declaration of Independence by George Gifford. Isis, 69(1):131–132, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230656.

VanHelden:1978:BRB

[1493] Albert Van Helden. Book review: Martinus Van Marum, Life and Work. Volume VI by E. Lefebvre; J. G. De Bruijn. Isis, 69(1):132–133, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230657.

Egerton:1978:BRBa

[1494] Frank N. Egerton, III. Book review: Linnaeus’s Öland and Gotland Journey, 1741 by Carl Linnaeus; Marie Åsberg; William T. Stearn. Isis, 69(1):133, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230658.

Shea:1978:BRB

[1495] William R. Shea. Book review: Isaac Newton filosofo della natura. Le lezioni giovanili di ottica e la genesi del metodo newtoniano by Maurizio Mamiami. Isis, 69(1):134, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230659.

Multhauf:1978:BRBb

[1496] Robert P. Multhauf. Book review: The Various and Ingenious Machines of Captain Agostino Ramelli by Agostino Ramelli; Martha Teach Gnudi. Isis, 69(1):134–135, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230660.

Marcil-Lacoste:1978:BRB

[1497] Louise Marcil-Lacoste. Book review: Dal sistema al senso comune. Studi sul newtonismo e gli illiministi britannici by Luigi Turco. Isis, 69(1):135– 137, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230661.

Stuewer:1978:BRB

[1498] Roger H. Stuewer. Book review: The Kind of Motion We Call Heat: A History of the Kinetic Theory of Gases in the 19th Century by Stephen G. Brush. Isis, 69(1):137–138, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230662.

Speziali:1978:BRB

[1499] Pierre Speziali. Book review: Oeuvres complètes d’Augustin Cauchy. Isis, 69(1):138–139, March 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230663.

Jenkins:1978:BRB

[1500] Reese V. Jenkins. Book review: The Birth of Photography: The Story of the Formative Years, 1800–1900 by Brian Coe. Isis, 69(1):139–140, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230664.

Sokal:1978:BRBa

[1501] Michael M. Sokal. Book review: Young John Dewey: An Essay in American Intellectual History by Neil Coughlan. Isis, 69(1):140, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230665.

Dauben:1978:BRB

[1502] Joseph W. Dauben. Book review: Richard Dedekind et les fondements des mathématiques by Pierre Dugac. Isis, 69(1):141–144, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230666.

Gutting:1978:BRBa

[1503] Gary Gutting. Book review: Knowledge and Error: Sketches on the Psychology of Enquiry by Ernst Mach; Brian McGuinness; Paul Foulkes; Thomas J. McCormack. Isis, 69(1):144–145, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230667.

Ikle:1978:BRB

[1504] Frank W. Iklé. Book review: Sir Aurel Stein, Archaeological Explorer by Jeannette Mirsky. Isis, 69(1):145–146, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230668.

Vandervliet:1978:BRB

[1505] Glenn Vandervliet. Book review: Prophetess of Health: A Study of Ellen G. White by Ronald L. Numbers. Isis, 69(1):146–147, March 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230669.

Filner:1978:BRBa

[1506] Robert E. Filner. Book review: Frédéric Joliot-Curie. A Biography by Maurice Goldsmith. Isis, 69(1):147–148, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230670.

Post:1978:BRB

[1507] John F. Post. Book review: Can Theories Be Refuted? Essays on the Duhem-Quine Thesis by Sandra G. Harding. Isis, 69(1):148–149, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230671.

MacKinnon:1978:BRB

[1508] Edward MacKinnon. Book review: Werner Heisenberg: 1901–1976 by Armin Hermann; Timothy Nevill. Isis, 69(1):149–150, March 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230672.

Kohler:1978:BRB

[1509] Robert E. Kohler. Book review: Half a Century of Medical Research by A. Landsborough Thomson. Isis, 69(1):150–151, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230673.

Jones:1978:BRB

[1510] Daniel P. Jones. Book review: A Minor Miracle. An Informal History of the National Science Foundation by Milton Lomask. Isis, 69(1):151– 152, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230674.

Keiner:1978:BRBa

[1511] Melvyn Keiner. Book review: Pioneers in Neuroendocrinology by Joseph Meites; Bernard T. Donovan; Samuel M. McCann. Isis, 69(1):152–154, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230675.

Anonymous:1978:BMa

[1512] Anonymous. Back matter. Isis, 69(1):155–158, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230676.

Anonymous:1978:FMb

[1513] Anonymous. Front matter. Isis, 69(2):159–161, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230426.

Brush:1978:NMS

[1514] Stephen G. Brush. Nettie m. Stevens and the discovery of sex determination by chromosomes. Isis, 69(2):162–172, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230427.

Schneer:1978:GTC

[1515] Cecil J. Schneer. The great taconic controversy. Isis, 69(2):173–191, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230428.

Daston:1978:BRP

[1516] Lorraine J. Daston. British responses to psycho-physiology, 1860–1900. Isis, 69(2):192–208, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230429.

Norberg:1978:SNE

[1517] Arthur L. Norberg. Simon Newcomb’s early astronomical career. Isis, 69(2):209–225, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230430.

Campbell:1978:SEC

[1518] W. A. Campbell. Some early chemical analyses of proprietary medicines. Isis, 69(2):226–233, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230431.

Stigler:1978:LEW

[1519] Stephen M. Stigler. Laplace’s early work: Chronology and citations. Isis, 69(2):234–254, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230432.

Scriba:1978:EHK

[1520] Christoph J. Scriba. Éloge: Hans Kangro, June 12, 1916–September 15, 1977. Isis, 69(2):255–256, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230433.

Schutt:1978:EBS

[1521] Hans-Werner Schütt. Éloge: Bernhard Sticker: August 2, 1906–August 30, 1977. Isis, 69(2):257–258, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230434.

Yadegari:1978:UMI

[1522] Mohammad Yadegari. The use of mathematical induction by Abū Kāmil Shujā’ Ibn Aslam (850–930). Isis, 69(2):259–262, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230435.

Ghiselin:1978:BRB

[1523] Michael T. Ghiselin. Book review: Ontogeny and Phylogeny. Isis, 69(2): 263–264, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230436.

Appel:1978:BRB

[1524] Toby A. Appel. Book review: The Spirit of System, by Richard W. Burkhardt. Lamarck and Evolutionary Biology. Isis, 69(2):265–266, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230437.

Russell:1978:BRB

[1525] J. L. Russell. Book review: Colloques d’histoire des sciences by Centre d’Histoire des Sciences et des Techniques de l’Université Catholique de Louvain; J. Ladrière; W. Voisé. Isis, 69(2):266–267, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230438.

Reingold:1978:BRB

[1526] Nathan Reingold. Book review: The Nature of Science: The History of Science in Western Culture since 1600 by David M. Knight. Isis, 69(2): 267–268, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230439.

Steneck:1978:BRB

[1527] Nicholas H. Steneck. Book review: The History of Ideas: A Bibliographical Introduction. Volume II: Medieval and Early Modern Europe by Jeremy L. Tobey. Isis, 69(2):268–269, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230440.

Koertge:1978:BRBa

[1528] Noretta Koertge. Book review: Methodological and Historical Essays in the Natural and Social Sciences by Robert S. Cohen; Marx W. Wartofsky. Isis, 69(2):269–270, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230441.

Hemmendinger:1978:BRBa

[1529] David Hemmendinger. Book review: Motion and Time, Space and Matter: Interrelations in the History of Philosophy and Science by Peter K. Machamer; Robert G. Turnbull. Isis, 69(2):270–271, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230442.

Oberdan:1978:BRB

[1530] Thomas Oberdan. Book review: Logic, Methodology and Philosophy of Science IV by Patrick Suppes; Leon Henkin; Athanese Joja; Gr. C. Moisil. Isis, 69(2):271–272, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230443.

Eklund:1978:BRB

[1531] Jon Eklund. Book review: Scales and Weights. An Historical Outline by Bruno Kisch. Isis, 69(2):272–273, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230444.

Kauffman:1978:BRB

[1532] George B. Kauffman. Book review: Scientific Quotations: The Harvest of a Quiet Eye by Alan L. Mackay; Maurice Ebison. Isis, 69(2):273– 274, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230445.

Field:1978:BRB

[1533] J. V. Field. Book review: The Historical Roots of Elementary Mathematics by Lucas N. H. Bunt; Phillip S. Jones; Jack D. Bedient. Isis, 69 (2):274, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230446.

Konijn:1978:BRB

[1534] H. S. Konijn. Book review: On the History of Statistcs and Probability by D. B. Owen. Isis, 69(2):275, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230447.

Rabkin:1978:BRB

[1535] Yakov M. Rabkin. Book review: Istoriia organicheskoi khimii by Georgii Vladimirovich Bykov. Isis, 69(2):275–277, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230448.

Mauskopf:1978:BRB

[1536] Seymour H. Mauskopf. Book review: Crystal Form and Structure by Cecil J. Schneer. Isis, 69(2):277–278, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230449.

Egerton:1978:BRBb

[1537] Frank N. Egerton, III. Book review: Des révolutions vertes. Histoire et principles de l’agronomie by François Dagognet. Isis, 69(2):278–279, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230450.

Brozek:1978:BRBa

[1538] Josef Brožek. Book review: The Development of Traditional Psychopathology: A Sourcebook by Mark D. Altschule; Patterns of Psychological Thought: Readings in Historical and Contemporary Texts by James R. Averill. Isis, 69(2):279–280, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230451.

Rosenberg:1978:BRB

[1539] Charles Rosenberg. Book review: The English Malady (1733) by George Cheyne; Eric T. Carison; A Treatise of the Hypochondriack and Hysterick Diseases (1730) by Bernard Mandeville; Stephen H. Good. Isis, 69(2): 280, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230452.

Triolo:1978:BRB

[1540] Victor A. Triolo. Book review: Theory and Practice in American Medicine: Historical Studies from the Journal of the History of Medicine and Allied Sciences by Gert H. Brieger. Isis, 69(2):280–281, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230453.

Stieb:1978:BRB

[1541] Ernst W. Stieb. Book review: Kremers and Urdang’s History of Pharmacy by Glenn Sonnedecker. Isis, 69(2):281–282, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230454.

Cohn-Haft:1978:BRB

[1542] Louis Cohn-Haft. Book review: The Healing Hand: Man and Wound in the Ancient World by Guido Majno. Isis, 69(2):282–283, June 1978.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230455.

Bullough:1978:BRBb

[1543] Vern L. Bullough. Book review: The Double Face of Janus and Other Essays in the History of Medicine by Owsei Temkin. Isis, 69(2):283– 284, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230456.

Rousseau:1978:BRB

[1544] G. S. Rousseau. Book review: The Royal College of Physicians of London. Portraits. Catalogue II by Gordon Wolstenholme; John F. Kerslake. Isis, 69(2):284, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230457.

Feldman:1978:BRB

[1545] Gerald D. Feldman. Book review: Moderne Technikgeschichte by Karin Hausen; Reinhard Rürup. Isis, 69(2):284–285, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230458.

Kibre:1978:BRB

[1546] Pearl Kibre. Book review: Codices Vossiani Chymici by P. C. Boeren. Isis, 69(2):285–286, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230459.

Hansen:1978:BRB

[1547] Bert Hansen. Book review: Le signe zodiacal du Scorpion dans les traditions occidentales de l’antiquité gréco-latine à la renaissance by Luigi Aurigemma. Isis, 69(2):286–287, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230460.

Potter:1978:BRBa

[1548] Paul Potter. Book review: Knidische Medizin. Teil I: Die Testimonien zur ältesten knidischen Lehre und Analysen knidischer Schriften im Corpus Hippocraticum by Hermann Grensemann. Isis, 69(2):287–289, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230461.

Longrigg:1978:BRBb

[1549] James Longrigg. Book review: The Origins of Stoic Cosmology by David E. Hahm. Isis, 69(2):289–290, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230462.

Smith:1978:BRBa

[1550] A. Mark Smith. Book review: Diocles on Burning Mirrors: The Arabic Translation of the Lost Original by G. J. Toomer. Isis, 69(2):290–292, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230463.

McVaugh:1978:BRB

[1551] Michael McVaugh. Book review: Historia social de la medicina en la España de los siglos XIII al XVI. Volumen I: La minoría musulmana y morisca by Luis García Ballester. Isis, 69(2):292, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230464.

Kren:1978:BRB

[1552] Claudia Kren. Book review: Traités sur les éclipses de soleil de 1333 et 1337 by Barlaam de Seminara; Daniel Donnet. Isis, 69(2):293, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230465.

Gingerich:1978:BRB

[1553] Owen Gingerich. Book review: On the Revolutions of the Heavenly Spheres by Copernicus; A. M. Duncan. Isis, 69(2):293–295, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230466.

Copenhaver:1978:BRB

[1554] Brian Copenhaver. Book review: Essays on the Life and Work of Thomas Linacre c. 1460–1524 by Francis Maddison; Margaret Pelling; Charles Webster. Isis, 69(2):295–297, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230467.

Finocchiaro:1978:BRBa

[1555] Maurice A. Finocchiaro. Book review: Ulisse Aldrovandi: Scienza e natura nel secondo cinquencento by Giuseppe Olmi; La formazione scientifica e il “Discorso naturale” di Ulisse Aldrovandi by Sandra Tugnoli Pattaro. Isis, 69(2):297, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230468.

Jayawardene:1978:BRBa

[1556] S. A. Jayawardene. Book review: The Italian Renaissance of Mathematics: Studies on Humanists and Mathematicians from Petrarch to Galileo by Paul Lawrence Rose. Isis, 69(2):298–300, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230469.

Frank:1978:BRB

[1557] Robert G. Frank, Jr. Book review: The Correspondence of Marcello Malpighi by Howard B. Adelmann. Isis, 69(2):300–301, June 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230470.

Truesdell:1978:BRB

[1558] C. Truesdell. Book review: Die Berliner und die Petersburger Akademie der Wissenschaften im Briefwechsel Leonhard Eulers. Teil III: Wissenschaftliche und Wissenschaftsorganisatorische Korrespondenzen 1726–1774 by A. P. Juskevic; E. Winter. Isis, 69(2):301–303, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230471.

Cannon:1978:BRB

[1559] Susan F. Cannon. Book review: Benjamin Martin: Author, Instrument-Maker, and “Country Showman” by John R. Millburn. Isis, 69(2):303– 304, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230472.

Kangro:1978:BRB

[1560] Hans Kangro. Book review: Physics circa 1900: Personnel, Funding, and Productivity of the Academic Establishments by Paul Forman; John L. Heilbron; Spencer Weart. Isis, 69(2):304–305, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230473.

Herbert:1978:BRB

[1561] Sandra Herbert. Book review: The Works of Charles Darwin: An Annotated Bibliographical Handlist by R. B. Freeman. Isis, 69(2):305–306, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230474.

Bynum:1978:BRB

[1562] W. F. Bynum. Book review: Antivivisection and Medical Science in Victorian Society by Richard French. Isis, 69(2):306–307, June 1978.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230475.

Mortimer:1978:BRB

[1563] Joanne S. Mortimer. Book review: Prozesse wissenschaftlicher Differenzierung. Eine organisations- und wissenschaftssoziologische Fallstudie by Rainald von Gizycki. Isis, 69(2):307, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230476.

Ostrowska:1978:BRB

[1564] Teresa Ostrowska. Book review: Historia patologii w XIX wieku by Romuald Wieslaw Gutt. Isis, 69(2):307–309, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230477.

Gutting:1978:BRBb

[1565] Gary Gutting. Book review: Method and Appraisal in the Physical Sciences: The Critical Background to Modern Science, 1800–1905 by Colin Howson. Isis, 69(2):309–310, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230478.

Daub:1978:BRB

[1566] Edward E. Daub. Book review: Applications of Energy: Nineteenth Century by R. Bruce Lindsay. Isis, 69(2):310–311, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230479.

Multhauf:1978:BRBc

[1567] Robert P. Multhauf. Book review: Seiyu roku: On Oil Manufacturing by Okura Nagatsune; Eiko Ariga; Carter Litchfield. Isis, 69(2):311, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230480.

Greene:1978:BRB

[1568] John C. Greene. Book review: Creation by Natural Law: Laplace’s Nebular Hypothesis in American Thought by Ronald L. Numbers. Isis, 69 (2):312, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230481.

Manier:1978:BRB

[1569] Edward Manier. Book review: Darwin in America: The Intellectual Response, 1865–1912 by Cynthia Eagle Russett. Isis, 69(2):313–314, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230482.

Egerton:1978:BRBc

[1570] Frank N. Egerton, III. Book review: Last of the Naturalists: The Career of C. Hart Merriam by Keir B. Sterling. Isis, 69(2):314, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230483.

Badash:1978:BRB

[1571] Lawrence Badash. Book review: Independence and Deterrence. Britain and Atomic Energy 1945–1952. Volume I: Policy Making by Margaret Gowing; Lorna Arnold. Isis, 69(2):314–315, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230484.

Trenn:1978:BRB

[1572] Thaddeus J. Trenn. Book review: Die Jahrhundertwissenschaft: Werner Heisenberg und die Physik seiner Zeit by Armin Hermann. Isis, 69(2): 315–316, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230485.

Wessels:1978:BRB

[1573] Linda Wessels. Book review: The Uncertainty Principle and Foundations of Quantum Mechanics. A Fifty Years’ Survey by William C. Price; Seymour S. Chissick. Isis, 69(2):316–317, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230486.

Allen:1978:BRBb

[1574] Garland E. Allen. Book review: The Quarterly Review of Biology: Special Issue 50th Anniversary, 1926–1976. Isis, 69(2):317–319, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230487.

Sokal:1978:BRBb

[1575] Michael M. Sokal. Book review: Particulars of My Life by B. F. Skinner. Isis, 69(2):319–320, June 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230488.

Sopka:1978:BRBa

[1576] Katherine R. Sopka. Book review: John C. Slater: Solid-State and Molecular Theory: A Scientific Biography. Isis, 69(2):320–321, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230489.

Anonymous:1978:BMb

[1577] Anonymous. Back matter. Isis, 69(2):322–326, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230490.

Anonymous:1978:FMc

[1578] Anonymous. Front matter. Isis, 69(3):327–329, September 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231038.

Kohler:1978:WFF

[1579] Robert E. Kohler. Walter Fletcher, F. G. Hopkins, and the Dunn Institute of Biochemistry: A case study in the patronage of science. Isis, 69 (3):330–355, September 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231039.

Turner:1978:VCB

[1580] Frank M. Turner. The Victorian conflict between science and religion: A professional dimension. Isis, 69(3):356–376, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231040.

Machamer:1978:ANP

[1581] Peter K. Machamer. Aristotle on natural place and natural motion. Isis, 69(3):377–387, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231041.

Campbell:1978:QPF

[1582] Ian Campbell and David Hutchison. A question of priorities: Forbes, Agassiz, and their disputes on glacier observations. Isis, 69(3):388–399, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231042.

Wallace:1978:CFS

[1583] William A. Wallace. Causes and forces in Sixteenth-Century physics. Isis, 69(3):400–412, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231043.

Pyne:1978:MGG

[1584] Stephen J. Pyne. Methodologies for geology: G. k. Gilbert and t. c. chamberlin. Isis, 69(3):413–424, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231044.

Pickering:1978:E

[1585] George Pickering, James Longrigg, and Michael T. Walton. To the Editor. Isis, 69(3):425–426, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231045.

Anonymous:1978:Na

[1586] Anonymous. News. Isis, 69(3):427–428, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231046.

Donahue:1978:BRB

[1587] W. H. Donahue. Book review: The Ash Wednesday Supper–La Cena de le Ceneri. Isis, 69(3):429–430, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231047.

Eisele:1978:BRB

[1588] Carolyn Eisele. Book review: Complete Published Works, including Selected Secondary Materials, by Charles Sanders Peirce; Contributions to the Nation. Part One: 1869–1893 by Charles Sanders Peirce; Kenneth Laine Ketner; James Edward Cook. Microfiche Edition, Boxed. With A Comprehensive Bibliography and Index of the Published Works of Charles Sanders Peirce with a Bibliography of Secondary Studies. Isis, 69(3): 431–434, September 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231048.

Graham:1978:BRB

[1589] Loren R. Graham. Book review: 200 Nauchnykh Biografii: Bibliograficheskii spravochnik (200 Scientific Biographies: A Bibliographical Reference Book) by Z. K. Sokolovskaia. Isis, 69(3):434–435, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231049.

Brozek:1978:BRBb

[1590] Josef Brožek. Book review: Historisches Worterbuch der Philosophie by Joachim Richter; Karlfried Grunder; Die Psychologie des 20. Jahrhunderts (Psychology of the 20th Century) by Gerhard Strube; Heinrich Balmer; Dieter Eicke; Hans Zier. Isis, 69(3):435–437, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231050.

Finocchiaro:1978:BRBb

[1591] Maurice A. Finocchiaro. Book review: Towards a Rational Philosophical Anthropology by Joseph Agassi. Isis, 69(3):437–438, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231051.

Hemmendinger:1978:BRBb

[1592] David Hemmendinger. Book review: Chance, Cause, Reason: An Inquiry into the Nature of Scientific Evidence by Arthur W. Burks. Isis, 69(3): 438–439, September 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231052.

Koertge:1978:BRBb

[1593] Noretta Koertge. Book review: The Structure of Scientific Inference by Mary Hesse. Isis, 69(3):439, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231053.

Menkes:1978:BRBa

[1594] Joshua Menkes and Evan Vlachos. Book review: Science, Technology and Society: A Cross-Disciplinary Perspective by Ina Spiegel–Rösing; Derek de Solla Price. Isis, 69(3):439–441, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231054.

Berendzen:1978:BRB

[1595] Richard Berendzen. Book review: The Historical Supernovae by David H. Clark; F. Richard Stephenson. Isis, 69(3):441, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231055.

Aaboe:1978:BRB

[1596] Asger Aaboe. Book review: A History of Ancient Mathematical Astronomy by O. Neugebauer. Isis, 69(3):441–445, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231056.

Lyon:1978:BRB

[1597] John Lyon. Book review: Fossils and Progress. Paleontology and the Idea of Progressive Evolution in the Nineteenth Century by Peter J. Bowler; History of the Earth Sciences during the Scientific and Industrial Revolutions, with Special Emphasis on the Physical Geosciences by D. H. Hall. Isis, 69(3):445–446, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231057.

Theodorides:1978:BRB

[1598] Jean Théodoridès. Book review: Insects, Hygiene, and History by J. R. Busvine. Isis, 69(3):446–447, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231058.

Egerton:1978:BRBd

[1599] Frank N. Egerton, III. Book review: Dictionary of British and Irish Botanists and Horticulturists, including Plant Collectors and Botanical Artists by Ray Desmond. Isis, 69(3):447–448, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231059.

Ihde:1978:BRB

[1600] Aaron J. Ihde. Book review: The Early History of Agriculture by Joseph Hutchinson; Grahame Clark; E. M. Jope; R. Riley. Isis, 69(3):448–449, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231060.

Abir-Am:1978:BRB

[1601] Pnina Abir-Am. Book review: A Century of DNA: A History of the Discovery of the Structure and Function of the Genetic Substance by Franklin H. Portugal; Jack S. Cohen. Isis, 69(3):449–450, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231061.

Burns:1978:BRB

[1602] Chester R. Burns. Book review: Retrospectroscope: Insights into Medical Discovery by Julius H. Comroe. Isis, 69(3):451, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231062.

Maulitz:1978:BRB

[1603] Russell C. Maulitz. Book review: Essays on the History of Medicine by Saul Jarcho. Isis, 69(3):451–452, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231063.

Abrams:1978:BRB

[1604] John W. Abrams. Book review: The Masterbuilders. A History of Structural and Environmental Design from Ancient Egypt to the Nineteenth Century by Henry J. Cowan. Isis, 69(3):452–453, September 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231064.

Norberg:1978:BRB

[1605] Arthur L. Norberg. Book review: History of Technology. First Annual Volume, 1976 by A. Rupert Hall; Norman Smith. Isis, 69(3):453–454, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231065.

Schipperges:1978:BRB

[1606] Heinrich Schipperges. Book review: Alchemy: The Medieval Alchemists and Their Royal Art by Johannes Fabricius. Isis, 69(3):454, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231066.

Potter:1978:BRBb

[1607] Paul Potter. Book review: Allgemeine Krankheitsbezeichnungen im Corpus Hippocraticum. Gebrauch und Bedeutung von Nousos und Nosema by Gert Preiser. Isis, 69(3):455, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231067.

Leighton:1978:BRB

[1608] Albert C. Leighton. Book review: Medieval French Bridges: A History by Marjorie Nice Boyer. Isis, 69(3):455–456, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231068.

Kreutz:1978:BRBb

[1609] Barbara M. Kreutz. Book review: The Medieval Machine: The Industrial Revolution of the Middle Ages by Jean Gimpel. Isis, 69(3):456–457, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231069.

Hall:1978:BRB

[1610] Robert E. Hall. Book review: An Annotated Bibliography of Islamic Science. Volume I by Seyyed Hossein Nasr; William C. Chittick. Isis, 69 (3):457–461, September 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231070.

Bowden:1978:BRB

[1611] Mary Ellen Bowden. Book review: Folie et déraison à la Renaissance. Isis, 69(3):461–462, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231071.

Struik:1978:BRB

[1612] D. J. Struik. Book review: Die Werke von Jakob Bernoulli. Band III. Isis, 69(3):462–463, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231072.

Mahoney:1978:BRBa

[1613] Michael S. Mahoney. Book review: Ciencia y religión en el siglo XVII by Elías Trabulse. Isis, 69(3):463–465, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231073.

Numbers:1978:BRB

[1614] Ronald L. Numbers. Book review: Protestants in an Age of Science: The Baconian Ideal and Antebellum American Religious Thought by Theodore Dwight Bozeman. Isis, 69(3):465–466, September 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231074.

Shapiro:1978:BRB

[1615] Alan E. Shapiro. Book review: Étienne Louis Malus (1775–1812) et la théorie corpusculaire de la lumière. Traitement analytique de l’optique géométrique, polarisation de la lumière et tentative d’explication dynamique de la réflexion et de la réfraction by André Chappert. Isis, 69 (3):466–467, September 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231075.

Limoges:1978:BRB

[1616] Camille Limoges. Book review: Correspondance entre Charles Darwin et Gaston de Saporta. Précédée de “Histoire de la paléobotanique en France au XIXe siècle.“ by Yvette Conry. Isis, 69(3):467–468, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231076.

Paul:1978:BRB

[1617] Erich Robert Paul. Book review: Die Berliner Akademie der Wissenschaften in der Zeit des Imperialismus. Teil I: Von den neunziger Jahren des 19. Jahrhunderts bis zur Grossen Sozialistischen Oktoberrevolution by Conrad Grau. Isis, 69(3):469–470, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231077.

Bates:1978:BRB

[1618] Donald G. Bates. Book review: Cholera 1832: The Social Response to an Epidemic by R. J. Morris. Isis, 69(3):470–471, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231078.

Mathias:1978:BRB

[1619] Simão Mathias. Book review: Beginnings of Brazilian Science. Oswaldo Cruz, Medical Research and Policy, 1890–1920 by Nancy Stepan. Isis, 69 (3):471–473, September 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231079.

Filner:1978:BRBb

[1620] Robert E. Filner. Book review: The Spaceflight Revolution: A Sociological Study by William Sims Bainbridge. Isis, 69(3):473–474, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231080.

Kopperl:1978:BRB

[1621] Sheldon J. Kopperl. Book review: Fighting Infection: Conquests of the Twentieth Century by Harry F. Dowling. Isis, 69(3):474–475, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231081.

Olby:1978:BRB

[1622] Robert Olby. Book review: The Professor, the Institute, and DNA. Oswald T. Avery, His Life and Scientific Achievements by Rene J. Dubos. Isis, 69(3):475–476, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231082.

McMullin:1978:BRB

[1623] Ernan McMullin. Book review: The Solvay Conferences on Physics: Aspects of the Development of Physics since 1911 by Jagdish Mehra. Isis, 69(3):476–477, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231083.

Anonymous:1978:BMc

[1624] Anonymous. Back matter. Isis, 69(3):478–480, September 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231084.

Anonymous:1978:FMd

[1625] Anonymous. Front matter. Isis, 69(4):481–483, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231086.

Multhauf:1978:E

[1626] Robert P. Multhauf. Editorial. Isis, 69(4):484–487, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231087.

Westfall:1978:NP

[1627] Richard S. Westfall. A note from the President. Isis, 69(4):488, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231088.

Burnyeat:1978:PST

[1628] M. F. Burnyeat. The philosophical sense of Theaetetus’ mathematics. Isis, 69(4):489–513, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231089.

Lorch:1978:DNN

[1629] Jacob Lorch. The discovery of nectar and nectaries and its relation to views on flowers and insects. Isis, 69(4):514–533, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231090.

Dauben:1978:GCP

[1630] Joseph W. Dauben. Georg cantor: The personal matrix of his mathematics. Isis, 69(4):534–550, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231091.

Grant:1978:PIB

[1631] Edward Grant. The principle of the impenetrability of bodies in the history of concepts of separate space from the middle ages to the seventeenth century. Isis, 69(4):551–571, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231092.

Woodward:1978:AGF

[1632] William R. Woodward. From association to gestalt: The fate of Hermann Lotze’s theory of spatial perception, 1846–1920. Isis, 69(4):572–582, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231093.

Rousseau:1978:LSS

[1633] G. S. Rousseau. Literature and science: The state of the field. Isis, 69 (4):583–591, December 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231094.

Wilson:1978:EVA

[1634] Leonard G. Wilson. Éloge: Victor Ambrose Eyles, 1895–1978. Isis, 69 (4):592–594, December 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231095.

Rocke:1978:GLD

[1635] Alan J. Rocke. Gay-lussac and dumas: Adherents of the Avogadro– Ampère hypothesis? Isis, 69(4):595–600, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231096.

Anonymous:1978:Nb

[1636] Anonymous. News. Isis, 69(4):601, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231097.

Finocchiaro:1978:BRBc

[1637] Maurice A. Finocchiaro. Book review: Perception, Theory and Commitment: The New Philosophy of Science. Isis, 69(4):602–604, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231098.

Burian:1978:BRB

[1638] Richard M. Burian. Book review: Progress and Its Problems: Towards a Theory of Scientific Growth by Larry Laudan. Isis, 69(4):604–606, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231099.

Schlagel:1978:BRB

[1639] Richard H. Schlagel. Book review: The Concepts of Space and Time: Their Structure and Their Development by Milic Čapek. Isis, 69(4):607, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231100.

Forman:1978:BRB

[1640] Paul Forman. Book review: Wissenschaftsentwicklung und Wissenschaftssteuerung. Einführung und Material zur Wissenschaftsforschung by Ina-Susanne Spiegel–Rösing. Isis, 69(4):607–608, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231101.

Dam:1978:BRB

[1641] Cecil F. Dam. Book review: The History of Meteorology: To 1800 by H. Howard Frisinger. Isis, 69(4):608–609, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231102.

Jayawardene:1978:BRBb

[1642] S. A. Jayawardene. Book review: Introduction bibliographique à l’histoire de la biologie. Isis, 69(4):609–611, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231103.

Keiner:1978:BRBb

[1643] Melvyn Keiner. Book review: Advances in American Medicine: Essays at the Bicentennial by John Z. Bowers; Elizabeth F. Purcell. Isis, 69 (4):611–613, December 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231104.

Menkes:1978:BRBb

[1644] Joshua Menkes. Book review: Research in Philosophy and Technology: An Annual Compilation of Research by Paul T. Durbin; Carl Mitcham. Isis, 69(4):613–614, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231105.

Plotkin:1978:BRB

[1645] Howard Plotkin. Book review: Retrospective Technology Assessment– 1976 by Joel A. Tarr. Isis, 69(4):614–615, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231106.

McCluskey:1978:BRBb

[1646] Stephen C. McCluskey. Book review: Native American Astronomy by Anthony F. Aveni. Isis, 69(4):615–616, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231107.

Weisheipl:1978:BRB

[1647] James A. Weisheipl. Book review: Speculum astronomiae by Alberto Magno; S. Caroti; M. Periera; S. Zamponi; Paola Zambelli. Isis, 69 (4):616–618, December 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231108.

Smith:1978:BRBb

[1648] A. Mark Smith. Book review: The Translation of the Elements of Euclid from the Arabic into Latin by Hermann of Carinthia (?), Books VII-XII by H. L. L. Busard. Isis, 69(4):618–619, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231109.

Toomer:1978:BRB

[1649] G. J. Toomer. Book review: The Exhaustive Treatise on Shadows by Abu al-Rayh.an Muh.ammad b. Ah.mad al-Biruni; E. S. Kennedy. Isis, 69 (4):619–622, December 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231110.

Atiyeh:1978:BRB

[1650] George N. Atiyeh. Book review: Les manuscrits médicaux et pharmaceutiques dans les bibliothèques publiques d’Alep by Salmane Kataye. Isis, 69 (4):622–623, December 1978. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231111.

Mahoney:1978:BRBb

[1651] Michael S. Mahoney. Book review: Il pitagorismo di Niccolò Copernico by Bronislaw Biliński. Isis, 69(4):623–624, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231112.

Schmitt:1978:BRB

[1652] Charles B. Schmitt. Book review: Girolamo Cardano (1501–1576). Arzt, Naturphilosoph, Mathematiker, Astronom und Traumdeuter by Markus Fierz. Isis, 69(4):624–625, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231113.

Duncan:1978:BRB

[1653] A. M. Duncan. Book review: De revolutionibus libri sex by Nicolaus Copernicus. Isis, 69(4):625–626, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231114.

Kargon:1978:BRB

[1654] Robert H. Kargon. Book review: The Correspondence of Henry Oldenburg. Volume VIII: 1671–1672 by A. Rupert Hall; Marie Boas Hall; The Correspondence of Henry Oldenburg. Volume X: June 1673–April 1674. Isis, 69(4):626, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231115.

Lenoir:1978:BRBb

[1655] Timothy Lenoir. Book review: Hegel und die Chemie. Studie zur Philosophie und Wissenschaft der Natur um 1800 by Dietrich von Engelhardt. Isis, 69(4):626–628, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231116.

Albury:1978:BRB

[1656] William Randall Albury. Book review: A System of Complete Medical Police: Selections from Johann Peter Frank by Johann Peter Frank; Erna Lesky. Isis, 69(4):628–629, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231117.

Brush:1978:BRB

[1657] Stephen G. Brush. Book review: Briefwechsel zwischen Alexander von Humboldt und Carl Friedrich Gauss by Kurt-R. Biermann. Isis, 69(4): 629, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231118.

Glick:1978:BRB

[1658] Thomas F. Glick. Book review: Charles Darwin: “Storia” ed “economia” della natura by Giuliano Pancaldi. Isis, 69(4):629–630, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231119.

Multhauf:1978:BRBd

[1659] Robert P. Multhauf. Book review: Svedenstierna’s Tour of Great Britain, 1802–3. The Travel Diary of an Industrial Spy by Eric T. Svedenstierna; E. L. Dellow. Isis, 69(4):630, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231120.

Stanton:1978:BRB

[1660] William Stanton. Book review: Science in America since 1820 by Nathan Reingold. Isis, 69(4):631, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231121.

Althouse:1978:BRB

[1661] James W. Althouse. Book review: Schlumberger: The History of a Technique by Louis A. Allaud; Maurice H. Martin. Isis, 69(4):631–632, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231122.

Darnell:1978:BRB

[1662] Regna Darnell. Book review: Contributions to Anthropology: Selected Papers of A. Irving Hallowell. Isis, 69(4):632–633, December 1978. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231123.

Moyer:1978:BRB

[1663] Albert E. Moyer. Book review: The Physicists: The History of a Scientific Community in Modern America by Daniel J. Kevles. Isis, 69(4): 634, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231124.

Allen:1978:BRBc

[1664] Garland E. Allen. Book review: Thomas Hunt Morgan, Pioneer of Genetics by Ian Shine; Sylvia Wrobel. Isis, 69(4):635–636, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231125.

Wheaton:1978:BRB

[1665] Bruce R. Wheaton. Book review: The Compton Effect: Turning Point in Physics by Roger Stuewer. Isis, 69(4):636–638, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231126.

Anonymous:1978:BMd

[1666] Anonymous. Back matter. Isis, 69(4):639–640, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231127.

Anonymous:1978:VI

[1667] Anonymous. Volume information. Isis, 69(4):641–660, December 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231085.

Anonymous:1978:FMe

[1668] Anonymous. Front matter. Isis, 69(5):1–4, ???? 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230251.

Neu:1978:OHT

[1669] John Neu. One hundred third critical bibliography of the history of science and its cultural influences (to January 1978). Isis, 69(5):5–15+17– 137+139–163+165–183, ???? 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230252.

Anonymous:1979:FMa

[1670] Anonymous. Front matter. Isis, 70(1):1–5, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230873.

Thackray:1979:BFG

[1671] Arnold Thackray. Le but d’Isis. (French) [the goal of Isis]. Isis, 70(1): 6–10, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230874.

Bogaard:1979:HWA

[1672] Paul A. Bogaard. Heaps or wholes: Aristotle’s explanation of compound bodies. Isis, 70(1):11–29, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230875.

Breiner:1979:CC

[1673] Laurence A. Breiner. The career of the cockatrice. Isis, 70(1):30–47, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230876.

Tamny:1979:NCP

[1674] Martin Tamny. Newton, creation, and perception. Isis, 70(1):48–58, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230877.

Guralnick:1979:CFE

[1675] Stanley M. Guralnick. The contexts of Faraday’s electrochemical laws. Isis, 70(1):59–75, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230878.

Illy:1979:AEP

[1676] József Illy. Albert Einstein in Prague. Isis, 70(1):76–84, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230879.

Stothers:1979:AA

[1677] Richard Stothers. Ancient aurorae. Isis, 70(1):85–95, March 1979. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230880.

French:1979:JMR

[1678] R. K. French. De juvamentis membrorum and the reception of galenic physiological anatomy. Isis, 70(1):96–109, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230881.

Christianson:1979:TBG

[1679] J. R. Christianson and Tycho Brahe. Tycho Brahe’s German treatise on the comet of 1577: A study in science and politics. Isis, 70(1):110– 140, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230882.

Kauffman:1979:ERO

[1680] George B. Kauffman. Éloge: Ralph E. Oesper, June 14, 1886–December 10, 1977. Isis, 70(1):141–143, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230883.

Schiffeler:1979:QCB

[1681] John Wm. Schiffeler. Query: Chinese brass water-spouting bowl. Isis, 70(1):144–145, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230884.

Kohlstedt:1979:NP

[1682] Sally Gregory Kohlstedt, E. J. Aiton, and Michele L. Aldrich. News of the profession. Isis, 70(1):146–151, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230885.

Lenoir:1979:BRBa

[1683] Timothy Lenoir. Book review: The Sociology of Science in Europe. Isis, 70(1):152–153, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230886.

Forman:1979:BRBa

[1684] Paul Forman. Book review: Wissenschaftspolitik im Wilhelminischen Deutschland, by Lothar Burchardt. Vorgeschichte, Gründung und Aufbau der Kaiser-Wilhelm-Gesellschaft zur Förderung der Wissenschaften. Isis, 70(1):154–155, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230887.

Kleinert:1979:BRB

[1685] Andreas Kleinert. Book review: Scientists under Hitler. Politics and the Physics Community in the Third Reich by Alan D. Beyerchen. Isis, 70(1):156–157, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230888.

Kevles:1979:BRB

[1686] Daniel J. Kevles. Book review: Sputnik, Scientists, and Eisenhower: A Memoir of the First Special Assistant to the President for Science and Technology by James R. Killian. Isis, 70(1):157–158, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230889.

Bailes:1979:BRB

[1687] Kendall E. Bailes. Book review: The Technological Level of Soviet Industry by R. Amann; J. M. Cooper; R. W. Davies. Isis, 70(1):158–159, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230890.

Multhauf:1979:BRBa

[1688] Robert P. Multhauf. Book review: Essays and Papers in the History of Modern Science by Henry Guerlac. Isis, 70(1):160, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230891.

Jayawardene:1979:BRBa

[1689] S. A. Jayawardene. Book review: Isis Cumulative Bibliography. A Bibliography of the History of Science Formed from Isis Critical Bibliographies 1–90, 1913–65. Volume III: Subjects by Magda Whitrow. Isis, 70(1):160– 163, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230892.

Finocchiaro:1979:BRBa

[1690] Maurice A. Finocchiaro. Book review: Le origini moderne della storiografia filosofica by Maria Assunta Del Torre. Isis, 70(1):163–164, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230893.

Maziarz:1979:BRB

[1691] Edward A. Maziarz. Book review: Graph Theory: 1736–1936 by N. L. Biggs; E. K. Lloyd; R. J. Wilson. Isis, 70(1):164–165, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230894.

Callery:1979:BRB

[1692] Bernadette G. Callery. Book review: An Illustrated History of the Herbals by Frank J. Anderson. Isis, 70(1):165–166, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230895.

George:1979:BRB

[1693] Wilma George. Book review: Zoological Illustration: An Essay towards a History of Printed Zoological Pictures by David Knight. Isis, 70(1):166– 167, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230896.

Egerton:1979:BRBa

[1694] Frank N. Egerton, III. Book review: Nature’s Economy: The Roots of Ecology by Donald Worster. Isis, 70(1):167–168, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230897.

Sokal:1979:BRB

[1695] Michael M. Sokal. Book review: The History of Psychology and the Behavioral Sciences: A Bibliographic Guide by Robert I. Watson. Isis, 70(1):168–169, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230898.

Theodorides:1979:BRBa

[1696] Jean Théodoridés. Book review: Edward Jenner’s Cowpox Vaccine: The History of a Medical Myth by Peter Razzell. Isis, 70(1):169, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230899.

Reiser:1979:BRB

[1697] Stanley Joel Reiser. Book review: The Development of Anaesthetic Apparatus. A History Based on the Charles King Collection of the Association of Anaesthetists of Great Britain and Ireland by K. Bryn Thomas. Isis, 70(1):170, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230900.

Mayr:1979:BRBa

[1698] Otto Mayr. Book review: Die deutsche Räderuhr: Zur Kunst und Technik des mechanischen Zeitmessers im deutschen Sprachraum. Volume I by Klaus Maurice. Isis, 70(1):170–171, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230901.

Reingold:1979:BRBa

[1699] Nathan Reingold. Book review: America by Design: Science, Technology, and the Rise of Corporate Liberalism by David F. Noble. Isis, 70 (1):171–173, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230902.

Beaver:1979:BRB

[1700] Donald deB. Beaver. Book review: Greek and Roman Sundials by Sharon L. Gibbs. Isis, 70(1):173–175, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230903.

Sivin:1979:BRB

[1701] N. Sivin. Book review: La tradition scientifique chinoise by Joseph Needham. Isis, 70(1):175, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230904.

Solnick:1979:BRB

[1702] Bruce B. Solnick. Book review: Historia natural y moral de las Indias by José de Acosta. Isis, 70(1):176, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230905.

Mahoney:1979:BRBa

[1703] Michael S. Mahoney. Book review: Prolegomena alle Vite dei matematici di Bernardino Baldi (1587–1596). Manoscritti Rosminiani-Celli già Albani-Boncompagni by Bronislaw Biliński. Isis, 70(1):176–177, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230906.

Kupsch:1979:BRB

[1704] W. O. Kupsch. Book review: Albrecht von Haller by Heinz Balmer. Isis, 70(1):177–178, March 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230907.

Jaki:1979:BRB

[1705] Stanley L. Jaki. Book review: Le savant et philosophe mulhousien Jean-Henri Lambert (1728–1777): Etudes critiques et documentaires by Roger Jaquel. Isis, 70(1):178, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230908.

Beddall:1979:BRB

[1706] Barbara G. Beddall. Book review: Darwin and His Flowers: The Key to Natural Selection by Mea Allan. Isis, 70(1):179, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230909.

Risse:1979:BRBa

[1707] Guenter B. Risse. Book review: La anatomía española en la primera mitad del siglo XIX by Juan Arechaga Martínez. Isis, 70(1):179–180, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230910.

Koppelman:1979:BRB

[1708] Elaine Koppelman. Book review: Vermischte Philosophische und Physikalische Schriften 1832–1848 by Bernard Bolzano. Isis, 70(1):180– 181, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230911.

Rocke:1979:BRBa

[1709] Alan J. Rocke. Book review: Naturwissenschaft, Staat und Industrie in Baden (1848–1914) by Peter Borscheid. Isis, 70(1):181–182, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230912.

Maulitz:1979:BRBa

[1710] Russell C. Maulitz. Book review: Organized Medicine in the Progressive Era by James G. Burrow. Isis, 70(1):182–183, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230913.

Sheets-Pyenson:1979:BRB

[1711] Susan Sheets-Pyenson. Book review: An Annotated Calendar of the Letters of Charles Darwin in the Library of the American Philosophical Society by P. Thomas Carroll. Isis, 70(1):183, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230914.

Lankford:1979:BRB

[1712] Nelson D. Lankford. Book review: Sir John Richardson, Arctic Explorer, Natural Historian, Naval Surgeon by Robert E. Johnson. Isis, 70(1):183– 184, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230915.

Adams:1979:BRB

[1713] Mark B. Adams. Book review: Pis’ma (1863–1916 gg.) [Letters (1863– 1916)] by I. I. Mechnikov; A. E. Gaisinovich; B. V. Lëvshin. Isis, 70(1): 184–185, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230916.

Forman:1979:BRBb

[1714] Paul Forman. Book review: Die Akademie der Wissenschaften der DDR. Ein Beitrag zu ihrer Geschichte by Werner Hartkopf. Isis, 70(1):185, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230917.

Wolfenstein:1979:BRB

[1715] Lincoln Wolfenstein. Book review: In at the Beginnings: A Physicist’s Life by Philip M. Morse. Isis, 70(1):185–186, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230918.

Anonymous:1979:BMa

[1716] Anonymous. Back matter. Isis, 70(1):187–189, March 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230919.

Anonymous:1979:FMb

[1717] Anonymous. Front matter. Isis, 70(2):191–195, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230786.

Hirsh:1979:RGN

[1718] Richard F. Hirsh. The riddle of the gaseous nebulae. Isis, 70(2):196– 212, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230787.

Malley:1979:DAT

[1719] Marjorie Malley. The discovery of atomic transmutation: Scientific styles and philosophies in France and Britain. Isis, 70(2):213–223, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230788.

Desmond:1979:DDR

[1720] Adrian J. Desmond. Designing the dinosaur: Richard Owen’s response to Robert Edmond Grant. Isis, 70(2):224–234, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230789.

Larson:1979:VFR

[1721] James L. Larson. Vital forces: Regulative principles or constitutive agents? A strategy in German physiology, 1786–1802. Isis, 70(2):235– 249, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230790.

Voigts:1979:ASP

[1722] Linda E. Voigts. Anglo-Saxon plant remedies and the Anglo-Saxons. Isis, 70(2):250–268, June 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230791.

Hill:1979:NGW

[1723] David K. Hill. A note on a Galilean worksheet. Isis, 70(2):269–271, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230792.

Roland:1979:E

[1724] Charles G. Roland. To the Editor. Isis, 70(2):271, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230793.

Anonymous:1979:NPa

[1725] Anonymous. News of the profession. Isis, 70(2):272, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230794.

Westfall:1979:BRBa

[1726] Richard S. Westfall. Book review: Galileo at Work, by Stillman Drake; Operations of the Geometric and Military Compass by Galileo Galilei; Stillman Drake. His Scientific Biography. Isis, 70(2):273–275, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230795.

Edsall:1979:BRB

[1727] John T. Edsall. Book review: Heraclitean Fire: Sketches from a Life before Nature by Erwin Chargaff. Isis, 70(2):276–277, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230796.

Jayawardene:1979:BRBb

[1728] S. A. Jayawardene. Book review: The Catalogue of the History of Science Collections of the University of Oklahoma Libraries by Duane H. D. Roller; Marcia M. Goodman. Isis, 70(2):278, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230797.

Hull:1979:BRB

[1729] David L. Hull. Book review: A Rational Animal and Other Philosophical Essays on the Nature of Man by Antony Flew. Isis, 70(2):278–279, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230798.

Kourany:1979:BRB

[1730] Janet A. Kourany. Book review: The Essential Tension: Selected Studies in Scientific Tradition and Change by Thomas S. Kuhn. Isis, 70(2):279– 280, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230799.

Bedini:1979:BRBa

[1731] Silvio A. Bedini. Book review: Images des sciences: Les anciens instruments scientifiques vus par les artistes de leur temps by Henri Michel. Isis, 70(2):280–281, June 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230800.

Larson:1979:BRB

[1732] James L. Larson. Book review: A History of Uppsala University, 1477– 1977 by Sten Lindroth; Neil Tomkinson; Jean Gray. Isis, 70(2):281– 282, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230801.

Baesel:1979:BRB

[1733] Don R. Baesel. Book review: The Naturalist in Britain: A Social History by David Elliston Allen. Isis, 70(2):282, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230802.

Bugliarello:1979:BRB

[1734] George Bugliarello. Book review: Technology as a Social and Political Phenomenon by Philip L. Bereano. Isis, 70(2):282–283, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230803.

Seidel:1979:BRB

[1735] Robert W. Seidel. Book review: Scientific Elite. Nobel Laureates in the United States by Harriet Zuckerman. Isis, 70(2):283–284, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230804.

Swinburne:1979:BRB

[1736] Richard Swinburne. Book review: Science, Chance, and Providence by Donald M. Mackay. Isis, 70(2):284–285, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230805.

DeVorkin:1979:BRB

[1737] David H. DeVorkin. Book review: Man Discovers the Galaxies by Richard Berendzen; Richard Hart; Daniel Seeley. Isis, 70(2):285–286, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230806.

Gosselin:1979:BRB

[1738] Edward A. Gosselin. Book review: This Wild Abyss: The Story of the Men Who Made Modern Astronomy by Gale E. Christianson. Isis, 70(2): 286–287, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230807.

Becker:1979:BRBa

[1739] Stanley L. Becker. Book review: From Animal Chemistry to Biochemistry by Noel G. Coley. Isis, 70(2):287, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230808.

Lindberg:1979:BRB

[1740] David C. Lindberg. Book review: Origins in Acoustics: The Science of Sound from Antiquity to the Age of Newton by Frederick Vinton Hunt. Isis, 70(2):287–288, June 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230809.

Pedersen:1979:BRB

[1741] Kurt Møller Pedersen. Book review: Kosmos af Chronos. Traek af en udviklingshistorie for vort astronomiske verdensbillede. (Episode of a History of the Development of Our Astronomical World Picture) by Kristian Peder Moesgaard. Isis, 70(2):288, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230810.

Hamlin:1979:BRB

[1742] Christopher Hamlin. Book review: Geology at MIT 1865–1965: A History of the First Hundred Years of Geology at Massachusetts Institute of Technology. Volume I: The Faculty and Supporting Staff by Robert Rakes Shrock. Isis, 70(2):288–289, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230811.

Millbrooke:1979:BRB

[1743] Anne Millbrooke. Book review: Essays on History of Geology by George W. White. Isis, 70(2):289–290, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230812.

Portugal:1979:BRB

[1744] Franklin Portugal. Book review: A History of Biochemistry: Part IV. Early Studies on Biosynthesis by Marcel Florkin. Isis, 70(2):290–291, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230813.

Egerton:1979:BRBb

[1745] Frank N. Egerton, III. Book review: An Introduction to Population Ecology by G. Evelyn Hutchinson. Isis, 70(2):291, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230814.

Burnham:1979:BRBa

[1746] John C. Burnham. Book review: R. I. Watson’s Selected Papers on the History of Psychology by Josef Brožek; Rand B. Evans. Isis, 70(2): 292, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230815.

Woodward:1979:BRB

[1747] William R. Woodward. Book review: The Origin of Consciousness in the Breakdown of the Bicameral Mind by Julian Jaynes. Isis, 70(2):292– 293, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230816.

Brozek:1979:BRB

[1748] Josef Brožek. Book review: Language, Man and Society: Foundations of the Behavioral Sciences by Robert W. Rieber; Significant Contributions to the History of Psychology, 1750–1920 by Daniel N. Robinson. Isis, 70(2):293–295, June 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230817.

Mandelbaum:1979:BRB

[1749] Dorothy Rosenthal Mandelbaum. Book review: Women in Medicine: A Bibliography of the Literature on Women Physicians by Sandra L. Chaff; Ruth Haimbach; Carol Fenichel; Nina B. Woodside. Isis, 70(2): 295–296, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230818.

Farley:1979:BRB

[1750] John Farley. Book review: Tropical Medicine and Parasitology: Classic Investigations by B. H. Kean; Kenneth E. Mott; Adair J. Russell. Isis, 70 (2):296, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230819.

Reich:1979:BRB

[1751] Leonard S. Reich. Book review: Turning Points in American Electrical History by James E. Brittain. Isis, 70(2):297, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230820.

Bedini:1979:BRBb

[1752] Silvio A. Bedini. Book review: European Clocks and Watches in the Near East by Otto Kurz. Isis, 70(2):297–298, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230821.

Becker:1979:BRBb

[1753] Stanley L. Becker. Book review: A Long, Deep Furrow: Three Centuries of Farming in New England by Howard S. Russell. Isis, 70(2):298, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230822.

Jackson:1979:BRB

[1754] Donald C. Jackson. Book review: Man and Water. A History of Hydro-Technology by Norman Smith. Isis, 70(2):299–300, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230823.

Maurice:1979:BRB

[1755] Klaus Maurice. Book review: Zur Geschichte der Löttechnik by Jochem Wolters. Isis, 70(2):300, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230824.

Bruins:1979:BRB

[1756] E. M. Bruins. Book review: In Search of Ancient Astronomies by E. C. Krupp. Isis, 70(2):300–301, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230825.

Hall:1979:BRBa

[1757] Bert S. Hall. Book review: Engineering in the Ancient World by J. G. Landels. Isis, 70(2):301–302, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230826.

Cadden:1979:BRBa

[1758] Joan Cadden. Book review: On the Elements by Marius; Richard C. Dales. Isis, 70(2):302–303, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230827.

Kren:1979:BRB

[1759] Claudia Kren. Book review: Jordanus de Nemore and the Mathematics of Astrolabes: De plana spera by Ron B. Thomson. Isis, 70(2):303, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230828.

Thomson:1979:BRB

[1760] Ron B. Thomson. Book review: Witelonis Perspectivae Liber Primus. Book I of Witelo’s Perspectiva by Sabetai Unguru. Isis, 70(2):304, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230829.

Kaplan:1979:BRBa

[1761] Barbara Beigun Kaplan. Book review: On the Causes of Symptoms. Maqalah fi Bayan ba’d al-a’rad wa-al-jawab ‘Anha. Ma’amar Hahå\underline{k}ra’ah. De causis accidentium by Moses Maimonides; J. O. Leibowitz; Shlomo Marcus; M. Beit–Arié; E. D. Goldschmidt; F. Klein-Franke; E. Lieber; M. Plessner. Isis, 70(2):304–305, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230830.

Yoke:1979:BRBa

[1762] Ho Peng Yoke. Book review: A Critical Study of the Yang Hui Suan Fa, A Thirteenth-Century Chinese Mathematical Treatise by Lam Lay Yong. Isis, 70(2):305–306, June 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230831.

Nakayama:1979:BRB

[1763] Shigeru Nakayama. Book review: Science and Civilisation in China. Volume V, Part 2: Spagyrical Discovery and Invention: Magisteries of Gold and Immortality by Joseph Needham; Lu Gwei-djen. Isis, 70(2): 306–307, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230832.

Yagi:1979:BRB

[1764] Eri Yagi. Book review: Science and Technology in East Asia by Nathan Sivin. Isis, 70(2):308, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230833.

Gingerich:1979:BRBa

[1765] Owen Gingerich. Book review: Gesamtausgabe. Band I: De revolutionibus. Faksimile des Manuskriptes by Nicolaus Copernicus; Heribert M. Nobis. Isis, 70(2):308–309, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230834.

Burke:1979:BRB

[1766] Michael E. Burke. Book review: Medicina, ciencia, y minorías marginadas: Los Moriscos by Luis García Ballester. Isis, 70(2):309– 310, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230835.

Reeds:1979:BRBa

[1767] Karen Reeds. Book review: Shakespearean Plant Names: Identifications and Interpretations by Mats Rydén. Isis, 70(2):310, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230836.

Garcia-Ballester:1979:BRB

[1768] Luis Garcia-Ballester. Book review: The Royal College of San Carlos: Surgery and Spanish Medical Reform in the Late Eighteenth Century by Michael E. Burke. Isis, 70(2):310–311, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230837.

Drake:1979:BRB

[1769] Stillman Drake. Book review: New Perspectives on Galileo by R. E. Butts; J. C. Pitt. Isis, 70(2):311–312, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230838.

Bernstein:1979:BRB

[1770] Howard R. Bernstein. Book review: Règles utiles et claires pour la direction de l’esprit et la recherche de la vérité by René Descartes. Isis, 70(2): 312–313, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230839.

Donovan:1979:BRBa

[1771] Arthur Donovan. Book review: Early American Science by Brooke Hindle. Isis, 70(2):313–314, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230840.

Langer:1979:BRB

[1772] Bernard Langer. Book review: Robert Boyle and the English Revolution: A Study in Social and Intellectual Change by J. R. Jacob. Isis, 70(2): 314–315, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230841.

Albury:1979:BRB

[1773] William Randall Albury. Book review: The Philosophy of Medicine: The Early Eighteenth Century by Lester S. King. Isis, 70(2):315, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230842.

Hujer:1979:BRB

[1774] Karel Hujer. Book review: Bibliografie Knizních Komenian 1945–1975 by Jan Amos Komensky. Isis, 70(2):315–316, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230843.

Hatch:1979:BRB

[1775] Robert A. Hatch. Book review: Cosmological Letters on the Arrangement of the World-Edifice by J. H. Lambert; Stanley L. Jaki. Isis, 70(2):316– 317, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230844.

Multhauf:1979:BRBb

[1776] Robert P. Multhauf. Book review: Correspondance du P. Marin Mersenne, Religieux Minime. Volume XIII: 1644–1645 by Marin Mersenne; Armand Beaulieu. Isis, 70(2):317, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/230845.

Burchfield:1979:BRB

[1777] Joe D. Burchfield. Book review: The Making of Geology: Earth Science in Britain, 1660–1815 by Roy Porter. Isis, 70(2):317–318, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230846.

Goodwin:1979:BRB

[1778] George Goodwin, Jr. Book review: Oak from an Acorn: A History of the American Philosophical Society Library, 1770–1803 by Murphy D. Smith. Isis, 70(2):318–319, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230847.

Bynum:1979:BRBa

[1779] W. F. Bynum. Book review: Boerhaave’s Men, at Leyden and After by E. Ashworth Underwood. Isis, 70(2):319–320, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230848.

Shklar:1979:BRB

[1780] Gerald Shklar. Book review: Der Zahnarzt in der niederländischen Malerei des 17. Jahrhunderts by Maria Elisabeth Wasserfuhr. Isis, 70(2): 320–321, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230849.

Olson:1979:BRB

[1781] Richard G. Olson. Book review: Hermeticism and the Scientific Revolution: Papers Read at a Clark Library Seminar, March 9, 1974 by Robert S. Westman; J. E. McGuire. Isis, 70(2):321, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230850.

Castrodeza:1979:BRB

[1782] Carlos Castrodeza. Book review: La introducción de la teoría celular en España by Juan Aréchaga; Guillermo Olagüe; Luis García Ballester; Mariano López Mateos. Isis, 70(2):322, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230851.

Oppenheimer:1979:BRBa

[1783] Jane M. Oppenheimer. Book review: Perepiska Karla Bera po Problemam Geografii. (Correspondence of Karl von Baer on Problems of Geography.); Karl Ber i Peterburgskaya Akademiya Nauk. (Karl von Baer and the Petersburg Academy of Sciences.); Pis’ma Karla Bera uchenym Peterburga. (Karl von Baer’s Letters to Members of the Petersburg Academy of Sciences.); Iz Epistolyarnogo Naslediya K. M. Bera v Arkhivakh Evropy. (From the Epistolary Legacy of Karl von Baer in the Archives of Europe.) by Karl Ernst von Baer; T. A. Lukina. Isis, 70(2):322–324, June 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230852.

Leavitt:1979:BRB

[1784] Judith Walzer Leavitt. Book review: The Medicine Show: Patients, Physicians and the Perplexities of the Health Revolution in Modern Society by Patricia Branca. Isis, 70(2):324–325, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230853.

Kish:1979:BRBa

[1785] George Kish. Book review: Friedrich Ratzel: Leben und Werk eines deutschen Geographen, 1844–1904 by Günther Buttmann. Isis, 70(2): 325, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230854.

Galatzer-Levy:1979:BRB

[1786] Robert M. Galatzer-Levy. Book review: Freud in Germany. Revolution and Reaction in Science 1893–1907 by Hannah S. Decker. Isis, 70(2): 325–326, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230855.

Robinson:1979:BRJ

[1787] Stephen M. Robinson. Book review: J. M. Dubbey: The Mathematical Work of Charles Babbage. Isis, 70(2):326–327, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230856.

Paul:1979:BRB

[1788] Erich Robert Paul. Book review: Scientific Materialism in Nineteenth Century Germany by Frederick Gregory. Isis, 70(2):327–328, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230857.

Bromberg:1979:BRB

[1789] Joan Bromberg. Book review: Heinrich Hertz: Memoirs, Letters, Diaries by Heinrich Hertz; Johanna Hertz; Charles Susskind; Lisa Brinner; Mathilde Hertz. Isis, 70(2):328–329, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230858.

Shafer:1979:BRB

[1790] Glenn Shafer. Book review: I. J. Bienayme: Statistical Theory Anticipated by C. C. Heyde; E. Seneta. Isis, 70(2):329, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230859.

Roberts:1979:BRB

[1791] Gerrylynn K. Roberts. Book review: Science in Victorian Manchester: Enterprise and Expertise by Robert H. Kargon. Isis, 70(2):329–330, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230860.

Rocke:1979:BRBb

[1792] Alan J. Rocke. Book review: Pflanzenchemie zwischen Lavoisier und Liebig by Reinhard Löw. Isis, 70(2):330–331, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230861.

Cannon:1979:BRB

[1793] Susan Faye Cannon. Book review: The Young Darwin and His Cultural Circle: A Study of Influences Which Helped Shape the Language and Logic of the First Drafts of the Theory of Natural Selection by Edward Manier. Isis, 70(2):331–332, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230862.

Rudwick:1979:BRB

[1794] Martin Rudwick. Book review: Henry De la Beche: Observations on an Observer by Paul J. McCartney. Isis, 70(2):332–333, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230863.

Turner:1979:BRB

[1795] R. Steven Turner. Book review: Die Hochschule auf dem Wege zum wissenschaftlichen Grossbetrieb. Die Universitat Heidelberg und das badische Hochschulwesen 1860–1914 by Reinhard Riese. Isis, 70(2):333–334, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230864.

Gale:1979:BRB

[1796] Barry G. Gale. Book review: Henslow of Hitcham. Botanist, Educationalist, Clergyman by Jean Russell-Gebbett. Isis, 70(2):334, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230865.

Myers:1979:BRB

[1797] George S. Myers. Book review: Francis Day (1829–1889) and His Collections of Indian Fishes by P. J. P. Whitehead; P. K. Talwar. Isis, 70(2): 334–335, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230866.

Ihde:1979:BRB

[1798] Aaron J. Ihde. Book review: Lucius Polk Brown and Progressive Food and Drug Control by Margaret Ripley Wolfe. Isis, 70(2):335–336, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230867.

Mayr:1979:BRBb

[1799] Ernst Mayr. Book review: Rudolf Leuckart. Weg und Werk. Biographien bedeutender Biologen. Band 2 by K[laus] Wunderlich. Isis, 70(2): 336, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230868.

Ellenberger:1979:BRB

[1800] François Ellenberger. Book review: Tectonics of Asia by Emile Argand; Albert V. Carozzi. Isis, 70(2):336–337, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230869.

Forman:1979:BRBc

[1801] Paul Forman. Book review: Correspondance. Choix de Lettres (1905– 1934) by Marie Curie; Irène Curie; Gilette Ziegler. Isis, 70(2):337–338, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230870.

Higham:1979:BRB

[1802] Robin Higham. Book review: Lunar Impact: A History of Project Ranger by R. Cargill Hall. Isis, 70(2):338, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230871.

Anonymous:1979:BMb

[1803] Anonymous. Back matter. Isis, 70(2):339–342, June 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230872.

Anonymous:1979:FMc

[1804] Anonymous. Front matter. Isis, 70(3):343–347, September 1979. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231372.

Calinger:1979:KNS

[1805] Ronald Calinger. Kant and Newtonian science: The pre-critical period. Isis, 70(3):348–362, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231373.

Hatfield:1979:SCM

[1806] Gary C. Hatfield and William Epstein. The sensory core and the Medieval foundations of early modern perceptual theory. Isis, 70(3):363–384, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231374.

Zetterberg:1979:HGJ

[1807] J. Peter Zetterberg. Hermetic geocentricity: John Dee’s celestial egg. Isis, 70(3):385–393, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231375.

Konstan:1979:PEP

[1808] David Konstan. Problems in epicurean physics. Isis, 70(3):394–418, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231376.

Tropp:1979:EKM

[1809] Henry S. Tropp. Eloge: Kenneth O. May, July 8, 1915–December 1, 1977. Isis, 70(3):419–422, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231377.

Eastwood:1979:AFE

[1810] Bruce Eastwood. al-Fārābī on extramission, intromission, and the use of platonic visual theory. Isis, 70(3):423–425, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231378.

Donovan:1979:NP

[1811] Arthur L. Donovan. News of the profession. Isis, 70(3):426–428, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231379.

Klein:1979:BRB

[1812] Martin J. Klein, Abner Shimony, and Trevor J. Pinch. Book review: Black-Body Theory and the Quantum Discontinuity, 1894–1912. Isis, 70 (3):429–440, September 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231380.

Vucinich:1979:BRB

[1813] Alexander Vucinich. Book review: Ocherki razvitiia estestvennonauchnykh i tekhnicheskikh predstavlenii na Rus v X-XVII vv. (Essays on the Development of Scientific and Technical Ideas in Russia in the 10–17th Centuries.) by V. K. Kuzakov. Isis, 70(3):441, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231381.

Barnes:1979:BRB

[1814] Barry Barnes. Book review: Varieties of Interpretation by Joseph Anthony Mazzeo. Isis, 70(3):441–442, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231382.

Bedini:1979:BRBc

[1815] Silvio A. Bedini. Book review: The Chamber of Physics. Instruments in the History of Science Collections of the Royal Swedish Academy of Sciences, Stockholm by Gunnar Pipping. Isis, 70(3):442–443, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231383.

Weingart:1979:BRB

[1816] Peter Weingart. Book review: Die gesellschaftliche Orientierung des wissenschaftlichen Fortschritts by Gernot, Böhme; Wolfgang van den Daele; Rainer Hohlfeld; Wolfgang Krohn; Wolf Schäfer; Tilman Spengler. Isis, 70(3):443, September 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231384.

Reingold:1979:BRBb

[1817] Nathan Reingold. Book review: Toward a Metric of Science: The Advent of Science Indicators by Yehuda Elkana; Joshua Lederberg; Robert K. Merton; Arnold Thackray; Harriet Zuckerman. Isis, 70(3):443–445, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231385.

Lakoff:1979:BRB

[1818] Sanford A. Lakoff. Book review: Scientists and World Order: The Uses of Technical Knowledge in International Organizations by Ernst B. Haas; Mary Pat Williams; Don Babai. Isis, 70(3):445–446, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231386.

Woolf:1979:BRB

[1819] Patricia Woolf. Book review: Perspectives on the Emergence of Scientific Disciplines by Gérard Lemaine; Roy MacLeod; Michael Mulkay; Peter Weingart. Isis, 70(3):446–447, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231387.

Lenoir:1979:BRBb

[1820] Timothy Lenoir. Book review: Wissenschaftssoziologie. Studien und Materialien by Nico Stehr; René König. Isis, 70(3):447–448, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231388.

Reeds:1979:BRBb

[1821] Karen Reeds. Book review: On Aesthetics in Science by Judith Wechsler. Isis, 70(3):448–449, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231389.

Heffner:1979:BRB

[1822] John Heffner. Book review: Studies in Perception: Interrelations in the History and Philosophy of Science by Peter K. Machamer; Robert G. Turnbull. Isis, 70(3):449–450, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231390.

Householder:1979:BRB

[1823] Alston S. Householder. Book review: A History of Numerical Analysis from the 16th through the 19th Century by Herman H. Goldstine. Isis, 70 (3):450–451, September 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231391.

Aiton:1979:BRB

[1824] E. J. Aiton. Book review: Prismata: Naturwissenschaftsgeschichtliche Studien. Festschrift für Willy Hartner by Y. Maeyama; W. G. Saltzer. Isis, 70(3):451–452, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231392.

Gingerich:1979:BRBb

[1825] Owen Gingerich. Book review: Supplement to the Catalogue of the Crawford Library of the Royal Observatory Edinburgh by Mary F. I. Smyth; Michael J. Smyth. Isis, 70(3):452, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231393.

Hutchinson:1979:BRB

[1826] G. E. Hutchinson. Book review: Survivals of Greek Zoological Illuminations in Byzantine Manuscripts by Zoltán Kádár. Isis, 70(3):452–453, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231394.

Eyler:1979:BRB

[1827] John M. Eyler. Book review: A Social History of Medicine by Frederick F. Cartwright. Isis, 70(3):453–454, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231395.

Theodorides:1979:BRBb

[1828] J. Théodoridès. Book review: The Virus: A History of the Concept by Sally Smith Hughes. Isis, 70(3):454–455, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231396.

Pitts:1979:BRB

[1829] John A. Pitts. Book review: Epidemics: The Story of Mankind’s Most Lethal and Elusive Enemies–From Ancient Times to the Present by Geoffrey Marks; William K. Beatty. Isis, 70(3):455–456, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231397.

Davis:1979:BRB

[1830] Audrey B. Davis. Book review: Medicine and the Reign of Technology by Stanley Joel Reiser. Isis, 70(3):456–457, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231398.

Risse:1979:BRBb

[1831] Guenter B. Risse. Book review: Konzepte der Medizin in Vergangenheit und Gegenwart by Karl E. Rothschuh; Iatromagie: Begriff, Merkmale, Motive, Systematik by Karl E. Rothschuh. Isis, 70(3):457–458, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231399.

Avrith:1979:BRB

[1832] Gale Avrith. Book review: Histoire des techniques by Bertrand Gille. Isis, 70(3):458, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231400.

Donovan:1979:BRBb

[1833] Arthur Donovan. Book review: The Steam Engine of Thomas Newcomen by L. T. C. Rolt; J. S. Allen; Steam Power and British Industrialization to 1860 by G. N. von Tunzelmann. Isis, 70(3):459–460, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231401.

Friedel:1979:BRB

[1834] Robert Friedel. Book review: Perspectives on Technology by Nathan Rosenberg. Isis, 70(3):460–461, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231402.

Gibbs:1979:BRB

[1835] Sharon Gibbs. Book review: The Stone Circles of the British Isles by Aubrey Burl; Science and Society in Prehistoric Britain by Euan W. MacKie. Isis, 70(3):461–462, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231403.

McCullough:1979:BRB

[1836] E. J. McCullough. Book review: Philo of Byzantium. Pneumatica. The First Treatise on Experimental Physics: Western Version and Eastern Version by Frank David Prager. Isis, 70(3):462–463, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231404.

Annas:1979:BRBa

[1837] Julia Annas. Book review: Aristotle on Mind and the Senses. Proceedings of the Seventh Symposium Aristotelicum by G. E. R. Lloyd; G. E. L. Owen. Isis, 70(3):463, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231405.

Kaplan:1979:BRBb

[1838] Barbara Beigun Kaplan. Book review: Marbode of Rennes’ (1035–1123) “De Lapidibus” by John M. Riddle. Isis, 70(3):463–464, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231406.

Saliba:1979:BRB

[1839] George Saliba. Book review: Le livre du qarast.un de \underline{T}abit Ibn Qurra: Étude sur l’origine de la notion de travail et du calcul du moment statique d’une barre homogène by Khalil Jaouiche. Isis, 70(3):464–465, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231407.

Yoke:1979:BRBb

[1840] Ho Peng Yoke. Book review: Pacing the Void: T’ang Approaches to the Stars by Edward H. Schafer. Isis, 70(3):465–466, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231408.

Tanner:1979:BRB

[1841] R. C. H. Tanner. Book review: Einführung in die Neue Algebra by François Viète; Karin Reich; Helmuth Gericke. Isis, 70(3):466, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231409.

Kreiling:1979:BRB

[1842] Frederick C. Kreiling. Book review: Gottfried Wilhelm Leibniz by Ronald Calinger; Bernard Wunderlich; Caryn Wunderlich. Isis, 70(3):466–467, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231410.

Finocchiaro:1979:BRBb

[1843] Maurice A. Finocchiaro. Book review: Metafisica del senso e scienze della vita: Tommaso Cornelio by Franco Crispini. Isis, 70(3):467–468, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231411.

Mahoney:1979:BRBb

[1844] Michael S. Mahoney. Book review: Register zu Gottfried Wilhelm Leibniz: Mathematische Schriften und Der Briefwechsel mit Mathematikern (herausgegeben von C. I. Gerhardt) by Joseph Ehrenfried Hofmann. Isis, 70(3):468, September 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231412.

Leicester:1979:BRBa

[1845] Henry M. Leicester. Book review: Lomonosov. Statei i materialov. VII. (Articles and Materials). Isis, 70(3):468–469, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231413.

Oppenheimer:1979:BRBb

[1846] Jane M. Oppenheimer. Book review: I Giornali delle Sperienze e Osservazioni Relativi alla Fisiologia della Generazione e alla Embriologia Sperimentale by Lazzaro Spallanzani; C. Castellani; V. G. Leone. Isis, 70(3):469, September 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231414.

Fellmann:1979:BRB

[1847] E. A. Fellmann. Book review: Geschichte der mechanischen Prinzipien und ihrer wichtigsten Anwendungen by István Szabó. Isis, 70(3):469–471, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231415.

Berkeley:1979:BRB

[1848] Edmund Berkeley. Book review: In the Presence of Nature by David Scofield Wilson. Isis, 70(3):472, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231416.

Mazumdar:1979:BRB

[1849] Pauline M. H. Mazumdar. Book review: Wilhelm Pfeffer. Apotheker, Chemiker, Botaniker, Physiologe 1845–1920 by Erwin Bunning. Isis, 70 (3):472–473, September 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231417.

Badash:1979:BRBa

[1850] Lawrence Badash. Book review: William Henry Bragg 1862–1942: Man and Scientist by G. M. Caroe. Isis, 70(3):474, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231418.

Brieger:1979:BRB

[1851] Gert H. Brieger. Book review: Charles Krafft (1863–1921), ein Pionier der Appendektomie und der Krankenpflege in Europa by Paul Gasser. Isis, 70(3):474–475, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231419.

White:1979:BRB

[1852] George W. White. Book review: The Science of Minerals in the Age of Jefferson by John C. Greene; John G. Burke. Isis, 70(3):475–476, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231420.

Ackerknecht:1979:BRB

[1853] Erwin H. Ackerknecht. Book review: Physiologie und Psychoanalyse im Leben und Werk Josef Breuers by Albrecht Hirschmuller. Isis, 70(3):476, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231421.

Craig:1979:BRB

[1854] Gordon Y. Craig. Book review: The Wernerian Theory of the Neptunian Origin of Rocks by Robert Jameson. Isis, 70(3):476–477, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231422.

Numbers:1979:BRB

[1855] Ronald L. Numbers. Book review: American Medical Education: The Formative Years, 1765–1910 by Martin Kaufman. Isis, 70(3):477, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231423.

Bowler:1979:BRB

[1856] Peter J. Bowler. Book review: T. H. Huxley: Man’s Place in Nature by James G. Paradis. Isis, 70(3):477–478, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231424.

VanHelden:1979:BRB

[1857] Albert Van Helden. Book review: Joseph von Fraunhofer: Handwerker, Forscher, Akademiemitglied, 1787–1826 by Günter D. Roth. Isis, 70(3): 478, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231425.

Daub:1979:BRB

[1858] Edward E. Daub. Book review: The Concepts and Logic of Classical Thermodynamics as a Theory of Heat Engines, Rigorously Constructed upon the Foundation Laid by S. Carnot and F. Reech by C. Truesdell; S. Bharatha. Isis, 70(3):478–479, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231426.

Knight:1979:BRB

[1859] David Knight. Book review: Album of Science: The Nineteenth Century by L. Pearce Williams. Isis, 70(3):479, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231427.

Stocking:1979:BRB

[1860] George W. Stocking, Jr. Book review: The Triumph of Evolution: American Scientists and the Heredity-Environment Controversy, 1900–1941 by Hamilton Cravens. Isis, 70(3):479–480, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231428.

Burnham:1979:BRBb

[1861] John C. Burnham. Book review: Science and Patterns of Child Care by Elizabeth M. R. Lomax; Jerome Kagan; Barbara G. Rosenkrantz. Isis, 70 (3):480–481, September 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231429.

Stevens:1979:BRBa

[1862] Rosemary A. Stevens. Book review: Almost Persuaded: American Physicians and Compulsory Health Insurance, 1912–1920 by Ronald L. Numbers. Isis, 70(3):481, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231430.

Anonymous:1979:BMc

[1863] Anonymous. Back matter. Isis, 70(3):482–486, September 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231431.

Anonymous:1979:FMd

[1864] Anonymous. Front matter. Isis, 70(4):487–491, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230718.

Cawood:1979:MCS

[1865] John Cawood. The magnetic crusade: Science and politics in early Victorian Britain. Isis, 70(4):492–518, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230719.

Rocke:1979:RCA

[1866] Alan J. Rocke. The reception of chemical atomism in Germany. Isis, 70 (4):519–536, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230720.

Pycior:1979:BPL

[1867] Helena M. Pycior. Benjamin Peirce’s linear associative algebra. Isis, 70 (4):537–551, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230721.

Drake:1979:EGR

[1868] Stillman Drake. Éloge: Guglielmo Righini, 1908–1978. Isis, 70(4):552– 554, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230722.

Unguru:1979:HAM

[1869] Sabetai Unguru. History of ancient mathematics — some reflections on the state of the art. Isis, 70(4):555–565, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230723.

Knorr:1979:MPP

[1870] Wilbur Knorr and M. F. Burnyeat. Methodology, philology, and philosophy. Isis, 70(4):565–570, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230724.

Saliba:1979:FNP

[1871] George Saliba. The first non-ptolemaic astronomy at the maraghah school. Isis, 70(4):571–576, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230725.

Gough:1979:CO

[1872] J. B. Gough. Charles the Obscure. Isis, 70(4):576–579, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230726.

Anonymous:1979:NPb

[1873] Anonymous. News of the profession. Isis, 70(4):580–583, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230727.

Farrall:1979:BRB

[1874] Lyndsay A. Farrall. Book review: Chinese Science: An Informal and Irregular Journal Dedicated to the Study of Traditional Chinese Science, Technology, and Medicine. Isis, 70(4):584–587, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230728.

Webster:1979:BRB

[1875] Charles Webster. Book review: The Chemical Philosophy: Paracelsian Science and Medicine in the Sixteenth and Seventeenth Centuries by Allen G. Debus; Man and Nature in the Renaissance by Allen G. Debus; Der sächsische Paracelsist Georg Forberger. Mit bibliographischen Beitragen zu Paracelsus, Alexander von Suchten, Denys Zacaire, Bernardus Trevirensis, Paolo Giovio, Francesco Guicciardini und Natale Conti by Rudolph Zaunick; Hans-Heinz Eulner; Kurt Goldammer. Isis, 70 (4):588–592, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230729.

Corsi:1979:BRB

[1876] Pietro Corsi. Book review: Science in Culture: The Early Victorian Period by Susan Faye Cannon. Isis, 70(4):593–595, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230730.

Jayawardene:1979:BRBc

[1877] S. A. Jayawardene. Book review: Die Mathematischen S. Marcohandschriften in Florenz by Axel Anton Björnbo. Isis, 70(4):596, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230731.

Finocchiaro:1979:BRBc

[1878] Maurice A. Finocchiaro. Book review: Empirismo e metafisica alle origini della scienza moderna by Antonina M. Alberti. Isis, 70(4):596–597, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230732.

Manier:1979:BRB

[1879] Edward Manier. Book review: What Is This Thing Called Science? An Assessment of the Nature and Status of Science and Its Methods by A. F. Chalmers. Isis, 70(4):597–598, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230733.

Rehbock:1979:BRB

[1880] Philip F. Rehbock. Book review: The Aquatic Explorers. A History of the Fisheries Research Board of Canada by Kenneth Johnstone. Isis, 70 (4):598–599, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230734.

Lubrano:1979:BRB

[1881] Linda L. Lubrano. Book review: Nauchno-tekhnicheskaia revoliutsiia by S. V. Shukhardin; V. I. Gukov. Isis, 70(4):599, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230735.

Konstan:1979:BRB

[1882] David Konstan. Book review: The Myth of Invariance: The Origin of the Gods, Mathematics and Music from the Rg Veda to Plato by Ernest G. McClain; The Pythagorean Plato: Prelude to the Song Itself by Ernest G. McClain. Isis, 70(4):599–600, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230736.

Leicester:1979:BRBb

[1883] Henry M. Leicester. Book review: Istoriya khimii belka (History of Protein Chemistry) by A. N. Shamin; Istoriya organicheskoi khimii. Otkrytie vazhneishikh organicheskikh soedinenii (History of Organic Chemistry. Discovery of the Most Important Organic Compounds) by G. V. Bykov; Ocherki po istorii organicheskoi khimii (Sketches on the History of Organic Chemistry) by G. V. Bykov. Isis, 70(4):600–601, December 1979.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230737.

Bennett:1979:BRB

[1884] J. A. Bennett. Book review: The Invention of the Telescope by Albert Van Helden. Isis, 70(4):601–602, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230738.

Davies:1979:BRB

[1885] Gordon L. Herries Davies. Book review: A Source Book in Geography by George Kish. Isis, 70(4):602–603, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230739.

Hall:1979:BRBb

[1886] Thomas S. Hall. Book review: The Pursuit of Nature: Informal Essays on the History of Physiology. … Written as Part of the Celebrations of the Centenary of the [British] Physiological Society in 1976 by A. L. Hodgkin; A. F. Huxley; W. Feldberg; W. A. H. Rushton; R. A. Gregory; R. A. McCance. Isis, 70(4):603–604, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230740.

Clark:1979:BRB

[1887] Linda Loeb Clark. Book review: Toward the Final Solution: A History of European Racism by George L. Mosse. Isis, 70(4):604, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230741.

Cadden:1979:BRBb

[1888] Joan Cadden. Book review: Birds with Human Souls: A Guide to Bird Symbolism by Beryl Rowland. Isis, 70(4):604–605, December 1979. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230742.

Farber:1979:BRB

[1889] Paul Lawrence Farber. Book review: La notion d’organisation dans l’histoire de la biologie by Joseph Schiller. Isis, 70(4):605, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230743.

Blumenthal:1979:BRB

[1890] Arthur L. Blumenthal. Book review: Problemi di teoria e storia della psicologia by Nino Dazzi; Marcello Tricarico. Isis, 70(4):605–606, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230744.

Weiner:1979:BRB

[1891] Dora B. Weiner. Book review: Le sujet de la folie: Naissance de la psychiatrie by Gladys Swain. Isis, 70(4):606, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230745.

Cowen:1979:BRB

[1892] David L. Cowen. Book review: Pharmazie und Geschichte: Festschrift für Günter Kallinich zum 65. Geburtstag by Werner Dressendörfer; Reinhard Löw; Annette Zimmermann. Isis, 70(4):607, December 1979. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230746.

Maulitz:1979:BRBb

[1893] Russell C. Maulitz. Book review: Classics in Modern Otology by Luis García–Ballester; Guillermo Olagüe; Miguel Ciges. Isis, 70(4):607–608, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230747.

Pelling:1979:BRB

[1894] Margaret Pelling. Book review: Mosquitoes, Malaria and Man: A History of the Hostilities since 1880 by Gordon Harrison. Isis, 70(4):608, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230748.

Stevens:1979:BRBb

[1895] Rosemary Stevens. Book review: Sickness and Health in America: Readings in the History of Medicine and Public Health by Judith Walzer Leavitt; Ronald L. Numbers. Isis, 70(4):608–609, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230749.

Bynum:1979:BRBb

[1896] W. F. Bynum. Book review: Historical Explorations in Medicine and Psychiatry by Hertha Riese. Isis, 70(4):609, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230750.

Annas:1979:BRBb

[1897] Julia Annas. Book review: Aristotle’s De motu animalium by Martha Craven Nussbaum. Isis, 70(4):610–611, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230751.

Folkerts:1979:BRB

[1898] Menso Folkerts. Book review: Archimedes in the Middle Ages. Volume II: The Translations from the Greek by William of Moerbeke by Marshall Clagett. Isis, 70(4):611–612, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230752.

Sylla:1979:BRB

[1899] Edith D. Sylla. Book review: Der alchemistische Traktat “Von der Multiplikation” von Pseudo-Thomas Von Aquin by Dietlinde Goltz; Joachim Telle; Hans J. Vermeer. Isis, 70(4):612, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230753.

Rankin:1979:BRB

[1900] F. K. C. Rankin. Book review: Ein italienisches Rechenbuch aus dem 14. Jahrhundert: (Columbia X 511 A13) by Kurt Vogel. Isis, 70(4):612–613, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230754.

Hansen:1979:BRB

[1901] Bert Hansen. Book review: Medieval Religion and Technology: Collected Essays by Lynn White. Isis, 70(4):613–615, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230755.

Scarborough:1979:BRB

[1902] John Scarborough. Book review: Palladius Kommentar zu Hippokrates ’De fracturis’ und seine Parallelversion unter dem Namen des Stephanus von Alexandria by Dieter Irmer. Isis, 70(4):615, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230756.

Dold-Samplonius:1979:BRB

[1903] Yvonne Dold-Samplonius. Book review: The Arithmetic of Al-Uqlidisi by Abu al-H .asan Ah.mad ibn Ibrahim al-Uqlidisi; A. S. Saidan. Isis, 70 (4):615–617, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230757.

Gingerich:1979:BRBc

[1904] Owen Gingerich. Book review: The Cosmographical Glass: Renaissance Diagrams of the Universe by S. K. Heninger. Isis, 70(4):617, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230758.

Riddle:1979:BRB

[1905] John M. Riddle. Book review: On Minerals and Mineral Products: Chapters on Minerals from His “Kreutterbuch” by Eucharius Rösslin the Younger; Johanna Schwind Belkin; Earle Radcliffe Calev. Isis, 70 (4):617–618, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230759.

Pippin:1979:BRB

[1906] Robert B. Pippin. Book review: Kant’s Theory of Science by Gordon G. Brittan. Isis, 70(4):618–619, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230760.

Taylor:1979:BRB

[1907] Kenneth L. Taylor. Book review: À l’aube de la géologie moderne: Henri Gautier (1660–1737) by François Ellenberger. Isis, 70(4):619, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230761.

Shapiro:1979:BRB

[1908] Alan E. Shapiro. Book review: The Library of Isaac Newton by John Harrison. Isis, 70(4):619–620, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230762.

Michler:1979:BRB

[1909] Markwart Michler. Book review: Albrecht von Hallers Briefe an Auguste Tissot, 1754–1777 by Erich Hintzsche. Isis, 70(4):620–621, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230763.

Shea:1979:BRB

[1910] William Shea. Book review: Contributo alla interpretazione scientifica dell’opera astronomica di Galileo by Guglielmo Righini. Isis, 70(4):621– 622, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230764.

Konvitz:1979:BRB

[1911] Josef Konvitz. Book review: The Compleat Plattmaker: Essays on Chart, Map and Globe Making in England in the Seventeenth and Eighteenth Centuries by Norman J. W. Thrower. Isis, 70(4):622–623, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230765.

Westfall:1979:BRBb

[1912] Richard S. Westfall. Book review: Newton and Newtoniana, 1672–1975. A Bibliography by Peter Wallis; Ruth Wallis. Isis, 70(4):623, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230766.

Neve:1979:BRB

[1913] Michael Neve. Book review: Social Change and Scientific Organization: The Royal Institution, 1799–1844 by Morris Berman. Isis, 70(4):623–625, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230767.

McKinney:1979:BRB

[1914] H. Lewis McKinney. Book review: The Collected Papers of Charles Darwin by Charles Darwin; Paul H. Barrett. Isis, 70(4):625, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230768.

Kish:1979:BRBb

[1915] George Kish. Book review: Elisée Reclus: Historian of Nature by Gary S. Dunbar. Isis, 70(4):625–626, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230769.

Rosenberg:1979:BRB

[1916] Charles E. Rosenberg. Book review: Edward Jarvis and the Medical World of Nineteenth-Century America by Gerald N. Grob. Isis, 70(4): 626–627, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230770.

Guralnick:1979:BRB

[1917] Stanley M. Guralnick. Book review: Science and Religion in America 1800–1860 by Herbert Hovenkamp. Isis, 70(4):627–628, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230771.

Pyenson:1979:BRB

[1918] Lewis Pyenson. Book review: Medizin, Naturwissenschaft, Technik und das Zweite Kaiserreich: Vorträge eines Kongresses vom 6. bis 11. September 1973 in Bad Nauheim by Gunter Mann; Rolf Winau. Isis, 70 (4):628–629, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230772.

Jones:1979:BRB

[1919] Charles V. Jones. Book review: The New Elements of Mathematics by Charles S. Peirce; Carolyn Eisele. Isis, 70(4):629–631, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230773.

Parssinen:1979:BRB

[1920] Terry M. Parssinen. Book review: The Medical Profession in Mid-Victorian London by M. Jeanne Peterson; Birth Control in Nineteenth-Century England by Angus McLaren. Isis, 70(4):631–632, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230774.

Bers:1979:BRB

[1921] Lipman Bers. Book review: Pafnutii L’vovich Chebyshev 1821–1894 by N. E. Prudnikov. Isis, 70(4):633, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230775.

Layton:1979:BRB

[1922] Edwin T. Layton, Jr. Book review: The Britannia Bridge: The Generation and Diffusion of Technological Knowledge by Nathan Rosenberg; Walter G. Vincenti. Isis, 70(4):633–634, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230776.

Staudenmaier:1979:BRB

[1923] John M. Staudenmaier. Book review: American Engineers of the Nineteenth Century: A Biographical Index by Christine Roysdon; Linda A.

Khatri. Isis, 70(4):634, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230777.

Allen:1979:BRB

[1924] Garland E. Allen. Book review: Eugenics and Politics in Britain, 1900– 1914 by Geoffrey Searle. Isis, 70(4):634–635, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230778.

Meikle:1979:BRB

[1925] Jeffrey L. Meikle. Book review: Hawthorne’s Mad Scientists: Pseudoscience and Social Science in Nineteenth-Century Life and Letters by Taylor Stoehr. Isis, 70(4):635–636, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230779.

Norberg:1979:BRB

[1926] Arthur L. Norberg. Book review: Astronomy Transformed: The Emergence of Radio Astronomy in Britain by David O. Edge; Michael J. Mulkay. Isis, 70(4):636–637, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230780.

Bearman:1979:BRB

[1927] David Bearman. Book review: The Excitement and Fascination of Science: Volume II: Reflections by Eminent Scientists by William C. Gibson. Isis, 70(4):637–638, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230781.

Badash:1979:BRBb

[1928] Lawrence Badash. Book review: Why I Left Canada: Reflections on Science and Politics by Leopold Infeld; Helen Infeld; Lewis Pyenson. Isis, 70(4):638, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230782.

Kasack:1979:BRB

[1929] Wolfgang Kasack. Book review: Sowjetmacht und Wissenschaft: Dokumente zur Rolle Lenins bei der Entwicklung der Akademie der Wissenschaften by Günter Kröber; Bernhard Lange. Isis, 70(4):638–639, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230783.

Lesch:1979:BRB

[1930] John E. Lesch. Book review: La vie et l’oeuvre de Frédéric Battelli (1867– 1941). L’école genevoise de physiologie de 1899 à 1941 (J. L. Prevost, F. Battelli, L. Stern) by Georges de Morsier; Marcel Monnier. Isis, 70 (4):639–640, December 1979. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230784.

Anonymous:1979:BMd

[1931] Anonymous. Back matter. Isis, 70(4):641–664, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230785.

Anonymous:1979:VI

[1932] Anonymous. Volume information. Isis, 70(4):645–663, December 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230717.

Anonymous:1979:FMe

[1933] Anonymous. Front matter. Isis, 70(5):1–4, ???? 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230065.

Neu:1979:OHF

[1934] John Neu. One hundred fourth critical bibliography of the history of science and its cultural influences (to January 1979). Isis, 70(5):5–17+19– 153+155–185+187–211, ???? 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230066.

Anonymous:1979:BMe

[1935] Anonymous. Back matter. Isis, 70(5):212, ???? 1979. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230067.

deSollaPrice:1974:GGAa

[1936] Derek de Solla Price. Gears from the Greeks. The Antikythera mechanism: a calendar computer from ca. 80 B.C. Transactions of the American Philosophical Society, New Series, 64(7):1–70, November 1974. CO-

DEN TAPSAY. ISSN 0065-9746. URL http://www.jstor.org/stable/
1006146.

Price:1974:GGAb

[1937] Derek J. de Solla (Derek John de Solla) Price. Gears from the Greeks: the Antikythera mechanism, a calendar computer from ca. 80 B.C., volume 64(7) of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1974. ISBN 0-87169-647-9. ISSN 0065-9746. 70 pp. LCCN QB107 .P74. US$5.00.